A Complete Bibliography of Publications in Communications in Statistics: Theory and

Methods: 1990–1999

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
             WWW URL: http://www.math.utah.edu/~beebe/

26 December 2021 Version 1.03

Title word cross-reference

(h, ϕ) [SMMP93]. (πps) [Wri90]. (q − 1) [YB95]. (r, k) [MVM90]. 1 [Gup92a]. 2 [CP94, Gup92a, LS94, PRP92, YC99a]. 2^{7} [Gup92a]. 2^{k−p} [GJ97, Kro94a]. 2^{m} [Gup91, Gup92a]. 2^{n} [LS97b]. 2 × 2 [BD92, Ng94]. 2 × k [Myo95]. 3^{m} [KA91, KA97a]. 4 [BMK93]. ^{m} [Yua92]. _{n} [SAP92]. _{pk} [ZSW90]. _{pmk} [Wri98]. r [Jai91]. _{S} [Hud91]. A [GP96, Kab91b, KG90, Stu91]. AR(1) [And97, Cha91a, Dia90, GCD94, Gri97, SS94a, Sil94]. AR(∞) [Kar95]. AR(p) [Lee99]. ARIMA(p, 1, 0) [SS94c]. B [Fei93, She94]. X̄ [Cos98]. β [Awa92, KK91c, SH94]. c [Her96, JS95, MN98]. C_{pk} [CP97b]. C_{pmk} [CH95a, Joh96]. χ^{2} [Dou97]. D [Cha92b, WN92, HY97, Kro94b, McM97]. E [Dut93]. ϵ [MP93]. F [CL90, DJ98, EJ93, HC94, KAKW99, KMBTG99, Leu92b, Per91, QCL98, SA96, Ter90, UNM92]. G [Bor95a, Mac92]. Γ [BR92]. h [Mac92]. IMA(1, 1) [Sil94]. j [Sal90]. K [AP91, Liu95b, CGP92, God90, Gup92a, GM96, Kar94, KM97, MJ92b, PS99a, TTG92, TP97, YB94b].

K(≥ 2) [GL90]. k = 3 [BMK93]. L [Cha93b, RL94, Ryc93, Xia95]. L_{1} [CW93, CW94b, GP90, MR95b, PM93, SG94, Shi95]. L_{2} [WC96a, ZZ99]. LU [Cho90b]. M [GL90, Kar94, Luc97b, YV95, BWCM90, CW98, DAM99, HSS93, Mar96, Wan94c]. MA(1) [HK97a]. MA(q) [SS94c]. n [FTV99, HW90]. p [BS98a, DRB97, IK97, KS98, MSL95, Tho96a, Tho96b]. P (X > Y) [RFG92]. P [X > Y] [JN94]. p^{∗} [MD94b]. ϕ [TZF97]. P P [Par93c]. Pr{Y > c|x} = p [Fuj90]. Q [CT92, Pre98]. R [HB91, Cho91b, Hua90]. R^{2} [KW93]. r × c [Gru93]. RV [CTT90]. s [Cho91b, Fin90]. σ [KKK91]. T [Cho92, Nan92, ACK95, CL90, CM95c, GC98a, HK90b, HL94, KS98, Koz94, Luc97b, MMS91, Nag97, Pos92b, PP90, SH94, Smi92a, Wan96, cL91]. T^{2} [FS99, JR91]. U [CGE99, JV92, SS92]. V [BW93]. X [SAS91a, Fuj90]. Z [Woo92, Ang91].

-and- [Mac92]. -block [PRP92]. -chart [MN98]. -class [Cha93b]. -coefficient [CTT90]. -Consistency [GP90, MR95b]. -contaminated [MP93]. -content [KK91c]. -criterion [Bor95a]. -D [YC99a]. -distribution [ACK95, Hua90, Cho92, QCL98]. -divergence [TZF97]. -divergences [Sal90]. -Efficient [GP96]. -entropies [SMMP93]. -error [KMBTG99, ZZ99]. -estimates [BWCM90, Ryc93, YV95]. -estimation [Mar96, Kar94, RL94]. -estimator [Luc97b]. -estimators [CW98, DAM99, Wan94c]. -Expectation [SH94]. -family [HL94]. -fold [YB95]. -method [Fin90]. -minimax [BR92]. -norm [SG94]. -optimal [Cha92b, Dut93, Hud91, Stu91, WN92]. -plot [CGE99]. -robust [She94]. -sample [CP94, Her96, Kar94]. -samples [KS98]. -scores [Woo92]. -sided [GL90]. -star [HY97]. -statistic [Ang91, Smi92a, Nan92]. -statistics [JV92, SA96, SS92, Xia95]. -sufficient [Awa92]. -system [Per91]. -test [Pos92b, cL91, DJ98]. -tests [UNM92]. -th [HSS93, PS99a]. -type [Ter90]. -value [DRB97, Tho96a]. -values [BS98a, Tho96b]. -variate [MSL95].

/1 [Jai91].

1 [Den91, Gov90b, Jai91]. 105 [SA93].

2SHI [HC90, HC93, HC97].

3 [RZ98]. 3-designs [RZ98].

Abel [Cha90a]. absence [SB97]. absolute [BLBE99]. absolutely [HK91, KM94a, SK91b]. absorption [New99]. accelerated [Lu90, Mad93a, Men92a, RDA92, Rao92, RBLN93, YK90]. accelerating [SH90]. acceptable [Gov90b]. acceptance [Ami96, AS94b, Fan91a, LV98, STR97, Str90]. accessible [SSCS99]. accidents [Jan98b]. account [ML91]. accounting [SM96]. accumulating [Ho98]. accuracy [Nag96, PLC98]. Accurate [HRS93, Llo95, Won95]. achieving [Gou95]. actions [BR92]. actual [CW94a]. acyclic [MAPV96]. ad [SS94d]. ad-hoc [SS94d]. Adaptation [Lai98]. adaptations [MSL95]. Adaptive [BS91, Gao97, Her96, MK90a, BH92a, DU91, GT91, HH94, HL94, KP93a, LD99a, PM90, PFU91a, Phi96]. Added [SP99]. Adding [Blo97]. additional [Kab91a]. additive [Han91, Lee99, Luc97a, SSM90, WDH90]. adjacent [TL91]. adjusted [BL90, HB90, KAKW99, PD92, SB90, WW91b]. adjusting [Ree96]. adjustment [Che98a, JB93, Oht96b, RDLT98, Sri96, SW99]. Admissibility [Cy95, YP93, Gro99, MGLF98, PF93, PN95, Tho96b]. Admissible [Kab97, KM94a, Kro97, ZK90]. adopted [Wes98]. advantages [Ste91]. affine [Xu90]. after [Ano98n, Gil93, HS90b, KH90, Oht91]. against [BB90a, BDT98, CL90, El 97, FN92, GD99, Han97b, HR98, KF97a, Kro94b, KGM94, Lis92, MT92, Mat99, Olu93, SJ90, SJ91c, SGB98, Sil94, SGD90, Szr95]. age [AW92, DR96b, Fre92]. aggregate [BC90]. aggregated [Bol90]. agreement [LWH98]. agricultural [Mar90b]. Aiala [KJ90]. AIC [Ati90]. aid [CT92]. al [HD92]. alarm [BK98]. Alcoa [SWD93]. alcohol [Moh90]. algebra [EF92]. algorithm [AL96, ACK95, MH98, Ozt91, Pai95, PC98, Ree96, TL91]. algorithms [AAB97, DG92]. alignment [FS93]. allocation [Bil97, Fan98, Kre92, Lin93, Wan91]. allocations [Zou94]. allowing [KD95, NS99, RL97]. Almost [Wan94c, AK95]. Alternate [Pro90]. Alternating [CJ98a, LP94, Pap90, PS96]. Alternative [DAM99, KA97b, LG98, MPB90, Wan93b, BS90a, BG96b, CH96, CS94a, CC95b, CS96b, Edw94, GD99, JSM98, Kro94b, Olu93, Per91, SJ90, SJ91b, SJ91c, SKS93, SD90, WH99, WB93, dlH91]. alternatives [Arn98, BB90a, FN92, Fro96, GSD96, GGG98, Han97b, HR98, He90, Her96, HY97, KGM94, MT92, NHY95, Pan96, SGB98, Szr95, TK94]. Amemiya [GT97]. among [Bel93, CFRT92, CWH93, HS90a, JS91a, Kim94c, Men97, PB93, PPS97, SC97]. Amplitude [Lop93, Kat95]. analogue [New99]. Analysing [SD95]. Analysis [BW93, FM92, GGL96, HCW91, HL93a, KP92a, LT92, PPS97, Pei90, QT98, RAP95, SSCS99, Sub92, AG96, AC99, Ang91, AB91b, BDG90, Bar95, Beh99, BR95, BY92, BF90, BS98a, BB90c, BC92, BG96b, Byr96, CR93, CTT90, CTT91, CT92, CT99, Cha92a, Cha99, CHA95, CHA97a, Che92a, Che98d, CT91, Chi99a, CK95, CBG90, Dag99, DG92, DM91, Dra90, Dra92, Dra94, Dra98, DH99, FS93, FTS94, Fre92, FMV98, Fun96a, Fun96b, FT97, Fun98, Gal94, GYH95, GJO90, GAN90, GM94b, Gra98, HV91, Hun90, HCL94, HD94, Hwa94, JH99, Jel95, JI97, JS90, Kan90, KAKW99, Kla96, KA97a, KKK90, KF97b, LW90, LTvC^{+}97, LP98a, LSR94, LD99b, Liu91, Luc92a, Mar90b, Mar95, MM96a, Men91, Mi90, ML91, MD92, Moh90, MS90b, MM96b, MK94, NPA93, NHY95]. analysis [Ng90, OT91, Pat94b, PT93, Pos92a, QSC94, RRAB93, RT96b, Rog92, Ros96, Rug91, Rug93, SS91b, Sco95a, Sco95b, SCSG94, SK97b, SCXG96, Sna94, SS99c, SLL97, SFT99, TK99a, TCT90, TC90, TWM91, TW92, Tiw94, TCSJ94, Tsi99, Tur90, VRL93, WL93, Wan98a, WS94, Wen94, WL92, Wil94, Wun91, Yip91, YB94b, vdMvdM90]. Analytic [WP92]. Analyzing [LM97, TM95, Ent90, FY95]. Anderson [And91b, RV99, Shy98, SSA90]. angle [SS90a]. angular [Law97]. angular-radial [Law97]. Animal [ML93, AZYK98, KCM94]. annotated [Ham91]. ANOVA [BK91, HK94, LC91, LLL92, SS98, WG96, Zha92]. ante [MA94]. ante-dependence [MA94]. antedependence [MA95]. antedependent [Alb92]. any [Liu95b]. applicable [Mis96]. Application [AAB97, BB90b, QL92, SAK97, AK96, Ahn92, AKP93, Ati90, BM90, Boz90, Buo90, CR93, Coa94, Dav91, Dem92, Fre92, Fri94, Mof95, MJ92b, MH96, NTS91, Pep90, Per91, Rie99, Rod92, SG90, SGN93, Str91b, TCT90, TWM91, VI91, VL92, ZZ99, Lai99]. Applications [BS99b, Che90, Bal94, BL98b, Bro90, CKH95, Cha90a, Cho92, CL98, Con90c, DS97, DM96, ESDC96, ES95, FMV98, GC98a, Gue94, GGT95, GS98b, Hoo90, JC98a, KC91a, KTTL99, Kot90, Lee96, LBB92, Liu92a, Liu93a, Luc92a, Mae98, MSMP97, Ong90, PGT98, PH92, PS93b, QCL98, RT96b, SW96, Sco99, SSS97, YB94b, ZM93b]. Applied [Dra90, Dra92, Dra94, Dra98, AHJ94, Bol90, JS90, LTvC^{+}97, Mae90, PE92, RWP95]. appraisal [MMS91]. approach [AK99, Ali90a, AC99, Arn96, BL93, BG96a, Bro95, Cer97, CS94a, CM97, CM95a, Che92a, CD98, CB94, De 92, DP97b, DM94, Edw94, EGM95, GJM96, Gou95, GPA98, Haj90, HL93a, Hoc90, HK99, Hwa92, JS93, Kar94, Kea91, KP92a, KY91b, KM94b, LY94, Liu95a, MM96a, ML91, NM96, PS95b, RK96, Rod98, Rog92, SM96, SJ97, SK99, Tiw94, Var98, Vel92, WLG99, WN97, YIK96, dP90, dlH91]. approaches [MPB90, MMPV98, Pet94, VCvdL97]. appropriate [Sob90]. Approximate [AHJ94, AC95, Fay99, JG95, KMM90, PH92, SM94, BS90c, GG92b, Mak91, vdMvdM90]. Approximating [ZM93a, FN95]. approximation [ATT99, BB90b, Cho94c, DG90, Gol90, Kat95, KY91b, Rod92, SW91, TSI98, Tsu99, VSS95]. Approximations [Mar92a, MD94a, MY91, HC93, JK91, TK97b, TK99b]. AQL [KG95]. AR-model [SK95]. arbitrarily [PL99]. arbitrary [TC90]. arch [Kun93, Kun97]. Archimedean [SY96]. area [Hug91, Isa92]. areas [DP97a]. ARIMA [Nie97]. arising [FKR91, BWCM90]. arithmetic [PL97]. arm [AFR94, HAKM93]. ARMA [CCC96, Cho91b, Cho91a, DRS90, EB98, FKY91, FKY95, Gle96, Gle98, Kum92, LS97a, Liu95a, Nas91, Pei90, PM93]. array [AK99]. arrays [Cho91b, GS98a, cH91b, Hyo94, YJF91, Yua92]. arrival [POGP95, Sin97]. arts [Moo96]. ASH [SW96]. aspects [AKP93, FR92, Iha93, SOSW94]. assays [CBRS90, SV97]. assess [US92]. Assessing [Gar96, TK94, vHL92, LW94]. Assessment [FT97, BDG90, CG92b, HL90, HL92, Jan95, LC92, MW98, WK93]. Assessments [Sob92a]. assigned [AR90]. assignment [Ste90, VI91]. assisted [CM94b]. associated [BSK95, Con94a, LWH98, Sal90, SA90b, Yam97b]. association [CLL95, Col98, CN99b, Fra91, Gup99, Kul94, Kul95, LG94, OW90, OW93, SA98]. assumed [ZK94]. assumption [Hu94, KP92b, Li99, Lin99, Pos92b, RL97]. assumptions [BDG90, Che94, YLB98, Yu97]. asymmetric [CP96b, CW98, GG93, HH94, MS90c, PF93, Rod98, RV95, TWB99b, Vän97b, VR96, YN99]. Asymptotic [Ang92, BGMT90, BR93b, BP99a, BS98b, CR92a, CM95a, CL96, CN99b, DRB97, Dik95, Don93, FMP98, FS99, Gao95a, Gao95b, GZ97, GJU94, HC93, Ino99, Iso92b, Jie90, Kar95, KB90, Kun91, KM96, LZ98a, Lin95, Luc97a, Mil91, MF97, Qum97, SMMP93, Tar90, Wan93a, WG96, Wu96, Ang91, BR91, BS99a, BS99b, CXZ90, Cha96a, CH95a, CP97b, Cho90a, Cho91b, Cho91a, CW98, Fal98, GJ92b, God92, Hen94a, Joh96, KL96, LC96a, Lou98, Mae98, Nai96, Sar97, Set90, SB95, Yan96b, Zha94a, ZZ99]. Asymptotically [GP93, LZ96b, LZZ99]. Asymptotics [GJM96, bL95, Yan98, Zha95b, Bos94, CFRT92]. at-a-time [GS95]. attraction [AG96]. attribute [KG95, RW94]. augmented [Kun97, MH98]. Australian [Moh90]. auto [BR91]. auto- [BR91]. autocorrelated [CT99, DN90, Hos90, Mae90, MPB90, OS93a, Rad97, SS94b, Wun91]. autocorrelation [Alw92, Cho90a, Sma93, SSCS99]. autocorrelations [And90, And91a, PM91b]. autocovariance [Chr97]. autopsy [AM99b]. autoregression [BLY92]. autoregressions [BC92, Luc97a]. autoregressive [AOA92, Ara90, BMM^{+}91, BR93b, Byr96, Cho90a, Cho94a, Dut99, FZ96, FM97, Gao95a, GS95, Hal96, Hon91, Li95, Mat98, Mat99, SS93a, SB95, TWB99a]. auxiliary [Ali91a, DB95a, FZ97, GM94b, PS92, TB91, Zha98]. availability [RW94]. Average [RWP95, AM93, Bor95b, BDT98, Cho94b, FN92, KF90, MMT97, SS93a, SJ91a, SW97, Sri99]. averaged [JU93, MD94a]. averaging [MAPV96].

B. [KJ91, Sar92a]. back [Rao90a, Rao92, RDA93]. Bahadur [AK94b, SS94e, Yon95]. balance [WN92]. balanced [AC95, AR90, BN91, CC90b, CS97c, CK97b, DMG91, DH95, GGO96, Goo91, Hyo94, KP95, Kim95a, PB98, PB93, PS91, RZ94, SA99, SSS97, SK91a, Sin92, SB90, SB91, SS99c, VP98, WL96, Wee92, YB95, YJF91, Yua92]. Balancing [HO90]. BAN [AN90]. BAN-estimators [AN90]. band [Ste91]. bandit [Yue99]. bands [CS95, Lin95, Mor91]. bandwidth [GP93]. Bansal [GC99]. Barr [KJ90]. Bartlett [BC97, Cor92, CF96b, CF98, KML91a]. Bartlett-type [CF96b]. based [Abd98, AF98, AH97, AAW93, AC99, ARS95, Ati90, BGP95, Boz90, Cer97, CH96, CBR91, Che99, CM98, CLL95, CN99b, CR92b, CA90b, DN92, Djo95, FKY91, FKY95, Fro93, GP90, GR90, GS90, GW94, GW99a, GT92, Hal99, HR98, HP92, HL94, HK95, HS96, Hor92, HY97, HH95, HK99, JH99, Kam99, Kan90, KS98, KKK91, KK94, Kra92, Kum97, KM93, Let97, LN90, LD99a, Li99, LZZ99, Lu90, Luc97b, LK98, MMD93, MLGD98, MK90a, MMPV96, Nai96, NM94, Ozt91, PMTP93, RS98, Rao90b, RRAB93, Rie95, RRrA91, SS99a, SR91, Sat90, Sch91, Shi99, SD95, TSI98, Tan94, TM96a, TC92, Vas96, VP98, Vos94, Wan98a, Wit90, WDH90, YT95, YK90, Zac91, ZK94, ZL91, Zog98, ZRM90]. basis [Ati96, Tho96b]. Bayes [AHJ94, BC90, Bro95, Bru95, CP90, CGP90, CP92, CP94, CP95, CP96a, CP97a, CAS97, CT91, Chi93, Dat91, Den98, GVSM92, HW90, HFP94, Hwa92, Iwa94, JP99, Kel96, iK99, Lia90, LL94, LR93, Lu93, LM94, Mar90a, MD94b, Muk94, Muk95, NS90, NS94, Oh98, PB90a, PSM96, Pep90, RL91, RT96a, Rod98, SJ91a, SC93, SC96, SUS95, Str91b, Stu91, SI97, Sug99, Tak99, Tar91, TZ90b, Ton96b, Ton96a, WG96, Yan90b, YLB98, You92, ZW99, vdMvdM90]. Bayes- [MD94b]. Bayesian [AHJ95, Ali90a, Ana97, BDG90, Bar95, BR95, BS98a, BR90b, BG96a, BC92, BG96b, BH99, CT99, CHA95, CHA97a, Che92a, Che92b, Che98d, CD98, CBG90, CB94, DP97a, Dag99, DLS97, DF94a, DY98, DP97b, DM94, Dia90, Dia91, Eva97, Fan91a, FT99, GYH95, Gen97, GAN90, GK91, GJM96, GG94b, Gri97, Gro91, HL93a, HH95, Jel95, Jel99, KP93a, LTvC^{+}97, LP98a, LLP98, Liu95a, LG93, MGR95, MAPV96, Mak91, MM96a, Men91, MW98, MD92, MGM98, MPQ91, Ng90, NAW96, NM92, PR92, PMD95, Pet94, PT93, QT98, RBC97, Rog92, Rug91, Rug93, San92, SAS91a, Sco95a, Sco95b, Sha92a, SK96, Ste92, SVM95, Tah91, Tiw94, Tsi99, Var98, Vel92, VRL93, WL93, Wan98b, WN97, YIK96, YK97, dlH91]. be [HJ90]. beaten [Tho96a]. Behavior [ACK95, AH93, Hen94a, PMSM97]. Behaviour [DRB93, DRB97]. Behnken [Bor95b]. Behrens [GW99b, LF91, NvdMM90]. Benchmarks [SM96]. Berkson [FKR91, FR93b]. Bernoulli [Hwa92, Olu94, SGS94, SKS93]. Berry [Lia94]. Best [BBK98, KWtB96, NS93, TS94, AB91a, AB94, Bof92, BD92, CZ97b, CT98, Dri91, GG94a, Gro99, HV96, LL90, LL94, Liu93b, Liu94b, Liu94a, MD93b, SK95, SS99b, TL94, YP93, ZS97]. Beta [MN93, CS90b, CMS99, Chu92, Den94, EJ93, Fan91b, Had98, KG92a, PG94, PGT98, RM90, SY93, SD97]. beta-binomial [Had98, SY93]. beta-compliance [CMS99]. better [SJ91b]. between [BS96, Bro95, CLL95, DN92, Den98, FN91, FMP98, GS90, Goo91, KMBTG99, LF93, LG94, Liu95b, Liu95a, MNZ90, MSMP97, Nag95, Ng93, Nor93, Par90, PK94, SS93b, SGS94, Tu99, US92, Vas96, WK97, Yan90b]. Bézier [KKP^{+}99]. Bhandary [GC99]. Bhattacharyya [CBR91]. BHNBUE [Han97b, SJ90]. Bias [ERO97, FZ96, GL97, MP97, PC95, TM96b, TG94, You93, AK95, Ali91b, Cor92, DS92, Dre96, Isa92, Kro97, LSR94, Mae90, Mor91, VR90, Wan96, WLG99, hZ93, ZR98]. biased [Bil97, GT90, GT92, Kej93, Llo99, MD90]. biases [CW98, Mae96]. BIBD [PS91]. bibliography [Dra90, Dra92, Dra94, Dra98, Ham91]. BIFR [SJ91c]. BIFRA [HR98]. bilinear [Che92a, Che92b, Gab92, Gab98, KB90]. binary [Buo90, Che98a, CW92a, Cza93, FWG94, Fuj94, GMR97, KP95, Kla96, Mee90, MF90, SGA99, SD97, WB97]. binary-valued [Mee90]. Binned [Chu94a]. Binomial [HB90, AOA92, SAP92, AP91, BL98b, Con90c, CF95, DF93, FS95b, Had98, HW90, IK97, Lia90, LD95, Mad93b, Mis96, MY91, New99, Olu94, Ong90, QSC94, SY93, Tap92, XD92]. bio [VP98]. bio-model [VP98]. bioassay [KML91a, KML91b]. bioassays [KY93]. bioavailability [LC92, Loc90, YR94]. bioavailability/bioequivalence [LC92]. bioequivalence [CL94, Cor90, LC92, LW94, pL98, Tsu99]. biology [JVM95]. biomedical [Zha95a]. bipartite [HY97]. Birnbaum [CT94, DY95, Rie95, Rie99]. birth [Cro96, Fre92, JVM95, KD95, PK91a]. birth-death [KD95]. bisexual [MGM98]. Bivariate [CS90b, GS98b, Ham92, MF90, AL96, AH97, AN91, Cho92, Con94b, DQS98, De 92, Fre92, GT96, HK91, HK92, Han92, Han97b, HR98, KK98b, KK90, KK91b, Koc95, KS94a, Kun93, LPR99, Lee96, LR96, Mis96, ML96, Olu94, Ong90, Ong93, RDA93, SJ90, SJ91b, SJ91c, SJ91d, SM97, SGJ95, SI97, Tah92, Tho96a, TK92, Ver96, WG95, YW98]. blended [SBS96]. Blends [HGM94]. Blinded [SL98a]. block [AC95, AC99, AR90, BN91, BY92, Cho90b, DGD92, DG97a, Dri91, GDK95, GP96, Jin97, KC98, KP95, PRP92, PS91, RRAB93, SLS96, SSS97, SK97a, Sin92, SB90, SS99c, Stu91, Wee92, YC99a, Zha92]. blocked [Pre98]. blocking [PJ99]. blocks [AAW93, CG91, CL97, Man91, Tar90]. blockwise [CG91]. blood [Moh90]. Blum [JZ96]. BLUP [MY95]. Board [Ano90a, Ano90b, Ano90c, Ano90d, Ano90e, Ano90f, Ano90g, Ano90h, Ano90i, Ano90j, Ano90k, Ano91a, Ano91b, Ano91c, Ano91d, Ano91e, Ano91f, Ano91g, Ano91h, Ano91i, Ano91j, Ano91k, Ano92c, Ano92d, Ano92e, Ano92f, Ano92g, Ano92h, Ano92i, Ano92j, Ano92k, Ano92l, Ano92m, Ano92n, Ano93a, Ano93b, Ano93c, Ano93d, Ano93e, Ano93f, Ano93g, Ano93h, Ano93i, Ano93j, Ano93k, Ano94a, Ano94b, Ano94c, Ano94d, Ano94e, Ano94f, Ano94g, Ano94h, Ano94i, Ano94j, Ano94k, Ano94l, Ano95a, Ano95b, Ano95c, Ano95d, Ano95e, Ano95f, Ano95g, Ano95h, Ano95i, Ano95j, Ano95k, Ano95l, Ano96a, Ano96b, Ano96c, Ano96d, Ano96e, Ano96f, Ano96g, Ano96h, Ano96i, Ano96j, Ano97a, Ano97b, Ano97c, Ano97d, Ano97e, Ano97f, Ano97g, Ano97h, Ano97i]. Board [Ano97j, Ano97k, Ano98b, Ano98c, Ano98d, Ano98e, Ano98f, Ano98g, Ano98h, Ano98i, Ano98j, Ano98k, Ano98l, Ano99b, Ano99c, Ano99d, Ano99e, Ano99f, Ano99g, Ano99h, Ano99i, Ano99j, Ano99k, Ano99l]. Bonferroni [Hoo90]. Bootstrap [BMM^{+}91, DG90, FM96, HK95, Mor90, BGMT90, BK97b, FS95a, FT95, Hor92, Ibr91, Jin97, KA92b, ZB93, cL91]. bootstrapped [VV97]. Bootstrapping [JV92, JGMGMR98, Lee90, Liu91, Qum94, Ver94, Ang92, DU91]. Borrowing [CM94a]. both [CHS98b, Che95b]. Bougaard [Lu90]. bound [BG95, BM93, iK99, Rug91, Yan90a]. Boundary [DS92, CZ97a, SAK97]. bounded [KS90a, Wil95]. bounding [Lav92]. Bounds [GL93, Koo93b, AHJ95, AF95, CGP92, CT94, DH90, DK95, GLL90, Hoo90, Iwa94, Lia94, Liu95b, Pap92, Rug93, Ryc93, SAS91a, Spu91, Spu95, Tak99, Tar91]. box [HC94, Bor95b, Kim96b, QL92, Yan96b]. branching [DK91, GS91b, KE93, KP93b, KKT98]. Breakdown [YV95, CR92b, MNS96, Zha96]. breaking [GP94]. Bridge [CC96]. Bridge-to-bridge [CC96]. brief [Gov95, Ham92, Yan90a]. broom [HY97]. Brownian [Bou90a, Liu97a]. BTIB [KM90]. Buckley [Hil93]. building [CL97]. Burr [AHJ94, AHJ95, GGL96, HN97, NAW96, PS99a, SAS91a, Tsa90a].

C [Wri98, ZSW90]. Calculating [Rab98]. calculation [LWH98, SAK97, TM96b]. calculations [Cha91a, DH90]. Calibration [WF95, BLS97, Dav91, KL99a, Sri95, Tak97b, TS90, WLG99]. can [DD97b, HJ90]. cancer [AKP93, CSK^{+}98]. Canonical [Ong93, Col98, Fou91, Leu94b, LL96, SKF94, SK97b, Shy98]. capability [CXZ90, CH95a, CP97b, CO90, DH99, Joh96, KKK91, KK91b, KK94, SWD93, ST93b, Sun96, Vän97a, Vän97b, Wri98, ZSW90]. capture [Gou95, GPA98, Yip96]. capture-recapture [Yip96]. carcinogenesis [Zhe99]. carcinogenicity [AZYK98, CM94c, ML93]. cardiograms [Kea96]. Carlo [KS90b, Lav92, NS90, SSSS90, SM91b]. carry [RL97]. carry-over [RL97]. cascading [Han91]. Case [Lin99, Age99, CM94a, CAS97, CZ97b, DS91, Her96, Kar95, iK99, Moh90, NvdMM90, OW93, RGU97, STR97, SA90b, Ull92, Vän97b, Wan93a, Xia95, YC99b]. case-control [OW93]. cases [Bur97, EJ93, Hwa92, KC90a, KC90b, VR96]. catastrophes [KD95]. catchability [YLB98]. categorical [BG96b, Che98d, HCW91, HAKM93, MM97, MML99, ML91, MD92, SK97b, SA98, ZB97]. categories [BR97]. categorized [GG94b, Koo93b]. Category [Dam93, Bou90b, OW93]. Category-specific [Dam93]. Cauchy [AA91, AK98]. caused [GL97]. cell [CH91a, CBRS90, FTS94, Fro96, Myo95, SS98, Zhe99]. cells [Liu94a, Mon99, UG95]. censored [Abd98, Ana92, BGP95, BE93, BDF95, CP90, CW94b, Chi99c, Coa94, DG90, DS93, EP91, FN92, Fer90, FBD97, FBD99, GG92a, GP90, GR90, Han92, Han97a, HR98, HH95, ID98, Kar99, KL99b, KSmS90, LZ98a, LY97, LTCC94, Luo94, MD92, PC98, Par93a, PB90b, PB91, PG92, PL99, Rie95, RSAJ90, Sha92a, SP99, SL94b, Tah91, TV91, Tur95, VT93, WB97, WS94, YB93, YW98]. censoring [Ano99m, BE93, CS97a, FK94, FKT95, Gas92, GK98, KY91b, LW95, Lou98, Men92a, MPQ91, PG95, PL97, TK91b, WC96b]. censorship [Her92, Her93, Stu94, Ver94]. center [MK90a, SL98a]. Centered [AA91, Sar95]. Central [HC96, Bor95a, Bor95b, EJ93, Fal98, Mae96, SKS93]. ceramics [KLH^{+}91]. Certain [Gov90a, CG90, FS95a, HC94, LBBCMR97, Rao91, SK99, TCT90]. CGMANOVA [CS97b]. chain [Ahm97a, Ber99, BR91, FT99, Gov90b, KM97, KM99a, KM99b, Raj91, SSSS94]. chains [BGMT90]. challenges [Zah90]. chance [Gar96, Sne96]. Change [JI99, Par93c, Zha93b, BS90b, Bou90a, BD93, Che98b, CS97c, Chu94b, DP97b, Fro90, GJ92b, GJM96, GL93, JW92, LP98b, ST99, VV94, WL93, Wes98, Zha95a]. change-over [CS97c]. Change-point [Zha93b, GJM96, JW92, VV94]. change-points [Chu94b]. changepoints [Wan98b]. changes [CT99, CJ98b, LLL92, Ng90]. changing [PG92, PG95]. character [BSS94, SMS93]. characteristic [CR93, GK92, KA97b, MK90a, Nai96]. characteristics [BS90a, CBRS90, DFT91]. Characterization [ALW91, GV90, Hyo94, Yua92, BH92b, GH92, Hür98, YSM92, ZRM90]. Characterizations [AK91, HL93b, Con90b, Ham92, SS91c]. Characterizing [DL95]. Charier [Nis94]. chart [Alw92, AJC99, Cha92a, DHW90, DJG94, MN98, Rad97, RAM96, Vau92b]. charts [ACR91, AH93, ARS95, AW99, Ami93, Cos98, NS96, Par98, RT96b, RWP95, Woo92, Zha95b]. check [PGC97, PGC98, RL97]. checking [Li99, SS94d]. chemical [WN92]. chi [Che95b, DS97, EJ93, KY91b, LB98, Mie91, Olu93, RG92, SBS96, Smi92a]. chi-square [DS97, EJ93, KY91b, LB98, Mie91, Olu93, RG92, SBS96]. chi-squared [Che95b, Smi92a]. Choice [ÉA95, RFG92, AG96, CM96b, DP93, LLL93, LS92, OW93, SGA99, TS95, WW90]. Cholesky [Kim94b, YC99a]. Choosing [Bel93, Bre93, Bro95, RA96, PR93, PS99b]. chow [GL93]. chronicle [Ley90]. ChSP [Gov90b]. ChSP-1 [Gov90b]. circulant [GDK95]. Circular [BP99b]. claims [Men92b]. class [Ahm97a, AH97, AAB96, BB90a, Bar95, Cha93b, CK96, CL98, Con90b, CR92b, DS97, Dut93, Gal94, Gro96, Han90, Hau91, HZ90, Hen92, Jan97, Jan98b, JA92, Kan93, Kej93, Kum97, LW95, MVM90, MR90, PS91, QL92, Rao92, RDA93, Sat91a, Sco97, SSS97, Sha90, SGB98, SSSS94, SPS94, SV97, Sob92b, SY96, TS92, Vän97a, Vän97b, VV94, YL92]. classes [MAPV96, NM92, NM94, Nor95, SP94, YLB98]. classical [PSM96, Sri95, VCvdL97]. Classification [KKK90, Leu94a, Bro95, Kim95a, KP93a, LG98, Shy98]. classified [GG94a, HCW91, WW91b]. clinical [AFR94, CMD94, Cha96b, CSK^{+}98, Dou97, Ent90, FWG94, HD94, Jai94, LRD94, LM97, Mor90, Pat94b, Ric90, Rog92, Sco99, SGJ95, SL98a, Sna94, SLL97, TY99, TM95, Wan91, Was94, Wen94]. clock [Rao90a, Rao92, RDA93]. closed [Wan98a]. closeness [BG91, Bos91, Dat91, DS91, Fou91, GK91, Kub91, MB91a, PK91b, SS91a, Sar91, SS91c]. closest [Hob91]. cluster [Ala96, Gro91, OS93a, WW91b]. clustered [GMR97, ML91, QT98]. clustering [Bri97, MS95]. coalescent [Yam97b]. Cochran [CM92, WH99, ZB97]. Cochrane [ST93a]. Coefficient [Cha93b, BM98, BH99, CTT90, CN99b, Hsi90, JS90, JS91b, JN91, KKS99, Liu95a, LK96, Mat98, Mat99, Mi90, PB90a, RS98, RBM95, SM91b, Tah92, WG95]. Coefficients [Lee95a, Bén90, Dia90, GM96, HW91, Jel99, Koo93b, Leu94b, LR96, LL96, Mil91, MF97, STS95, SFT99]. coherent [ED93]. Cohpasisons [Hoo90]. Coincidence [Nab93]. cointegrated [Kun93]. cointegration [Shi99]. Collapsibility [Dav90]. collecting [CS93]. Collinearity [LM93, KG90]. column [GC98b, PD92, PS95a, SB91]. Combination [CB90, AB91b, HCL94, PGT98]. combinations [Bar91, CR92a, DAM99]. Combined [AK99, BS92, TK91b, RSAJ90]. Combining [Ahm91, Dou93, VCvdL97, Sar92b]. Comments [Bly91, Dey91, Mas91, Nay91, Sen91, Str91a, CTT90, Has95b, HO90, KJ90]. common [CTP93, CZ97b, EP92b, Hod90, JC98a, JC98b, Kel96, KR90, MT90, MPH97, PM94, PT92, RS92, Sar91, ST90b, SS99d, Wan99, Wes98, Wit91a, YB94b]. commonly [LSR94]. commonly-used [LSR94]. commutativity [Rag90]. Comparative [CBRS90, Kan90, AG96, IY99]. compared [BB93b]. Comparing [Cer97, HAKM93, Iso92a, KCM94, RBLN93, Smi92a, Wil92, YC99b, CS95, GM94a, GP96, HL92, KC98, Kim95c, Liu96, MA92, QSC94, Sob96, Stu91]. comparisions [CL96]. Comparison [Ban90, DG92, KS98, Liu95a, MS90c, PMTP93, PK91b, RW96, SNK91, Sri95, SS94e, TS90, TS94, Tu99, Ver96, Arn98, BL98a, BD91, But98, CG92a, DS97, DY98, Fou91, GM94a, HT98, Hug91, JS93, KO94a, Kea96, KAKW99, KS90b, Kla96, KL96, LV98, MM93, MLGD98, NvdMM90, PB90a, SSSS90, SO95, SI97, US92, WSM91, Wik98, Wil91]. Comparisons [CW98, Sob95, Bof92, DM91, Dri91, Has95a, KSA98, LSU90, MB91a, SBS96, dMD95]. compartmental [BM98, KP93b, KKT98]. compatibility [CT93b]. compatible [SA93]. competing [DK95, Geo91, KD99, RZ98, Wan93a, Wan94b, YL92, YL93, ZK94]. Complete [Sob90, DRL97, Gou95, GC98b, KC98, KAHK94, MP91a]. Completeness [Bha95c]. complex [GM94b, GF91, Hun90, Kun91, Mor90, SA90b]. complexity [Boz90, MMD93]. compliance [CMS99, RT91, Rob94]. component [CTT90, HV91, Hoc90, KKK90, KM94c, Lu90, PB93, RRA92, SLS96, SJ93, TK99a, WJ90]. components [Bén90, BM94, BS90b, Bre90, CM95a, CG92c, DMG91, EB94, ÉA95, Fan91a, Hei93, HK94, KMM90, LC96b, LD99b, LSU90, MBR91, MR95a, Par98, Pre98, RA96, SCSG94, SG91a, Sub91, TC90, VP98, YB95, YB94b, ZK90]. Composite [NA94, Bor95a, Bor95b, Fri90a, Fri90b, Law93, RK96]. Compound [DC92, FM92, GN95, PG92]. compound-symmetric [FM92]. Compromise [Yan90b]. Computation [Ong95, HS96, PZ99]. computer [EF92, Ste90, WS91b]. Computing [CW97]. concavity [FR93a]. Concentration [STS95, Ree96]. concentration-controlled [Ree96]. concept [Eva97]. concerning [DA98, Sco97]. concomitant [CR92a, JB93, MD90, Tsu98]. condition [CW93, Mac98, MQZ99, PGC97, SK99]. Conditional [BF90, BR92, CH95b, Chi99b, Gov97, Koo93a, Liu93a, SA92a, Tak94, TY99, WA95, Yip91, ALW91, Age99, ACS95, CT98, DH90, DQS98, Dut99, GV90, KK98a, Nay96, Ong93, ST90a, VV96, VV97, ZRM90]. conditionals [AA91, CS90b, Sar95]. Conditions [POGP95, EPdR99, Kun93, QCL98, RG96, Sha92b, WD97]. cone [TK99a, Tan98]. Confidence [DK95, FTV99, GF91, Hol95, KK94, Kra92, Mor91, PB98, PB93, SG91a, WSM91, YB94a, YB95, AM99a, BS96, CH96, CM96a, CS95, DH92, DG97b, EB94, Fay99, FS95a, Fuj99a, GLL90, HB90, Huw93, KMM90, KML91b, Lin95, Liu95b, Llo95, MD93b, Mor90, Nag96, Nag97, Pap92, RS92, SM94, SL94a, Ste91, SS99d, Tah92, Tu99, WK97, Wit91b, YL91, dlH91]. Confidentiality [CS93]. confirmatory [TWM91]. confounding [GJ97, HB90]. conjecture [FKR91]. conjectures [ZM93b]. conjugate [GCD94, HB99, Nag99]. conjunction [TM95]. connected [Dou93, IY99, KG92a]. connectedness [PS95a]. connections [CFRT92]. Consequences [SJ93]. Considerations [Hol91, CG92a, CL94, CG90, DW90]. Consistency [CW94b, GP90, Hau91, Kun95, Str95, YW98, cL91, BOR92, CW93, GS90, LYS96, MR95b, PM93].

Consistent [MEC93, BH92b, HZ90, Hon99, LY97, MP91b, TZ90a, YW98, ZK94, ZL91]. constancy [Lom99]. constituent [Smi92a]. constrained [JBY90, KI94, SO95]. constraints [Gru93, HL93a, MG93, MP93, PS93a, Sar99, SL99, TW92, Wan96, WK91b, Yua92]. Constructing [KAHK94, BMK93, LV98, Was94]. Construction [AS94b, BN91, GT91, KKS99, PS91, Sin92, AR90, GP92, Liu95b, MMS91]. Constructions [CS97c, DS91]. Consul [Sap94]. consumer [Str90]. containing [JN94, LL91]. contaminated [Don93, Gle98, JS95, MP93]. contamination [CW98, FZ96, Sat91b, WG99]. contemporaneous [MC91]. content [KK91c, LR92, Moh90]. context [BK98, BC90, DX90, Gar96]. Contingency [SP90, Cer97, Dav90, DY96, GG94b, LS94, MGLF98, Myo95, OW90, PS93a, STL91, UG95]. continuation [Ley90]. continued [NS99]. Continuous [WS99, Ano98n, Car93, CC95b, GPA98, GL94, HK91, KM94a, LG94, Leu94a, MC91, OS93b, Par94, Par93c, SK91b, WK93, WK91b]. continuum [Liu93a]. contoured [NS92]. contrast [Cha92b]. contrasts [HY97, MG91, Sug99]. contributions [RC96]. Contributors [RAM96]. control [ACR91, Alw92, ARS95, AW99, Ami93, AJC99, BB93b, BD91, Cha92a, Che95a, DS97, DHW90, DJG94, DGMR96, DFT91, Dou93, Fuj90, HL92, ID98, KC98, Luc92b, MR90, NA90, NS96, OW93, PL96, RA96, RT96b, Ros96, RWP95, RAM96, SS90b, SWD93, SW91, SW97, Sri99, Var96, Vau92b, Woo92, WB93, Zha95b]. controlled [Ree96, Rog92]. controls [KY91a]. conundrum [Smi98]. Convergence [ACK95, LS94, Nis94, OT91, Ton96b, Ton96a, Vio95, GN91, Let97, SG94, Sun97, VV96, VV97, Yu97, Zha93a, ZW99]. converges [Blo97]. Converting [Won92]. convex [BR92, CM95a, DAM99, LAC90, TK99a, Tan98]. Cook [Ban98, Hid91, McM97]. coordinate [HS90b, dP90]. Coping [AG90]. copula [GM90, LK96, ZK94]. copulas [Car94, SY96]. corner [PE92]. corollary [Kea91]. correct [Bof90, Liu93a]. Corrected [dSC99, Alb92, hZ93]. Correcting [BT93, RT91, Rob94]. Correction [Sar92a, Cor92, CF98, DS92, Dre96, Has95b, HC94, KML91a, PC95, WLG99]. corrections [BC97, Cor92, CF96b]. correlated [AA92, Che98d, DF94b, FM96, FS99, Kla96, LD95, Mar92b, SS93b, SD97, Tak97a, Wan93b, WG95, ZB97]. correlation [BC90, BH99, CC95a, Cho91a, CD98, CLL95, Fuj99b, Goo91, Jel99, JA92, Kab90, KS94a, Leu94b, LL96, LK96, Men97, Mi90, MJ92a, Ong93, PB90a, RDM92, RM93, Sch90, Set90, Tah92, TS92, WG95]. correlations [LF93, Mak94, SKF94]. Correspondence [Beh99]. corresponding [Wil92]. Corrigendum [Ano92a, Ryc94]. cost [Boy95]. count [BJ94, FY95, Lee97, Yip91]. counter [Sap94]. counter-example [Sap94]. counterexample [Lis92]. counts [NPA93, Tha92]. course [Gar96, Moo96, Sne96]. courses [Var96, Zah90]. covariables [Che97]. covariance [Abt92, AM99a, Boz90, Byr96, Cal94, Cho94a, DJ98, FM92, Gal94, JH99, JB93, Kim94b, Kim95c, Kri91, MD94a, MEC93, PK94, SE97, ST96, SI97, TA96, TWM91, TW92, TK93, VV94, Wan93b, Yok96, YSM92, YB94b, ZB93, ZS97]. covariances [BR91, VR90]. Covariate [RDA92, Dam93, Fuj94, Gra98, HD94, Koo93b, LG94, Li99, PPS97, RDLT98]. covariates [BC97, Coa94, GZ97, IH94, Leu92a, RI93, SPQ98, Zou94]. coverage [Gou95, Isa92]. Cox [Kim96b, QL92, XB97, Yan96b]. CPIT [Gau98]. Cpm [ST93b]. Cramér [DM96, DRL97, Gho91, RV95, RV99, Sco99, VR96]. Cramér-von [RV95]. credibility [Gro91, vdMvdM90]. credible [DF94a]. criteria [Gen97, Hau91, Hon94, Kc91b, Kar95, MA95, US92, Zha93a]. Criterion [Jah91, Ati90, Bor95a, CN99a, CH91c, HC97, Kub91, Mad92, NC97, PK91b, Sar91, SNK91, SLL97, Tan94, YY93]. Critical [Dun95]. criticism [MNS96]. Cross [LR98, SH91, Sob92a, Chu94a, DG97a, HCW91, RW94, RL97, SS99a, WP91, WW91b]. cross-classified [HCW91, WW91b]. cross-over [RL97]. cross-sectional [WP91]. Cross-Validated [Sob92a]. Cross-validation [SH91, Chu94a, SS99a]. crosses [GDK95, GC98b, PS91, VP98]. Crossing [Kul99]. crossings [SR91]. crossover [Cor91b, KAHK94, ZS98]. Crow [Fuj99a]. crude [DQS98]. cubic [Djo95, Dut91, HH90]. cum [CM95a]. Cumulants [NSI90, ZM93b]. cumulative [Che95a, PZ99, POGP96]. Current [Zah90, CP96a]. curtailing [Li98]. curtailment [DH92]. curve [Bha95a, CR93, Dem96, Fol92, Fuj99a, GK92, Gup99, Kim94b, KKP^{+}99, LTvC^{+}97, TK93, Von91, Yok96, ZR98]. curves [CHS98a, JU93, MP91b, PC98, RI93, Vas96, VV94]. cuscore [NA90]. cusum [Boy95, Edg96, RWP95, Sri97]. cycles [Wan99].

D [Hud91, SA93, YC99a]. d’Agostino [Kro94b]. damage [DX90, POGP96]. Darling [RV99, SSA90]. data [Abd98, AAB96, AAB97, AC99, Ana92, Arn96, Ati90, AM99b, BS94, Bar98, Bar89, BBK98, Beh99, Bel93, BC90, Ber98a, BB93a, BSSB90, BJ94, BDF95, Bro97, Cal91, CS97a, CM94a, Che99, CG92b, Che98d, CM98, CW92a, CHC98, Chi99c, CK96, Coa94, CM96b, CA90b, DG92, DJ98, DG90, DS93, Don93, Dru90, DRL97, ERO97, Ent90, FS93, Fay99, FN92, FBD97, FBD99, FY95, Fre92, Fri94, Fuj94, Fuj95, Fuj99b, Gal94, Gao94, GG94a, GJO90, GMR97, GP90, GR90, GG94b, GC99, Gle98, Gra98, Gro94, GNR98, GP94, GT92, GGL96, GGG98, HCW91, HP92, Ho98, HH95, ID98, Jin97, KO94a, Kim96b, KL99b, Kla96, Kri91, KM95, Law97, LT92, Lee95b, Lee97, LLL93, LZ98a, LY97, Luc92a, LTCC94, Luo94, LK98, MH98, Mad92, MMPV98, MEC93, MM97, MML99]. data [MD92, Mon99, MPQ91, MH96, MS99, NPA93, NT92, OC99, PC98, PFU91b, Par93a, PLW93, PS95b, QT98, RS98, RD96, SB97, SS90b, SS94b, SK97b, SY94, SCXG96, SP99, SM91b, SSCS99, SL94b, SD95, SD97, Tah91, Tiw94, Tre94, TK93, TM95, WW91a, WG90, WS94, WW91b, Wun91, YN99, Yip91, YW98, Zha95a, ZB97, ZK94]. data-based [Ati90, HP92]. datums [SSCS99]. Deal [Ino99]. death [KD95, PK91a, WP92]. deciding [FN91]. Decision [LLL93, LL96, AB91a, GdWG91, Kim94c, WB93]. decision-making [GdWG91]. decision-support [Kim94c]. decomposable [MGLF98]. Decomposition [BP96, TC90, Cho90b, HV91, RV99, Wit90, YC99a]. deconvolution [TZ90a]. decreasing [BB90a, Bru95]. defective [PM92]. defectives [GT91]. deferred [KG92b]. deficiency [GR90]. definitions [Pei92, Ree96]. degrees [HC94, KAKW99, Oht96b]. delay [Zha90]. Delta [SK99, NTS91]. densities [BS99b, CW92b, ES94, GV90, PL96, RS90, SD90, WW91a]. Density [HS93, Ati96, Bru95, Car90, Chu94b, CBRS90, DS92, GP90, GP93, GM90, GNR98, GP94, GV90, Hon91, Jon94, KHJY97, KK98c, Let97, Lou98, MR95b, Oht93a, Oht93b, RGU97, Rod92, S99, SK91b, TZ90a, Tro93, Ull92, WS99, Wri98, Wu96, Yu97, Zha98, ZZ99]. Department [Bra90, Ley90]. departure [SJ93]. departures [Lee98]. depend [SJ93]. dependence [De 92, EPdR99, GC98a, Hon91, Jai94, KS95, KM95, Law97, MA94, Myo95, RG96, ST90b, Sun99, Tiw94, Yu97, Zog98, dR94]. dependent [BT93, Chu94a, DK95, Hon99, Jin97, Kla96, KK91c, KG92b, LW95, Li99, Mar95, Per91, Ryc93, Smi92a, SU95, SC92, Tsu98, TM95, WA96, ZK94]. derivable [Yua92]. derivation [Ati90, CN99a, Cho90a, MY95]. derivations [Lei90]. derivative [Jon94]. derivatives [Abd93, Lou98]. derived [Eva97, GG96, Hyo94, RV99, SM96]. Describing [LD95, JU93]. descriptive [GS93, GS94]. Design [Par98, Adh91b, AB98, AFR94, AR90, Arn98, BY92, Bre90, BL90, CG92a, Cha92b, CM95a, CG92c, Cor91b, GS90, Hai93, Jai94, KC98, KAHK94, LRD94, LTvC^{+}97, LM97, Lin93, MM96a, Nor95, Pas94, QSC94, Rag90, Ras90, San92, Sha92b, SK97a, SMS93, SLL97, Tsa90b, Tsa90a, Ull92, Vau93, VL92, WK91a, WK91b]. design-based [GS90]. design-cum-model [CM95a]. Designing [Men92a, SB93]. Designs [Par90, Abt92, Afs90, AK99, AZYK98, AC95, AVN97, BN91, BL98a, BMK93, BW93, Bor95a, Bor95b, BD92, CMD94, CC90b, Cha96b, CHS98b, CS97c, Cor91a, Dam93, DGD92, DG97a, DLW96, DL95, Dri91, Dut91, Dut93, GT91, Goo91, GP92, GDK95, GP96, GC98b, Han91, HGM94, Hud91, Hyo94, JP93, JN96, Joh95, JP92, KM90, KP95, KA91, Kro94a, Lee94a, LS97b, Lia99, Liu97b, MG94, MK90b, NSS99, OS93a, OS93b, Özt99, PD92, PS95a, PK98, PJ99, PL96, PS91, Pre98, RW94, RZ98, RRAB93, RL97, Sas91b, Sat91a, SSS97, SM91a, SM92, SK91a, Sin92, SW98, SGD90, SB90, SB91, SS99c, Stu91, Sub92, Tem94, Udd98, Wee92, WN92, WK91a, Wij93, Yua92, Zac91, Zha92, ZAW98]. destructively [Men97]. detect [Ali90a, SD90]. Detecting [Hid91, Lee98, MBR91, Zha95a, Che95a, LPK95, OC99, SS91b, Wan98b]. Detection [Abd95, BB93a, Bha95a, Bha95b, Hos90, Mad92, TCSJ94, CZ97a, GC99, Nai90, Pet94, SGA99, Var98]. deterioration [NS99]. determinant [Cho91a, Mar92a]. determination [LC99, PL97]. determine [KA92a]. Determining [Gle98, SS91b, FT99, Sha98]. deterministic [OC99]. Development [YK90]. deviation [ATT99, CS90c, You93]. device [SMS93]. DFR [JSM98]. diagnostic [CS95, Fun98, Hoc90, KK98c, Mat98, RS96, SV97, SS94d, WS91a]. Diagnostics [KC90a, KC90b, Bro97]. Diagonal [SY96, CJ98a, CG91]. dial [GC98b, VP98]. diallel [GDK95]. dice [GL90]. dichotomous [CC95b, Cus96, Leu94a, PGT92, WN97]. Dickey [KS90c]. Dickey-fuller [KS90c]. dietary [LF93]. difference [Dou97, KMBTG99, Lai99, Par90, PT93, SS93b, TA96, Vas96, WK97, Won95, YB94a]. differenced [Swi98]. differences [CWH93, DN92, DG97b, GW94, Nor93]. differencing [Li95, Swi98]. different [TA96, TM91]. Difficulties [MA95]. digraphs [MAPV96]. dilution [CBRS90]. dimension [Abt92, Ati96]. dimensional [Car90, DS91, Ete91, GS91b, Koo90, SK91b]. dimensionality [Hwa94, SS94e]. dimensions [Ein96]. diminishing [JR90]. direct [CM95a, RL97, UG95]. Direction [KG90, YY95]. directional [Fin90, PK98]. directions [AVN97, Dun95]. directly [Fay99]. Dirichlet [Den91, FBD97, MMPV96, Swa93]. Dirichlet-multinomial [MMPV96]. disasters [KKT98]. discontinuous [GPW98]. discordancy [He90]. discovering [Nay96]. discovery [GL94]. discrepancy [TS95]. Discrete [Koc95, KM97, ATT99, Bar89, BR97, BB90c, Con90b, DY96, FMP98, Hal96, Kra92, NS97, Par94, YW98, ZM93a]. Discriminant [Wil94, BB90c, Buo90, Cha99, CBR91, DG92, Fun96a, Fun96b, Fun98, Hwa94, KL96, KKK90, LW90, Pos92a, VSS95, WL92]. Discriminating [SGS94]. Discrimination [CCC96, DN92, KI94, Law90, LG93, QL92, SC97]. discussion [KJ91]. disease [CSK^{+}98, KCM94, Wen94]. disease-free [CSK^{+}98]. disease-state [Wen94]. disjoint [GP96]. disparity [MMPV98]. Dispersion [JR90, Kim94a, AS94a, BC97, CF96b, Cor91b, Cro96, GW94, Kre92, Lee94a, MM96a, PMD95, RAP95, SNK91, Wik98]. Dispersion-diminishing [JR90]. dispersions [Jor99]. dispersive [MM93]. distance [Ano99m, Ban98, CBR91, CA90b, DH96, DRL97, Hid91, Liu95b, MNZ90, Tan94, Yan98]. Distances [PK94, Liu93b, Luo94, MD93b]. distantly [Tho96a]. distinct [Mar97]. distributed [BP93, GS92, Kam94, Nag97, PE92, Sei98, WC96b]. Distribution [CS90a, ES94, FR93a, GS94, Guo96, JA92, KGM94, Man91, Oht93a, Pan96, PGT98, Sin91, SL94b, Vän97a, Abd93, AL96, AK96, AB92b, AK91, Ahs96, ATT99, AK94a, AS94a, ACK95, AK98, Ati96, BCA93, BA94, BGP95, BR97, BP99a, Bho97, Bou90a, Bou90b, BS90c, BSK95, BS98b, CW90, CP94, Car93, CW94a, CT94, Cha96a, CHA95, CHA97a, CH95a, CP97b, CM98, Chi92, Chi93, Chi95, Cho90a, Cho91a, Cho92, Chu92, CG90, Con90a, Con90c, Con90d, CF95, DG94, Dah95, Den94, DY95, DY98, Dix91, DM91, DF93, DF94b, DH99, DA98, DP93, ESDC96, EJ93, Fam93b, Fan91b, Fan91c, FKT95, Fri90a, Fri90b, Fro93, FS95b, FS99, Fuj95, FMV98, GB98, GH92, Ghi98, GJ92b, Gil92, God92, GN91, Gru93, GC98a, Gue94, GP94, GGL96, GS98b, GK98]. distribution [HH94, HK91, Han96, HD99a, HK90b, HSS93, HN97, Hsi90, Hua90, HB91, HFP94, JP90, JK91, JT92, Joh96, JS90, JS91b, JN91, Kab90, Kab91b, KK98a, KK91a, KH90, KS90a, KK92, Kem92, Ker95, Kib99, KK90, KK91b, KS94a, KSmS90, KR90, KM94b, LC96a, LPR99, LM96, Lei90, Leu92b, LD99a, LH99, LBB92, LR98, LSU90, Llo99, MP91a, Men91, Men95, Mis96, MPR98, MPQ91, McM97, Nag99, NS93, New99, NAW96, NS92, OR98, Oh98, Oht93b, Ong90, PZ99, PM90, PFU91b, PR92, PMD95, PSM96, Pan97, PG95, PT92, PL97, Per91, PP90, QCL98, RRA92, Rao90a, RKN91, Rie95, Rie99, RRrA91, STR97, SMMP93, Sar95, Sca92, Sch97, Sco95a, Set90, SH94, ST90b, Sha92a, Sha90, SGB98, SNM97, SGN93, SUS95, Smi92a, Sna94, SKS93, Sun96, Swi98]. distribution [Tah92, TM91, Tho93, TP97, Tsa90a, Tsi99, Tsu98, Wit91b, Wun91, YY95, YSM92, YP93, YW98, ZQ97]. Distribution-free [GS94, Guo96, KGM94, Man91, Pan96, DM91, Sha90, SGB98]. Distributional [Bur97, MLH92, PLC98]. Distributions [SR91, ALW91, AH97, AD98, AAG98, AC95, AS92, AA91, AP91, AN91, BS91, BM98, BLF99, Bar91, BCLBMR99, BP99a, BS90c, CR93, CW97, Cap92, CS90b, CXZ90, Cha90a, CTP93, Cha93b, CP99b, Cho91b, CC95b, Con90b, Con94b, Con94a, CF96a, CT97, DG94, DLS97, DS97, DR96a, DQS98, DM96, Den91, DY91, DC92, DH96, DX90, DT98, ES95, EP91, Epp95, FN91, FK94, Fer90, FMP98, Gar92, GJ92a, GG96, GSD96, God90, GG92b, GGVM98, GK90, GT92, GGT95, GT96, Han92, Iso92b, JN95, Jan97, Jan98a, Jan98b, JP91a, JP91b, JC98a, JS95, KHR97, Ker95, KG92a, KM94a, Koc95, KCM94, KM97, Koz94, Kra92, Kul90, KM93, Lee96, Lei90, LL90, LC99, Lin95, LL93, LR93, LBBCMR97, Luc92a, MT90, Mad93b, MM93, MB91b, Min98, Min99]. distributions [MAS94, MS98, NN91, NR99, NvdMM90, NSI90, NS97, Olu94, Ong93, Ong95, PS99a, RG92, RL91, Rao92, RDA93, RS92, SA90b, SJ91b, SKF94, SD93, SK91b, Tah91, TV91, TK92, TTG92, TK99b, WSM91, WP91, WA95, Won95, Yu91, YB94a, YK90, ZM93a]. disturbance [Nag95, Nag96, Nag97]. disturbances [HC93, Kib96, Mae90, MPB90, Sil94, SPS94, Tak97a]. disturbed [Wil98]. dithered [Was99]. Divergence [MSMP97, ZFP90, Est96, FMP98, KL99a, LS94, MMPV96, Mie91, TZF97]. divergence-based [MMPV96]. divergences [Sal90]. diverse [SCXG96]. divisible [BMK93, MK90b, Sas91b]. DNA [Gol90]. do [Mil96, RL97]. Dolby [Huw93]. domain [But98, CM94a, CM95a, CAS97]. domains [AG96]. dominance [ESY93, MC92]. Dominating [ST96]. Donnelly [Yam97b, Yam97a]. Dose [CG92b, Pat94a, AKP93, BSSB90, HGM94, Jan95, KAHK94, LTvC^{+}97, MF90, Ree96]. dose-adjusting [Ree96]. dose-ranging [MF90]. Dose-response [CG92b, Pat94a, AKP93, HGM94]. dose-response-curve [LTvC^{+}97]. Double [Buo90, Ahm97a, AS94b, Ber99, DB95a, FBD99, HK90a, KI94, KG95, SA92a, SA92b, SA93, Tim96, TB91]. doubly [BE93, BDF95, ES95, Fer90, FBD97, LY97, PB91]. doubly-censored [LY97]. down [Wil91]. drawn [Adh90]. drift [Liu97a, OC99]. dropout [MM97]. dropouts [QSC94]. drug [AB91b, HCL94, Tiw94]. DS [WL92]. due [KM95, MP97, TM96b]. Dunnett [CM95c]. duration [WP91]. Durbin [Kul90]. during [ZS98]. dynamic [Bol90, KK92, Mae90, Mak91, SK96, Wan98b, Wen94].

E-optimal [DGD92]. E-optimality [Dut91]. each [Bou90b]. early [BB90c, Tem94]. earthquake [Iha93]. easy [Tur90]. ecological [BC90, RDM92]. economic [NS99]. Economical [Sri96, SW99]. Edgeworth [Cha91b, CHA95, CHA97a, GS98a, Qum94]. editor [Ano92o, Ano95m, Ano97l, Ano98m, mC96, GC99, Joh96, Kle94, Kum98, Sap94, She98]. Editorial [Ano92b, Ano98a, Ano99a, Smi92b, Smi94, Smi95, Smi96, Smi97, Smi98, Ano90a, Ano90b, Ano90c, Ano90d, Ano90e, Ano90f, Ano90g, Ano90h, Ano90i, Ano90j, Ano90k, Ano91a, Ano91b, Ano91c, Ano91d, Ano91e, Ano91f, Ano91g, Ano91h, Ano91i, Ano91j, Ano91k, Ano92c, Ano92d, Ano92e, Ano92f, Ano92g, Ano92h, Ano92i, Ano92j, Ano92k, Ano92l, Ano92m, Ano92n, Ano93a, Ano93b, Ano93c, Ano93d, Ano93e, Ano93f, Ano93g, Ano93h, Ano93i, Ano93j, Ano93k, Ano94a, Ano94b, Ano94c, Ano94d, Ano94e, Ano94f, Ano94g, Ano94h, Ano94i, Ano94j, Ano94k, Ano94l, Ano95a, Ano95b, Ano95c, Ano95d, Ano95e, Ano95f, Ano95g, Ano95h, Ano95i, Ano95j, Ano95k, Ano95l, Ano96a, Ano96b, Ano96c, Ano96d, Ano96e, Ano96f, Ano96g, Ano96h, Ano96i, Ano96j, Ano97a]. Editorial [Ano97b, Ano97c, Ano97d, Ano97e, Ano97f, Ano97g, Ano97h, Ano97i, Ano97j, Ano97k, Ano98b, Ano98c, Ano98d, Ano98e, Ano98f, Ano98g, Ano98h, Ano98i, Ano98j, Ano98k, Ano98l, Ano99b, Ano99c, Ano99d, Ano99e, Ano99f, Ano99g, Ano99h, Ano99i, Ano99j, Ano99k, Ano99l]. education [Mil96]. educational [BB90c]. Effect [GJ92a, GT90, PFU91a, SGJ95, Yip96, Aga90, Bil97, BKS99, CP99a, GJO90, GL93, Gup91, Gup92a, OW93, PJ99, PP92, Pep90, PL97, RDA92]. Effective [Cus96, Ksh98, RT96b]. effectiveness [NA94]. Effects [Alw92, DRS90, Fuj99b, Tak97a, BK91, BW93, BH99, CC95a, CC90b, CK95, CG90, Dam93, Gov95, Hei93, HS99, Isa92, JP92, KY91a, Lai98, Lee94a, LC96b, Lee97, LTCC94, LK98, MW96, MM96a, Mon99, PB98, PPS97, RW94, RZ98, RDLT98, RL97, SS93b, SM96, SKF94, SPQ98, SL98a, aS99, SO98, SS99c, WL96, Wik98, Yok96]. efficacy [CB98, FWG94, MF90, Yip98]. efficiencies [MPH97]. Efficiency [AHAEK92, HW91, Lin93, Yan96a, AN90, Boi91, Bol90, CC90a, CC90b, CL96, CN99b, Kar95, KP95, LP94, Pap90, Ree96, SK91a, SY93, Tur95, Wan93a, Yan90a, YB93]. efficiency-balanced [KP95, SK91a]. Efficient [AZYK98, GP96, HS96, RL97, DD97b, KLH^{+}91, LZZ99, MNS96, SSM90, Stu91]. Efron [GP98]. Eggenberger [Jan98a, SJ97]. Eicker [Ein96]. eigenvalues [HM99, Leu92b, TCT90, YB94b]. eigenvectors [TCT90]. element [ZM90]. Elementary [SET96, Kot90]. elicitation [AAG98]. elicited [Ste92]. Eliciting [MGR95]. Eliminating [Fre92]. elimination [Liu94b]. ellipsoidal [Cy95]. Elliptical [AVB96, JA92, TK99b, WH99]. elliptically [ML96, NS92]. EM-type [AAB97]. Empirical [Chi93, HW90, HFP94, Hwa92, LW90, Lia90, LL94, Bro95, Cam93, Car93, CAS97, Ein96, Kmi95, Lu93, LM94, Mar90a, MK90a, Nai96, NPA93, NS90, Qum94, RT96a, SL99, SC93, SC96, Sun97, Ton96b, Ton96a, Vio95, ZW99]. employers [HJ90]. EMS [PC98]. end [LM97]. endpoint [CSK^{+}98]. endpoints [Che98a, SM96, TK94, TM95, Wan98a]. energies [MJ92b]. engineering [Var96]. enhance [MGR95]. enhancement [NKN99]. enrichment [Tem94]. ensure [Blo97]. entropic [DH99]. entropies [SMMP93]. Entropy [GW99a, Sal90, Cor95, Lei90, Liu91, PMTP93, PMSM97, PN95, PH92, Yan90b, YK97]. Entropy-based [GW99a]. entry [Lin93]. environment [CH97, DFT91]. Environmental [Jan95]. EOR [Tak99]. epidemic [Bri97]. equal [CS94b, PK98, PT92]. equality [And91b, CGP92, CO91, Fro96, GT97, GM96, HL90, Koz94, MJ92a, NN91, Olu93, TW92, TV91, VT93, YL92, YL93]. equation [BWCM90, CP96c, HK90c, HL93a, LT92, PS95b]. equations [Bha95c, Cho90b, Dem92]. equi [AA92]. equi-correlated [AA92]. Equiblock [SK91a]. Equiblock-sized [SK91a]. equilibrium [GS98b]. equireplicate [SS99c]. Equivalence [BB93b, BB96, Par93b, Rao92, Fun98, Hil93, LWH98, MAPV96, Was94]. Equivalent [LLL96, AAW93, CKH95, Ksh98, Pei92]. Equivariant [JB93, Kub91, gWyJ97, ZS97]. ergodicity [Lim93]. Erlang [PS96]. Erratum [Ano91l, Ano96k, Che96, Pra91]. error [Abd93, AK94b, Ati96, BLBE99, BOR92, Cal91, CHS98b, Che97, DH90, DD97b, Edw94, GBC99, Gri97, Gri92, GT90, Güv95, HK90b, HK97a, Has95a, Hol91, HO90, Isa92, KSA98, KF90, KMBTG99, Lia94, LZ98b, LH99, MD93a, Muk94, Muk95, Oht96b, Oht96a, PB93, SH94, Sin91, SG91b, SW99, TL99, TC91, VSS95, Wan94a, Wan96, Wan93b, WLG99, WJ90, WL92, YB95, ZZ99, as97]. errors [Abd95, Alb92, Ban90, BT93, CC95a, CT99, Cha91a, Cha93a, CMS99, Chu94a, DN90, Dut99, FM96, Fin90, GCD94, HW91, Hos90, KC91a, Kim93, LS97a, Nag97, Pas94, PR93, Pro90, SS93a, SS94b, SW98, Wan93b, WZ94, Zha94b]. errors-in-variables [Abd95, Ban90, PR93, Pro90]. escape [Tem94]. Escoufier [CTT90]. Esséen [Lia94]. estimability [SSM90]. Estimate [SS92, AW92, BS92, Bru95, Dem96, GJ92b, Kej93, LSR94, LLH90, Liu97a, Ng93, RD94, WC96a, WL92, Yan96a]. estimated [Bak91, CC95a, CP97b, JV92, Ker95, PMSM97, Pas94, Vio95]. Estimates [KSmS90, Lee94a, Abd93, BWCM90, Bar95, Cha91a, CW93, CW94b, Cho91a, DD97b, ERO97, FT95, GG96, Hob91, Hon91, LP98a, LD99a, Liu92c, LR93, Mar90a, Nag99, NT92, PM93, RL91, Ryc93, SG94, Wan93b, Wit90, Wit91a, YV95, YLB98]. Estimating [Ana92, Bil97, Bou90b, CH94, CH97, Fuj90, GM90, GL94, IY99, JN94, JS91b, KF90, LF93, Llo99, MW96, MH95, PC98, RC95, S99, SG90, SO98, SC92, Str90, TWB99a, WG95, ZR98, Bha95c, CHS98a, Fuj99b, GJM96, Hug91, Hwa94, KO94a, Kel96, KA91, KM94b, LH93, MD95, MP97, Nay94, Nay96, NM96, Par90, PS95b, PL97, Sar91, SV97, SH90, Son92, Tar91, TM96b, Tre94, UI99, WK97, WN92, YY93, Yip98, Yu98]. Estimation [AL96, AB92a, AHAEH90, AAB96, SAP92, AM91, Anr94, Arn91, Arn92, Bal98, BGP95, BCLBMR99, Bho97, BSK95, CPK92, CM95b, Che98c, CW92a, Chu94b, CK96, DK91, DW90, DL90, DU91, Dru90, EP92a, FKT95, Gil93, GBC99, Gri92, Gup92b, GGHA95, Han97a, HD99a, Her92, HD92, HN97, Jai91, Jan99, JP90, JP91a, JI97, JW92, KE93, KHR97, KS90a, KKK91, Kri91, KY93, KM93, KM94c, Leu92b, LM94, MT90, Men97, MD93a, MAS94, MJ92b, Oht91, Qiu94, RRA92, RS98, RDM92, Sca92, SGN93, SMS97, SE97, Sun96, TK92, TB91, TK97a, VP98, Wil96, Won93, YB93, Abd98, AF98, AB92b, Ahm98, AK95, AKP93, Ala96, Ali91b, AM93, Ano99m, Ara90, Arn94, Ati96, BS91, BMM^{+}91, BR93b, BC90, Bev92, BLY92, BE93, BDF95, But98, CP90, CGP90]. estimation [CP96b, CTP93, Cha90b, CA90a, CM94b, CAS97, CM98, Che90, Chi92, Chi93, CH95b, Chi99b, Cho94c, CK97b, CD98, Chu98, Cip98, CBFCN99, CT97, CJH95, CW92b, Dah95, DLS97, DN90, DMG91, Dav91, De 92, DY91, Dix91, DS92, Edw94, ES92, EPdR99, EHL93, FTS94, FK91, Fer90, FBD97, FBD99, Fri92, Fro90, Fyg91, Gab92, Gab98, GPW98, GW91, GG93, GPA98, GS92, Gro99, Gru93, GNR98, GP94, Güv95, HEB92, HW90, Hei93, Her93, Hoc90, Hon99, HSS93, HK94, HH95, HS93, HPS98, Hwa92, HS99, HM99, Iha93, Isa92, IU93, JP99, Jel99, JB93, Jin95, JT92, JG95, Jon94, KD95, KK98b, Kap92, Kar94, KM94a, KHJY97, KK98c, KS94a, KR90, Kro97, Kub91, Kub99, Lai99, LM96, sL98b, Law90, LAC90, LS97a, Leu92a, Leu94b, Lia90, Lia94]. estimation [LZ96b, LZZ99, LH99, Lin97, LL96, Lop93, LB98, LBBCMR97, Lu90, MQZ99, Mak94, Mak91, MG93, Mar96, Mar92a, MN93, MY95, MK90a, MMT97, MM97, Mie91, MU97, MP91b, MPQ91, Muk97, Myo95, Nag95, NS93, NS94, NMB98, OSL91, OSL92, PB90a, PM90, PFU91a, PR92, PM94, PSM96, Par93b, Pas94, PB90b, PB91, PG92, PB92, Pep90, Phi96, PS93b, RKN91, RTS92, RG94, RL94, Rie95, RZ94, RW96, SS99a, ST90a, SA99, Sar92b, Sha92a, Sha99, SPQ98, SCSG94, SL98a, SS94b, SY94, SK96, SG91b, SU95, ST93a, Stu94, Sub91, Sun98, Tah91, TK99a, Tap92, TL99, TZ90b, TC92, TY99, Tro93, Vau92a, Vau93, Vel93, VS91, WS99, WW91a, WJ90, WC96b, YIK96, Yu91, YK97, YR94, ZSW90, Zha94b, Zha94a, Zha98, as97, dlH91]. estimation* [DC92]. Estimations [Lu93, XB97, FK94, HFP94, SSM90]. estimator [Adh91a, AK95, AP93, AM99b, BS90a, Car93, Cha91b, Cha96a, CM92, CM96a, Che94, CT98, CP99b, Chi99c, Chr97, CC95b, Cor95, DJD96, DU91, DY96, DD96, DSTW91, DT98, FZ96, FRS97, Fir99, GP93, GJU94, GN91, GT97, HH94, HK90a, HL94, HC97, HK90d, HB91, IT93, Ino99, KSA98, KC91a, KFAH92, KO94b, KA92b, KM96, KJ91, Let97, LZ98b, Luc97b, Mae90, MB91a, MGLF98, MR95b, MPR98, NS90, NM94, Oht93a, Oht93b, Oht96b, Oht96a, PFN92, Rai96, RG96, Sar92b, SA90c, SG91b, Str95, ST96, SI97, SD92, Tak97b, TZ90a, VV96, Wan96, WM99, Wu96, gWyJ97, Yan98, Yon95, YP93, Yu97, ZS97, ZZ99, ZK94, ZL91]. Estimators [FKY91, Ahm97a, AN90, AHJ94, AK98, BPT90, Bas93, Bol90, BS98b, BS99a, BB90d, Bre90, BH99, CT93a, Cy95, Cha93a, Cha91c, CM92, CP96c, Cho94a, CL96, CW98, CH91c, CR92b, CC91b, CC91c, CBRS90, Cza93, DB95a, DAM99, Den98, DH96, Dou93, Dre96, Fou91, Fro93, Fuj95, Gao95a, GP90, GR90, GS90, GGO96, Gri90, Gro94, Gro96, Hal99, Has95a, Hil93, HC90, HC93, Hür98, IK97, Jia99, Kc91b, Kan90, KKS99, KS98, KS97, Ki92, Kib96, KI94, Klo90, KK94, KA97b, LY97, LLH90, Liu90, Lou98, LK98, Mae96, MVM90, MEC93, Mor91, MM96b, MPH97, Muk95, NS90, PM91a, PMD95, PF93, PG95, PK91b, PR93, Pra89, PS90, Pra91, PS92, Rao91, RS95, SS91a, SK93, Sch97, SJ91a, SNK91, SS91c, SC93, SC96, SSSS94, SPS94, SUS95, Sob95]. estimators [SY93, STS95, SO95, Sri95, Swi98, Ton96b, Ton96a, TS90, TS94, TS92, Ull92, Wan94a, Wan94c, Wik98, WP92, Yan90b, You93, YW98, YT95, ZW99, hZ93, ZK90]. Euclidean [KL96]. Euler [Kem92]. Evaluating [Bak91, Ibr91, Gar96, JP93, JN96, KAHK94, PJ99, Tho96b]. Evaluation [ACR91, AM93, LWH98, LW94, ML91, SW97, Boi91, Boz90, Cor90, Gen97, pL98, Moh90]. event [GJO90, Lin99, LL91]. events [DF94b]. everywhere [Tho96a]. EWMA [Luc92b, Par98, SW97]. Exact [CB98, Chu92, GG92b, Oht96a, RD96, SS99d, US92, vdMvdM90, Abt92, AC95, Buo90, BD93, CG92c, GGO96, God92, HL93a, JK91, Kab90, Kan93, KT91, LN90, Liu95b, NvdMM90, Oht93b, Smi92a, SS94e, Szr95, Ull92]. Exact-and [vdMvdM90]. example [FKR91, FR93b, Sap94, Tho96a]. examples [SSCS99]. exhaust [LdBPR92]. exhibiting [Ng90]. Existence [WD97, Dem96, GG96, Gup92a, KK98b, gWyJ97]. existing [Koo93a]. expansion [Cha91b, Nis94, Qum97]. expansions [FS99, GS98a, KL96, Ong93, Qum94, TCT90, TM96a]. expectancy [RDA92]. Expectation [KK91c, SH94, LZZ99]. expectations [ES95, GS91a, Rag90, Ryc93, ZRM90]. expected [Gar92, NM92, Nor93, RC95]. experiment [Ali90a, KLH^{+}91, MM96a, SS94d, Tur95, Yip96]. Experimental [RW94, Vau93, BW93, CG92c, KH90, Ras90, Sub92, WK91b]. experiments [AZYK98, Cha92b, Cro91, DG97a, Den98, FS95a, GJ97, Han91, IY99, KCM94, KA92a, KA97a, Lai98, Llo95, ML93, Mon99, PMTP93, WS91b]. experts [GdWG91]. explanatory [EGM95]. explicit [Bou90a, CK97a]. Exploratory [NPA93, CTT91, Kan90]. explosive [And90, And91a]. Exponential [Lai99, YB94a, Age99, AK91, Ana92, Bal94, Boy95, CP90, CP96b, CPK92, CH94, CS97a, CTP93, CW92a, Chi92, Chi93, CH95b, Chi95, CBFCN99, Dem96, DG97b, EP91, FK94, FTV99, GZ97, GCD94, GGVM98, GK98, HK91, HK92, Han92, Han93, HD99a, Heb94a, HB99, JP99, JN94, JP91a, JP91b, JC98a, JC98b, JN91, Kun95, KM93, LL90, LL94, LBBCMR97, MT90, MH95, MSMP97, MAS94, NN91, Pan97, PP92, PG92, PG95, PC95, Sat90, SJ91b, Sha92a, SGN93, SUS95, SH90, SD95, Tah91, TV91, Ton96b, VT93, VS91, VR96, Won95, YK90, ZW99, dSC99]. Exponential-type [Lai99]. exponentiality [Asc90, BB90a, BH92a, FN92, GW99a, Han97b, HR98, Hen92, SJ90, SJ91c]. Exponentially [Ami93]. exponentials [SS99d]. exponentiated [MH96]. exposures [Nor93]. expression [Bou90a, Hid91]. expressions [KO94b, Mat91, MP91a, WP92]. Extended [NS97, Yok96, Dem92, Fuj99a, LM94]. Extension [CS96a, Dha95, Ein96, TL91, MS90b, PS95b, YJF91]. extensions [Kem92, SLS96]. extinction [DK91, KE93]. extra [JG95, LD95, ML91]. extra-binomial [LD95]. extra-Poisson [JG95]. extra-variation [ML91]. extrapolation [And97, Jan95]. extremal [BP96]. extreme [Ahn94, AG96, AS92, BCA93, BGP95, Bho97, Chi99b, DM96, DM94, DH96, DD96, DT98, FKT95, Hug91, KSmS90, LR98, MD94b, Olu94, RW96, Sca92, SY91, Tsu98]. extreme-value [AS92, Chi99b, DM96, DD96, DT98, FKT95]. extremes [Ang92, De 92, Hüs96, TM91, Tho93].

F [Ahn92]. F-probability [Ahn92]. factor [Aga90, CTT91, CC90b, GVSM92, HC94, Kan90, KA91, KWtB96, KML91a, KF97b, LP98a, Mon99, PB98, aS99, TWM91, TL99]. factorial [AF95, BW93, BD92, Bre90, GP92, Hyo94, JP92, KA91, KA92a, KA97a, Lee94a, LS97b, MM96a, Ng94, Yua92]. factorials [Gup91, Gup92a]. factorization [Kab91b, PS93b]. factorized [PZ99]. factors [AB98, BKS99, Gal94, GL97, Iwa94, KG90, PM92, PS99b, WG96, You92]. fail [Fan91a]. Failure [YB94a, AHJ95, Cha96a, DQS98, FN92, GZ97, GJ92b, GGL96, GK98, GGG98, JP90, Kul99, Mad93a, Nab93, PG92, PG95, Rao92, RT91, RZ94, SJ91a, Tah91, WC96b, Won95, YL92, YL93]. failure-times [WC96b]. failures [CP95, KM99a, TZ90b]. fair [HJ90]. fairest [GL90]. fallacy [BC90]. familial [BB93a, GC99, Mak94]. families [AKH91, Car94, CS97a, Con94b, Fer90, HFP94, Hür98, Mac92, PB91, PB92, Sat91b, Ton96b, ZW99]. Family [RG92, AN90, BP99a, Bri97, CF96a, CBFCN99, DQS98, GCD94, GGVM98, HL94, Jan98a, KM94a, Lee96, LD99a, LP98b, LB98, LBBCMR97, Luc92a, MSMP97, MK94, MH96, NS97, Olu94, PB90b, PC95, RC94, SD95, TK92, dSC99]. Farlie [DF95]. Fat [GSD96]. fatalities [Jan98b]. feasible [Oht93a, SG91b]. features [YLB98]. feedback [Luc92b]. FGM [SJ93]. field [Mar90b, YN99]. fields [But98]. Fieller [Dav91]. fifteen [Bra90, Ley90]. filter [Cho94c, Dem92, Tap98, YAM91]. filtering [CM94a, CAS97, YC99a]. filters [TM96a]. Finding [Bor95a]. Finite [BR90b, ALW91, Arn91, Arn92, BGMT90, CT97, FT99, Fir99, Gro91, KM94b, KS94b, KJ91, Liu90, LR93, MBR91, Muk94, Muk95, Pra89, PS90, Pra91, PS92, RL91, SSSS94, Ste91, TC90, TC92, TZF97]. First [AOA92, AM93, Ara90, AR90, BMM^{+}91, Bra90, Cho94b, FZ96, Fre92, Haj90, Hal99, KB90, SJ97, WK91a]. first-order [AM93, Ara90, KB90, WK91a]. Fisher [Ang91, Awa92, BG91, Bro98, GW99b, LF91, NvdMM90, TZF97, Tur95, Zog98]. fit [AAW93, AS92, BR97, Che99, CHC98, CGE99, Djo95, Epp95, EHL93, FM96, FKY95, Fro96, Gau98, Gho91, GW99a, HK95, Hon98, KP92b, KHJY97, KF97a, Koz90, Kul90, LZ96a, LN90, MMPV96, OR98, Ozt91, RV95, RRrA91, Sch91, SBS96, Sun97, Swa99, VR96, WS94, ZFP90]. fitted [LF91]. Fitting [CM96b, Gao94, Mon99, Kab97, hK97b]. five [KS90b]. Fixed [HEB92, Tah92, AM99a, Che95b, Cos98, EB98, FTS94, Hol95, LY94, MY91, Ull92, WK91b]. fixed-size [AM99a]. Fixed-width [Tah92]. flats [LS97b, Lia99]. flexible [Hen92]. floating [Cal91]. flood [MH96]. Florida [Bra90, Ley90]. fluctuations [Cha90a]. fold [Güv95, Mis96, PB93, TC91, YB95]. folded [PP90]. foldover [Joh95]. Follmann [Wan98a]. follow [TZ90b]. follow-ups [TZ90b]. following [CH95b, Chi99b, Liu97a, TY99, Vel93]. food [LF93]. footrule [CN99b]. forced [bL95]. forced-in [bL95]. Forecasting [Lee95b, MO91, Che92b, Dia90, FR92, MLGD98, Wan98b]. forecasts [Liu95a]. form [EJ93, Fou91, Shy98, Vän97a, Wil96]. forms [AM93, FMV98, MS90a, Min99, QCL98]. formula [CM95c, PZ99, SK97a, SKS93, Yam97b, Yam97a]. formulae [Fri91]. formulation [CZ97b]. formulations [Ong90]. Four [Pei92, BS90c, BKS99, Mis96, Sca92]. four-fold [Mis96]. fourfold [Lev92]. Fourier [BK97b, Che94, Hon91]. Fourth [KT91, Kun97, TK91a]. Fourth-order [Kun97]. fractional [AF95, BW93, Hyo94, KA91, KA92a, Li95, Mat91, Yua92]. fractionally [Swi98]. fractions [LRD94]. framework [DH99]. Franklin [SS93c]. Fréchet [HSS93]. free [CSK^{+}98, CJ98b, DM91, GS94, Guo96, KGM94, Man91, Pan96, SM91a, Sha90, SGB98, dP90]. freedom [HC94, KAKW99, Oht96b]. frequency [But98, DY96, Kat95, KM99b, Llo99, Lop93, Mof95, Swi98]. frequentist [BS98a]. Freund [KM94c]. Friedman [BL90]. full [CC90b, CG90, SK95]. fuller [KS90c]. fully [EGM95, HCW91]. Function [RG94, Abd93, Abd98, Age99, SAP92, Ana92, BB95, BR90b, BDF95, Car93, Car90, Cha96a, CBR91, Chi99c, Cho90a, Chu94b, CK97b, De 92, DK95, EJ93, EPdR99, EHL93, FK91, GH92, GM90, GG93, Has95a, HS90c, HB91, HB99, KK91a, KS90a, Koc95, KA92a, KKK90, Let97, Liu96, Lou98, MK90a, Mor91, Nai96, NPA93, PZ99, PR92, PSM96, Pan97, PFN92, PF93, PN95, Par93c, PGC98, Rai96, Rie99, RZ94, RV95, RRrA91, RG96, SFT99, Szr95, TZ90b, Tsa90b, VSS95, VR96, WW90, Wri98, YY93, YP93, YW98, Zhe99]. Functional [AM91, NTS91, Abd95, Ban91, BLS97, BP99b, Edw94, Iso92a, Iso93, Iso92b, KM94b, Mat98, Mat99, Pro90, Rod92, TS95]. functional-coefficient [Mat98]. functionals [DR96b]. functions [Abt92, Blo97, BE93, CP96b, Car90, CR92a, CHS98b, Cha90a, ES95, Enc92, EJ93, FR93a, GPW98, GP94, JI99, Kan93, KD99, KS90a, KL96, Kot90, KM94b, LW90, Lee99, Leu94b, LH93, LY97, LL96, MM96b, New99, Oht93a, Oht93b, PB92, Qiu94, RDA92, RTS92, Rei90, Rod98, Rug90, She94, SSW93, Sob90, SK91b, TCT90, Tsa90a, WL92, ZK94]. Further [FR93b, HC90, KF97b, MT92, HO90, Pun97, Sat91b, Tom92]. futility [SLL97]. future [CWH93, Mil96, Raq96].

G [Sar92a]. gain [HC93]. Gains [Hod90]. Galton [ES92, MGM98]. games [KK92]. gamma [ALW91, AK96, Ahn96, BSK95, BS99b, Bro97, CH94, CM98, Dix91, FN91, Ghi98, JSM98, JK91, KG92a, PT92, RRA92, RKN91, RM90, SY91, Son93, TM91, Tho93, Wit91b]. gamma-exponential [CH94]. Gasser [AP93]. Gauss [Gro96, Gro99]. Gaussian [AHAEH90, AHAEK92, Ami93, Ami96, Bar91, But98, Cho94c, Chr97, DN92, DY98, Edg96, GN95, HK90d, Hsi90, JS90, JS91b, Kar95, KM95, Kum92, Mar92a, PM90, PMD95, SS99b, SE97, Sun96, VR96, WF95]. GEE [ERO97, Hal99]. Gee-based [Hal99]. Geeta [Con90a, Con90d]. Gehan [LL98]. General [ACS95, Gal94, TK97b, TK99b, Ahm97a, AC99, AF95, BM90, BR93b, BT93, Gro96, Kim94c, KO94b, Loc90, Meh93, Ozt91, PGT98, SAK97, SK97a, SSSS94, SKS93, UNM92, Vän97a, Vän97b]. Generalised [MR95a, HC97, MY95, Mis96, Sin97, SG91b, YB93]. generalization [Bri94, FS95b, Ghi98, GGVM98, KS94b, Yam97a]. generalizations [CN99b]. Generalized [Ahs96, CF92, CF98, Jan98a, Kre92, Mad93b, MK94, SY94, SS93c, Sug99, SD97, Yue99, ZB97, AK96, AK95, Arn98, BPT90, BCA93, BA94, BL98b, BC97, Cha96a, CQ91, CF95, CB94, CW92b, Dem92, DL90, DF93, DT98, ESDC96, Fam93a, Fam93b, Fir99, Fro93, GT97, HK95, Hol91, Kab90, KK91a, KP95, LC96b, LZ96a, LM93, LD99a, LZ96b, LM94, Mar95, MB91b, NR99, NP92, Oht93a, Ong95, PMTP93, PS95b, PS99a, PC99, RRA92, RKN91, RDLT98, RC94, SR91, Sar92a, Sca92, Seg92, SNK91, SJ97, SPS94, SK96, SS99b, SFT99, Sun97, Sun98, Tak97b, Vos94, WLG99, WS91a, Yan90a, Yan90b]. Generalizing [CN99a]. generated [Iha93, Lee99, Olu94, Ong93]. Generating [Heb94a, DX90, HS90c, NPA93, New99, Rie99, RRrA91]. generator [PGC98]. geometric [AOA92, BS90b, DR96a, Yeh92]. Geometry [LK96, PRP92]. Gibbs [Bru95, Che92a, GS91a, Men95]. Gill [GP98]. Gini [CN99b, Gup99]. given [Gov90b, Rug91, Rug93]. GM [MU97]. GMANOVA [KKS99, MEC93]. good [LY94]. Goodness [BR97, CHC98, KHJY97, RRrA91, AAW93, AS92, Che99, CGE99, EHL93, FM96, FKY95, Fro96, Gau98, Gho91, GW99a, HK95, Hon98, KP92b, KF97a, Kul90, LZ96a, MMPV96, Ozt91, RV95, Sch91, SBS96, Sun97, Swa99, VR96, ZFP90]. Goodness-of-fit [KHJY97, AAW93, AS92, Che99, CGE99, FKY95, Gau98, GW99a, KF97a, Kul90, LZ96a, MMPV96, Sch91, SBS96, Sun97, Swa99, VR96]. GPED [Jan99]. Grade [Mie91]. gradient [Gil92, Vau93]. Gram [Nis94]. grand [SC96]. Graphical [Fun95, BL98a, CT92, JP93, JN96, MGR95, Ozt91, Ste91]. Graybill [Ino99]. great [DR96b]. Green [Dik95, Her92, Her93, Jie90]. grid [OS93b]. Griffiths [Yam97b, Yam97a]. Group [CHS98b, Dou97, Li99, Pat94b, Zha90, BMK93, Cha96b, GW92, GAN90, GT91, KK92, MSS97, MK90b, NS96, PB93, PG92, Ric90, Sas91b, SAK97, SMS97, TY99, Xu90, YC99b]. group-dynamic [KK92]. Grouped [Mof95, CM98, DRL97, Par93a, Tre94, WG90, YB93]. grouping [CC95a, MMPV98]. groups [GM94a, PPS97, Pos92a, Ros96, Wil92, You92]. Growth [Fol92, Bha95a, Dem96, JU93, RI93, TK93, Vas96, VV94, Von91, Yok96, Zhe99]. guaranteed [CH97, CS93]. Gumbel [DF95]. Gurland [Lee95a].

Haenszel [ZB97]. Handa [Sar92a]. hardware [Zha93b]. harmonic [SSCS99]. Hartley [CM92]. having [MBR91, Rao92, WJ90]. hazard [BE93, DG90, EPdR99, GJM96, Gil92, HH90, HO90, HPS98, sL98b, RTS92, WG90, Zhe99]. hazards [CP99b, DQS98, Gho91, GP93, Li99, Lin99, MS99, RDA92, Ver94]. Healy [Aos94]. Heine [Kem92]. heterogeneity [AG90, Sin91]. heterogeneous [TS94]. heteroscedastic [Aos94, Cor92, DGD92, Dut99, HK90c, HK94, Kib99, SW98, Wan93b, Wil91]. heteroscedasticity [GL93, Wil96]. heteroskedasticity [BD93, SA90c]. heuristic [MS95]. Hierarchical [NS94, Pep90, Ste92, Str91b, CC93, Dag99, DMG91, Dat91, GJO90, Had98, HC90, Jia99, QT98, Rod98, WZ94]. high [CR92b, MNS96]. high-breakdown [CR92b]. Higher [EF92, ATT99, Ein96, MC92, ZM93b]. highly [SD90]. histograms [Ati90]. historical [Che98d, KY91a]. HNBUE [POGP95, POGP96]. hoc [SS94d]. Hodges [Fri90a, Fri90b]. Hogg [HC96, RC96]. Homogeneity [ST99, CG91, Fam93b, Gil93, GD99, HD99b, KGM94, Lev92, MT92, Pos92b, RR97, SGB98, SD93, TSI98]. homogeneous [TS94]. homoscedastic [HS90a, RS92]. Honda [Lai98]. honor [HC96]. Horvitz [Adh91a]. Hotelling [FS99]. house [Mof95]. hull [LAC90]. Huntsberger [KFAH92]. Hybrid [GK98, DT98]. hyper [Den91]. hyper-Dirichlet [Den91]. hyper-Liouville [Den91]. hyperbinomial [Fer96]. hypercube [Bor95a, Han91]. hypercubes [Tan94]. hypercubic [Dut93]. hypergeometric [DP93, GT91, God90, LC99, SFT99]. hypotheses [AKH91, Ban91, CKH95, CK97a, Est96, Fri90a, GVSM92, Meh93, MRS92, SA90b, WM99, ZB93]. hypothesis [Fri90b, Gas92, Kab91a, Ksh98, Oh98, Tan98, TK92, TC91, UNM92, Wij93].

i.d [Xia95]. I.I.D. [Loy90]. identical [Bal94, BB95]. identically [Kam94, Sei98]. Identifiability [DX90]. Identification [DFT91, Kum92, Li95, DW90, DRS90, OT91, RS99, SK95]. identify [JP92]. Identifying [Gle96, PE92, Cho91b, GT91, Hon94, KS90b, KC90a, KC90b]. identities [JS91a, Tho93]. identity [CKH95, Gas92, JC98a]. ig [MH95]. ig-exponential [MH95]. ignorable [ZR98]. Ihara [Kan90]. ihe [Rie95]. II [Ana92, BGP95, CHS98b, DH90, FK94, FKT95, Fer90, Goo91, HO90, HH95, KY91b, KSmS90, Kul95, Men92a, PB90b, PB91, Pos92b, RSAJ90]. iid [SS93a]. III [BS90c, Fuj99a]. illustrations [Ste90]. illustrative [RT96b]. Image [KP93a, DL95]. images [NKN99]. imaging [EE93]. Impact [CM94c, Ng94]. Impacts [Ali91a, AA92]. imperfect [GZ97, LLP98]. Implications [Lye91, Hoc90, JBY90, MC92]. Improper [Had98, CT91]. improve [Wij93]. Improved [Ahm98, BB90d, DMG91, FS95a, Leu94b, Spu91, Stu94, Sun98, TC92, Hoo90, JK91, Lev92, NS90, PM94, PC95, Pra89, PS90, Pra91, RS92]. improvement [BK97a, Cor91a, Gro99, HS90c, Ino99, Mye91, Nag95, WS91b]. improvements [NS99]. Improving [DJG94, MN98, Tho96b, Rao91, SA90c]. impurity [SA98]. imputation [Bel93, SSCS99]. inadequacy [Ali90b]. Inadmissibility [Fun91, HK90d, MD95, PF93]. Inagaki [Has95a]. inc.Ome [Lei90]. incapability [Che98c]. incidence [WP92]. inclusion [AR90, DSTW91, Swi98]. income [DX90, Lei90]. incomplete [AAB96, AC95, AR90, BN91, BBK98, Bel93, BDF95, CP90, Cy95, CK96, Dri91, FBD99, Gra98, Gro94, Kri91, KM95, LT92, Lee95b, Leu94a, MML99, MPQ91, PD92, PS91, Sco97, SSS97, SK97a, Sin92, SB91, Wee92]. Incorporating [KY91a, Che98d]. Incorporation [SLL97]. increasing [FN92]. increment [CW94a]. indefinite [EJ93]. independence [AH97, CG91, CN99b, DF95, DM96, HK91, HK92, JZ96, KK98a, LS94, Lye91, PGC97, PGC98, SJ93, US92, YSM92]. independence-distribution-preserving [YSM92]. Independent [Var96, ATT99, CGP92, Den94, Ete91, Fan91b, Gru93, Koo93a, KMBTG99, Koz94, MJ92b, Sat90, SC93, Tsu98, Wil92]. Index [Sha98, AS94a, CH95a, CP97b, Che98c, CO90, DD96, Joh96, KKK91, KK91b, MLH92, MPR98, SB97, ST93b, Sun96, Wri98, Zha90, ZSW90, ZL91]. index-lag [Zha90]. indexed [KG95, SA93, SB93]. indicator [Kai96]. indices [CXZ90, PLC98, Vän97a, Vän97b]. indirect [CM95a]. individual [BC90, pL98, RDM92, YR94]. individuals [Woo92]. inequalities [Bro90, Ete91, Kot90, Lai99, MS90a]. inequality [JBY90, Wan96]. Inference [Rod98, SS90b, Sco97, ST93b, Zac91, Aos94, CP97a, CH96, DY98, ES94, Eva97, GMR97, GT92, Han92, Han93, HK90c, HK97a, Koo93a, Kot90, KMBTG99, Li94, Mar96, MF97, MGM98, NK99, RD96, Sha92a, SOSW94, Str91b, SVM95, Won95, Yeh92, Yip91]. Inferences [Hsi90, Che92b, LLL96, SM91b]. inferential [Cer97, PLC98]. inferior [Nor93]. infinite [Gle95, Gle96, Nas91]. Inflated [GGT95, MS98]. Influence [Cam93, Kim94b, TWM91, You92, BL93, Bro97, CTT91, Fuj94, Fun95, Fun96a, Fun96b, FT97, Kim95c, Kim95b, Kim96b, Kim96a, KKT98, KF97b, LW90, LD99b, Lin99, Mar92b, She94, TC90, Vos94, WL92, Wil95, vHL92]. Influential [Vel92, Abd95, Hid91, Hos90, KC90a, KC90b, RS99, TCSJ94]. Information [Gen97, Hon94, KM91b, KM95, LRD94, PGT92, Wan94b, AH97, Ali91a, Boz90, CP90, CN99a, CT93b, CS93, Che98d, CJH95, Fan98, FY95, Gue94, Hau91, HC90, HS90c, HK99, HD94, Hyo94, KI94, LR92, Mad92, MGR95, MS90c, NC97, PG94, PS92, SLS96, Sha99, SK97a, Ste92, TB91, TCSJ94, TZF97, Tur95, Wee92, YIK96, Zha98, Zog98]. information-based [Boz90]. informative [Ali90a, GM98b]. informatively [MD92]. infrared [NKN99]. inhomogeneous [Cy95]. inner [GS93, GS94]. innovations [Kum92, Pei90, TWB99a, TWB99b]. insensitive [WW90]. inspection [WZ94]. Institute [HC96]. Instrumental [TL99, as97]. intake [LF93]. integral [Che94, GP98]. integrated [Abd93, CMS99, CZ97b, FM97, Sri99]. intensities [Won95, YB94a]. intensity [HB99, JSM97, Law93]. inter [SLS96, Sin97, Wee92]. inter-arrival [Sin97]. inter-block [SLS96, Wee92]. interact [RL97]. Interaction [CM96b, Jär99, MGLF98, Ng94, SS98, SSM90, SS94d]. interactions [KA91, SS91b, Zha92]. interblock [SK97a]. intercept [SA90a, WL93, Wil98]. intercepts [Lim93]. interclass [Fuj99b]. interim [SLL97]. intermediate [Fal98]. Internal [FK94]. interpenetrating [Adh90, Koo93a]. interpolability [BB97]. Interpolation [PM91b, Nie97]. Interpreting [RBM95]. Interval [Dav91, ZSW90, Che99, CH95b, Chi99b, GW91, HEB92, IK97, KI94, Meh93, Nag95, Nag96, Nag97, PC98, PL99, SL94a, Sha99, SO95, Sri96, Ste91, Tah92, Vau92b, Wan94a, WK97]. interval-censored [PL99]. interval-hypotheses [Meh93]. interval-scaled [Che99]. intervals [AH93, AW99, BS96, Bro95, CGP90, CP92, CH96, CM96a, CWH93, DH92, DY95, DG97b, EB94, Fay99, FTV99, FS95a, Fuj99a, KMM90, KK94, Kra92, KK91c, KML91b, Liu93b, Llo95, MG91, MD93b, Mor90, PB98, PB93, SA90a, SET94, SG91a, SS99d, Tak94, Tu99, WL96, WSM91, Wit91b, YL91, YB94a, YB95, vdMvdM90]. intervened [Sco95a]. intervention [DW90, Moh90]. interviewer [Sin90]. intraclass [BH99, CD98, Jel99, KS94a, PB90a]. Intrinsic [De 92]. Introduction [Hog96a]. introductory [Moo96]. invariant [GM98b, HZ90, Hu94, SE97, Sub91, Sun99, YP93]. Inverse [Ami96, Min99, RK96, AHAEH90, AHAEK92, Ami93, Bar91, CP94, DY98, Edg96, HK90d, Hsi90, JS90, JS91b, PM90, PMD95, PM91b, RBC97, Sar97, Sar99, SNM97, SS99b, Sri95, Sun96, Tak97b, TP97, WF95, ZL91]. Inverse-Gaussian [Ami96, Ami93, SS99b]. inverted [Heb94b]. Invertibility [DB95b]. investigating [RS96]. Investigation [CSK^{+}98, WL92]. investigations [FR93b]. invex [Han90]. Invexity [Rei90]. involving [Arn93, Fer90, GG92a, Tho93, Zou94]. IQL [SA93]. irregular [MD94a, NT92, SS94b, SK99]. Ising [Jär99]. isotonic [Cha96a, Shi95, WC96a]. issues [Ano96l, Li98]. iterated [cH91b]. Iterative [hK97b, AF98, AM93, HK99, Kan90, Par93b]. iteratively [Jor99]. IV [GP92].

Jaccard [MLH92]. jackknife [Cha90b, LLH90, Mae96, TG94]. Jackknifed [CC91b, CC91c]. Jacobi [Let97]. Jaeschke [Ein96]. James [Bas93, Hil93, MG93, ST96, YT95]. James-stein [MG93]. Jennen [AP93]. Johnson [Mac92, Nan92]. Joint [CC95a, Cho91a, DK95, FS99, GV90, JZ96, KMBTG99, Ng93, PZ99, PE92, Tsu98]. joint-risk [Ng93]. Jolly [Bur97]. jump [JI97, OT91, Qiu94]. jumps [Qiu94].

Kallinen [Gle96]. Kalman [CM94a, CAS97, Dem92, Tap98, YAM91]. Kano [Kan90]. Kaplan [Cha91b, GP98, GJU94, VV96, VV97]. Karber [GN91]. Karlin [Kea91]. Katz [CF96a, Jan98a]. kept [Nor93]. Kernel [GNR98, MQZ99, Abd93, Che90, Chu94b, CC91b, CC91c, DY96, GPW98, GP90, GR90, GP93, Hon98, Jon94, Lou98, MLGD98, MR95b, Tro93, Yon95, Zha98]. kernel-based [MLGD98]. kernel-type [MR95b]. kernels [HS93]. Kiefer [JZ96, Yon95]. kind [CS90b, Haj90, Jan97]. knots [Cha91a, RD94]. knowledge [FBD97, Res96, YT95]. known [Cal94, GK92, JS90, JS91b, JN91, KS90a, Lei90, ZM93a]. Kolmogorov [CC96, GG92a, KF97a, RR91, YP93]. Kolmogorov-type [KF97a]. Koreisha [Gle96]. Koziol [Dik95, Her92, Her93, Jie90]. Kruskal [LC91]. Kullback [HK99, KL99a, Rod92, Sha99, YIK96]. kurtosis [BS99a, Hen94a, Hen94b].

l [Lim93]. laboratory [SET96, Wen94]. Lachin [Fri91, Fri94]. lack [Djo95, LN90]. LAD [Liu92c]. lag [LTCC94, PE92, Zha90]. lagged [SU95, WA96]. lagged-dependent [SU95]. Lagrange [Jan97, Min98, Min99]. Lagrangean [CF96a]. Lagrangian [Con94b, Jan98a, SNM97]. Laplace [BB90b, LBB92, RD94, STR97]. Large [CS90c, Dut99, HK91, HK92, Kul94, Kul95, LH99, Lin97, PMSM97, SP94, ATT99, Car94, Sob96, Xia95, Zah90]. large-deviation [ATT99]. large-lecture [Zah90]. larger [CPK92, EP92a, KHR97]. largest [MAS94]. last [EB92, Luc92b]. latent [Dag99, SV97]. latest [LL93]. latin [Han91, Mar97, Tan94]. lattice [AAB96, CK96, Epp95, NMB98]. law [CGP90, CGP92, CP97a, CGE99, Gau98, cH91b, Xia95]. Lawless [KO94b]. layout [CG92a, Sha90, Tha92, WDH90]. lead [DD97b, Kmi95]. learning [Gar96]. Least [Bev92, Lu90, SPQ98, BSSB90, CM97, CH91a, CP96c, CT98, Dem96, GT97, IT93, Jor99, Kab97, KM96, Luc97a, Mar95, MM96b, PRP92, Qum97, Sar92b, SY94, SO95, Str95, Wan96]. least-squares [Mar95]. lecture [Zah90]. Lee [Lee95a]. left [CP96b, GJU94, JP99, PC98]. left-truncated [CP96b]. legislation [Moh90]. Lehman [Fri90b, GGG98]. Lehmann [Fri90a]. lei [GC98b, VP98]. Leibler [HK99, KL99a, Rod92, Sha99, YIK96]. length [Bar98, Cha92a, DRB93, GT90, GT92, KM97, WK97]. length-biased [GT90, GT92]. lengths [RWP95, SW97, Zha95b]. leptokurtosis [MN93]. less [Mil96]. Letter [Ano92o, Ano95m, Ano97l, Ano98m, GC99, Joh96, Kle94, Kum98, Sap94, She98, mC96]. level [CJ98b, Gov90b, Hol95, HD94, Hyo94, JP92, Lia99, MM96a, NHY95, Tsu99]. levels [AB91b, Rab98, Won92]. Leverage [Mar92b, ST93a]. Lévy [OR98]. liberal [Moo96]. life [AB94, BB90a, Cha93b, DL90, DR96b, FK91, GH92, KM91a, LP98b, Lin97, Lu90, Mad93a, Men92a, MPQ91, NK99, PFU91b, Rai96, RDA92, Rao92, RDA93, RBLN93, RG96, SJ91d, SJ93, YK90]. life-test [YK90]. lifelength [GB98]. lifetesting [Rao92]. Lifetime [FBD97, BE93, CP96a, NAW96]. lifetimes [CH97]. like [BLBE99, GJO90, KO94a]. Likelihood [FTS94, He90, SS99a, SS94a, Vos94, AF98, Age99, AKP93, Alb92, AK98, Bar95, Bha95c, BS98b, BS99a, BS99b, CP96a, Cha91a, CO91, Chu98, CBFCN99, Cza93, Dav90, DY98, ES92, FN91, FRS97, Fuj95, GB98, GMR97, Gru93, Güv95, Had98, HP92, HK94, HPS98, Hür98, HS99, JS93, JK91, Jia99, JT92, JG95, KD95, KK98b, Klo90, KS94a, KML91a, LM96, Law90, LAC90, LS97a, LH93, Lis92, Liu97a, LM94, MH98, MA95, Mak94, Mar92a, MGLF98, MM96b, Myo95, Nis94, Pai95, PB90a, Par93b, PG92, PG95, SS93b, SL99, SSW93, Vas96, WP92, Won92, YY95, You93, Zha94a, hZ93]. Likelihood-based [SS99a, Vos94]. likelihoods [Mar96]. Limit [Bri97, NR99, CS90c, ESDC96, Kul90, Ng93, NM92, Yon95]. limitations [KJ91]. Limiting [KK98a, ZK90, CBRS90, Gov90b, Min99, Sma93]. limits [BK91, CT94, GS93, GS94, Gro91, GF91, HB90, LWH98, Pap92, SM94, Str90]. line [Ber98a, BB96, Dru90, KY93, Sri99]. Linear [AM91, BLS97, BP93, CK97a, Pas94, RKN91, Raq96, Abt92, ACR91, Ali91b, And97, Arn98, Ban98, Ban91, Bar91, BBK98, Bha95b, BR90b, Bol90, BC97, Boz90, Bre93, BT93, CKH95, CR92a, Cy95, CB90, Cha93a, CHA95, CHA97a, CBR91, CW93, CW94b, Che98b, CT98, CT91, CK95, CBG90, CB94, Cox92, DN90, DAM99, Dat91, DB95b, DSTW91, Enc92, FM96, Fou91, Fri94, Fuj90, Fun96a, Gao95a, Gao95b, GPW98, Geo91, Gri92, Gro99, Gru93, Guo96, Güv95, HK90b, HK97a, Has95a, Hau91, HC90, HC93, Hon98, Hon99, Iso92a, Iso93, Iso92b, Iwa94, Jah91, Jel95, Jel99, JT90, Kab91a, KO94a, Kej93, KP92a, Ki92, KS90b, Kib96, Kib99, hK97b, KW93, KL96, Koz90, KWtB96, Ksh98, Kun91, Le94, LW95, LC96b, LZ96a, LM93, LN90, LZ96b, LZ98a, LZ98b]. linear [LZZ99, Liu97b, LG98, LM94, Lye91, MY95, MGLF98, MU97, Mor91, MP97, Nag95, Nag97, Ng90, PB93, Par93a, PP92, Pei92, PR93, PL96, PGT98, Pun96, Pun97, Qum94, Qum97, RS99, RS96, RL94, RDLT98, Rod92, RBC97, SP94, Sar99, Seg92, SD91, SG94, SL99, Sil94, SPS94, SSM90, SP99, SM91b, SK96, Sri95, ST93a, ST99, SD97, Tak97b, TM96b, Ter90, Tim96, TC91, Ton96a, TS90, TS94, TL94, Tu99, UNM92, Var98, VSS95, Vos94, WL93, WLG99, WS91a, Wil96, Won93, Yan98, YB94b, Zhe99]. linear-plateau [Gri92]. linearity [BDT98, Mat99]. linearly [Bol90]. lines [Wil92]. LINEX [SAP92, PR92, Pan97, PFN92, Rai96]. link [Cza93]. Liouville [Den91]. Listing [Ano96l]. litter [SM96]. Liu [AK95]. LIVEX [PSM96]. LMP [Gov95]. load [CP92]. load-strength [CP92]. loading [CTT91]. Local [BL93, Bro97, Fal98, GPW98, Gro99, Kim95b, Kim96b, Kim96a, SY93, WX99, CS90c, EF92, Fun96a, FT97, GSD96, Kat95, LD99b]. Locally [HP92, RSAJ90, CG90, KP93a, RS99, Tho96a]. location [AK98, BS91, BGP95, BS92, BB90d, CT93a, CPK92, CH94, CTP93, CH95b, CL96, Con90b, DAM99, DS91, DG97b, EP91, GLL90, HL94, Her96, Hu94, Hür98, IH94, JP91a, JP91b, JC98a, JC98b, Kar94, Kum97, LL94, Liu95b, Lom99, MH95, MU97, MAS94, ML96, PK91b, RP90, RKN91, Sat91b, SM97, SGB98, SS99d, TV91, VT93, YV95, JR91]. location-parameter [Con90b]. location-scale [Hür98]. location/scale [Sat91b]. locations [MT90]. LOD [Rab98]. Log [Ahn96, Cha92b, FR93a, Fro93, Fun96a, HP92, hK97b, LR96, LW94, MGLF98, Nis94, Ong95]. log-concavity [FR93a]. Log-gamma [Ahn96]. log-linear [hK97b, MGLF98]. log-normal [LR96]. log-odd [Fun96a]. log-series [Fro93, Ong95]. log-transformation [LW94]. logarithm [cH91b]. logarithmic [GT96, KK90]. Logistic [Arn93, SV97, AS92, CMS99, ESDC96, FT95, Lee90, LD99a, PK91a, SGA99, SET94, SCXG96, SC92, TL99, Zha94b]. logit [Col98, Tre94, ZAW98]. lognormal [AK94a, LR98, LK98, Men91, PL97]. lognormality [AKH91]. Lomax [CR93]. long [Gup92b, OC99]. Longitudinal [Von93, Wil98, Ban98, Dam93, ERO97, Gal94, HAKM93, MM97, PS95b, SD97]. loop [Sug99]. loop-ordered [Sug99]. Lorenz [CHS98a, Gup99, MP91b]. loss [SAP92, BG91, CP96b, CK97b, DY91, FY95, Gar92, GG93, GGO96, Has95a, KA92a, KM91b, KM95, PR92, PSM96, Pan97, PFN92, PF93, PN95, PGT92, Rai96, RZ94, Rod98, Sob90, Sun98, Tsa90b, Tsa90a, WW90, Wan94b, Wan94a, Yan90b, YP93]. losses [BR92]. lost [KK92]. lost-games [KK92]. Lot [SA92b, Gov90a]. lottery [LdBPR92]. Low [Mar97, Jan95, Swi98].

Low-order [Mar97]. Lower [Iwa94, SK91b, Tak99, Tar91, AK91, GLL90, iK99, Liu95b, NM92, Rug91]. LQL [KG95]. LRT [LV98]. LSDSP [SA92b]. LSE [FZ96]. LSI [AJC99]. lying [Tan98].

M [Jai91]. M/E [Jai91]. MA [CCC96]. made [CK97a]. magnetic [EE93]. mail [SMS93]. main [Aga90, BW93, BKS99, CC90b, Gup91, Gup92a, Mon99, aS99]. maintenance [BK98]. Making [RT96b, GdWG91]. Mallows [AF98]. manifolds [Chi99a]. Mann [FR98, PC99, Ibr91]. MANOVA [KC90a, KC90b, aS99, SS94e, Tim96]. Mantel [ZB97]. manufacturing [DGMR96, DFT91]. many [HV96]. many-one [HV96]. MAPD [SB93]. Mardia [Hen94b]. marginal [Bha95c, Car90, Col98, GS91a, Rod92, ST90b, SD93, SK91b, ZK94]. Marginally [DLW96]. marginals [AOA92, GM98b]. margins [Che95b, Hal96]. mark [Ana97]. mark-resighting [Ana97]. markers [CS95]. market [GC98a, MN93]. Markov [Adh91a, AB92a, BGMT90, Ber99, BR91, FT99, Fyg91, Gro96, Gro99, Hoo90, KE93, KM97, KM99a, KM99b, KP93b, KKT98, Lee94b, MAPV96, MR95b, RR97, RC95, SGS94, SS99b]. Markovian [Arn93, ZM93b]. marriage [Fre92]. martingale [Lai99]. matched [BG96b]. matching [Gol90]. Mathematical [HC96, Ree96]. matrices [Cho90b, Hyo94, KG92a, Kim95c, Kri91, Kro94a, Mee90, MJ92a, PK94, RM93, SE97, TA96, YB94b, ZB93]. matrix [AM99a, Cal94, CTT91, CKH95, Cho94a, DJ98, IT93, JA92, Kab90, KKS99, Kim94b, Leu92b, MEC93, NS93, Pun96, SNK91, SH94, ST96, SI97, TK93, TZF97, Tur95, Yan96a, ZS97, Zog98]. matrix- [SH94]. maxima [HSS93]. Maximal [Ete91, Set90]. maximized [Rab98]. Maximum [CP96a, ES92, Gru93, Güv95, Hal96, HK94, HPS98, HS99, JT92, KD95, KK98b, Klo90, KS94a, LAC90, LS97a, Mak94, MM96b, PG92, hZ93, AF98, AKP93, Ano98n, AK98, Bor95a, Bor95b, BS98b, BS99a, BS99b, CH94, Cha91a, Chu98, CBFCN99, Cza93, FN91, FRS97, Fuj95, Hür98, JS93, Jia99, JG95, LM96, Law90, Lei90, Liu97a, MH98, MH95, Mar92a, MGLF98, Myo95, PB90a, Par93b, PK98, PG95, SS99a, VR90, WP92, Woo92, YSM92, You93, Yua92, Zha94a]. Maximum-likelihood [ES92]. maxmin [Fri94]. Mean [Dah95, KSA98, Abd93, AB92a, AB92b, Ara90, Arn92, BB90a, BLBE99, CC93, CK97b, CBRS90, Dou97, EP92b, Fri90a, Fri90b, GJ92a, GN95, HH94, JB93, Jin95, JS90, JS91b, Kar95, Kel96, Ker95, KS97, KI94, KR90, KM91b, Kun93, Lee94a, LP98b, LSU90, Lin97, Liu90, LK98, NK99, Nag99, Ng90, Oh98, Oht91, Oht96b, Oht96a, PB98, PM90, PR92, PK91a, PL97, Pra89, Pra91, Rai96, RS92, Rob94, RZ94, RG96, Sar91, Sat90, SJ91d, SGJ95, SC96, SMS97, Son92, SG91b, ST96, SD92, Tan98, TB91, TS92, TC92, TK97a, Wan96, YY95, YT95]. means [Ahm91, Ahm97a, Ali90b, Bar91, BR91, BB93b, CWH93, FTS94, Fri94, KS98, Koz94, Lee95a, Liu96, MT92, Men91, MA92, NS94, SS93b, SC93, SSW93, SSSS94, SG91a, Str91b, Sub92, Sug99, TA96, TL99, WK97, Yan90b, Yu98]. measure [Boz90, CTT90, Fra91, JP93, Kea91, Kim96a, MMD93, MB91a, PMD95, SA98, Zog98]. measured [Che97, Men97]. measurement [BT93, CC95a, DD97b, Edw94, GBC99, Gri92, KF90, Muk94, Muk95, SW99, TL99, Wan93b, WLG99, YIK96, as97]. measurements [Afs90, FWG94, FMV98, KAKW99, Lai98, Par94, PPS97, Pat94b, Udd98, VV94, WK91b]. Measures [GC98a, PJ99, RS99, Boi91, CTT91, DJ98, Fun98, Kre92, KF97b, LLH90, MSMP97, NSS99, OW90, OW93, PMTP93, PMSM97, PLW93, RRAB93, RD96, Sal90, Vos94]. Measuring [BSSB90]. mechanical [CP99a]. mechanism [PLW93]. median [CW90, ID98, JGMGMR98, Luc97a, MS90b, SL99, Str95, Tho96a]. medians [Kum97, Men91, Rao90b]. medium [DRB93]. meeting [HC96, HC96]. Meier [Cha91b, GP98, GJU94, VV96, VV97]. meiosis [ZS98]. memories [Hog96b]. memory [Gup92b]. Menten [SW98]. messy [SD95]. meta [BS98a, HD94, LSR94, Mi90, Rog92]. meta-analysis [BS98a, HD94, LSR94, Rog92]. method [AAG98, BK97b, CHS98a, CGE99, DJD96, DN90, FT99, FN91, FZ97, Fin90, HL92, Ibr91, JP93, JN96, JP92, Kan90, Kea96, Kel96, LV98, NS99, NTS91, SK95, Sha98, SK99, Vau92a, Wil95, YJF91]. methodologies [MA92, SAK97]. methodology [Boh96, HD94, Mye91, Yan96b]. Methods [BMK93, Ana97, BC90, But98, DRL97, EB92, FM92, Hug91, HCL94, LV98, Luo94, Mar90b, MLGD98, MY95, OT91, Rao91, RW96, Sar92b, TCT90, TC90, Vas96, WK93, ZB93]. metric [Che99]. mice [JU93]. Michaelis [SW98]. mid [PM91b]. Midzuno [KS94b]. migration [GS91b]. MIL [SA93]. MIL-STD-105 [SA93]. minimal [Mig98, Shi95]. Minimax [GG94a, BR92, JS93, MD94b, Muk94, Muk95]. Minimaxity [SC93, SC96, AB91a, PFN92]. Minimization [Ali90b, Arn93]. Minimum [Bor95b, DH96, DRL97, Fer90, Luo94, Abd93, Ali91b, Ano99m, Boy95, CH94, CJ98b, DY96, Fri92, HK90d, HB91, IU93, KI94, Kro97, MH95, Oht96b, Oht96a, SG94, gWyJ97, Yan98]. MINQUE [Bre90, LC96b]. mirror [DL95]. mirror-image [DL95]. mis [Gil91, Gil93]. mis-specified [Gil91, Gil93]. misclassification [CK95, Fun96b, LW90]. misclassified [CHC98]. Mises [DM96, DRL97, Gho91, RV95, RV99, Sco99, VR96, YY95]. Mises-type [DM96]. mismodelled [Che97]. Missing [SD91, Bev92, DG92, FTS94, FWG94, JS93, KS94a, KM91b, MEC93, Mon99, NM96, PLW93, SK97a, SY94, SGD90, Sub92, TK93]. misspecification [AK94a, Hal99, Kro94a]. Misspecified [JR91, ERO97, Kab97]. Mixed [PP92, Aga90, BBK98, Boi91, Bre93, CT93b, Che98d, CBG90, CA90b, DMG91, Dat91, HPS98, IT93, JP90, Jel95, Jel99, KAKW99, KP92a, Klo90, MY95, SPQ98, STS95, SO95, SS99c, WB97, WLG99]. mixed-linear [Jel95]. mixed-model [KAKW99]. mixing [LC96a, LR93, RL91, S99, Tar91, Yan98, Yu91]. Mixture [Cor91a, sL98b, LP94, Ong90, ALW91, AHAEH90, AHAEK92, CS97a, CT91, Cro91, FY95, JN96, Kim94a, Kro94b, LBB92, LR93, MBR91, MML99, Phi96, Pre98, RRA92, RL91, Sei98, WX99, Yip91, Yip98, Yu91]. mixtures [CH94, Cha92b, CT97, Heb94a, Heb94b, Ker95, Leu92a, MH95, Ong93, SE97, Wit91a]. ML [GG96, MY95]. MLE [Fun91, Mi90]. MLEs [BLF99]. mode [ST90a]. Model [AM91, CS97b, DH95, EE93, Gle95, Ham91, Isa92, Jor99, Abd95, Adh91a, AF98, Aga90, Ahm98, Ahn94, AHAEH90, AHAEK92, AHJ94, AHJ95, Alb92, Ali90b, Ban90, Bar98, Bar89, BBK98, Ber99, Bha95b, BK91, Boi91, BR90b, Bol90, Bre93, Buo90, BT93, BH99, Bur97, Byr96, CP92, CT93b, CT99, Cha91a, Cha92b, CHA95, CHA97a, Cha91c, CM94b, CM95a, CM95b, CT98, CP99b, Cho94a, Cho94b, CG90, Col98, CF92, CBG90, CA90b, DGD92, DF95, Dem92, DW90, Dha95, DG90, Dik95, Dut99, ÉA95, Fam93a, FZ96, FTS94, FM96, FY95, Fri92, Fuj99a, Gao97, GYH95, GJO90, GS95, Gho91, GP93, GS90, Gil91, Gil93, GBC99, Gle98, Gri97, Gri92, Gro96, Gro99, Guo96, GL94, GN95, Had98, HEB92, HK90b, HK90c, Has95a, Hau91, Hei93]. model [HK95, Her92, Her93, HH90, Hon98, Hon91, Hos90, HD92, HK94, Huw93, HS99, Iwa94, Jah91, Jan95, Jan98b, Jär99, Jel95, Jel99, Jie90, JT90, JU93, Kar95, KKS99, KK92, KAKW99, KP92a, Kib96, KY91a, KB90, Kla96, Klo90, Koz90, Kun95, KM93, KM94c, Kun91, KM96, Law90, Law97, Lee94b, LC96b, Lee98, LZ96b, LZ98a, LZ98b, LZZ99, LH99, LP94, LBB92, LLP98, Loc90, Lu90, Lu93, LM94, Mad93a, MAPV96, Mae90, MSS97, MC91, MN93, MNS96, MMPV96, MEC93, MML99, MU97, Muk97, MP97, Nag95, Nag97, NC97, PB98, Pai95, PB93, Par93a, Par98, Pas94, PP92, PG92, PR93, Pep90, PL96, POGP95, POGP96, Pre98, Pro90, Pun96, QT98, RZ98, Rag90, RAP95, Rod92, SLS96, SK95, SAS91a, SJ97, SH94, SJ93, Sha92b, SG94, SNM97, SB95]. model [Sil94, Sin91, SS93c, Sin97, SM91b, SK96, SW98, SY93, STS95, Szr95, TM96b, Tim96, Tom92, TC91, Tre94, TS95, TK93, UNM92, VP98, Ver94, Von91, Wan93a, Wan94b, Wan96, Wan93b, WL96, WB97, Wil96, WP92, WK91b, gWyJ97, WX99, Yip91, YL92, YL93, Yip98, YB93, Zha93a, ZM93b, ZAW98, vdMvdM90]. model-and [GS90]. Model-robust [Jor99]. Modeling [GGG98, MC91, Ahn96, Bre93, CG92b, GS95, GT90, GK90, MS99, OC99, SSCS99, Ste92, ZS98]. Models [STL91, Sob92a, AB94, AKP93, AM93, AVB96, Arn91, Arn92, ACS95, Arn98, BM98, Ban98, BS94, BR93b, BF90, Bha95a, BLS97, BM94, BS90b, BC97, Boz90, BDT98, Bri97, CC93, CKH95, CCC96, Cy95, Cha93a, CMS99, Che92a, Che92b, CW93, CW94b, CP96c, Che98b, CL97, Con90d, CB94, Cor92, CF96b, CBFCN99, CM96b, Cza93, DP97a, Dag99, DMG91, Dat91, Dav91, DB95b, Dem96, DL90, Dia90, DX90, Djo95, DH95, DSTW91, ERO97, EB98, EGM95, FRS97, FKY91, FKY95, Fol92, FM97, FM92, Fro90, Gab92, Gab98, Gao95a, Gao95b, Gen97, Geo91, GZ97, Gle96, Gov95, GLL90, Gup92b, Hai93, HK92, Han93, Han91, HK97a, Har90, HL93a, HC90, HC93, Hon94, Hon99, HPS98, Hüs96, JA92, Kab97, KO94a, Kib99, Kim94a, hK97b, KW93, KMBTG99, Kum92, KKT98]. models [sL98b, Lee90, LAC90, LT92, LPK95, LZ96a, LS97a, Lee97, LN90, Li95, Lia94, Liu95a, LD95, Lye91, LK98, MA94, MA95, MGR95, Mar90b, Mar92a, Mar92b, Mat98, Mat99, MPB90, MY95, MR95a, MGLF98, MMT97, Mof95, Muk94, Muk95, MF90, Nab93, Nai90, NT92, Ng90, PE92, Par93b, Par94, PK98, PC95, Pei92, Per91, PM93, PS93b, PL99, Pun97, Qum94, Qum97, RDA92, Rao92, RL94, RT91, Rob94, RDLT98, RBLN93, SA90a, SGA99, SET94, SP94, Sco95b, Seg92, Sei98, SD91, SPQ98, She94, SAK97, SL99, SS94a, SS94b, SY94, SSM90, SV97, SG91a, SU95, SC92, SS99c, Sun97, SD97, SS94e, Ter90, TL99, TWB99a, TWB99b, Tom92, Ton96a, TL94, Tu99, UG95, Var98, VV94, Von93, Vos94, Wan98b, WLG99, WJ90, WS91a, WS94, WG99]. models [WK91a, WA96, WDH90, Won92, XD92, Yan98, Yok96, YB95, Zha94b, Zhe99, as97, dMD95, dSC99]. Modification [Nai96, Bro98, Cha97b, HD92, Kub99, Nan92]. Modified [Chu98, CBFCN99, KG95, LM96, Liu90, SSA90, Chu94a, Fin90, GT92, GGT95, HK90a, KSA98, LZ98b, MS98, NN91, Olu93, YY95]. modulated [CP97a]. modules [ED93]. Mohanty [Sar92a]. Moment [MP91a, Wit91a, BS90c, BS99b, CC95a, ESDC96, GH92, Hal99, KT91, Rie99, SY93, TK91a]. moment-generating [Rie99]. Moments [Heb94b, LBB92, New99, RS90, Sei98, SA96, Son93, VI91, ALW91, BCA93, BA94, Car93, CJ98a, DJD96, Gil92, Heb94a, Her92, Her93, JS91a, KO94b, Kun97, LR98, Mae96, Mat91, PZ99, PS99a, SKS93, Spu91, Spu95, TM91, Tho93, Tsu98, Vän97a, ZM93a]. monetary [KA92a]. monitored [Wen94]. monitoring [Ber98a, CMD94, LRD94, Ric90, Sri97, Zha95a]. monotone [Fuj95, JT92]. monotonic [WG90]. Monte [KS90b, Lav92, NS90, SSSS90, SM91b]. Moors [MSS97]. Morgenstern [DF95]. most [CG90, GT91, Hu94, LL91, RSAJ90, Sin97, Tho96a]. motion [Bou90a, Liu97a]. moving [AM93, BDT98, Cho94b, KF90, MMT97, Rad97, SS93a, Sri99]. moving-range [Rad97]. MRBP [KT91, TK91a]. MSE [AK95, IT93]. MSE-matrix [IT93]. Multi [BSS94, CH96, CTP93, GW92, PT92, AFR94, Cer97, CS94b, Ete91, JW92, KAHK94, KM97, KM99a, KM99b, Mon99, Sri95, Tak97b, Ton96b]. multi-arm [AFR94]. Multi-character [BSS94]. multi-dose [KAHK94]. multi-parameter [Ton96b]. multi-path [JW92]. Multi-sample [CH96, PT92]. Multi-stage [CTP93, GW92]. multi-state [KM97, KM99a, KM99b]. multi-step [CS94b]. multi-univariate [Sri95, Tak97b]. multi-way [Cer97]. multicenter [KP92a]. multicollinear [KO94a]. multicollinearity [AA92, WS91a]. multicomponent [KM93]. multiconsequence [DW90]. Multidimensional [CC91a, BR91, Dav90, She94, SB90, Zha90]. multilinear [SH94]. multinomial [AN90, BS94, Bha97, CH91a, Dha95, Fan98, Fro96, Fun91, GG96, LL91, Liu94a, MMPV96, Olu93, SET94, TSI98, ZFP90, Kul95]. multinormal [Dah95, Gil92, TA96]. multinormality [KTTL99]. Multiparameter [DY91]. Multiple [Bof92, Dri91, EB92, Ki92, Mad93a, ZAW98, BS90a, CT99, CB90, CS96a, Che98a, CK95, Dia91, DM91, FK94, Fun95, Fun96b, Gao94, GYH95, GM94a, HD99b, Jah91, Kab90, KA97b, KG92b, KML91b, LM97, LG93, Lop93, Mak91, Qum94, RC94, Sug99, Szr95, TK94, VI91, Wan98a, WJ90, WK91a, Wil91, ZW99]. Multiple-objective [ZAW98]. multiple-parameter [ZW99]. multiple-recapture [GYH95]. Multiplicative [Tom92, BOR92, Koc95]. multipliers [TA96]. multiply [BGP95, FKT95]. multiprocess [SK96]. multiresponse [LN90, Wij93]. multisample [MS99]. multiscale [Wu96]. multisource [Sat91b]. multistage [AKP93]. Multistate [ED93]. multitype [DK91, KE93]. Multivariate [AAW93, Byr96, DR96a, GS95, Gue94, Hüs96, ID98, JH99, JT90, Min98, Par94, RP90, SW96, AB91a, AB92b, Ahm98, Ahn92, AAG98, Alb92, AP91, AJC99, ACK95, AK98, BP96, Bar98, Bar89, BR95, Bel93, Bil97, Bof92, Boi91, BSSB90, Boz90, Bri94, Car90, CB90, CK97b, Col98, DS97, DM96, DJ98, DFT91, Dun95, Fuj95, FT97, GJ92a, Gil92, GGVM98, GC98a, Ham92, Han93, Han96, HK90b, HK97a, HS90b, HZ90, Hen94a, Hen94b, Iso92a, Iso93, JN95, Jel95, Jel99, JT92, JBY90, Kab91a, KK91a, KAKW99, KS98, KP92a, Kib99, Kim95b, Kim96b, Koz94, Kre92, KR90, KY93, KML91a, KML91b, Leu92b, LR98, LSU90, Lu90, MW96, Men95, Nag99, Nag97, NvdMM90, NS93, Oh98, Ozt91, PZ99, Pei90, PL96, Per91, RA96, RS92, RAM96]. multivariate [SA90b, SJ97, SNM97, SY94, SFT99, SK91b, TCT90, TC90, TB91, TP97, TK93, Tsi99, TK97b, Var98, Wan96, Wan94c, WA95, Wes98, WK91a, YL93, YSM92, ZS97, ZM93a]. mutual [AH97, Gue94].

n [Hon99]. N.S. [QCL98]. natural [AB91a, HB99]. NBUE [Kan93]. near [Ano98n, LN90]. near-records [Ano98n]. nearest [Adh91b, Yu97]. Nearly [Pre98, Rak92a, Rak92b]. nearness [CH91c, HC97, SNK91]. necessary [CW93, FR98]. negative [AOA92, AP91, BL98b, CF95, Dha95, GG96, KS95, MY91, Myo95, New99, Ong90, TS92, WS99, XD92]. neighbor [Yu97]. neighborhoods [NKN99]. neighbours [Goo91]. nested [BL90, FM97, Güv95, Kim94c, MPB90, PB93, Rob94, Ter90, TC91, WJ90, YB95]. networks [CL97]. Neumann [Bri94]. neural [CL97]. neutrino [MJ92b]. new-better-than-used [SJ91b]. Neyman [Nag96]. no [CLL95, Mon99, SC92, WDH90]. noise [Cho94a, Cus96, Iha93, SS94c]. noisy [PGT92]. Non [GM98b, MS98, ST90a, SSM90, AF98, And97, Bal94, BB95, Cha93a, CM95b, CS97c, DN90, DB95b, FM97, GG96, GT91, Hau91, Kam94, Kan90, Kib96, Ksh98, Kum92, Kun91, LBB92, MPB90, MRS92, PC95, Pei90, RT91, Rob94, RM93, SS93a, SSSS90, Sei98, SA96, SK96, TK99a, Tem94, TWB99a, TK97a, WS99, Wan94a, WDH90, Won93, Xia95, YL92, ZR98]. non-adaptive [GT91]. Non-additive [SSM90, WDH90]. non-compliance [RT91, Rob94]. non-convex [TK99a]. non-existence [GG96]. non-exponential [PC95]. non-Gaussian [Kum92]. non-i.d [Xia95]. non-identical [Bal94, BB95]. non-identically [Kam94, Sei98]. non-ignorable [ZR98]. non-iid [SS93a]. Non-informative [GM98b]. non-iterative [AF98, Kan90]. non-linear [And97, DN90, DB95b, Hau91, Kun91, SK96]. non-negative [WS99]. non-nested [FM97, MPB90]. non-normal [Kib96, TWB99a, Won93]. non-normality [SA96]. non-oblique [MRS92]. non-occurrence [CM95b]. non-optimality [Wan94a]. non-overlapping [LBB92]. Non-parametric [ST90a, YL92]. non-responders [Tem94]. non-response [TK97a]. non-spherical [Cha93a]. non-stationary [Pei90]. non-strongly-balanced [CS97c]. non-testable [Ksh98]. non-uniform [SSSS90]. Non-zero [MS98]. Noncentral [EJ93, CS90a, Cho92, DG94, Kab90, Kab91b]. noncentrality [TM96b]. nonconsecutive [SS94a]. Nondifferentiable [CMS99]. nonhierarchical [hK97b]. nonhomogeneous [HB99, Kim93]. nonincreasing [WW91a]. noninformative [ÉA95]. noninvasive [CS93]. Nonlinear [Hun90, RBC97, TM96a, BOR92, BDT98, BT93, Cip98, Dem96, Djo95, FM92, Geo91, Hai93, JU93, KM96, Lee99, NT92, PC95, Ras90, SS93a, SY94, Str95, as97]. Nonnegative [Arn94, Hei93]. nonnormal [Has95a, TK92]. nonnormality [CL90, SKF94]. Nonparametric [Abd98, ARS95, BE93, CG92a, CZ97a, Cor91b, EPdR99, FK91, Fro90, GW94, HY97, JS95, KC98, Kat95, Liu92c, MLGD98, Par93a, Swa99, TZ90b, WW91a, WP91, WDH90, Yeh92, Ahm98, AM99b, Boh96, CB90, Chu94b, CLL95, Cor90, DD97a, DS92, EGM95, EHL93, Fre92, GPW98, GG94a, GP94, Hol91, Hor92, JI97, JI99, KK98c, Kot90, Koz90, Lai99, LW95, Lin97, MR90, NM94, SD92, TZ90a, Tiw94, Ull92, WP92, YL93, Zha96, dR94]. Nonparametrie [RG96]. nonpooled [KAKW99]. nonrandom [FP90, MM97]. nonregular [BLF99, KHR97, KM94a]. nonreplicated [SS94d]. nonresponse [KS97]. nonsmooth [Rei90]. nonstandard [Ker95]. Nonstationary [WA96, GG96, Nas91]. nonzero [Kar95]. Norm [Cza93, Bro90, MG93, PM93, SG94]. normal [AB91a, AB92b, AAG98, AS92, BR95, Bof92, BB93b, Buo90, CS90a, CC93, Car90, CZ97b, CT91, CO91, CK97b, Cox92, Dia90, DH99, EP92b, ES94, Fri90a, Fri94, Fuj90, Fuj95, GW91, GM96, Ham92, Hoo90, IU93, JB93, JT92, Kab91a, Kel96, Kib96, KY91a, KI94, KK91b, KS94a, KR90, Kri91, Kro94b, LC91, Lee95a, LR96, LSU90, Lin95, Liu94b, Liu96, Lu93, MBR91, MG91, MT92, Mat91, Men95, Nag99, Nai90, NS93, Oh98, Oht91, Oht93b, PM91a, PR92, Pap92, PK94, Pep90, Phi96, RGU97, RS92, Sar95, Sar91, SC93, SSW93, SPS94, Smi92a, ST96, SI97, Sug99, ST99, SB93, Tah92, Tan98, Tho96a, TWB99a, UI99, Var98, WK97, WG95, Won93, WX99, ZS97, Zhe99]. normality [Ahn92, AKH91, CR92a, CS96b, Dik95, Dun95, Fal98, FS99, HS90b, HZ90, Hen94b, Kro94b, LZ98a, Lou98, MSL95, Nai96, Per91, RM90, SA96, Ver96, Wu96, ZZ99]. normalization [NR99]. normalizing [TSI98, TK97b]. normally [GS92, Kar99]. norms [Vio95]. Note [Abd93, BK97a, Muk97, Adh90, And91b, AVN97, AP91, AK98, BPT90, Bas93, Bos94, CM92, EP92b, Enc92, FY95, Fri92, Fuj95, GK92, HK90a, HC94, IU93, JP91b, JC98b, KP95, KS95, Kot91, LC96a, Lav92, LC99, Lis92, Liu96, LB98, LL98, Mae96, Nas91, Nie97, NS92, OR98, Phi96, PS99b, SS93a, Sat90, Sat91b, SL94a, TSQ91, Tim96, Vas96, Wee92, WG99, Woo92, Yu91, ZQ97, Zha98, ZW99]. Notes [Mi90]. notions [KS95]. NP [Vau92b]. nuisance [DF94a, Liu92c, Rab98]. null [Cal94]. number [Ano98n, AS94b, CP95, Jan98b, LY94, MBR91, NM92, NM94, Nor93, PS99b, Qiu94, TC90, Xia95, YLB98, Yua92, YC99b]. numbers [Haj90, LdBPR92, Mad92, Sob96]. Numerical [Bre90].

OA [Tan94]. OA-based [Tan94]. objective [ZAW98]. objects [Nor95]. oblique [MRS92]. observable [FBD97]. observation [Ano98n, Cho94b, Kam99, Kra92, Luc92b, SSSS90, SS98, YB93]. observations [Abd95, Ali90a, Ali90b, BE93, EP91, FWG94, Fun95, Fun96a, Fun96b, Gov90a, Hid91, Hon99, Hos90, JN94, Kab91a, KS94a, KK91c, LR92, LR98, LG93, Mar92b, Mar95, NM96, PL99, RS99, S99, SS93b, Sei98, SS94a, SGD90, Sub92, TC92, TCSJ94, Vel92, WB97, WG95, WN97, You92]. observed [SS94c, Won92]. observing [Nor95]. obtain [POGP95]. Obtaining [GJ97, WL96]. occasion [EB92]. Occupancy [SD93]. occurrence [CM95b]. odd [Fun96a, Hyo94]. odds [Anr94, El 97, IY99, Lev92, MML99, OW93]. oil [Lom99]. OLS [BG91, Mae90]. omnibus [BH92a, Bro98, Hen92]. On-line [Sri99, Ber98a]. One [BK91, Cal94, Car90, OS93a, Özt99, STR97, Abt92, Baj95, BP99a, Bri93, BH99, Car94, CG92a, CH96, CG90, CBFCN99, DS91, DG97b, DM91, ÉA95, FK94, GS95, GD99, Gov95, Gup91, Güv95, HEB92, Hei93, HV96, KM94a, KA91, KM95, Lai98, Liu91, MG91, MN98, Mor91, Nan92, Oh98, Oht91, OSL91, PB93, RW94, Rao90b, RS90, Sri99, SLL97, TC91, WL96, dSC99, vdMvdM90]. one- [Rao90b]. One-and [Özt99, FK94]. One-cluster [OS93a]. One-dimensional [Car90, DS91]. one-fold [PB93, TC91]. one-parameter [BP99a, CBFCN99, KM95, dSC99]. one-sample [CH96]. One-sided [BK91, Cal94, Baj95, Bri93, DG97b, GD99, Lai98, MG91, MN98, Oh98, Oht91]. one-way [BK91, BH99, CG92a, CG90, DM91, ÉA95, WL96, vdMvdM90]. online [SW91]. only [Dou93, TA96]. onset [KCM94]. operating [CR93, GK92]. operational [Oht96a, TS94]. opiate [Gro94, Jai94]. Optimal [Cha96b, DG97a, Fyg91, GC98b, Hai93, Hon91, Joh95, KA91, LTvC^{+}97, LS97b, LH93, PL96, RTS92, Sob96, SW98, Tro93, WK97, Wij93, WZ94, Zou94, Abt92, AK99, AW92, Cha92b, CS97c, DGD92, Dra95, Dut93, Fan98, GP93, Hol91, Hud91, Jah91, KA92a, Lin93, Liu97b, MVM90, Meh93, Pre98, Rak92a, Rak92b, San92, SM91a, SM92, She94, Sma93, Sob92b, Stu91, TZF97, Udd98, VL92, WW90, WN92, ZAW98]. optimality [Adh91a, Bha95c, Dut91, RZ98, Tar90, Wan94a]. Optimum [MG94, Ras90, SW97, WL92, YIK96]. Option [MB91b, Hüs96]. Orcutt [ST93a]. Order [BB95, Mig98, Sch91, Ahs96, AN90, ATT99, AOA92, AK94b, AM93, AVN97, Anr94, AP91, AM99a, Ara90, BN91, Bal94, BMM^{+}91, BP99a, Bha97, BLY92, CW97, CJ98a, CR92a, CFRT92, Cox92, ES95, EP92b, EF92, ESY93, Fal98, Fan91c, FT99, FZ96, FRS97, FTV99, Gle98, God90, JP91b, JS91a, KB90, Kun97, LBB92, LG98, MA94, Mar97, Nag90, NR99, Olu93, Par93a, PK98, Raq96, RS90, Sat90, Son93, SW91, Sri99, TTG92, TP97, WK91a, ZM93b]. Order-restricted [Sch91]. ordered [BR97, DG97b, GD99, GGHA95, JS95, KGM94, LR96, MD95, NHY95, PPS97, PS93a, PN95, SGB98, Sug99, TK94, VS91]. ordering [KF97a, Koc90, SSW93, Yan96a]. orders [AR90, CJ98b, MC92, TM91]. ordinal [BS94, Bar98, DY96, HAKM93, QT98, dMD95]. ordinary [KO94b, MM96b, Seg92]. ordinates [Swi98]. orthant [NS92]. Orthogonal [DS93, Lia99, BCLBMR99, BKS99, GP92, GDK95, SB90, Wit90, Zha92]. orthogonality [Age99, PD92]. orthogonally [Pre98]. other [LV98]. outcome [Kla96]. outcomes [GS92]. outer [GS93, GS94]. Outlier [NMB98, BR93a, DRS90, Dia91, Gle98, SGA99, Var98]. Outlier-robust [NMB98]. outliers [Ali91a, BB93a, Bha95a, Bha95b, Dix91, GC99, HN97, IT93, KS90b, Luc97a, Mad92, Nai90, Pet94, PM91b, TCSJ94, UG95, WG99, YAM91]. over-dispersion [RAP95]. Overdispersed [XD92, Kim94a, Sco95b]. overdispersion [CS97a, LPK95]. overlapping [God92, LBB92]. oversaturated [DLW96]. Overview [BS94].

P [Che98a]. P-value [Che98a]. paired [Mor90, SS90b, TC92]. pairwise [CWH93, DG97b, LSU90, SS93b]. Panel [Ano90l, MC91]. paper [KJ90]. paradox [DD97b]. parallel [KY93, LS97b, Lia99]. parallelism [NSS99, SSSS90]. Parameter [Age99, CT99, CT97, Li94, Rie95, SS94b, Ahs96, SAP92, BP99a, BS92, BSK95, BB90d, Car94, CC90a, CTP93, CM98, CH95b, Chi99b, Con90b, CBFCN99, DJD96, DN90, DY98, Dix91, FK94, FKT95, Fun91, Gil93, GG93, GLL90, HW90, Hon91, Iha93, Jär99, JC98a, JC98b, Jor99, JN91, KM94a, KM95, Kun95, LL90, Lia90, LL94, Liu92c, Lom99, MAS94, NA90, PSM96, Pan97, PN95, PB90b, PB91, PB92, PT92, PK91b, RR97, RT96a, RS90, SS99a, Sca92, Sch97, SGN93, SUS95, Swi98, Tap92, Tar91, TC91, Ton96b, WW90, WC96b, Wit91b, XB97, Yu91, ZW99, dSC99]. parameters [AF98, AHAEH90, AK98, BGP95, Ban90, BR93b, Bha95b, Bha97, Bho97, CPK92, CH94, CW92a, DF94a, DL90, DG97b, DS93, EP91, EP92a, EB98, Ete91, Fer90, Fol92, Fro93, GD99, Gru93, GGHA95, HD99a, HK90c, HD99b, Hob91, HD92, HN97, Jan99, JV92, JP90, JP91a, JP91b, KHR97, Kap92, Kar94, KSmS90, KF90, Law90, Liu95b, LR93, MH95, MD93a, MU97, MD95, NN91, PM90, PMSM97, PF93, Pas94, PG95, PT92, Rab98, RRA92, RKN91, Rug90, Sch97, SPQ98, SG90, SGN93, Sob96, TV91, TWB99a, Ton96b, Ton96a, UI99, VT93, Ver96, VS91, WM99, Zha94a, ZW99]. Parametric [FBD99, BDF95, Cza93, Gao97, MP91b, OT91, PB92, Ric90, ST90a, She94, SAK97, Sun97, WS94, YL92]. Paretian [MPR98]. Pareto [BA94, DA98, DT98, Han96, PSM96, PS99a, Sob92b, WA95]. Pareto-optimal [Sob92b]. part [ST96, Jah91, Pos92b]. partial [AF98, And90, And91a, BSSB90, Cho90a, CL98, Ete91, GDK95, Koc90, LZZ99, LR92, PRP92, PS91, SS90b, Yan96a, Zha90]. Partially [BD91, GG94a, GG94b, GT97, HCW91, Hon99, KAHK94, LZ98a, LD95, NHY95, Phi96, SSS97, SB91]. particle [VL92]. particles [KP93b, KKT98]. particular [Hal96]. Partition [Har90, Awa92, Kai96, Liu93a]. partitioned [Dah95, NP92, Pun96]. partitions [Cer97, RG94, TZF97]. partly [Gao95a, Gao95b, LZ98b, SG94]. pass [Fan91a]. pass-fail [Fan91a]. passage [SJ97]. past [CM94a]. path [JW92]. pattern [DJ98, GJ97, MML99]. pattern-mixture [MML99]. Patterned [Kro94a, TK93]. PBIB [Sat91a]. PCS [GLL90]. peaked [SD90]. Pearson [LZ96a]. Peculiar [Mae90, GM98b]. penalised [Bar95]. penalized [HPS98, MA95, SS99a]. Percentile [CT94, Sch97, FK91, RDA92]. Perfect [SGA99, UG95]. Performance [DHW90, HC97, Luc92b, SK93, Adh91b, Alw92, DJG94, Jin97, MGR95, Men95, SA90c, TK91a]. performances [NS90]. period [Cor91b, Udd98]. periodic [CM95b]. periodically [FM97]. periodicity [WN97]. periodogram [Has95b, Swi98]. Periodograms [AD98]. permutation [Kul94, LC96a, RD96]. permuted [CP99b]. personal [Lei90]. perspective [BR90b]. perturbation [TCT90]. perturbations [Lee99]. pharmacokinetics [Pat94a]. Phase [Wan99, Kim93]. Phase-shifting [Wan99]. phenomenon [Fei93]. physics [VL92]. Pickands [Dre96]. piece [MP93]. piece-wise [MP93]. piecewise [Zhe99]. pitfalls [AKH91]. Pitman [BG91, Bos91, CH91c, Dat91, DS91, Fou91, GK91, HC97, Kea91, Kub91, MB91a, PFN92, PK91b, SS91a, Sar91, SNK91, SS91c]. Pitman-closeness [DS91, SS91a]. pivotal [SVM95]. placement [BB90c]. placements [GW94]. plan [Gov90a, GK92, KG95, SA92b]. plans [Ami96, BKS99, Gov90b, Gup91, Gup92a, KG92b, Raj91, SA93, SB93, WL96, WB93, YK90]. plant [Gou95, GPA98, Yip96]. plant-capture [Gou95, GPA98, Yip96]. plateau [Gri92]. play [AFR94]. play-the-winner [AFR94]. plot [Ahn92, CGE99, Par93c, SD91, SK97a]. plots [CT92, SP99]. plug [dR94]. plug-in [dR94]. plus [Gup91, Gup92a]. Point [CP96b, GW91, Bou90a, Cal91, Che98b, DP97b, Fro90, GJ92b, GJM96, IU93, JP99, JW92, Law97, Nag95, Ng90, Sma93, SW98, ST99, VV94, Wes98, YIK96, Zha93b]. points [Chu92, Chu94b, JI97, JI99, Kul99, LM97, MG94, MNZ90, YV95, Zha95a]. Poisson [AL96, Ahm91, BK98, CF92, Cro96, Dag99, DC92, DF94b, Fam93a, Fam93b, GJO90, GG96, Gol90, Gru93, HK95, HB99, HL93b, Hwa92, JSM97, JG95, KK91a, Kem92, KP92b, LPR99, LPK95, Lee98, LP94, Mof95, NS94, Pap90, PS96, PN95, Pet94, RRrA91, SM94, Sco95a, Sco95b, Sco97, Tsi99, WSM91, Wil98, Wun91, Yan90b]. Poisson-distribution [Wun91]. Poisson-like [GJO90]. policies [BK98]. policy [AW92]. Pólya [Jan98a, SJ97, TP97]. polygon [DY96]. Polynomial [HM99, Cha91a, CLC97, CM96b, ES94, FP90, Kat95, Liu97b, MA95, RTS92, SM92]. polynomial-normal [ES94]. polynomials [BCLBMR99, Let97, Mon99]. polytomous [LT92, Von93]. pool [Ahm97b, Rai96, TL91]. pool-adjacent-violators [TL91]. pooled [MMS91]. pooling [Chi95]. poor [SV97]. population [AB91a, AB94, Arn91, Arn92, BR90b, Cha91c, CZ97b, CC95b, Fan98, GG94a, Gou95, GPA98, Gro91, IU93, KS98, KS97, KM94b, KA97b, KJ91, LL94, Liu90, LK98, Muk94, Muk95, OS93a, Oht91, PM91a, PK91a, Pra89, PS90, Pra91, PS92, SSSS94, SMS97, Son92, SD92, Tap92, TC92, TK99b, WG95, YLB98, ZS97]. populations [Bof92, CO91, DM94, EP92b, GM96, JC98b, Kel96, KS94b, Liu94b, Liu95b, MG91, MSMP97, MD93b, MD94b, ML96, MY91, Olu93, PM94, Pap92, PK94, SC97, Sar91, SSSS90, Sob92b, SH90, TSI98, VT93, WK97, WG90, YC99b, Zou94]. portmanteau [Kul90]. Positive [ST90b, Fro90, Kap92, NS93, RM93, SR91, ST96, TS92]. positive-part [ST96]. positivities [SV97]. Possible [CS94b]. possibly [GG92a, QSC94]. Posterior [Bof90, Rug90, Bos91, CC93, DRB97, GS91a, Rod92]. posteriors [Had98]. potency [KY93]. Power [CGP92, HL90, SA99, Age99, BB95, BG95, BD93, CGP90, CP97a, CB98, CG90, CGE99, DH90, DY91, Dix91, FN95, FR98, Gau98, GGVM98, GT92, GGT95, HL92, Ibr91, Kro94b, LS94, Lev92, LLL92, Liu96, MP97, MS98, NR99, SSW93, Sma93, Szr95, TSI98, TM96b, Tho96b, TK91a, US92, Wij93]. power-divergence [LS94]. power-law [CGE99, Gau98]. powerful [CG90, Hu94, RSAJ90, SD90, Tho96a]. powers [Dun95, JSM97, LS94, SS94e]. practical [PS95b]. pragmatic [Cha97b]. Prasad [KJ91]. Pre [Gil91, AR90, CH95b, Chi99b, Gil93, GGO96, Oht91, Wan94a]. pre-assigned [AR90]. pre-test [CH95b, Chi99b, Gil93, GGO96, Oht91, Wan94a]. Pre-testing [Gil91]. Precedence [Liu92a, LS92]. precise [GVSM92]. precision [Kri91, Ng90, PLC98, Sun98]. predetermined [Hol95]. Predicting [BB90c, LLL92, PL96]. Prediction [Ber98b, HK90b, Res96, AHJ95, Ber98a, BR90b, Bor95b, CGP90, CP94, CP95, CP96a, CS94a, CM94a, CWH93, CA90b, DP97a, DH92, DY95, DA98, HH95, KL99a, Kib99, KMBTG99, Liu93b, NAW96, SA90a, SET94, SAS91a, Tak94, Tak99, TL94, Zha94b]. predictions [GJO90]. Predictive [HK97a, Liu92b, dMD95, CHA95, DRB97, Gen97, LY94, MGR95, Str91b, YL91]. predictor [BBK98, CK95]. predictors [CM95a, Dat91, Gri97, KWtB96, Raq96]. Preface [Ano91m, HC96]. preference [CS94b]. Preliminary [BLY92, DB95a, GG93, Moh90, Kar94, Kib96, KA92b, LL90, Ng94, SS91a, SK93]. Preliminary-test [GG93]. Preposterior [Gar92]. presence [BD93, DF94a, Dix91, FP90, Gri92, Gru93, HN97, Iha93, IT93, JG95, KS97, KF90, LS97a, MM96a, PT92, SB97, SM92, TK97a, WG95, YAM91, ZR98]. presence-absence [SB97]. present [IH94, Mad92]. preserving [RT91, YSM92]. prespecified [BR91]. pretest [Ahm98, SA90c]. prevalence [VRL93]. price [Hüs96]. pricing [MB91b]. Principal [TK99a, Bén90, BM94, CTT90, KKK90, LD99b, RA96, YB94b]. principle [Lei90, Lis92, Was94]. principles [Wes98]. Prior [Fan98, BR95, CT93b, CHA95, CHA97a, CS94b, Che98d, CT91, CD98, CJH95, DF95, DP93, FBD97, Gar92, HB99, KY91a, KP93a, MGR95, MS90c, PG94, Res96, Ste92, SI97, TCSJ94, Wil94, YLB98, YT95]. priors [ÉA95, GM98b, Had98, Lia90, MP93, Rug90, SP94, Swa93]. privacy [MSS97]. Probabilistic [Tsa90a, Haj90, Tsa90b]. Probabilities [Bri93, AR90, CM95c, Fro96, GW91, HRS93, HO90, KK98b, Liu92a, Myo95, NHY95, Ong95, Wil94, ZM90]. Probability [MS90a, Ste90, Wri90, Ahn92, Bof90, CP92, CHS98b, CM95b, Con90b, Con94b, Con94a, DK91, FRS97, FR92, Fun96b, Gro94, HB91, KE93, KK91a, KG92a, KSm96, Liu93a, Liu97a, Luc92a, MD90, NPA93, Nay96, NS97, NS92, PFU91b, RGU97, RRrA91, Rug91, Rug93, SJ97, SNM97, Sob95, Wri98]. probable [CH91a, LL91]. probit [Bar98]. problem [AB94, Ano99m, Baj95, BR93a, BG96a, Bro95, Cho94b, Dia91, Edw94, FTV99, Fre92, Fun91, GW99b, Her96, Jin95, Kab91a, Kre92, LF91, MR90, ML96, QL92, RP90, RSAJ90, SM97, SGN93, Ste90, SL94b, TZ90a, Tsu99, WK91b, YC99a, Yue99]. problems [BSK95, DP97b, DA98, Han90, JW92, Koo90, Kub99, Tak97a, Tak99, Zha93b]. procedure [AF98, Aos94, BD91, Bri93, CS93, CH91a, Chi95, Dra95, Fri90a, GG92a, GAN90, Gov97, Liu94b, MNS96, MY91, PFU91a, Rak92a, Rak92b, RT96a, Ric90, SA90c, SK96, SW91, Sri97, UI99, Wan98a, Wil91, Won95]. procedures [CAS97, CS94b, Cor90, Ent90, Eva97, GG94a, GM94a, HS96, Jah91, Kab97, KO94a, KC98, KS90b, KY91b, KKK90, Muk94, NA90, RS99, RC94, Sob90, Sob96, SH90, Sri99, Ste91, TA96, WDH90, WX99]. Process [KK91b, Adh91a, Ami93, And90, And91a, Ara90, BK98, CGP90, CP97a, CXZ90, CH95a, CP97b, Che98c, Che95a, Cho90a, CO90, Cho91b, Cho91a, Cho94a, Cho94b, CGE99, Cus96, DGMR96, DK91, DJ98, Ein96, FBD97, Fer96, Gau98, GS91b, HB99, HL93b, JSM97, Joh96, KE93, KD95, KKK91, KK94, KM95, Law97, LP94, NS99, Pap90, PS96, PK91a, PLC98, POGP95, PGT92, RA96, RC95, RAM96, Sri96, SW99, Sri99, Sun96, Swa93, VV97, Vau92b, Vio95, Wri98, ZSW90, Zha94a, ZS98]. processes [AAB96, And97, Arn93, Bal98, BP96, BK96, BMM^{+}91, CGP92, Cha90a, Chr97, CK96, Cro96, Edg96, ES92, Fyg91, GG96, GCD94, HK97a, Hoo90, JVM95, KM91b, Kun97, LC96a, Lee99, bL95, Mak91, MD94a, MR95b, MGM98, Nas91, Nie97, NMB98, OT91, Pei90, Rad97, Sun97, WD97, WW91a, Wes98, Yeh92, Zha90]. Procrustes [Chi99a]. product [BA94, CC95a, CJ98a, Cha91c, CLC97, Cor91a, Fan91b, Liu90, Men91, Ng93, PS99a, SSSS94, TS92, Yon95]. product-limit [Ng93, Yon95]. product-type [TS92]. production [NS99]. productivity [Ano90l]. products [Den94]. professional [RC96]. Profile [NSS99, Age99, BS99b]. program [BB90c]. programming [Han90]. progressive [PG95]. progressively [PG92]. projected [DL95]. Projection [Pos92a, HS93, ZL91]. projects [Moo96, Var96]. Proof [Fei93, HO90, WH99, Yan90a]. proper [CF96b, Had98, KP95, SP94]. Properties [BK97b, DC92, Lee96, OS93b, AM99a, AJC99, AN91, Bos91, Con90a, Con90c, Don93, ESDC96, Fir99, Gao95a, GJU94, HS90a, Kam94, KB90, Kun91, Mac92, Mae90, MD94a, MLH92, MH96, Nag90, Oht96a, PLC98, PH92, SK93, Tak94, WG96, ZM93b, ZFP90]. property [Fan91c, Kim94c, LR96, MVM90, NP92, POGP96, Rao90a, Rao92, RDA93]. Proportion [Leu92a, Ahm97b, Ala96, Nay94]. proportional [Adh91b, CP99b, DQS98, DG90, Gho91, GP93, HPS98, hK97b, KP93a, KSm96, Li99, Lin99, MS99, RDA92, SA98, Ver94, Wri90]. proportional-reduction-in-impurity [SA98]. proportionally [BE93, MPQ91]. proportions [CO91, Den98, Dru90, Gro91, GF91, HL90, MS90c, Nor93, PT93, RL91, SG91a]. Protective [MM97]. proximity [YN99]. pruned [Kim94c]. Pseudo [GMR97, Gup99]. Pseudo-likelihood [GMR97]. PTC [PS91]. public [Mof95]. Pukkila [Gle96]. pure [Cro96]. purely [WD97]. purposes [GS93, GS94]. pursuit [Pos92a].

Quadratic [TCT90, AM93, BG91, Cha92b, EJ93, FMV98, HP92, Kim95a, Liu91, Liu97b, Luo94, Mat91, MS90a, NS93, Pre98, QCL98, Sub91, Tsa90b]. qualitative [AB98, Arn96, Per91]. quality [ARS95, Ano90l, Cha92a, DFT91, Gov90b, Mye91, NS99, NA90, Var96, WS91b, WB93]. quantile [Bos94, CC91b, CC91c, Kc91b, KC91a, MP93, Par93c, VV97, Yon95]. quantiles [CQ91, CP96c, Don93, FN95, GR90, Hug91, Rug90, SVM95, VV96, Ver94]. quantitative [AB98, CG92b, SM96, SMS97]. quantities [CFRT92]. quartic [VI91]. Quasi [Hor92, Con90c, JS93, LH93, LM94, Pai95, RBC97]. quasi-Bayesian [RBC97]. quasi-binomial [Con90c]. quasi-likelihood [LH93, LM94, Pai95]. Quasi-nonparametric [Hor92]. questionnaires [LF93]. queueing [Jai91]. quick [AS94b, MPH97]. quota [Kot91]. Quotas [KM99a, KM99b, LL93]. quotient [CS90a].

R [Sar92a]. radial [Law97]. radii [YSM92]. Random [Kai96, Lee97, ZM90, ATT99, Bal94, BB95, Bar98, BK91, BH99, CS90a, CH97, CW94a, Chu92, CG90, Con94a, Cox92, ÉA95, Ete91, FBD97, Fol92, Gov95, HEB92, Hei93, Her92, Her93, HK94, cH91b, HS99, Kam94, Kap92, KY91a, Kre92, Lai98, Lee90, LC96b, Lee99, Liu95a, Loy90, LK98, MW96, MSS97, MNZ90, NAW96, NSI90, NM94, PB98, Pep90, PGC97, PGC98, QSC94, Rao92, RM90, SR91, Sar92a, Sat90, ST90b, Sin91, SM91b, SO98, SS99c, Stu94, Ver94, WL96, WK97, Wil98, Xu90, Yok96]. random-effects [HS99, KY91a, WL96, Yok96]. Randomization [Ent90, GL97, Mar95, SMS93, Sna94]. Randomized [Arn96, Ahm97b, AFR94, Arn94, BSS94, BY92, CA90a, CC95b, KC98, Man91, Muk97, Nay94, Rag90, Ree96, RT91, Rob94, SS93c, Tar90]. randomly [GP90, MJ92b]. randomness [MMD93]. range [Loy90, Rad97]. ranges [Rug90]. ranging [MF90]. Rank [Cap92, Che97, RRAB93, AC95, AC99, CN99b, Geo91, Gov95, LW95, LYS96, Lom99, ML96, Par93a, RP90, RT91, SM97, Spu91, Spu95, Tar90, Wil95, WDH90, ZQ97]. Rank-based [RRAB93, WDH90]. rank-order [Par93a]. Ranked [MD90, Beh99, Bho97, Boh96, KB96, Liu95b, Özt99]. ranked-set [Boh96, KB96]. ranking [KY91b, Sob90]. rankings [AF98]. ranks [CLL95, FKY91, FKY95, KSm96, SL94b, Wri90]. Rao [CM92, Liu91]. rased [Han92]. Rate [GN91, BE93, Cha96a, Fay99, FN92, GJO90, GJ92b, GJM96, GL94, Mad93a, Nab93, Rao92, WL92, Zha93a]. Rates [Let97, BE93, DH90, Fay99, Kul99, LW90, sL98b, LS94, PG92, PG95, QSC94, SJ91a, SG94, Tah91, Ton96b, Ton96a, VSS95, WG90, YL92, YL93, Yu97, ZW99]. rather [Tho96b]. ratio [Ahm97a, Alb92, BM93, BS99b, Bri94, CS90c, Cha91c, Cha90b, CM96a, CO91, DJD96, Dav90, FN91, GB98, He90, JK91, KMM90, Kap92, KS97, KK94, KML91a, Lev92, Liu97a, Mae98, MML99, Nis94, OW93, Pra89, PS90, Pra91, QCL98, Rao91, SS93b, SS94a, Sin91, SSW93, SSSS94, Tah91, UI99, YY95]. ratio-type [PS90]. rational [PE92]. ratios [Anr94, Den98, Dru90, El 97, Fay99, IY99, SG90, YR94]. Rayleigh [HH95, LM96]. re [GS92, Jor99, SL98a]. re-estimation [GS92, SL98a]. re-weighted [Jor99]. real [Cal91]. realising [AR90]. realized [SO98]. recapture [EB92, GYH95, GL94, Llo95, Yip96]. receiver [CR93]. reconstruction [GG94b, KP93a]. record [AK91, BCA93, BA94, BB90b, Ber98b, GP94, Kam94, Koc90, PS99a, TZ90b]. record-breaking [GP94]. records [Ano98n, LF93]. records/true [LF93]. Recovery [SLS96, SK97a, Wee92, WP91]. rectangles [MNZ90]. rectification [MSS97]. Recurrence [BCA93, BA94, ES95, Hua90, HB91, PS99a, SFT99, TM91, Fro93, KK91a]. Recursive [CT93a, Gab92, Cip98, NM96, RG96, TSQ91]. redescending [YV95]. reduced [Pun97]. Reduction [RS95, SKS93, SA98, TG94]. referees [Ano96l]. Reference [SI97, Bak91, CD98]. Refined [Dre96]. reflections [SR91]. region [AM99a, KML91b, LV98, RS92, SH94]. Regional [HC96, MS95]. regions [DF94a, Hor92, Hud91, MD94a]. Regression [GJO90, HV91, LTCC94, NC97, OSL91, PL99, SS90a, Wun91, Ahm98, Ahn94, Ahn96, Alb92, Ali91a, Ali91b, Ati96, BDG90, BF90, BG91, Bha95b, BOR92, Blo97, BLS97, BM94, BJ94, BP93, Bro97, Buo90, CT92, CT93b, CT99, Cha91a, CB90, CM97, Cha93a, CHA95, CHA97a, CQ91, CW93, CW94b, CP96c, Che98b, CT98, Che90, CW92a, CT91, Cho94b, CK95, Chu94b, CF92, Cor92, CA90b, Cza93, DN90, DD97a, Djo95, Dra90, Dra92, Dra94, Dra98, DSTW91, Dut99, Enc92, Fam93a, Fir99, FT95, Fuj90, FT97, Gao97, GPW98, GJ92a, Gil91, Gil93, GG93, GGO96, GBC99, Gra98, Gri92, Guo96, Güv95, Hai93, Hal99, HK90b, Has95a, Has95b, Hau91, HW91, HC90, HC93, Hol91, Hon94, Hon98, Hos90, Hun90, Iwa94, Jah91, JU93, JI97, JI99, Kab97, Kab91a]. regression [KSA98, KO94a, Kam99, KKS99, Kej93, Ki92, KS90b, Kib96, Kim93, Kim95b, Kim96b, Koo93b, KO94b, Kun91, KM96, Lai99, Lee90, LYS96, LM93, Lia94, LLH90, Lu93, Lye91, Mar92b, MB91a, Mof95, Nag95, Nag97, NP92, Oht93a, OSL92, Pai95, PB93, Par93b, Par94, PRP92, Qiu94, Qum94, Qum97, RS98, RS99, Rao91, Ras90, RBC97, SA90a, SH91, SK93, SET94, Sar97, Sar99, Sar92b, SS93a, Sco95b, SW96, Seg92, SGJ95, Sha92b, SPQ98, SA96, SY91, Shi95, SS94b, Sil94, SCXG96, SPS94, SV97, SP99, SM91b, SG91b, STS95, SO95, SC92, ST93a, Str95, ST99, Szr95, Tak97a, Tak97b, Ter90, TC91, Ton96a, TS94, TCSJ94, Tu99, Ull92, VI91, Wan96, WL93, WLG99, WJ90, Wil92, Wil95, Wil96, WM99, Wil98, WA96, WF95, XB97, XD92, Yan98]. regression [You93, YB93, YB95, Zha94b, ZL91, dR94]. regression-based [Kam99]. regressions [Kar99, KMBTG99, Liu97b]. regressors [Lee90]. Regular [BLF99, Li94]. regularization [WW90]. rejection [CH95b, Chi99b]. Rejoinder [Ano91n]. related [BE93, CW90, DG94, Ent90, God90, KM97, Koo90, Kub99, MMPV98, MPQ91, Ong95, Pun96, Pun97, TK91a, YL91]. relation [Fro93, Jan98a, Ng93]. relation-based [Fro93]. relations [BCA93, BA94, ES95, Hua90, HB91, KK91a, PS99a, SFT99, TM91]. Relationship [AM91, Iso93, Iso92b, MC91, Min99, Nag95, Zha90]. relationships [Ban91, BP99b, HS90a, Iso92a, JT90, Won93]. Relative [GR90, Tur95, Abd98, Boi91, CL96, Eva97, Jin97, KY93, Lei90, NS90, RDM92, Yan90a]. Reliability [Kam94, AK96, Ana92, Chi92, Chi93, DL90, FBD99, GGHA95, GS98b, Han97a, KM93, KM94c, MS99, Nag90, Rod98, SB97, Sha92a, Zha93b]. remark [Sap94]. Remarks [GC99, AP93, Che95b, KP92b]. Reml [TL94, MY95]. renewal [BK96, BK98, Fyg91, RC95, WW91a]. Rényi [Set90]. repair [CP99a, LLP98]. repairable [CP99a, Ho98]. reparametrizations [Pei92]. repeat [DL95]. Repeated [Afs90, DH92, Ho98, Boi91, DJ98, FWG94, FMV98, Gal94, HCW91, KAKW99, Lai98, LLH90, NSS99, PLW93, Par94, PPS97, Pat94b, RRAB93, RD96, Udd98, VV94, WK91b]. replacement [Adh90, AW92, EB92, Kim96a]. replacing [DG92]. replicate [GDK95, KA92a]. replicated [Wan93b]. replicates [LN90]. replication [Iso93, MW96]. replications [Guo96, WDH90]. Representation [AF95, AK94b, Cal91, DH95, Mae98, Ozt91, Vio95, Wan94c, YN99]. reproduces [Car93]. Resampling [Chr97, SA99]. residence [BM98]. Residual [DR96b, MMT97, WS94, BB90a, FK91, GH92, HS90a, Her93, JI97, KM91a, LP98b, Lin97, NK99, RDA92, RG96, SJ91d, Shi99, Zha94a]. residuals [BP93, CG92c, JH99, Mar92b, PC95, RWP95, SY91]. resighting [Ana97]. resistance [ZQ97]. Resistant [KM90]. resolution [BW93, GP92, Hyo94, Lia99, SV97]. resonance [EE93]. resources [OS93b]. respect [Cy95, JN96, JSM97, MD90, Zha92]. respondent [MSS97]. responders [Tem94]. responding [CBRS90]. Response [mC96, Mye91, AB98, Ahm97b, AKP93, Arn94, Arn96, BSS94, BSSB90, CA90a, CG92b, CC95b, HGM94, HAKM93, JP93, KD99, Kro97, LTvC^{+}97, Muk97, Nay94, PK98, PJ99, Pat94a, Rag90, SS93c, TK97a, Vau93, You93]. responses [AA92, BG96b, HCW91, ML91, Par90, SSSS90, Von93, WB97]. Restricted [Ali91b, Fam93a, Liu97b, Cza93, EP92b, Fun91, HH90, KSA98, KW93, Kun95, MM96b, Olu93, RS95, Sar92b, Sch91, SGN93, Wan94a]. restriction [GGHA95, Tu99]. restrictions [Anr94, Bha97, Cy95, FRS97, Fol92, GL97, JP91b, Kib96, LG98, Myo95, Pei92, Tim96, Tom92]. result [Ein96, Sar92a, VL92]. results [Bur97, CQ91, Edg96, Fun91, GZ97, Gup91, Han92, Han93, HC90, Jie90, KT91, KK98c, Kmi95, Koc90, Kul94, Kul95, MT92, Mar95, MD95, NS90, PG95, Pun96, Pun97, Rei90, TK91a, VV96, VRL93, Wil95, Yan96b, Zha94a]. reverse [Tap98]. reversed [ESDC96]. review [Afs90, Boh96, DGMR96, KG90, Kub99, Von91]. rhg [BS90a]. Ridge [BG91, Cha93a, MB91a, BPT90, CJH95, Fir99, KSA98, Kib96, KO94b, NS90, Oht93a, SK93, Sar92b, Seg92, SG91b]. right [Ano99m, Bal94, GP90, GR90, Hug91, ID98, LZ98a, Par93a, SL94b, YW98]. right-censored [LZ98a, YW98]. right-truncated [Bal94]. risk [Abd98, CG92b, Has95a, IU93, Jan95, iK99, Ng93, RS95, SI97, Tak99, Tar91, WC96a, WK93, gWyJ97, ZK94]. risks [DK95, Geo91, GGO96, KD99, Nag99, RDM92, Wan93a, Wan94b, YL92, YL93]. Robert [HC96, RC96]. Robust [AB98, AM91, BY92, CM97, CT91, Cho94c, Cip98, Gab98, Kap92, Kul90, Mar96, OSL92, Rug91, Rug93, SL99, WK91a, BS91, Bas93, CL96, Gle96, HW91, Jor99, Kla96, KK94, LD99a, NMB98, NM92, She94, aS99, SCXG96, VR90, Wil92, WM99, YR94]. Robustness [CC93, LP98a, Luc97b, MK90b, Pos92b, SGD90, CL90, DF94a, Kro94a, Luc97a, Nan92, Zha92, Zha96]. ROC [CR93, ZR98]. role [GK90, Sin90]. roles [YLB98]. Root [Hon99, AD98, Kim94b, SS94a, SS94c]. Root-n-consistent [Hon99]. roots [BR93b, Li94]. Rosenblatt [JZ96]. rotatability [JP93]. rotatable [AVN97, BN91, PK98]. rotation [FZ97, MD93a]. rotationally [SE97]. Roundoff [Cal91]. Row [Mee90, GC98b, PD92, PS95a, SB91]. row-column [GC98b, PD92, PS95a, SB91]. rule [AB91a, Bil97, DSTW91, Gar92, Kim95a, SU95, WB93]. rules [Cha92a, LL94, LG98, MD94b]. Run [KM99b, Cha92a, CJ98b, DGMR96, OC99, RWP95, SW97, Zha95b]. Runs [NS96, Cha92a, DL95, God92, GG92b, KM97, KM99a, SR91].

s [Yua92, Was99]. S. [Sar92a]. sacrifice [ML93]. Saddlepoint [JK91, TK97b, TK99b]. safety [HO90]. Sample [FZ97, GS92, Nor95, PL97, Ano98n, Ano99m, Aos94, Baj95, BR93a, Bre90, Bri93, CP94, Car94, CH96, CT93b, CB98, CW94a, Cha90a, CM95c, CWH93, Cus96, DRB93, Dut99, Fir99, Fri91, FR98, GG92a, GW94, HK91, HK92, Her96, JGMGMR98, Jin95, JT92, JS95, Kar94, Kmi95, Kul94, Kul95, Kum97, Lee95a, LC99, LH99, LWH98, Lin97, Liu96, MY91, MP97, Nan92, NAW96, NM94, Oht96a, Özt99, PMSM97, Par93c, PT92, Pos92b, Pra89, PS90, Pra91, PS92, RP90, Rao90b, RD96, Ree96, RFG92, Ric90, RSAJ90, SA99, SS91b, Sco97, SM97, SKF94, Sha98, SL98a, SBS96, SGN93, Str91b, SL94b, TM96b, Ter90, TK91b, TZF97, VSS95, Wan93a, WK97, WP91, WL92, Wil95, WW91b, Wit90, YLB98, Zha96, cL91]. sampler [Bru95, Men95]. samples [BGP95, CP90, CP95, CP96b, DA98, Fer90, GF91, Gru93, Han92, HR98, JN94, KS98, KSmS90, Koo93a, Leu94a, MJ92b, PB90b, PB91, PG92, Pet94, Rie95, Ryc93, TV91, Var98, VT93, VRL93]. Sampling [Gue94, KH90, Pap90, WL96, Adh91b, Ahm97a, AHAEK92, Ala96, AH93, AW99, Ami96, AS94a, AR90, Arn91, AS94b, BS90a, BK97a, Bho97, Boh96, Buo90, CW94a, CM94b, Cha97b, Che92a, CC95b, CL98, Cos98, DB95a, DP93, EB92, Fan91a, FP90, GS91a, GS90, Gov90a, Gov90b, Gov97, GK92, Gro91, GT90, Koo90, Kot91, KB96, Kre92, Kul95, KS94b, KSm96, KJ91, KG92b, KG95, LP94, Liu90, Llo99, Mac98, MQZ99, MD93a, Mis96, MA92, Muk94, Muk95, MD90, Myo95, NA94, OS93a, OSL91, OSL92, OS93b, PS96, PB92, PGT92, Raj91, RK96, STR97, San92, SS93b, SSS97, SM91a, SM92, SS93c, SA92a, SA92b, SA93, Sri96, TB91, Vau92b, Wri90, WZ94, WB93, YC99b, YK90, ZFP90]. Sarmanov [Lee96]. Satterthwaite [KAKW99, SL94a]. saturated [Zac91]. Saunders [CT94, DY95, Rie95, Rie99]. scalar [GS95, TA96]. scale [Ano99m, BGP95, BB90d, CT93a, Cap92, CH94, CM98, Chi99b, CW98, CR92b, Dix91, EP91, EP92a, Gil93, GG93, GD99, Heb94a, Heb94b, Hu94, Hür98, JR91, JN91, Kap92, Leu92b, Lom99, MT90, MH95, MD95, ML96, NN91, Pan97, PF93, PT92, RKN91, Sat91b, Wit91a]. scaled [Che99]. scaling [CC91a]. scan [Kul97]. scatter [AK98]. Scheffe [BS96]. Scheffés [Boi91]. scheme [AW92, BS90a, BK97a, Che95a, Fan91a, KS94b, NA94]. schemes [AHAEK92, GK98, Myo95, SW97, YB93]. Schwarz [CN99a, NC97]. science [Ste90]. Score [Tha92, TK93, Che95a, CF96b, Cro96, Law93, SS99a, WG90, dP90]. scores [CB90, KWtB96, LC91, Rab98, TL99, TK91b, Woo92]. Scoring [PS93a, Sob92a, Gar92, MH98]. screen [Tho96b]. Screening [Liu92b, GW92, GS93, GS94, bL95, PM92, VRL93, WZ94]. seasonal [Li94]. Seber [Bur97]. Second [AN90, AM99a, AVN97, BN91, BLY92, CS90b, ESY93, GH92, Jan97, Ley90, Liu94a, PK98, SW91]. Second-order [AM99a]. sectional [WP91]. Security [BK96, BK98]. seemingly [KMBTG99, Par93b]. segmented [Gri92]. selected [Son92, SD92, Tap92]. Selecting [GL90, Liu94a, AB91a, AB94, CZ97b, LL94, LL91, Liu94b, SS99a, Sob92b, Tan94]. Selection [Ati96, BD92, Gov90b, KG92b, LL90, NKN99, PFU91b, SS99b, AS94b, Bof90, BM94, BD91, CS94a, CH91a, CS94b, DM94, Dra95, Dri91, EE93, GG94a, GP93, Gle95, Hau91, HV96, HK99, Jah91, Kab97, Kar95, KY91b, KG95, LY94, Liu93a, MAPV96, MD94b, MY91, MD90, NC97, PFU91a, SC97, SM91b, TS95, Vel93, VCvdL97, YL91, ZM90, Zha93a]. selective [Özt99]. Self [LY97, YW98, ZK94]. Self-consistent [LY97, YW98, ZK94]. semi [BDF95, Fyg91, KP93b, KKT98]. semi-Markov [Fyg91, KP93b, KKT98]. semi-parametric [BDF95]. semidefinite [NS93, RM93]. semimartingales [TSQ91]. Semiparametric [LZ98a, Gao97, Hon99, Lia94, LH99, SVM95]. semiPareto [Bal98]. Semirange [Van93]. Sen [KS94b]. sensitive [Bou90b, CS93, Gov90a, SMS93, SMS97, SA92b]. Sensitivity [Bén90, CTT91, DF95, LZ96a, Mar90a, Ros96, TC90, TW92, Bak91, CTT90, TCT90, YY93, hZ93]. separability [SB95]. separate [AKH91, Szr95]. sequence [Gol90, Hal96, JP99, RR97, SS99b]. sequences [AB92a, Lai99, OC99, SGS94]. Sequential [Ara90, Coa94, JS90, KA97a, MA92, Wan91, AW92, BMM^{+}91, BD91, Cha96b, CHS98b, CH91a, Dou97, Dra95, Fyg91, GAN90, Hwa92, iK99, LRD94, Li99, Liu94b, Liu97a, Mor90, Nag99, Pat94b, QSC94, Ric90, SAK97, Sna94, SH90, Tah91, Tah92, TY99, UI99, WX99, WB93, Zou94]. Serial [Goo91, Kea96, ML93, Wen94]. Series [KA91, AD98, AOA92, Ana92, BB97, Bev92, Bol90, BJ94, BK97b, CCC96, Cha90a, CHA95, CHA97a, CBR91, Che92a, Che92b, Cho94c, Cip98, DN92, DB95b, DY91, Dia91, Fro93, Gab92, Gab98, GS95, Gle95, Gup92b, GT92, GGT95, GT96, Hon91, KTTL99, KB90, KK90, KM96, Li94, Mar90b, MO91, MD93a, MS98, NT92, NC97, NM96, Ong95, PS93b, RAP95, RS96, RT96b, Sas91b, Swi98, TM96a, TWB99b, vHL92]. service [Ste90]. set [BR93b, Bho97, Boh96, CB90, Dem92, Huw93, KB96, Mad92, MD90, Özt99, Rug91, Rug93]. set-up [BR93b, CB90]. sets [AR90, Bar89, GP96, Hol95, MG91, US92]. Setting [CM96a, RDA93, Rao90a, Rao92]. settings [WH99]. setup [DK95]. several [CTP93, CWH93, GM94a, HS90a, Han93, HL90, HL92, JP91a, JC98b, Lei90, Liu94b, MT90, NvdMM90, OW90, OW93, RBLN93, SC97, Sar91, SS99d]. Shannon [YK97]. Shape [Mac92, BS90b, CC90a, DJD96, PSM96, PT92, Wit91b]. shapes [Tro93]. Shapiro [MSL95]. Sharp [Ryc93]. Shewhart [Cha92a]. shift [Ali90b, Fro90, GB98, JP99]. shifted [MT90]. shifting [Wan99]. shiftpoint [BG96a]. shifts [Che95a]. shock [POGP95, POGP96]. Short [EB94, DGMR96, RAP95]. Shortest [DY95]. shortfall [Van93]. shoulder [Mac98, MQZ99]. Shrinkage [Chi92, Jin95, Kub99, TK99a, Wit90, BR90a, Chi93, DS91, HK90a, IK97, KFAH92, KKS99, Ki92, PMD95, SS91a, SS91c, SPS94]. shrinking [SC96]. Shrunken [CH91c]. Sidak [Hoo90]. Sidak/Slepian [Hoo90]. sided [Baj95, BK91, Bri93, Cal94, CS96a, DG97b, GD99, GL90, Her96, Kar94, Lai98, LV98, MG91, MN98, Oh98, Oht91, Pap92, STR97]. sigma [GK92]. Sign [AW99, KB96, ARS95, HRS93, Özt99]. signal [Iha93, Kun95, RAM96, SS94c]. signalling [GZ97]. signed [Nis94, SM97]. Significance [AB91b, Kmi95, BMM^{+}91, Ho98, Kla96, Rab98, RS98, Won92]. significant [JP92]. similarity [MLH92, Mee90]. Simple [RS96, Ang91, BW93, CH96, GD99, Güv95, Hyo94, JP92, MT92, Oh98, PB93, PR93, Rao90a, SSW93, Sug99, TC91, WL93, Wil96, Won95]. simple-tree [CH96]. simpler [De 92]. simplified [Vän97a]. simulate [Bru95]. simulation [Bel93, KL96]. Simultaneous [BS96, CS95, CWH93, CK97b, DM94, DG97b, Fuj99a, GS93, Liu93b, MG91, MD93b, MD94b, SA90a, SET94, VS91, Aos94, GS94, HK90c, Mak91, NS94, SOSW94, Ter90]. Single [GDK95, BA94, DRS90, Gov97, GK92, Kra92, PS99a, SS98]. singular [Gro99, Pun96, Pun97, Yua92]. situation [Mor90]. situations [TWB99a, TS92]. six [LS94]. size [Adh91b, Aos94, AM99a, CB98, CW94a, Cus96, GPA98, GS92, GW99b, HS96, Hol95, KSm96, LWH98, MY91, MP97, MD90, NAW96, PL97, Ree96, RFG92, SS91b, Sco97, Sha98, SL98a, SMS97, TM96b, Wri90]. sized [SK91a]. sizes [Adh90, Bri93, CM95c, DRB93, Fri91, FR98, LC99, Liu96, WK97]. skewness [BS98b, MN93, RBM95, TG94]. Skorokhod [Vio95]. Slepian [Hoo90]. Sliced [Sar99, Sar97, ZL91]. slippage [He90, Sob95]. Slope [PK98, AVN97, BN91, Ban90, Guo96, Huw93, JP93, JN96, TC91, Won93]. Slope-rotatable [PK98]. slopes [SS94e]. Small [BKS99, CAS97, SBS96, Bre90, CM94a, CM95a, DP97a, Gro91, GF91, HC93, Isa92, Nor93, OW90, Oht96a, Ter90, Tsu99, VRL93, Wil95]. small-sample [Wil95]. smaller [CPK92, EP92a, KHR97]. Smirnov [CC96, GG92a, RR91, YP93]. Smooth [GP94, HH94, TK99a, Yon95]. Smoothed [NT92, Ami93]. smoothers [TSQ91]. Smoothing [KKP^{+}99, RI93, AAB96, Bos94, CK96, Djo95, Gao94, Hon91, SS99a]. software [Rod98, Zha93b]. sojourn [KP93b, RC95]. Solution [Zhe99, Dia91]. Solutions [Han90, Mak91, NvdMM90]. solvability [BWCM90]. solve [Kre92]. Solving [YC99a]. Some [Abt92, AKH91, AKP93, AP93, Bos91, CTT90, CQ91, Che95b, Con94b, DG94, DA98, ESDC96, FR92, Fro93, GM94a, Geo91, Gup91, GW99b, HS90a, Han92, Han93, JSM97, JZ96, JS91a, JBY90, KG92a, KP92b, KK98c, Koc90, Kot90, KJ90, Le94, Li98, Mar95, MD95, MSL95, Nag90, NS99, Nor93, PM94, PG95, Pra89, PS90, Pra91, Pun96, Pun97, SS93b, SM92, Sha92b, SOSW94, SB90, SB91, TTG92, Wes98, Wil95, Yan96b, Zha94a, ZM93b, AB92a, AS92, And97, AN91, BM98, Baj95, Bal94, BF90, CG92a, CXZ90, Chi99a, CL96, Con90c, Con94a, CF96b, EB98, Gao95a, GGO96, HK92, Has95a, HFP94, Hud91, HCL94, Iso93, Iso92b, JVM95, KS95, Kmi95, KM91b, Mar92a, MO91, MH96, MPH97, NS90, Rai96, SS91a, SK93, SA90b, Sco99, SS91c, Spu91, Spu95]. some [SGD90, SS94e, Tak97a, Tho93, TS90, YL91, Zha92, Zha96, Zhe99]. soonest [LL93]. source [Iha93]. sources [Gao94]. space [Dem92, DH95, Fun91, Kun95, SGN93, TZF97, TL94, Wit90]. spacing [Kai96, Mon99]. spacings [Sch91]. spanned [Bén90]. spanning [MS95]. sparse [Fay99, Fre92, Lev92]. spatial [AAB97, BM90, KM91b, KM95, Kul97, Law90, Mar92a, MD94a, MU97, NMB98, NKN99, OS93b, SB95, WD97, Zha94a]. spatially [Mar95, Mar97]. spatially-distinct [Mar97]. spatiotemporal [Ber98a]. Spearman [CN99b, GN91]. Special [Tem94, Bur97, Chi99a, EJ93, JVM95, Kot90]. species [Nay96, YC99b]. specific [Dam93]. specification [ACS95, DH95, Kam99]. specifications [WA95]. specificity [Bak91, hZ93]. specified [Gil91, Gil93, GW99b, HS96, Lia99, Rug90, SS91b, ZM90]. spectral [DN92, NMB98, PS93b]. spectrum [Lop93]. spending [CHS98b]. sphere [CC91a]. spherical [Bor95b, CW90, Cha93a, Hud91]. spline [Cha91a, Djo95, HH90, HS93, RI93]. splines [Bos94, CLC97, Gao94]. splitting [KH90]. Sprt [Edg96]. spurious [JN94]. square [DS97, EJ93, KY91b, LB98, Mie91, Olu93, PR92, RG92, SBS96, Wan96]. squared [Abd93, BLBE99, Che95b, KSA98, Oht96b, Oht96a, Smi92a, SG91b, Sun98, Wan94a, YSM92]. squared-radii [YSM92]. Squares [Lu90, Bev92, BSSB90, CM97, CP96c, CT98, Dem96, GT97, IT93, Jor99, Kab97, KM96, Luc97a, Mar95, Mar97, MM96b, PRP92, Qum97, Sar92b, SPQ98, SY94, SO95, Str95]. Sroka [AP93]. Stability [Kun93, Lim93, OW90, SCSG94]. Stable [PS93b, BS91, MPR98]. stably [BP93]. stage [Aos94, AM99a, Arn91, BK97a, BR90a, CMD94, CTP93, CP96c, GW92, HK90a, HC90, IH94, Jah91, KI94, LM97, MG94, Muk95, NA94, Par93b, QSC94, Raj91, Rak92a, Rak92b, SM91b, TA96, YT95, Yue99, Zhe99]. stages [Cha97b]. staggered [Lin93]. standard [CG92a, HW91, KC91a, Stu91, You93]. standardized [Ein96, Fay99]. standpoint [Tho96a]. star [HY97]. State [Bra90, Ley90, Cha92a, Dem92, DH95, FT99, KM97, KM99a, KM99b, KG92b, Lee94b, TL94, Wen94]. stationarity [Lee99]. stationary [AB92a, CBR91, Chr97, Gao95a, MR95b, Pei90]. Statistic [Nan92, AK94b, Alb92, ARS95, Ang91, Dou97, Gho91, HD99b, Kul97, KML91a, LZ96a, LL98, Mat98, McM97, PC99, RV99, Sco99, Shy98, Smi92a, VI91, Ver96, Was99, YSM92]. Statistical [AG96, AJC99, BS90b, CM98, CL94, CL97, GT92, Iha93, pL98, PH92, WK93, DGMR96, HGM94, HS96, Hob91, HCL94, JBY90, Kea96, Kot90, LWH98, Mar96, RA96, RAM96, Tak99, VCvdL97, Was94]. statistically [AAW93]. Statistics [HC96, Ahs96, Bal94, BB95, BP99a, BS90c, CW97, CS90c, CJ98a, CR92a, Cox92, ES95, EB98, Est96, Fal98, Fan91c, FTV99, FS99, GL97, Iso92b, JV92, JS91a, Kai96, Kim94a, Kul90, LC96a, LS94, LBB92, Mae98, MP91a, MMPV98, Mig98, Mil91, Mil96, MMS91, Nag90, Nai96, NR99, NvdMM90, Pro90, Raq96, RS90, RR91, Sat90, SA90b, SA96, SBS96, SS92, aS99, Son93, Spu91, Spu95, SET96, Ste90, Tho96a, Tho96b, TK97b, Var96, Xia95, Yan90a, Zah90, ZB97, ZFP90, dSC99, Bra90, Ley90]. STD [SA93]. Steady [Cha92a]. Steady-state [Cha92a]. stein [MG93, DY91, DSTW91, GGO96, HC93, KO94a, Nag96, Nag97, Oht93b, SU95, ST96, TK99a, YT95]. Stein-like [KO94a]. Stein-rule [DSTW91, SU95]. Stein-type [GGO96, Oht93b]. Steinmetz [AP93]. Step [PM92, CS94b, GW92, Mad93a, Wil91]. step-down [Wil91]. step-stress [Mad93a]. Step-wise [PM92, GW92]. stepping [ZM93b]. stepping-stone [ZM93b]. Stepwise [SK97b, KKK90, SH91]. Stirling [Haj90, NS97]. Stochastic [GM94b, MC92, ZS98, BM98, Bha95b, Cha90a, Con90d, DH92, DJ98, ESY93, GK90, HL93a, Jan98b, KF97a, Li98, Sin97, TCSJ94]. stock [GC98a]. stone [ZM93b]. stopping [Nag99, SLL97]. straight [BB96]. strata [MA92]. strategies [OSL92]. strategy [Gle96, MSS97]. stratification [MG94, TB91]. Stratified [Kla96, Arn91, CW94a, Gro91, Kre92, Kul95, NM94, San92, WG90, Wil94]. stratum [OSL91]. strength [CP92, CM94a, Han97a, KM93]. stress [Han97a, KM93, Mad93a]. Strict [Lee99]. Strong [BOR92, VV96, Xia95, Yan98]. strongly [CS97c, TZ90a]. Structural [HK90c, AVB96, Ban91, BD93, CP96c, GL93, HV91, HL93a, Hod90, LT92, MK94, RT91, Rob94, SP90, Wan98b, Won93]. Structurally [PD92, SB91]. structure [AM99a, GN95, Hol91, Ong93, PB93, SA93, TWM91, TW92, TC91, WJ90, YSM92, YB95]. Structured [Cha99, Ahn94, Den98]. structures [Gal94, Sun99, VV94, Yok96]. student [Gar96, Moo96, Var96, HK90b, KS98, Luc97b, Nag97, Smi92a, WA95]. Student- [HK90b, KS98, Nag97]. Studentized [aS99]. students [Moo96]. Studies [TC91, CB98, Cor91a, HGM94, HCL94, Kan90, LLH90, LC92, Loc90, OW93, RDM92, Tem94]. study [Bel93, Bre90, Cha91c, CM94a, CAS97, CO90, Dam93, HAKM93, HD94, KLH^{+}91, KAHK94, KL96, Moh90, Sin90, SM91b, Tem94]. study-level [HD94]. sub [CJ98a, Kum97, SMS97]. sub-diagonal [CJ98a]. sub-group [SMS97]. sub-sample [Kum97]. subclusters [QT98]. subfamiliy [Min98]. subject [Ali90b, MD93a]. Subjective [Swa93]. subpopulation [Yu98]. subranges [CR92b]. subsamples [Adh90, KKK91, Rao90b]. Subsampling [Kc91b, KC91a]. Subset [HV96, CH91a, LL91, SK95, Sob92b]. subsets [HS96, Jah91, ZM90]. subspaces [Bén90]. subsystem [Tsa90b, Tsa90a]. subtrees [Kim94c]. success [God92, KM97]. successes [KM99a]. succession [LL93]. Sufficient [EB98, Awa92, Mig98]. suitable [TS92]. sum [ATT99, Cho94c, Kot91, NSI90, SG90]. summaries [Fun95]. summary [MP91a]. summing [Son93]. sums [CJ98a, Ete91]. sup [Bro90]. sup-norm [Bro90]. super [Cha91c]. superconducting [KLH^{+}91]. superfluous [DSTW91]. superiority [IT93]. superpopulation [Arn91, Arn92, BR90b, GJO90]. supersaturated [BL98a, KA97a]. supplementary [Cha92a]. support [Kim94c]. supremum [Bou90a, Vio95, ZL91]. sure [Wan94c, gWyJ97]. Suresh [Vau92a]. surface [AB98, JP93, Mye91, PK98, PJ99, Vau93]. surfaces [Koo90, Kro97]. surprise [Eva97]. surrogate [CSK^{+}98, TM95]. Surveillance [Jär99, Wes98]. survey [Asc90, BSS94, CM94b, Cha97b, Che90, CL98, GS90, Gov95, Ham92, Hug91, Pra91, RS98, Rei90, SSS97, SMS93, Str91b]. surveys [Ahm97b, Ana97, Arn94, BR93a, Hun90, KA97b, Nay94, Pra89, PS90, PS92, Rag90, Sin90, WP91]. survival [Abd98, BE93, BDF95, CSK^{+}98, Chi99c, Coa94, DK95, EGM95, Fre92, Fri91, KK98b, Kan93, KM93, LG94, LY97, LTCC94, MPQ91, PC98, POGP95, TZ90b, TM95, WG90, ZK94]. surviving [CM94c]. Sweeping [FS93]. switching [AH93, AS94b]. symbol [Hyo94]. symmetric [FM92, GN95, HH94, Han93, Lia90, Liu97a, ML96, Nai90, Shi99, TWB99a]. symmetrical [AR90, CC90b, LD99a]. symmetrically [Rie95]. Symmetry [Cro91, AB92b, BM90, BH92b, HK92, Hu94, MK90a, Tom92]. system [Ana92, AS94b, CP96a, GM94b, KP93b, KM94c, Kun93, Lu90, Per91, PGC98, S99, SA92a, WS91b]. system-acceptance [AS94b]. Systematic [FP90, Koo90]. systems [BK96, CH97, CL98, DGMR96, ED93, Ho98, Jai91, SJ93].

table [KK98a, LS94, PE92, UG95]. Tables [Cox92, Cer97, Dav90, DY96, GG94b, Gup92a, KG95, Lev92, MGLF98, Myo95, OW90, PS93a, STL91, SS98, SP90]. Taguchi [NS99, SW91, Tsa90b, Tsa90a, Zac91]. tail [Cap92, CM95c, DJD96, GW91, GSD96, Gri90, Hug91, MPR98]. tail-area [Hug91]. tail-weight [Cap92]. tampered [Mad93a, Rao92]. tapping [OC99]. Tavaré [Yam97b, Yam97a]. Taylor [TM96a]. Teaching [SWD93, Sne96]. technique [BSS94, SD91, dR94]. techniques [Bel93, BSSB90, DGMR96, EHL93, KKP^{+}99, Kub99]. tell [Tur90]. Tensor [CLC97]. term [AK94b, Hal96, Mar92a]. terms [Has95a, Tur95, Wan96]. Test [AH97, Sob92a, YB94b, BM90, Bak91, BH92a, BMM^{+}91, BG95, BLY92, Boy95, BL90, Bro98, CGP92, Cal94, Car94, CFRT92, CS96a, Che95b, CH95b, Chi99b, CO91, CG92c, CG90, CLL95, Cro96, DB95a, DJ98, DRB93, Djo95, Epp95, ESY93, FM96, FKY95, Fri90b, Fri94, FR98, Fro96, GB98, GL97, Gao97, GG92a, Gas92, GL93, Gil93, GG93, GGO96, GD99, GC99, GW99a, Guo96, He90, Hen94b, HRS93, Hu94, Ibr91, ID98, Iso92b, JSM98, Kan93, KM91a, Kar94, KAKW99, Kib96, KHJY97, KB96, Koz90, Kro94b, Kul90, KA92b, KGM94, Lai98, LF91, Lee95a, LL90, Liu96, Liu97a, Lom99, Lu90, Mad93a, MMD93, Mat99, Mil91, MS90b, ML96, MSL95, NN91, Ng94, Oh98, Oht91, Olu93, PE92, PLW93, PT92, Pos92b, Pro90, RS98, Rao90b, RV95, RC94, SS91a, SK93]. test [SJ90, SJ91b, SJ91c, SM97, SBS96, SB95, SSA90, Stu91, SD90, Szr95, TSI98, Ter90, Tho96a, Tho96b, Tsu99, VR96, Wan94a, Wan98a, WG90, WW91b, WDH90, YY95, YL93, YK90, ZB97, cL91, dSC99]. testable [Ksh98]. testimation [BR90a, RGU97]. Testimator [Pan97, CC90a]. Testing [Aga90, BB97, BK98, Bha97, BDT98, CS97a, CG91, Che98b, El 97, EHL93, Fam93b, FWG94, FM97, Gas92, GM96, HS90b, Hon98, JN91, Kar99, Koz94, LN90, Mac98, MRS92, MJ92a, PM90, RR97, SS98, SS94c, Sil94, Tan98, TV91, VT93, Was94, ZB93, AB94, Ban91, Bro95, CP99a, CB90, CGE99, Dou97, Est96, Fri90a, Gho91, GT91, Gil91, GD99, GVSM92, GN95, HK92, JS90, JBY90, Kab91a, KP92b, LS94, MT92, MPB90, Olu93, PFU91b, RFG92, RT96a, RR91, Sat91b, SAK97, SGB98, SD93, TK92, TC91, Tur90, WN97, WX99, YK90]. Tests [CS96b, FN92, Han97b, HR98, KTTL99, LG94, LPK95, Lev92, MM93, SJ91d, WM99, Wit91b, AD98, AKH91, AAW93, AC95, AS92, Asc90, Baj95, BB90a, BH92b, BR93a, Bri93, Bri97, BD93, CC96, Cap92, CL90, CM94c, CHC98, Coa94, CN99b, CF96b, Cor91b, Dav90, DM96, DH95, Dun95, Dut99, Fri91, Gau98, Geo91, GW94, Gov95, GW99b, HK91, HK92, HL90, HL92, HZ90, Hen92, HK95, Her96, HT98, Ho98, HC94, HV96, HY97, IH94, Iso93, JSM97, JR91, JK91, JZ96, JS95, Kam99, KK98a, KAKW99, KT91, KF97a, Kla96, Kul94, Kul95, Kum97, KS90c, Law93, Le94, LW95, LLL92, LP98b, LS92, Liu92c, Man91, MR90, Meh93, MMPV96, Men92a, NS96, Nor93, Ozt91, Özt99, Pan96, Par93a, RP90, RSAJ90, SA99, Sch91, SS93b, SSSS90, Sha90, SGB98]. tests [SS94a, Shi99, SSW93, Sma93, Sug99, Sun97, SS94e, Tar90, Tha92, TK93, UNM92, Van93, Ver96, Was94, Wen94, Wij93, YL92, ZQ97, Zha96, ZFP90, Zou94, dP90]. th [HSS93, PS99a]. Thanks [Hog96b]. their [Bro90, Con90b, CN99b, DN92, DAM99, ES95, JSM97, Kul90, Liu92a, Ryc93, SK93, SR91, Swi98, TCT90, WS99]. theorem [And91b, Dav91, ED93, KA92b, Tan94, WH99]. theorems [Bri97, CS90c, ESDC96, Yon95]. theoretic [LL96, Mad92]. theoretical [Kmi95]. theory [Ang92, BWCM90, BR93b, ES94, Gao95b, Kab91b, KM96, LH99, Mof95, RBC97, RD94, RW96, Sar97]. there [Men92a, Wil96]. thermal [NKN99]. thin [GSD96]. Third [CFRT92, Mae96]. Thompson [Adh91a]. Three [CMD94, CC90b, DS91, Gro94, Raj91, Aga90, BSK95, Cer97, Cha97b, ED93, KT91, Lia99, Liu96, MLGD98, RT96b, Sch97, SS98, US92]. Three-factor [CC90b]. three-level [Lia99]. three-parameter [Sch97]. Three-stage [CMD94, Raj91]. three-way [SS98]. threshold [BC92, DY98, Lim93, POGP96, XB97]. tidal [SSCS99]. tightened [SB93]. tightened-normal-tightened [SB93]. tightening [AS94b]. Tightness [YL91]. Tiku [Vau92a]. Time [BJ94, CJ98b, TWB99b, AD98, AOA92, BM98, BB97, Bev92, Bol90, BK97b, CP99a, CCC96, CBR91, Che92a, Che92b, CM94c, Cho94c, Cip98, DN92, DB95b, DQS98, Dia91, Ete91, Gab92, Gab98, GS95, Gle95, GPA98, Gup92b, GL94, GGL96, GGG98, HAKM93, JP90, KTTL99, Kat95, KB90, KCM94, KP93b, KM96, LG94, LdBPR92, Li94, Li99, LL93, Mar90b, MD93a, MPQ91, NT92, NC97, Ng90, NM96, PS93b, RAP95, RR97, Rao90a, RS96, RT96b, RT91, RZ94, SSSS90, Sin97, TM95, WK91b, vHL92]. time-dependent [Li99, TM95]. time-series [Mar90b]. Time-trend [CJ98b]. time-varying [Kat95]. times [Cos98, Moh90, Nag99, Rai96, RC95, SSSS90, SJ93, WC96b]. TNT [SB93]. Toeplitz [Cho90b]. tolerance [BK91, CT94, GS93, GS94, Hor92, KK91c, MG91, Pap92, SH94, Tsa90b, Tsa90a, WL96]. tolerances [Vän97b]. tool [SS94d]. tools [RS96]. topological [Kea91]. total [Arn91, KP93b, Men92b, SG91a, WN92]. totals [KA97b]. toxicity [MF90]. toxicological [FS95a]. toxicology [SCXG96]. trace [CJH95, GM98a]. traces [AM93]. traditional [SA90c]. transect [Dru90, Mac98, MQZ99]. transform [RD94]. Transformation [RM93, HS90b, Kim96b, LW94, TSI98, Tre94, Ver96, Won92, Yan96b]. transformations [CC96, Che97, FN95, HP92, JR90, LLL92, LLL96, QL92, TK97b, Vel92]. Transformed [KS97, LYS96, LB98, MO91, RG92, RM90, cL91]. transforming [LLL93]. transforms [Cam93]. transition [Wen94]. transitive [Sob96]. treating [WK97]. treatment [Bil97, CS96a, CP99b, Dri91, Ent90, FWG94, HV96, Jai94, MR90, Nor93, Sin92, Stu91]. treatments [CG92a, CS94b, GM94a, GP96, HL92, KC98, KT91, LY94, Stu91]. Tree [Ahn94, CH96, Den98, HY97, Kim94c, MS95, MT92, SSW93]. Tree-structured [Ahn94, Den98]. trees [Ahn96]. trend [ACR91, CP99a, Cox92, CJ98b, El 97, Fro96, HT98, JSM97, KS90c, Le94, LP98b, SM91a, WG90, YB94b]. trends [CGP92, FP90, HAKM93, RS96, SS91b, SM92, Wan99]. trial [Ent90, KP92a, Mor90, Ng94, Pat94b, Ree96, SLL97, TY99]. triallel [DG97a, PS91]. trials [AFR94, Arn96, AB91b, CMD94, Cha96b, CS96a, CSK^{+}98, CL94, Dou97, Dou93, FWG94, HD94, Jai94, LRD94, LM97, Mar90b, MF90, Ric90, RT91, Rob94, Rog92, Sco99, SGJ95, SGS94, SL98a, Sna94, TM95, Wan91, Was94, YR94]. triangular [GS98a, OS93b]. Trimmed [CT98, CP96c, MMS91]. trimmed- [MMS91]. trinomial [SNM97, Tre94]. triple [CL98]. Triserial [Sch90]. true [LF93]. truncated [AL96, Bal94, BP99a, CP96b, Car90, DY91, Don93, ES95, Fun91, GG92a, Gil92, GLL90, HD99a, JP99, KL99b, LBB92, MP91a, MJ92b, PC98, Sha92a, SS99d, SK91b]. Truncation [Sun99, Fer90, KHR97, PB90b, PB91, PB92, RT96a, RS90]. Tukey [BS96, FZ96, Liu96]. tumor [Rao90a, WP92]. tumorigenicity [KCM94]. Two [CS94b, CP96c, Con90d, IH94, Kim93, MMPV98, NA90, Pap92, TA96, YT95, Ahm97a, Ahs96, Ano99m, Aos94, AM99a, AR90, Arn91, Baj95, BK97a, BR90a, Bre90, Bri93, Bro95, BD93, CPK92, CH94, CM92, CS96a, CO91, CG92c, Cor92, Cor91b, DB95a, Dou97, EP92a, EP92b, FK94, FKT95, Fer90, Fun98, Gal94, GT91, GW94, Gov95, GS91b, GP96, Her96, Hil93, HC90, HAKM93, Hyo94, Jah91, JP91b, Jin95, JP92, KMM90, KHR97, KK98a, Kar94, KS90a, Kel96, KT91, Klo90, Koo90, Koz94, Kum97, KM94c, Lee94b, Lee95a, LM97, LV98, LL90, Liu95b, Mae98, MG94, MH95, MM93, MM96a, Muk95, Nab93, NA94, Olu93, Özt99, PB98, PM94, Par90, Par93b, PB90b, PB91, PB92, PT93, PGT98, Pos92a, Pos92b, QCL98, QL92, QSC94, RRA92, Rak92a, Rak92b, RP90]. two [RS92, RD96, Ric90, RSAJ90, Ros96, STL91, SLS96, SA99, SA90b, SS93b, SSSS90, Sco95b, SM97, Sha90, SJ93, aS99, SG90, SGN93, SSSS94, SUS95, SM91b, SW98, SG91a, SL94b, Swi98, Tah91, Tha92, TV91, TK91b, Udd98, US92, Vas96, VCvdL97, VS91, Wan93a, Wan96, WK97, WSM91, Wik98, Wil92, WDH90, Won95, Yu91, YB94a, YC99b, Yue99, Zha95b, Zhe99, Zou94, cL91]. Two- [Pap92]. two-arm [HAKM93]. two-class [QL92]. two-component [RRA92]. two-dimensional [GS91b, Koo90]. two-factor [aS99]. two-level [Hyo94, JP92]. two-parameter [FK94, FKT95, LL90]. two-period [Cor91b, Udd98]. Two-phase [Kim93]. two-point [SW98]. two-sample [Ano99m, Baj95, GW94, Jin95, Kum97, Lee95a, Özt99, Pos92b, RP90, Ric90, SGN93, SL94b, TK91b, Wan93a, cL91]. two-sided [CS96a, Her96, Kar94, LV98]. Two-stage [CP96c, TA96, YT95, Aos94, AM99a, Arn91, BR90a, HC90, LM97, Muk95, Par93b, QSC94, Yue99, Zhe99]. two-state [Lee94b]. two-symbol [Hyo94]. two-truncation [PB90b, PB91, PB92]. two-way [Bre90, Gov95, STL91, Sha90, SG91a, Tha92, WDH90]. type [AHJ95, AAB97, Ana92, Bas93, BS90c, Chu92, CF96b, DM96, Dem96, Fer90, Fri92, GG94a, GR90, GGO96, HH95, JZ96, KFAH92, KS97, KF97a, KY91b, Lai99, LS97b, Men92a, MR95b, Nag96, Nag97, Oht93b, PFN92, PB90b, Pra89, PS90, Pra91, RV95, SAS91a, SS94a, Ter90, TS92, VR96, BGP95, CHS98b, DH90, Den91, FK94, FKT95, HO90, KSmS90, PB91, PG92, PG95, RSAJ90]. type-2 [Chu92]. Type-II [BGP95, FK94, FKT95, KSmS90].

UE [RZ98]. ultrastructural [Huw93, JT90]. umbrella [CS96b, Pan96]. UMVSU [Muk97]. UMVSU-estimation [Muk97]. Umvu [LK98, LB98, LBBCMR97, Vel93]. unbalanced [BK91, ÉA95, FS93, Hei93, Mon99, Sin91, SM91b, SG91a]. Unbiased [CJH95, KR90, PB90b, PB91, PS92, SG91b, Tap92, AK95, BBK98, Chi99c, Fer90, Gro99, HK90d, HB91, KJ91, NS93, Sub91, TS92, TL94, Tsu99]. Unbiasedness [Che94, EF92, Fri90b]. Unbiastld [PB92]. unblinding [GS92]. unblocked [Goo91]. unconditional [BF90, VSS95]. uncontrolled [SGJ95]. Uncorrelated [CG92c, HS90a]. uncorrelatedness [Lye91]. undamped [Kun95]. undergraduate [Var96]. unequal [Adh90, CM95c, LC91, Liu96, Mon99, PK94, SL98a, SC93, SC96, SSW93]. unification [Gri90]. unified [Arn96]. Uniform [CL90, VV96, Yu97, Ahs96, BCLBMR99, BR97, CT97, EP92a, Fan91c, FRS97, MD95, SSSS90, Son92]. uniformity [Dun95, Kc91b, SD90]. uniformly [HK90d, MP91b, WC96b, gWyJ97]. unimodal [DLS97]. unimodality [MP93]. Uniqueness [CP99b, KK98b]. unit [AD98, GM94b, Li94, OSL91, RM90, SS94a, SS94c]. unit- [RM90]. units [CP99a, KH90]. unity [RG94]. univariate [Jah91, NvdMM90, NM96, Ozt91, Sri95, Tak97b, Ver96]. Universal [RZ98]. universe [Koo93a]. University [Bra90, Ley90]. unknown [BB90d, Bro95, Cha91a, CZ97b, Fri90a, Fri90b, Fuj94, GS92, MT90, Ng90, PT92, Sch97, ST96, TA96, Ver96, Wil96]. unmeasured [SPQ98]. unobserved [NM92]. unpaired [Fuj99b]. unrelated [KMBTG99, Par93b]. unreplicated [JP92, Lee94a]. unstable [BMM^{+}91]. Untransformed [Cho94b]. unweighted [SG91a]. upcrossings [SR91]. update [Dra90, Dra92, Dra94, Dra98]. updated [Rei90]. upon [SA90c]. Upper [DH90, BG95, BM93, Hor92, MN98]. ups [TZ90b]. upward [SR91]. Use [KSm96, Loc90, Sna94, AR90, AB91b, GdWG91, Gro94, KA92a, MA95, Mar90b, Sch97, WS99, WW91a, Yip98]. used [LSR94, ML91, SJ91b, SK97a]. useful [NHY95]. user [Lia99]. user-specified [Lia99]. uses [Mar95]. Using [Bru95, FN95, LLH90, Moo96, PGC98, RD94, AL96, Ahm97a, SAP92, Bak91, BSS94, BCLBMR99, Bho97, BM94, BSSB90, CS97a, CHS98b, CW92a, Cho90b, CD98, Cor90, DP97a, DH92, FT95, Gle96, GNR98, Hid91, HL93a, HS93, HL93b, Ibr91, JB93, KP92a, KY91a, KKK90, LLL96, LWH98, LG93, LM94, Mad92, MB91a, NPA93, NC97, PZ99, PR92, Pan97, PE92, PF93, PK91b, PL96, POGP95, PS91, PS92, RG94, RT91, Rob94, Rod98, RS95, SA99, SS94b, SMS93, SG91a, Ste90, Tan94, TB91, TCSJ94, Tsa90b, Tsa90a, Was94, WM99, Wri90, YC99a, YR94, ZB93]. usual [GT97, Mar96, Nag99, Tom92]. utilities [GdWG91]. utilizing [HD94].

V [HC96, RC96]. vaccine [CB98, Yip98]. vague [Res96]. Validated [Sob92a]. validation [Chu94a, Ham91, SS99a, SH91]. validity [BGMT90]. Value [PG94, Ahn94, AG96, AS92, BCA93, BGP95, Bho97, Cal94, Che98a, Chi99b, DRB97, DM96, DH96, DD96, DT98, FKT95, Fuj90, KSmS90, LZZ99, RW96, SGA99, Sca92, Tho96a]. valued [Kun91, Mee90]. values [BCA93, BA94, BB90b, Ber98b, Bev92, BS98a, Bor95a, Buo90, Dun95, JS93, Kam94, Koc90, KM91b, PS99a, Tho96b]. variability [LC92, Sri97]. Variable [BM94, MA94, SC97, Vau92b, AH93, AW99, BT93, CS94a, FZ97, Fuj90, HB90, HK99, Kab97, MC91, MW98, Mor91, Rao92, SMS97, SP99, SM91b, SU95, SC92, TL99, Tsu98, WK97, WA96, as97]. variable-selection [Kab97]. variables [Abd95, ATT99, Ami96, Bal94, BB95, Ban90, CS90a, Cox92, Dag99, DB95a, DSTW91, EGM95, Fan91b, GK92, cH91b, JZ96, Kam94, Kap92, LV98, Leu94a, bL95, LR96, Loy90, Mat91, Men97, MD90, NKN99, PR93, Per91, PGT98, PGC97, PGC98, Pro90, RM90, STR97, Sat90, ST90b, SC93, Smi92a, Son93, SA98, WK93]. Variance [CA90a, CM94b, FT95, WJ90, AP93, Arn94, AG90, BY92, Bre90, CM92, CZ97b, CG92c, DMG91, DU91, DY96, DM91, EB94, ÉA95, FTS94, Fer90, Fri90a, Fri90b, Gle95, Gle96, GS92, Hei93, Hil93, Hoc90, Hod90, HK94, HK90d, HB91, HL93b, IU93, KMM90, KM91a, Kub99, LC91, LC96b, LSR94, Lia94, LZ98b, LLH90, Loy90, LK98, MR95a, MEC93, MMT97, Men92b, MS90b, MM96b, MP97, Nag95, Nag96, Nag97, NHY95, Nas91, NS93, Oht91, Oht93b, PM91a, PM94, PB93, Par98, PK98, Pos92b, PS90, PS92, SLS96, SS91b, SL94a, SCSG94, SS92, Sin91, SG91a, SC92, SB90, SB91, SS99c, Sub91, Tur90, VP98, Yan90a, YB95, ZK90, vdMvdM90]. variance-balanced [SS99c]. variance-covariance [MEC93]. variances [Ali90b, BM93, Bor95b, Bre90, CW98, EP92b, Lav92, Le94, Pep90, SL98a, SC93, SC96, Sin91, Str91b, ST99]. variant [Hen94a, Oht96a]. variants [NC97, TS94]. variate [AB91a, Bou90b, MSL95]. variates [Den94, US92]. variation [BM98, Cha93b, GM96, Hsi90, JG95, JS90, JS91b, JN91, LR96, LD95, Mil91, ML91, MF97, RR91]. variations [SG91a]. varieties [Dou93]. variety [Dou93]. various [Myo95]. varying [Kat95]. Vector [CLL95, YN99, AB92b, BB90d, Che99, Cho90b, Cho94a, FBD97, Gru93, GN95, JB93, Jin95, Kea96, LSU90, Oh98, RS92, TK97a]. vector-based [Che99]. vectors [CLL95, Dou97, Ete91, HS90a]. verification [hZ93, ZR98]. version [GP98, NTS91, WH99]. versus [BF90, BG91, CM96b, Fro96, Luc92b, Oh98, TK91b]. via [ALW91, CH91c, DF95, EHL93, Eva97, HV96, KL99a, KKP^{+}99, KL96, PC98, Phi96, QT98, WS91b]. Victorian [Moh90]. Violation [MSS97]. violations [Pos92b, SA90b]. violators [TL91]. visit [Mof95]. vitality [RG96]. vith [SSSS90]. volume [Ano96l, Chu92]. vs [AKH91, HO90, MR90, SJ91b]. VSSI [Cos98].

Waiting [LdBPR92, LL93]. Wald [HD99b, TK93, dSC99]. walk [SR91, Sar92a]. Walker [Cho90b]. walks [Con94a, Xu90]. Wallis [LC91]. Wang [KO94b]. warranty [Men92b]. Watson [ES92, MGM98]. wavelet [LZZ99, Wu96, ZZ99]. wavelets [WS99]. way [BK91, Bre90, BH99, Cer97, CG92a, CG90, DM91, ÉA95, Gov95, HEB92, Hei93, KK98a, Klo90, Liu91, STL91, SS98, Sha90, SG91a, Tha92, Tur90, WL96, WDH90, vdMvdM90].

Weak [Sun97, VV97]. weaker [Che94]. weakly [Gro99, Hon99]. Wedderburn [LH93]. Weibull [AHAEH90, AHAEK92, CP94, CP95, CC90a, FN91, FKT95, GBC99, Lu90, MP91a, MK94, MH96, Sch97]. weighing [WN92]. weight [Cap92, Koc95, RV95, SBS96, VR96, WN92]. Weighted [Bro90, DD97a, GT96, PM91a, RZ94, Shi99, SCXG96, AN91, CM97, Cha90b, FMP98, GNR98, GK90, JN95, Jor99, Koc95, LTvC^{+}97, Mar96, Qum97, RV99, Sco99, TTG92, Vio95]. weighting [CM95a]. weights [SSW93, SW97]. where [Nor93, Tho96a]. which [Kun93]. white [Cho94a]. whiteness [BB97]. Whitney [FR98, Ibr91, PC99]. wider [Jan97]. width [HEB92, Tah92]. Wiener [DJ98, YC99a]. Wilcoxon [LL98, FR98, PC99]. Wilk [MSL95]. Wilks [Kab91b, KG90]. will [Mil96]. window [Cha96a]. winner [AFR94]. wise [GW92, MP93, PM92]. Wishart [BR95, DG94, GM98a]. withdrawals [Ent90]. within [Bri97, GM94a]. without [Adh90, EB92, GS92, Lim93, MW96, MQZ99, ML93, SMS93, Tu99]. WLS [CB94]. Wolfowitz [Nag96]. women [Fre92]. worst [MD93b].

XII [AHJ95, HN97].

Yamada [HD92]. years [Bra90, Ley90]. yield [GM98b, Nai96]. Yule [Cho90b].

zero [Moh90, MS98, Rao90a, Rao92, RDA93, Sco97]. zeros [SP90]. zone [DHW90, DJG94].

References

Anderson:1991:CDC

[AA91] Dale N. Anderson and Barry C. Arnold. Centered distributions with Cauchy conditionals. Communications in Statistics: Theory and Methods, 20(9):2881–2889, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ali:1992:IEC

[AA92] M. A. Ali and M. Masoom Ali. Impacts of equi-correlated responses on multicollinearity. Communications in Statistics: Theory and Methods, 21(7):1965–1975, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Alonso:1996:ESI

[AAB96] F. J. Alonso, J. M. Angulo, and M. C. Bueso. Estimation and smoothing from incomplete data for a class of lattice processes. Communications in Statistics: Theory and Methods, 25(3):667– 681, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Alonso:1997:AET

[AAB97] F. J. Alonso, J. M. Angulo, and M. C. Bueso. Application of EMtype algorithms to spatial data. Communications in Statistics: Theory and Methods, 26(3):669–683, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Al-Awadhi:1998:EMM

[AAG98] Shafeeqah A. Al-Awadhi and Paul H. Garthwaite. An elicitation method for multivariate normal distributions. Communications in Statistics: Theory and Methods, 27(5):1123–1142, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Alam:1993:MGF

[AAW93] Khursheed Alam, Roger Abernathy, and Calvin L. Williams. Multivariate goodness-of-fit tests based on statistically equivalent blocks. Communications in Statistics: Theory and Methods, 22(6):1515–1533, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Abughalous:1991:MND

[AB91a] Mansour M. Abughalous and Naveen K. Bansal. On the minimaxity of the natural decision rule for selecting the best variate in a multivariate normal population. Communications in Statistics: Theory and Methods, 20(5–6):1557–1567, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ash:1991:SLU

[AB91b] Charles W. Ash and Robert H. Brown. Significance levels for use in the analysis of combination drug trials. Communications in Statistics: Theory and Methods, 20(10):3283–3292, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Adke:1992:EMS

[AB92a] S. R. Adke and N. Balakrishna. Estimation of the mean of some stationary Markov sequences. Communications in Statistics: Theory and Methods, 21(1):137–159, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ahmed:1992:EMV

[AB92b] S. E. Ahmed and S. O. Badahdah. On the estimation of the mean vector of a multivariate normal distribution under symmetry. Communications in Statistics: Theory and Methods, 21(6):1759– 1778, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Abughalous:1994:PSB

[AB94] Mansour M. Abughalous and Naveen K. Bansal. On the problem of selecting the best population in life testing models. Communications in Statistics: Theory and Methods, 23(5):1471–1481, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Aggarwal:1998:RRS

[AB98] M. L. Aggarwal and Anita Bansal. Robust response surface design for quantitative and qualitative factors. Communications in Statistics: Theory and Methods, 27(1):89–106, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Abdous:1993:NMM

[Abd93] B. Abdous. Note on the minimum mean integrated squared error of kernel estimates of a distribution function and its derivatives. Communications in Statistics: Theory and Methods, 22(2):603– 609, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Abdullah:1995:DIO

[Abd95] Mokhtar B. Abdullah. Detection of influential observations in functional errors-in-variables model. Communications in Statistics: Theory and Methods, 24(6):1585–1595, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Abdushukurov:1998:NES

[Abd98] A. A. Abdushukurov. Nonparametric estimation of the survival function from censored data based on relative risk function. Communications in Statistics: Theory and Methods, 27(8): 1991–2012, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Abt:1992:SEO

[Abt92] Markus Abt. Some exact optimal designs for linear covariance functions in one dimension. Communications in Statistics: Theory and Methods, 21(7):2059–2069, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Alvo:1995:AED

[AC95] Mayer Alvo and Paul Cabilio. Approximate and exact distributions of rank tests for balanced incomplete block designs. Communications in Statistics: Theory and Methods, 24(12):3073– 3121, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Alvo:1999:GRB

[AC99] Mayer Alvo and Paul Cabilio. A general rank based approach to the analysis of block data. Communications in Statistics: Theory and Methods, 28(1):197–215, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Arslan:1995:CBE

[ACK95] Olcay Arslan, Patrick D. L. Constable, and John T. Kent. Convergence behavior of the EM algorithm for the multivariate tdistribution. Communications in Statistics: Theory and Methods, 24(12):2981–3000, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Aerne:1991:ECC

[ACR91] Laura A. Aerne, Charles W. Champ, and Steven E. Rigdon. Evaluation of control charts under linear trend. Communications in Statistics: Theory and Methods, 20(10):3341–3349, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Arnold:1995:GCS

[ACS95] Barry C. Arnold, Enrique Castillo, and José Mariá Sarabia. General conditional specification models. Communications in Statistics: Theory and Methods, 24(1):1–11, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Akdi:1998:PUR

[AD98] Yilmaz Akdi and David A. Dickey. Periodograms of unit root time series: distributions and tests. Communications in Statistics: Theory and Methods, 27(1):69–87, 1998. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Adhikary:1990:NIS

[Adh90] Arun Kumar Adhikary. A note on interpenetrating subsamples of unequal sizes drawn with and without replacement. Communications in Statistics: Theory and Methods, 19(4):1475–1481, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Adhikary:1991:OHT

[Adh91a] Arun Kumar Adhikary. On optimality of the Horvitz–Thompson estimator under a Markov process model. Communications in Statistics: Theory and Methods, 20(7):2291–2297, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Adhikary:1991:PNP

[Adh91b] Arun Kumar Adhikary. On the performance of the nearest proportional to size sampling design. Communications in Statistics: Theory and Methods, 20(12):3933–3941, 1991. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anderson:1995:RBG

[AF95] D. A. Anderson and W. T. Federer. Representation and bounds for the general fractional factorial. Communications in Statistics: Theory and Methods, 24(2):363–373, 1995. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Adkins:1998:NIP

[AF98] Laura Adkins and Michael Fligner. A non-iterative procedure for maximum likelihood estimation of the parameters of Mallows’ model based on partial rankings. Communications in Statistics: Theory and Methods, 27(9):2199–2220, 1998. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Andersen:1994:RPW

[AFR94] John Andersen, Doug Faries, and Roy Ramura. A randomized play-the-winner design for multi-arm clinical trials. Communications in Statistics: Theory and Methods, 23(2):309–323, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Afsarinejad:1990:RMD

[Afs90] K. Afsarinejad. Repeated measurements designs — a review. Communications in Statistics: Theory and Methods, 19(11): 3985–4028, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Asiribo:1990:CVH

[AG90] Osebekwin Asiribo and John Gurland. Coping with variance heterogeneity. Communications in Statistics: Theory and Methods, 19(11):4029–4048, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Alves:1996:SCE

[AG96] M. I. Fraga Alves and M. Ivette Gomes. Statistical choice of extreme value domains of attraction — a comparative analysis. Communications in Statistics: Theory and Methods, 25(4):789– 811, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Agarwal:1990:TME

[Aga90] B. L. Agarwal. Testing a main effect in a three factor mixed model. Communications in Statistics: Theory and Methods, 19 (2):723–738, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ageo:1999:POC

[Age99] Gianna Ageò. Parameter orthogonality and conditional profile likelihood: the exponential power function case. Communications in Statistics: Theory and Methods, 28(8):1759–1768, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Amin:1993:SBC

[AH93] Raid W. Amin and Rohan Hemasinha. The switching behavior of charts with variable sampling intervals. Communications in Statistics: Theory and Methods, 22(7):2081–2102, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ahsanullah:1997:TIC

[AH97] M. Ahsanullah and Mohamed Habibullah. Test of independence of a class of bivariate distributions based on mutual information. Communications in Statistics: Theory and Methods, 26 (11):2803–2816, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Al-Hussaini:1990:EPI

[AHAEH90] Essam K. Al-Hussaini and Nagi S. Abd-El-Hakim. Estimation of parameters of the inverse Gaussian–Weibull mixture model. Communications in Statistics: Theory and Methods, 19(5):1607– 1622, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Al-Hussaini:1992:ESS

[AHAEK92] Essam K. Al-Hussaini and Nagi S. Abd-El-Kim. Efficiency of schemes of sampling from the inverse Gaussian–Weibull mixture model. Communications in Statistics: Theory and Methods, 21(11):3143–3169, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Al-Hussaini:1994:ABE

[AHJ94] Essam K. Al-Hussaini and Zeinhum F. Jaheen. Approximate Bayes estimators applied to the Burr model. Communications in Statistics: Theory and Methods, 23(1):99–121, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Al-Hussaini:1995:BPB

[AHJ95] Essam K. Al-Hussaini and Zeinhum F. Jaheen. Bayesian prediction bounds for the Burr type XII failure model. Communications in Statistics: Theory and Methods, 24(7):1829–1842, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ahmed:1991:CPM

[Ahm91] S. E. Ahmed. Combining Poisson means. Communications in Statistics: Theory and Methods, 20(3):771–789, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ahmed:1997:GCC

[Ahm97a] M. S. Ahmed. The general class of chain estimators for the ratio of two means using double sampling. Communications in Statistics: Theory and Methods, 26(9):2247–2254, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ahmed:1997:PPP

[Ahm97b] S. E. Ahmed. To pool or not to pool the proportion in randomized response surveys. Communications in Statistics: Theory and Methods, 26(10):2499–2525, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ahmed:1998:IPN

[Ahm98] S. E. Ahmed. Improved pretest nonparametric estimation in a multivariate regression model. Communications in Statistics: Theory and Methods, 27(10):2391–2421, 1998. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ahn:1992:FPP

[Ahn92] Sung K. Ahn. F-probability plot and its application to multivariate normality. Communications in Statistics: Theory and Methods, 21(4):997–1023, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ahn:1994:TSE

[Ahn94] Hongshik Ahn. Tree-structured extreme value model regression. Communications in Statistics: Theory and Methods, 23(1):153– 174, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ahn:1996:LGR

[Ahn96] Hongshik Ahn. Log-gamma regression modeling ’through regression trees. Communications in Statistics: Theory and Methods, 25(2):295–311, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ahsanullah:1996:GOS

[Ahs96] M. Ahsanullah. Generalized order statistics from two parameter uniform distribution. Communications in Statistics: Theory and Methods, 25(10):2311–2318, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Aparisi:1999:SPL

[AJC99] Francisco Aparisi, Jose Jabaioyes, and Andres Carrion. Statistical properties of the LSI multivariate control chart. Communications in Statistics: Theory and Methods, 28(11):2671–2686, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ahsanullah:1991:CED

[AK91] M. Ahsanullah and S. N. U. A. Kirmani. Characterizations of the exponential distribution through a lower record. Communications in Statistics: Theory and Methods, 20(4):1293–1299, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Al-Khalidi:1994:MLD

[AK94a] H. R. Al-Khalidi. On the misspecification of a lognormal distribution. Communications in Statistics: Theory and Methods, 23(8):2343–2350, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Alam:1994:ETB

[AK94b] Khursheed Alam and K. B. Kulasekera. On the error term in Bahadur’s representation of an order statistic. Communications in Statistics: Theory and Methods, 23(12):3361–3371, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Akdeniz:1995:AUG

[AK95] Fikri Akdeniz and Selahattin Kaçiranlar. On the almost unbiased generalized Liu estimator and unbiased estimation of the bias and MSE. Communications in Statistics: Theory and Methods, 24(7):1789–1797, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Agarwal:1996:GGD

[AK96] S. K. Agarwal and S. L. Kalla. A generalized gamma distribution and its application in reliability. Communications in Statistics: Theory and Methods, 25(1):201–210, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Arslan:1998:NML

[AK98] Olcay Arslan and John T. Kent. A note on the maximum likelihood estimators for the location and scatter parameters of a multivariate Cauchy distribution. Communications in Statistics: Theory and Methods, 27(12):3007–3014, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Aggrawal:1999:CAA

[AK99] M. L. Aggrawal and Renu Kaul. Combined array approach for optimal designs. Communications in Statistics: Theory and Methods, 28(11):2655–2670, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Al-Khalidi:1991:SPT

[AKH91] H. R. Al-Khalidi and L. J. Hwang. Some pitfalls of tests of separate families of hypotheses: normality vs. lognormality. Communications in Statistics: Theory and Methods, 20(8):2505–2528, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Al-Khalidi:1993:SAM

[AKP93] H. R. Al-Khalidi and A. R. Parsa. Some aspects of maximum likelihood estimation of multistage cancer dose-response models with application. Communications in Statistics: Theory and Methods, 22(12):3377–3392, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Adamidis:1996:ETB

[AL96] Konstantinos Adamidis and Sotirios Loukas. Estimation in a truncated bivariate Poisson distribution using the EM algorithm. Communications in Statistics: Theory and Methods, 25(10): 2215–2222, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Alalouf:1996:EPC

[Ala96] I. S. Alalouf. The estimation of a proportion in cluster sampling. Communications in Statistics: Theory and Methods, 25(2):325– 343, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Albert:1992:CLR

[Alb92] Jeffrey M. Albert. A corrected likelihood ratio statistic for the multivariate regression model with antedependent errors. Communications in Statistics: Theory and Methods, 21(7):1823– 1843, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ali:1990:BAD

[Ali90a] M. A. Ali. A Bayesian approach to detect informative observations in an experiment. Communications in Statistics: Theory and Methods, 19(7):2567–2575, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ali:1990:MMI

[Ali90b] M. A. Ali. Minimization of model inadequacy when observations subject to shift in means and variances. Communications in Statistics: Theory and Methods, 19(4):1467–1473, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ali:1991:IAI

[Ali91a] M. A. Ali. Impacts of auxiliary information on outliers in regression. Communications in Statistics: Theory and Methods, 20(10):3271–3281, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ali:1991:RMB

[Ali91b] M. A. Ali. Restricted minimum bias linear estimation in regression. Communications in Statistics: Theory and Methods, 20(12):3751–3760, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Adatia:1991:CMG

[ALW91] A. Adatia, A. G. Law, and Q. Wang. Characterization of a mixture of gamma distributions via conditional finite moments. Communications in Statistics: Theory and Methods, 20(5–6): 1937–1949, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Alwan:1992:EAC

[Alw92] Layth C. Alwan. Effects of autocorrelation on control chart performance. Communications in Statistics: Theory and Methods, 21(4):1025–1049, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Andrew:1991:REL

[AM91] Luong Andrew and K. Tak Mak. Robust estimation in a linear functional relationship model. Communications in Statistics: Theory and Methods, 20(2):721–733, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anderson:1993:EQF

[AM93] T. W. Anderson and R. P. Mentz. Evaluation of quadratic forms and traces for iterative estimation in first-order moving average models. Communications in Statistics: Theory and Methods, 22 (4):931–963, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Aoshima:1999:SOP

[AM99a] Makoto Aoshima and Nitis Mukhopadhyay. Second-order properties of a two-stage fixed-size confidence region when the covariance matrix has a structure. Communications in Statistics: Theory and Methods, 28(3–4):839–855, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Azcarate:1999:NEA

[AM99b] Cristina Azcárate and Fermín Mallor. A nonparametric estimator from autopsy data. Communications in Statistics: Theory and Methods, 28(2):479–494, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Aminzadeh:1993:ESC

[Ami93] M. S. Aminzadeh. Exponentially smoothed control charts for the inverse-Gaussian process. Communications in Statistics: Theory and Methods, 22(6):1665–1680, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Aminzadeh:1996:IGA

[Ami96] M. S. Aminzadeh. Inverse-Gaussian acceptance sampling plans by variables. Communications in Statistics: Theory and Methods, 25(5):923–935, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Aithal:1990:SOE

[AN90] B. U. Aithal and B. N. Nagnur. Second order efficiency of the BAN-estimators for the multinomial family. Communications in Statistics: Theory and Methods, 19(5):1763–1776, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Arnold:1991:SPB

[AN91] Barry C. Arnold and H. N. Nagaraja. On some properties of bivariate weighted distributions. Communications in Statistics:

Theory and Methods, 20(5–6):1853–1860, 1991. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ananda:1992:EER

[Ana92] Malwane M. A. Ananda. Estimating the exponential reliability function for a system in series with type II censored data. Communications in Statistics: Theory and Methods, 21(11):3219– 3227, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ananda:1997:BMM

[Ana97] Malwane M. A. Ananda. Bayesian methods for mark-resighting surveys. Communications in Statistics: Theory and Methods, 26 (3):685–697, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anderson:1990:PAE

[And90] Oliver D. Anderson. On the partial autocorrelations for an explosive process. Communications in Statistics: Theory and Methods, 19(9):3505–3526, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anderson:1991:PAE

[And91a] Oliver D. Anderson. On the partial autocorrelations for an explosive process. Communications in Statistics: Theory and Methods, 20(5–6):1951–1955, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Andrew:1991:NEA

[And91b] Soms P. Andrew. A note on equality in Anderson’s theorem. Communications in Statistics: Theory and Methods, 20(1):141– 145, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Andel:1997:ESN

[And97] Jiří Andél. On extrapolation in some non-linear AR(1) processes. Communications in Statistics: Theory and Methods, 26(3):581– 587, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Angus:1991:SAA

[Ang91] John E. Angus. A simple asymptotic analysis of Fisher’s zstatistic. Communications in Statistics: Theory and Methods, 20(8):2697–2700, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Angus:1992:ATB

[Ang92] John E. Angus. Asymptotic theory for bootstrapping the extremes. Communications in Statistics: Theory and Methods, 22 (1):15–30, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1990:EBa

[Ano90a] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 19(2):ebi, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1990:EBb

[Ano90b] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 19(3):ebi, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1990:EBc

[Ano90c] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 19(4):ebi, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1990:EBd

[Ano90d] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 19(5):ebi, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1990:EBe

[Ano90e] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 19(6):ebi, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1990:EBf

[Ano90f] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 19(7):ebi, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1990:EBg

[Ano90g] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 19(8):ebi, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1990:EBh

[Ano90h] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 19(9):ebi, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1990:EBi

[Ano90i] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 19(10):ebi, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1990:EBj

[Ano90j] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 19(11):ebi, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1990:EBk

[Ano90k] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 19(12):ebi, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1990:PQP

[Ano90l] Anonymous. Panel on quality and productivity. Communications in Statistics: Theory and Methods, 19(11):4435–4454, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1991:EBa

[Ano91a] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 20(5–6):ebi, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1991:EBb

[Ano91b] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 20(1):ebi, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1991:EBc

[Ano91c] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 20(2):ebi, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1991:EBd

[Ano91d] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 20(3):ebi, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1991:EBe

[Ano91e] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 20(4):ebi, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1991:EBf

[Ano91f] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 20(7):ebi, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1991:EBg

[Ano91g] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 20(8):ebi, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1991:EBh

[Ano91h] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 20(9):ebi, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1991:EBi

[Ano91i] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 20(10):ebi, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1991:EBj

[Ano91j] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 20(11):ebi, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1991:EBk

[Ano91k] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 20(12):ebi, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1991:E

[Ano91l] Anonymous. Erratum. Communications in Statistics: Theory and Methods, 20(12):4083, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1991:P

[Ano91m] Anonymous. Preface. Communications in Statistics: Theory and Methods, 20(11):3421–3422, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1991:R

[Ano91n] Anonymous. Rejoinder by. Communications in Statistics: Theory and Methods, 20(11):3693–3695, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1992:C

[Ano92a] Anonymous. Corrigendum. Communications in Statistics: Theory and Methods, 21(2):257, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1992:E

[Ano92b] Anonymous. Editorial. Communications in Statistics: Theory and Methods, 22(1):v, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1992:EBa

[Ano92c] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 21(2):ebi, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1992:EBb

[Ano92d] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 21(3):ebi, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1992:EBc

[Ano92e] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 21(4):ebi, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1992:EBd

[Ano92f] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 21(5):ebi, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1992:EBe

[Ano92g] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 21(6):ebi, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1992:EBf

[Ano92h] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 21(7):ebi, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1992:EBg

[Ano92i] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 21(8):ebi, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1992:EBh

[Ano92j] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 21(9):ebi, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1992:EBi

[Ano92k] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 21(10):ebi, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1992:EBj

[Ano92l] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 21(11):ebi, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1992:EBk

[Ano92m] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 21(12):ebi, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1992:EBl

[Ano92n] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 22(1):ebi, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1992:LE

[Ano92o] Anonymous. Letter to the editor. Communications in Statistics: Theory and Methods, 21(12):3627–3628, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1993:EBa

[Ano93a] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 22(2):ebi, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1993:EBb

[Ano93b] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 22(3):ebi, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1993:EBc

[Ano93c] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 22(4):ebi, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1993:EBd

[Ano93d] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 22(5):ebi, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1993:EBe

[Ano93e] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 22(6):ebi, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1993:EBf

[Ano93f] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 22(7):ebi, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1993:EBg

[Ano93g] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 22(8):ebi, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1993:EBh

[Ano93h] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 22(9):ebi, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1993:EBi

[Ano93i] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 22(10):ebi, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1993:EBj

[Ano93j] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 22(11):ebi, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1993:EBk

[Ano93k] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 22(12):ebi, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1994:EBa

[Ano94a] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 23(1):ebi, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1994:EBb

[Ano94b] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 23(2):ebi, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1994:EBc

[Ano94c] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 23(3):ebi, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1994:EBd

[Ano94d] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 23(4):ebi, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1994:EBe

[Ano94e] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 23(5):ebi, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1994:EBf

[Ano94f] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 23(6):ebi, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1994:EBg

[Ano94g] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 23(7):ebi, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1994:EBh

[Ano94h] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 23(8):ebi, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1994:EBi

[Ano94i] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 23(9):ebi, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1994:EBj

[Ano94j] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 23(10):ebi, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1994:EBk

[Ano94k] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 23(11):ebi, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1994:EBl

[Ano94l] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 23(12):ebi, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1995:EBa

[Ano95a] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 24(1):ebi, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1995:EBb

[Ano95b] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 24(2):ebi, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1995:EBc

[Ano95c] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 24(3):ebi, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1995:EBd

[Ano95d] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 24(4):ebi, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1995:EBe

[Ano95e] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 24(5):ebi, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1995:EBf

[Ano95f] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 24(6):ebi, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1995:EBg

[Ano95g] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 24(7):ebi, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1995:EBh

[Ano95h] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 24(8):ebi, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1995:EBi

[Ano95i] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 24(9):ebi, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1995:EBj

[Ano95j] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 24(10):ebi, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1995:EBk

[Ano95k] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 24(11):ebi, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1995:EBl

[Ano95l] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 24(12):ebi, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1995:LE

[Ano95m] Anonymous. Letter to the editor. Communications in Statistics: Theory and Methods, 24(11):2959–2962, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1996:EBa

[Ano96a] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 25(1):ebi, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1996:EBb

[Ano96b] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 25(2):ebi, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1996:EBc

[Ano96c] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 25(3):ebi, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1996:EBd

[Ano96d] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 25(4):ebi, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1996:EBe

[Ano96e] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 25(5):ebi, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1996:EBf

[Ano96f] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 25(6):ebi, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1996:EBg

[Ano96g] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 25(9):ebi, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1996:EBh

[Ano96h] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 25(10):ebi, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1996:EBi

[Ano96i] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 25(11):ebi, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1996:EBj

[Ano96j] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 25(12):ebi, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1996:E

[Ano96k] Anonymous. Erratum. Communications in Statistics: Theory and Methods, 25(5):1149, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1996:LRV

[Ano96l] Anonymous. Listing of referees for volume 25, issues 1–12. Communications in Statistics: Theory and Methods, 25(12):3247– 3250, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1997:EBa

[Ano97a] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 26(1):ebi, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1997:EBb

[Ano97b] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 26(2):ebi, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1997:EBc

[Ano97c] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 26(3):ebi, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1997:EBd

[Ano97d] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 26(4):ebi, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1997:EBe

[Ano97e] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 26(6):ebi, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1997:EBf

[Ano97f] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 26(7):ebi, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1997:EBg

[Ano97g] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 26(8):ebi, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1997:EBh

[Ano97h] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 26(9):ebi, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1997:EBi

[Ano97i] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 26(10):ebi, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1997:EBj

[Ano97j] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 26(11):ebi, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1997:EBk

[Ano97k] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 26(12):ebi, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1997:LE

[Ano97l] Anonymous. Letter to the editor. Communications in Statistics: Theory and Methods, 26(8):2073–2074, 1997. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1998:E

[Ano98a] Anonymous. Editorial. Communications in Statistics: Theory and Methods, 27(1):v–??, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1998:EBa

[Ano98b] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 27(1):ebi, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1998:EBb

[Ano98c] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 27(2):ebi, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1998:EBc

[Ano98d] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 27(3):ebi, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1998:EBd

[Ano98e] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 27(4):ebi, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1998:EBe

[Ano98f] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 27(5):ebi, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1998:EBf

[Ano98g] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 27(7):ebi, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1998:EBg

[Ano98h] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 27(8):ebi, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1998:EBh

[Ano98i] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 27(9):ebi, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1998:EBi

[Ano98j] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 27(10):ebi, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1998:EBj

[Ano98k] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 27(11):ebi, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1998:EBk

[Ano98l] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 27(12):ebi, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1998:LE

[Ano98m] Anonymous. Letter to the editor. Communications in Statistics: Theory and Methods, 27(3):775, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1998:NNR

[Ano98n] Anonymous. On the number of near-records after the maximum observation in a continuous sample. Communications in Statistics: Theory and Methods, 27(3):673–686, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1999:E

[Ano99a] Anonymous. Editorial. Communications in Statistics: Theory and Methods, 28(1):v, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1999:EBa

[Ano99b] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 28(3–4):ebi, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1999:EBb

[Ano99c] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 28(1):ebi, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1999:EBc

[Ano99d] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 28(2):ebi, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1999:EBd

[Ano99e] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 28(5):ebi, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1999:EBe

[Ano99f] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 28(6):ebi, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1999:EBf

[Ano99g] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 28(7):ebi, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1999:EBg

[Ano99h] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 28(8):ebi, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1999:EBh

[Ano99i] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 28(9):ebi, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1999:EBi

[Ano99j] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 28(10):ebi, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1999:EBj

[Ano99k] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 28(11):ebi, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1999:EBk

[Ano99l] Anonymous. Editorial Board. Communications in Statistics: Theory and Methods, 28(12):ebi, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anonymous:1999:MDE

[Ano99m] Anonymous. On minimum distance estimation in two-sample scale problem with right censoring. Communications in Statistics: Theory and Methods, 28(6):1461–1477, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anraku:1994:EOR

[Anr94] Kazuo Anraku. Estimation of odds ratios under order restrictions. Communications in Statistics: Theory and Methods, 23 (11):3257–3272, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Al-Osh:1992:FOA

[AOA92] Mohamed A. Al-Osh and Emad-Eldin A. A. Aly. First order autoregressive time series with negative binomial and geometric marginals. Communications in Statistics: Theory and Methods, 21(9):2483–2492, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Aoshima:1994:HSS

[Aos94] Makoto Aoshima. Healy’s sample size of two-stage procedure in heteroscedastic simultaneous inference. Communications in Statistics: Theory and Methods, 23(5):1297–1310, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Antzoulakos:1991:NMN

[AP91] Demetris L. Antzoulakos and Andreas L. Philippou. A note on the multivariate negative binomial distributions of order K. Communications in Statistics: Theory and Methods, 20(4):1389– 1399, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Altman:1993:SRA

[AP93] N. S. Altman and C. P. Paulson. Some remarks about the Gasser–Sroka–Jennen–Steinmetz variance estimator. Communications in Statistics: Theory and Methods, 22(4):1045–1051, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Arnab:1990:USB

[AR90] Raghunath Arnab and Debesh Roy. On use of symmetrical balanced incomplete block design in construction of sampling design realising pre-assigned sets of inclusion probabilities of first two orders. Communications in Statistics: Theory and Methods, 19(9):3223–3232, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Aras:1990:SEM

[Ara90] Girish Aras. Sequential estimation of the mean of a first-order autoregressive process. Communications in Statistics: Theory and Methods, 19(5):1639–1652, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Arnab:1991:EFP

[Arn91] Raghunath Arnab. Estimation of a finite population total in two-stage and stratified sampling under superpopulation models. Communications in Statistics: Theory and Methods, 20(8):2495– 2503, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Arnab:1992:EFP

[Arn92] Raghunath Arnab. Estimation of a finite population mean under superpopulation models. Communications in Statistics: Theory and Methods, 21(6):1717–1724, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Arnold:1993:LPI

[Arn93] Barry C. Arnold. Logistic processes involving Markovian minimization. Communications in Statistics: Theory and Methods, 22(6):1699–1707, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Arnab:1994:NVE

[Arn94] Ranhunath Arnab. Nonnegative variance estimation in randomized response surveys. Communications in Statistics: Theory and Methods, 23(6):1743–1752, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Arnab:1996:RRT

[Arn96] Raghunath Arnab. Randomized response trials: a unified approach for qualitative data. Communications in Statistics: Theory and Methods, 25(6):1173–1183, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Arnoldsson:1998:CDA

[Arn98] Göran Goran Arnoldsson. A comparison of design alternatives for generalized linear models. Communications in Statistics: Theory and Methods, 27(1):107–120, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Amin:1995:NQC

[ARS95] Raid W. Amin, Marion R. Reynolds, Jr., and Bakir Saad. Nonparametric quality control charts based on the sign statistic. Communications in Statistics: Theory and Methods, 24(6):1597– 1623, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Aly:1992:SGF

[AS92] Emad-Eldin A. A. Aly and Mohammed A. Shayib. On some goodness-of-fit tests for the normal, logistic and extreme-value distributions. Communications in Statistics: Theory and Methods, 21(5):1297–1308, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anderson:1994:SDI

[AS94a] Jana C. Anderson and M. M. Siddiqui. The sampling distribution of the index of dispersion. Communications in Statistics: Theory and Methods, 23(3):897–911, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Arumainayagam:1994:CSQ

[AS94b] S. Devaraj Arumainayagam and V. Soundararajan. Construction and selection of quick switching double sampling systemacceptance number tightening. Communications in Statistics: Theory and Methods, 23(7):2079–2100, 1994. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

as:1997:IVE

[as97] Jeffrey S. Buz as. Instrumental variable estimation in nonlinear measurement error models. Communications in Statistics: Theory and Methods, 26(12):2861–2877, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shiraishi:1999:SRS

[aS99] Taka aki Shiraishi. Studentized robust statistics for, main effects in a two-factor MANOVA. Communications in Statistics: Theory and Methods, 28(3–4):809–823, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ascher:1990:STE

[Asc90] Steven Ascher. A survey of tests for exponentiality. Communications in Statistics: Theory and Methods, 19(5):1811–1825, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Atilgan:1990:DAA

[Ati90] Taskin Atilgan. On derivation and application of AIC as a databased criterion for histograms. Communications in Statistics: Theory and Methods, 19(3):885–903, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Atilgan:1996:SDB

[Ati96] Taskin Atilgan. Selection of dimension and basis for density estimation and selection of dimension, basis and error distribution for regression. Communications in Statistics: Theory and Methods, 25(1):1–28, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Akahira:1999:HOL

[ATT99] Masafumi Akahira, Kunihiko Takahashi, and Kei Takeuchi. The higher order large-deviation approximation for the distribution of the sum of independent discrete random variables. Communications in Statistics: Theory and Methods, 28(3–4):705–726, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Arellano-Valle:1996:ESM

[AVB96] R. B. Arellano-Valle and H. Bolfarine. Elliptical structural models. Communications in Statistics: Theory and Methods, 25(10): 2319–2341, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Anjaneyulu:1997:NSO

[AVN97] G. V. S. R. Anjaneyulu, D. N. Varma, and V. L. Narasimham. A note on second order slope rotatable designs over all directions. Communications in Statistics: Theory and Methods, 26(6):1477– 1479, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Aras:1992:SSE

[AW92] Girish Aras and Lyn R. Whitaker. A sequential scheme to estimate the optimal age replacement policy. Communications in Statistics: Theory and Methods, 21(8):2325–2338, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Amin:1999:SCC

[AW99] Raid W. Amin and Oskar Widmaier. Sign control charts with variable sampling intervals. Communications in Statistics: Theory and Methods, 28(8):1961–1985, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Awad:1992:FSP

[Awa92] Adnan M. Awad. On the Fisher β-sufficient partition. Communications in Statistics: Theory and Methods, 21(9):2645–2664, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ahn:1998:EDA

[AZYK98] Hongshik Ahn, Wei Zhu, Joonsook Yang, and Ralph L. Kodell. Efficient designs for animal carcinogenicity experiments. Communications in Statistics: Theory and Methods, 27(6):1275– 1287, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Balakrishnan:1994:RRS

[BA94] N. Balakrishnan and M. Ahsanullah. Recurrence relations for single and product moments of record values from generalized Pareto distribution. Communications in Statistics: Theory and Methods, 23(10):2841–2852, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bajorski:1995:SOS

[Baj95] Piotr Bajorski. On some one-sided tests in the two-sample problem. Communications in Statistics: Theory and Methods, 24 (1):75–86, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Baker:1991:ENT

[Bak91] Stuart G. Baker. Evaluating a new test using a reference test with estimated sensitivity and specificity. Communications in Statistics: Theory and Methods, 20(9):2739–2752, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Balakrishnan:1994:OSN

[Bal94] N. Balakrishnan. On order statistics from non-identical righttruncated exponential random variables and some applications. Communications in Statistics: Theory and Methods, 23(12): 3373–3393, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Balakrishna:1998:ESP

[Bal98] N. Balakrishna. Estimation for the semiPareto processes. Communications in Statistics: Theory and Methods, 27(9):2307– 2323, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bansal:1990:CSP

[Ban90] Naveen K. Bansal. Comparison of the slope parameters in errorsin-variables model. Communications in Statistics: Theory and Methods, 19(7):2705–2716, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bansal:1991:TLH

[Ban91] Naveen K. Bansal. On testing linear hypotheses on linear functional and structural relationships. Communications in Statistics: Theory and Methods, 20(9):3001–3019, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Banerjee:1998:CDL

[Ban98] Mousumi Banerjee. Cook’s distance in linear longitudinal models. Communications in Statistics: Theory and Methods, 27(12): 2973–2983, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Barr:1989:MMD

[Bar89] Aiala Barr. A multivariate model for discrete data sets. Communications in Statistics: Theory and Methods, 18(2):445–459, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). See comments [KJ90].

Barr:1991:LCM

[Bar91] Aiala Barr. On the linear combinations of the means of inverse Gaussian distributions. Communications in Statistics: Theory and Methods, 20(9):2891–2905, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Barry:1995:BAC

[Bar95] Daniel Barry. A Bayesian analysis for a class of penalised likelihood estimates. Communications in Statistics: Theory and Methods, 24(4):1057–1071, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Barnhart:1998:PMM

[Bar98] Huiman X. Barnhart. A probit model for multivariate random length ordinal data. Communications in Statistics: Theory and Methods, 27(7):1693–1713, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Basak:1993:NJT

[Bas93] Indrani Basak. A note on James type robust estimators. Communications in Statistics: Theory and Methods, 22(4):1069– 1084, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bandyopadhyay:1990:CTE

[BB90a] Dipankar Bandyopadhyay and Asit P. Basu. A class of tests for exponentiality against decreasing mean residual life alternatives. Communications in Statistics: Theory and Methods, 19(3):905– 920, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Basak:1990:ALA

[BB90b] Prasanta Basak and Parthasarathy Bagchi. Application of Laplace approximation to record values. Communications in Statistics: Theory and Methods, 19(5):1875–1888, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Boothby:1990:PEE

[BB90c] Louise H. Boothby and James K. Brewer. Predicting early educational program placement with discrete discriminant analysis. Communications in Statistics: Theory and Methods, 19(11): 4049–4060, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bravo:1990:IEL

[BB90d] Gina Bravo and MacGibbon Brenda. Improved estimators of a location vector with unknown scale parameter. Communications in Statistics: Theory and Methods, 19(10):3657–3670, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bhandary:1993:DOF

[BB93a] Madhusudan Bhandary and Naveen K. Bansal. Detection of outliers in familial data. Communications in Statistics: Theory and Methods, 22(9):177–185, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bofinger:1993:ENM

[BB93b] Eve Bofinger and Mark Bofinger. Equivalence of normal means compared with a control. Communications in Statistics: Theory and Methods, 22(11):3117–3141, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Balakrishnan:1995:OSN

[BB95] N. Balakrishnan and K. Balasubramanian. Order statistics from non-identical power function random variables. Communications in Statistics: Theory and Methods, 24(6):1443–1454, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bofinger:1996:ESL

[BB96] Eve Bofinger and David Bofinger. Equivalence to a straight line. Communications in Statistics: Theory and Methods, 25(2): 435–455, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Baragona:1997:TTS

[BB97] Roberto Baragona and Francesco Battaglia. Testing time series for interpolability and whiteness. Communications in Statistics: Theory and Methods, 26(11):2623–2644, 1997. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Barroso:1998:BLU

[BBK98] Lúcia P. Barroso, Wilton O. Bussab, and Martin Knott. Best linear unbiased predictor in the mixed model with incomplete data. Communications in Statistics: Theory and Methods, 27 (1):121–129, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bendel:1990:BME

[BC90] Robert B. Bendel and Bradley P. Carlin. Bayes methods in the ecological fallacy context: estimation of individual correlation from aggregate data. Communications in Statistics: Theory and Methods, 19(7):2595–2623, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Broemeling:1992:BAT

[BC92] Lyle D. Broemeling and Peyton Cook. Bayesian analysis of threshold autoregressions. Communications in Statistics: Theory and Methods, 21(9):2459–2482, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Botter:1997:BCG

[BC97] Denise A. Botter and Gauss M. Cordeiro. Bartlett corrections for generalized linear models with dispersion covariates. Communications in Statistics: Theory and Methods, 26(2):279–307, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Balakrishnan:1993:RRM

[BCA93] N. Balakrishnan, P. S. Chan, and M. Ahsanullah. Recurrence relations for moments of record values from generalized extreme value distribution. Communications in Statistics: Theory and Methods, 22(5):1471–1482, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Barranco-Chamorro:1999:EUD

[BCLBMR99] T. Barranco-Chamorro, F. Lopez-Blazquez, and J. L. Moreno-Rebollo. Estimation in uniform distributions using orthogonal polynomials. Communications in Statistics: Theory and Methods, 28(5):1145–1167, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bristol:1991:PSS

[BD91] David R. Bristol and M. M. Desu. A partially sequential selection procedure for comparison with a control. Communications in Statistics: Theory and Methods, 20(2):703–710, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Borowiak:1992:SBF

[BD92] D. S. Borowiak and J. P. De Los Reyes. Selection of the best in 2 × 2 factorial designs. Communications in Statistics: Theory and Methods, 21(9):2493–2500, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Buse:1993:PTE

[BD93] A. Buse and N. K. Dastor. The power of two exact tests for structural change in the presence of heteroskedasticity. Communications in Statistics: Theory and Methods, 22(8):2239–2257, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bravo:1995:SPE

[BDF95] José I. Bravo, Íñigo De Fuentes, and Arturo J. Fernández. A semi-parametric estimation of a survival function from incomplete and doubly censored data. Communications in Statistics: Theory and Methods, 24(11):2735–2752, 1995. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bagchi:1990:BAA

[BDG90] Parthasarathy Bagchi, Norman Draper, and Irwin Guttman. Bayesian assessment of assumptions of regression analysis. Communications in Statistics: Theory and Methods, 19(3):921–934, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Brannas:1998:TLA

[BDT98] Kurt Brännäs, Jan G. De Gooijer, and Timo Teräsvirta. Testing linearity against nonlinear moving average models. Communications in Statistics: Theory and Methods, 27(8):2025–2035, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bravo:1993:NES

[BE93] José I. Bravo and M. Dolores Esteban. Nonparametric estimation of survival functions for doubly censored observations when the lifetime hazard rate is proportionally related to the censoring hazard rates. Communications in Statistics: Theory and Methods, 22(11):3237–3253, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Beh:1999:CAR

[Beh99] Eric J. Beh. Correspondence analysis of ranked data. Communications in Statistics: Theory and Methods, 28(7):1511–1133, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bello:1993:CAI

[Bel93] Abdul Lateef Bello. Choosing among imputation techniques for incomplete multivariate data: a simulation study. Communications in Statistics: Theory and Methods, 22(3):853–877, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Benasseni:1990:SCS

[Bén90] Jacques Bénasséni. Sensitivity coefficients for the subspaces spanned by principal components. Communications in Statistics: Theory and Methods, 19(6):2021–2034, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Berke:1998:SPL

[Ber98a] Olaf Berke. On spatiotemporal prediction for on-line monitoring data. Communications in Statistics: Theory and Methods, 27(9): 2343–2369, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Berred:1998:PRV

[Ber98b] Alexandre M. Berred. Prediction of record values. Communications in Statistics: Theory and Methods, 27(9):2221–2240, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Berchtold:1999:DCM

[Ber99] Andre Berchtold. The double chain Markov model. Communications in Statistics: Theory and Methods, 28(11):2569–2589, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Beveridge:1992:LSE

[Bev92] Steve Beveridge. Least squares estimation of missing values in time series. Communications in Statistics: Theory and Methods, 21(12):3479–3496, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Benjamini:1990:CVU

[BF90] Yoav Benjamini and Camil Fuchs. Conditional versus unconditional analysis in some regression models. Communications in Statistics: Theory and Methods, 19(12):4731–4756, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bertuzzi:1991:RRV

[BG91] Alessandro Bertuzzi and Alberto Gandolfi. Ridge regression versus OLS by Pitman’s closeness under quadratic and Fisher’s loss. Communications in Statistics: Theory and Methods, 20(11): 3581–3590, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bismarck:1995:UBP

[BG95] M. Bismarck and L. Gajek. An upper bound for the power of a test. Communications in Statistics: Theory and Methods, 24 (3):813–828, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Broemeiing:1996:BAS

[BG96a] Lyle D. Broemeiing and Mary Ann Gregurich. On a Bayesian approach for the shiftpoint problem. Communications in Statistics: Theory and Methods, 25(10):2267–2279, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Broemeling:1996:BAA

[BG96b] Lyle D. Broemeling and Mary Ann Gregurich. A Bayesian alternative to the analysis of matched categorical responses. Communications in Statistics: Theory and Methods, 25(7):1429–1445, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Basawa:1990:ABV

[BGMT90] I. V. Basawa, T. A. Green, W. P. McCormick, and R. L. Taylor. Asymptotic bootstrap validity for finite Markov chains. Communications in Statistics: Theory and Methods, 19(4):1493–1510, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Balakrishnan:1995:ELS

[BGP95] N. Balakrishnan, Shanti S. Gupta, and S. Panchapakesan. Estimation of the location and scale parameters of the extreme value distribution based on multiply Type-II censored samples. Communications in Statistics: Theory and Methods, 24(8):2105– 2125, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Baringhaus:1992:AOT

[BH92a] L. Baringhaus and N. Henze. An adaptive omnibus test for exponentiality. Communications in Statistics: Theory and Methods, 21(4):969–978, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Baringhaus:1992:CNC

[BH92b] L. Baringhaus and N. Henze. A characterization of and new consistent tests for symmetry. Communications in Statistics: Theory and Methods, 21(6):1555–1566, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Burch:1999:BEI

[BH99] Brent D. Burch and Ian R. Harris. Bayesian estimators of the intraclass correlation coefficient in the one-way random effects model. Communications in Statistics: Theory and Methods, 28(6):1247–1272, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bhandary:1995:DOG

[Bha95a] Madhusudan Bhandary. Detection of outliers in growth curve models. Communications in Statistics: Theory and Methods, 24(8):1923–1940, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bhandary:1995:DOL

[Bha95b] Madhusudan Bhandary. Detection of outliers in linear model when the regression parameters are stochastic. Communications in Statistics: Theory and Methods, 24(1):153–173, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bhapkar:1995:COM

[Bha95c] Vasant P. Bhapkar. Completeness and optimality of marginal likelihood estimating equations. Communications in Statistics: Theory and Methods, 24(4):945–952, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bhattacharya:1997:TMP

[Bha97] Bhaskar Bhattacharya. Testing multinomial parameters under order restrictions. Communications in Statistics: Theory and Methods, 26(8):1839–1865, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bhoj:1997:EPE

[Bho97] Dinesh S. Bhoj. Estimation of parameters of the extreme value distribution using ranked set sampling. Communications in Statistics: Theory and Methods, 26(3):653–667, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bilodeau:1997:EMT

[Bil97] Martin Bilodeau. Estimating a multivariate treatment effect under a biased allocation rule. Communications in Statistics: Theory and Methods, 26(5):1119–1124, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Brannas:1994:TSC

[BJ94] Kurt Brännäs and Per Johansson. Time series count data regression. Communications in Statistics: Theory and Methods, 23(10):2907–2925, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bhaumik:1991:OST

[BK91] Dulal K. Bhaumik and Pandu M. Kulkarni. One-sided tolerance limits for unbalanced one-way ANOVA random effects model. Communications in Statistics: Theory and Methods, 20(5–6): 1665–1675, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Barnett:1996:SSR

[BK96] Vic Barnett and Mike G. Kenward. Security systems and renewal processes. Communications in Statistics: Theory and Methods, 25(3):475–487, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bartos:1997:NIT

[BK97a] Jerzy Bartos and Marek Kaluszka. Note on an improvement in two stage sampling scheme. Communications in Statistics: Theory and Methods, 26(9):2129–2139, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Braun:1997:PFB

[BK97b] W. J. Braun and R. J. Kulperger. Properties of a Fourier bootstrap method for time series. Communications in Statistics: Theory and Methods, 26(6):1329–1336, 1997. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Barnett:1998:TPR

[BK98] Vic Barnett and Michael G. Kenward. Testing a Poisson renewal process in the context of security alarm maintenance policies. Communications in Statistics: Theory and Methods, 27(12): 3085–3094, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Burgess:1999:SOM

[BKS99] Leonie Burgess, Donald L. Kreher, and Deborah J. Street. Small orthogonal main effect plans with four factors. Communications in Statistics: Theory and Methods, 28(10):2441–2464, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Brits:1990:AFT

[BL90] S. J. M. Brits and H. H. Lemmer. An adjusted Friedman test for the nested design. Communications in Statistics: Theory and Methods, 19(5):1837–1855, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Billor:1993:LIN

[BL93] N. Billor and R. M. Loynes. Local influence: a new approach. Communications in Statistics: Theory and Methods, 22(6):1595– 1611, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lian:1995:AFV

[bL95] Ie bin Lian. Asymptotics of forced-in variables from screening processes. Communications in Statistics: Theory and Methods, 24(1):131–151, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Balkin:1998:GCS

[BL98a] Sandy D. Balkin and Dennis K. J. Lin. A graphical comparison of supersaturated designs. Communications in Statistics: Theory and Methods, 27(6):1289–1303, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bebbington:1998:GNB

[BL98b] M. S. Bebbington and C. D. Lai. A generalized negative binomial and applications. Communications in Statistics: Theory and Methods, 27(10):2515–2533, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bar-Lev:1999:MSE

[BLBE99] Shaul K. Bar-Lev, Benzion Boukai, and Peter Enis. On the mean squared error, the mean absolute error and the like. Communications in Statistics: Theory and Methods, 28(8):1813–1822, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bar-Lev:1999:RMN

[BLF99] Shaul K. Bar-Lev and Camil Fuchs. Regular MLEs for nonregular distributions. Communications in Statistics: Theory and Methods, 28(9):2037–2044, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Blodgett:1997:AFE

[Blo97] Robert J. Blodgett. Adding functions to ensure regression converges. Communications in Statistics: Theory and Methods, 26(5):1203–1214, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bolfarine:1997:LCF

[BLS97] Heleno Bolfarine, Claudia R. O. P. Lima, and Monica C. Sandoval. Linear calibration in functional regression models. Communications in Statistics: Theory and Methods, 26(10):2307– 2328, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Blyth:1991:C

[Bly91] R. Colin Blyth. Comments. Communications in Statistics: Theory and Methods, 20(11):3445–3452, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bohrer:1992:PTE

[BLY92] Robert Bohrer, Kunjung Lai, and Thomas A. Yancey. Preliminary test estimation for the second order autoregression. Communications in Statistics: Theory and Methods, 21(10):2797– 2807, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Baczkowski:1990:TSS

[BM90] A. J. Baczkowski and K. V. Mardia. A test of spatial symmetry with general application. Communications in Statistics: Theory and Methods, 19(2):555–572, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bofinger:1993:UBR

[BM93] Eve Bofinger and Kerrie Mengersen. An upper bound on a ratio of variances. Communications in Statistics: Theory and Methods, 22(7):1907–1922, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Boneh:1994:VSR

[BM94] Shahar Boneh and Gonzalo R. Mendieta. Variable selection in regression models using principal components. Communications in Statistics: Theory and Methods, 23(1):197–213, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bader:1998:CVR

[BM98] Julia O. Bader and James H. O. Matis. On the coefficient of variation for residence time distributions of some stochastic compartmental models. Communications in Statistics: Theory and Methods, 27(7):1757–1780, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Banerjee:1993:MCG

[BMK93] Shakti Banerjee, Ranjana Malhotra, and Sanpei Kageyama. Methods of constructing group divisible designs with k = 3 and 4. Communications in Statistics: Theory and Methods, 22(10):

2831–2842, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Basawa:1991:BTS

[BMM^{+}91] I. V. Basawa, A. K. Mallik, W. P. McCornick, J. H. Reeves, and R. L. Taylor. Bootstrap test of significance and sequential bootstrap estimation for unstable first order autoregressive processes. Communications in Statistics: Theory and Methods, 20(3):1015–1026, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Babu:1991:CSO

[BN91] B. Re. Victor Babu and V. L. Narasimhant. Construction of second order slope rotatable designs through balanced incomplete block designs. Communications in Statistics: Theory and Methods, 20(8):2467–2478, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bofinger:1990:PPC

[Bof90] Eve Bofinger. Posterior probability of correct selection. Communications in Statistics: Theory and Methods, 19(2):599–616, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bofinger:1992:MCB

[Bof92] Eve Bofinger. Multiple comparisons with ‘best’ for multivariate normal populations. Communications in Statistics: Theory and Methods, 21(4):915–941, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bohn:1996:RNR

[Boh96] Lora L. Bohn. A review of nonparametric ranked-set sampling methodology. Communications in Statistics: Theory and Methods, 25(11):2675–2685, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Boik:1991:SMM

[Boi91] Robert J. Boik. Scheffés mixed model for multivariate repeated measures: a relative efficiency evaluation. Communications in Statistics: Theory and Methods, 20(4):1233–1255, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bolstad:1990:EDL

[Bol90] William M. Bolstad. The efficiency of dynamic linear model estimators applied to a linearly aggregated time series. Communications in Statistics: Theory and Methods, 19(1):83–90, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bhattacharyya:1992:SCN

[BOR92] B. B. Bhattacharyya, Y. Otsuka, and G. D. Richardson. Strong consistency in nonlinear regression with multiplicative error. Communications in Statistics: Theory and Methods, 21(10): 2825–2831, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Borkowski:1995:FMC

[Bor95a] John J. Borkowski. Finding maximum G-criterion values for central composite designs on the hypercube. Communications in Statistics: Theory and Methods, 24(8):2041–2058, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Borkowski:1995:MMA

[Bor95b] John J. Borkowski. Minimum, maximum, and average spherical prediction variances for central composite and Box–Behnken designs. Communications in Statistics: Theory and Methods, 24(10):2581–2600, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bose:1991:SPP

[Bos91] Sudip Bose. Some properties of posterior Pitman closeness. Communications in Statistics: Theory and Methods, 20(11): 3697–3712, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bosch:1994:NAQ

[Bos94] Ronald J. Bosch. A note on asymptotics for quantile smoothing splines. Communications in Statistics: Theory and Methods, 23(11):3075–3083, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Boukai:1990:EED

[Bou90a] Benzion Boukai. An explicit expression for the distribution of the supremum of Brownian motion with a change point. Communications in Statistics: Theory and Methods, 19(1):31–40, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bourke:1990:EDE

[Bou90b] Patrick D. Bourke. Estimating a distribution for each category of a sensitive variate. Communications in Statistics: Theory and Methods, 19(9):3233–3241, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Boyd:1995:MCE

[Boy95] Amy Hancock Boyd. A minimum cost exponential cusum test. Communications in Statistics: Theory and Methods, 24(11): 2929–2958, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bozdogan:1990:IBM

[Boz90] Hamparsum Bozdogan. On the information-based measure of covariance complexity and its application to the evaluation of multivariate linear models. Communications in Statistics: Theory and Methods, 19(1):221–278, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Brorsen:1993:LRS

[BP93] B. Wade Brorsen and Paul V. Preckei. Linear regression with stably distributed residuals. Communications in Statistics: Theory and Methods, 22(3):659–667, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Balkema:1996:DME

[BP96] A. A. Balkema and E. I. Pancheva. Decomposition for multivariate extremal processes. Communications in Statistics: Theory and Methods, 25(4):737–758, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bhatt:1999:ADO

[BP99a] M. B. Bhatt and S. R. Patel. Asymptotic distribution for the order statistics of a truncated one-parameter family of distributions. Communications in Statistics: Theory and Methods, 28(8):1823–1855, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bowtell:1999:CFR

[BP99b] Phil Bowtell and Mike Patefield. Circular functional relationships. Communications in Statistics: Theory and Methods, 28(9):2103–2118, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Baksalary:1990:NGR

[BPT90] Jerzy K. Baksalary, Pawel R. Pordzik, and Götz Trenkler. A note on generalized ridge estimators. Communications in Statistics: Theory and Methods, 19(8):2871–2877, 1990. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bhattacharya:1990:TSS

[BR90a] C. G. Bhattacharya and B. L. S. Prakasa Rao. On two-stage shrinkage testimation. Communications in Statistics: Theory and Methods, 19(10):3527–3538, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bolfarine:1990:FPP

[BR90b] Heleno Bolfarine and Josemar Rodrigues. Finite population prediction under a linear function superpopulation model: a Bayesian perspective. Communications in Statistics: Theory and Methods, 19(7):2577–2594, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Boender:1991:MMC

[BR91] C. G. E. Boender and H. E. Romeijn. The multidimensional Markov chain with prespecified asymptotic means, and (auto- )covariances. Communications in Statistics: Theory and Methods, 20(1):345–359, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Betro:1992:CMA

[BR92] Bruno Betrò and Fabrizio Ruggeri. Conditional Γ-minimax actions under convex losses. Communications in Statistics: Theory and Methods, 21(4):1051–1066, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Barnett:1993:POT

[BR93a] Vic Barnett and Donna Roberts. The problem of outlier tests in sample surveys. Communications in Statistics: Theory and Methods, 22(10):2703–2721, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Basu:1993:ATE

[BR93b] A. K. Basu and S. Sen Roy. Asymptotic theory of estimation of parameters in autoregressive models under general set-up of the roots. Communications in Statistics: Theory and Methods, 22(8):2223–2237, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bekker:1995:BMN

[BR95] A. Bekker and J. J. J. Roux. Bayesian multivariate normal analysis with a Wishart prior. Communications in Statistics: Theory and Methods, 24(10):2485–2497, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Best:1997:GFO

[BR97] D. J. Best and J. C. W. Rayner. Goodness of fit for the ordered categories discrete uniform distribution. Communications in Statistics: Theory and Methods, 26(4):899–909, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bradley:1990:FSU

[Bra90] Ralph A. Bradley. The Florida State University Department of Statistics: The first fifteen years. Communications in Statistics: Theory and Methods, 19(11):3935–3960, 1990. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bremer:1990:NSS

[Bre90] Ronald Bremer. Numerical study of the small sample variances of the MINQUE estimators of variance components in the twoway factorial design. Communications in Statistics: Theory and Methods, 19(4):1261–1279, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bremer:1993:CMY

[Bre93] Ronald H. Bremer. Choosing and modeling your mixed linear model. Communications in Statistics: Theory and Methods, 22(12):3491–3521, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bristol:1993:PSS

[Bri93] David R. Bristol. Probabilities and sample sizes for the two onesided tests procedure. Communications in Statistics: Theory and Methods, 22(7):1953–1961, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Brien:1994:MGN

[Bri94] Peter C. O’ Brien. A multivariate generalization of von Neumann’s ratio. Communications in Statistics: Theory and Methods, 23(1):239–247, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Britton:1997:LTT

[Bri97] Tom Britton. Limit theorems and tests for within family clustering in epidemic models. Communications in Statistics: Theory and Methods, 26(4):953–976, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Brown:1990:WSN

[Bro90] Mark Brown. Weighted sup-norm inequalities and their applications. Communications in Statistics: Theory and Methods, 19(11):4061–4081, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Brouaye:1995:CBT

[Bro95] F. Brouaye. Choosing between two unknown intervals: an empirical Bayes testing approach to a classification problem. Communications in Statistics: Theory and Methods, 24(12):3163–3181, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Brown:1997:LIR

[Bro97] Gary C. Brown. Local influence regression diagnostics for gamma data. Communications in Statistics: Theory and Methods, 26(12):3051–3067, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Brostrom:1998:MFO

[Bro98] Göran Brostrom. A modification of Fisher’s omnibus test. Communications in Statistics: Theory and Methods, 27(11):2663– 2674, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Brunner:1995:UGS

[Bru95] Lawrence J. Brunner. Using the Gibbs sampler to simulate from the Bayes estimate of a decreasing density. Communications in Statistics: Theory and Methods, 24(1):215–226, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bansal:1990:AEM

[BS90a] M. L. Bansal and Ravindra Singh. An alternative estimator for multiple characteristics in rhg sampling scheme. Communications in Statistics: Theory and Methods, 19(5):1777–1784, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bookstein:1990:SMG

[BS90b] Fred L. Bookstein and Paul D. Sampson. Statistical models for geometric components of shape change. Communications in Statistics: Theory and Methods, 19(5):1939–1972, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bowman:1990:ADF

[BS90c] K. O. Bowman and L. R. Shenton. The approximate distribution of four moment statistics from type III distributions. Communications in Statistics: Theory and Methods, 19(4):1511–1579, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Badahdah:1991:ARE

[BS91] Saleh O. Badahdah and M. M. Siddiqui. Adaptive robust estimation of location for stable distributions. Communications in Statistics: Theory and Methods, 20(2):631–652, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bhattacharya:1992:CEL

[BS92] C. G. Bhattacharya and K. R. Shah. Combined estimate of a location parameter. Communications in Statistics: Theory and Methods, 21(7):1929–1934, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Barnhart:1994:OMM

[BS94] Huiman X. Barnhart and Allan R. Sampson. Overview of multinomial models for ordinal data. Communications in Statistics:

Theory and Methods, 23(12):3395–3416, 1994. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Boyle:1996:SCI

[BS96] Madeline J. Boyle and John D. Spurrier. Simultaneous confidence intervals between Tukey and Scheffe. Communications in Statistics: Theory and Methods, 25(7):1583–1593, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Blyth:1998:BMA

[BS98a] Stephen Blyth and Adrian F. M. Smith. Bayesian meta-analysis of frequentist p-values. Communications in Statistics: Theory and Methods, 27(11):2707–2724, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bowman:1998:ASD

[BS98b] K. O. Bowman and L. R. Shenton. Asymptotic skewness and the distribution of maximum likelihood estimators. Communications in Statistics: Theory and Methods, 27(11):2743–2760, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bowman:1999:AKM

[BS99a] K. O. Bowman and L. R. Shenton. The asymptotic kurtosis for maximum likelihood estimators. Communications in Statistics: Theory and Methods, 28(11):2641–2654, 1999. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bowman:1999:AMP

[BS99b] K. O. Bowman and L. R. Shenton. The asymptotic moment profile and maximum likelihood: Applications to gamma and gamma ratio densities. Communications in Statistics: Theory and Methods, 28(10):2497–2508, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bowman:1995:EPA

[BSK95] K. O. Bowman, L. R. Shenton, and Chris Karlof. Estimation problems associated with the three parameter gamma distribution. Communications in Statistics: Theory and Methods, 24(5): 1355–1376, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bansal:1994:MCS

[BSS94] M. L. Bansal, Sarjinder Singh, and Ravindra Singh. Multicharacter survey using randomized response technique. Communications in Statistics: Theory and Methods, 23(6):1705–1715, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bookstein:1990:MDR

[BSSB90] Fred L. Bookstein, Paul D. Sampson, Ann P. Streissguth, and Helen M. Barr. Measuring “dose” and “response” with multivariate data using partial least squares techniques. Communications in Statistics: Theory and Methods, 19(3):765–804, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Buonaccorsi:1993:CNM

[BT93] John P. Buonaccorsi and Tor D. Tosteson. Correcting for nonlinear measurement errors in the dependent variable in the general linear model. Communications in Statistics: Theory and Methods, 22(10):2687–2702, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Buonaccorsi:1990:DSE

[Buo90] John P. Buonaccorsi. Double sampling for exact values in the normal discriminant model with application to binary regression. Communications in Statistics: Theory and Methods, 19 (12):4569–4586, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Burnham:1997:DRS

[Bur97] Kenneth P. Burnham. Distributional results for special cases of the Jolly–Seber model. Communications in Statistics: Theory and Methods, 26(6):1395–1409, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Butler:1998:FDC

[But98] Neil Butler. A frequency domain comparison of estimation methods for Gaussian fields. Communications in Statistics: Theory and Methods, 27(9):2325–2342, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bonett:1993:ASM

[BW93] Douglas G. Bonett and J. Arthur Woodward. Analysis of simple main effects in fractional factorial experimental designs of resolution V. Communications in Statistics: Theory and Methods, 22(6):1585–1593, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Bai:1990:SEA

[BWCM90] Z. D. Bai, Y. H. Wu, X. R. Chen, and B. Q. Miao. On solvability of an equation Arising in the theory of m-estimates. Communications in Statistics: Theory and Methods, 19(1):363–380, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ben:1992:RAV

[BY92] Marta García Ben and Víctor J. Yohai. Robust analysis of variance for a randomized block design. Communications in Statistics: Theory and Methods, 21(6):1779–1798, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Byrne:1996:MAA

[Byr96] Philip J. Byrne. Multivariate analysis with an autoregressive covariance model. Communications in Statistics: Theory and Methods, 25(3):555–569, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chaudhuri:1990:VER

[CA90a] Arijit Chaudhuri and Arun Kumar Adhikary. Variance estimation with randomized response. Communications in Statistics: Theory and Methods, 19(3):1119–1125, 1990. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cuadras:1990:DBR

[CA90b] C. M. Cuadras and C. Arenas. A distance based regression model for prediction with mixed data. Communications in Statistics: Theory and Methods, 19(6):2261–2279, 1990. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Calvetti:1991:REF

[Cal91] Daniela Calvetti. Roundoff error for floating point representation of real data. Communications in Statistics: Theory and Methods, 20(8):2687–2695, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Calvin:1994:OST

[Cal94] James A. Calvin. One-sided test of a covariance matrix with a known null value. Communications in Statistics: Theory and Methods, 23(11):3121–3140, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Campbell:1993:IET

[Cam93] Edward P. Campbell. Influence for empirical transforms. Communications in Statistics: Theory and Methods, 22(9):730–752, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Caperaa:1992:RTS

[Cap92] Philippe Capéraà. Rank tests for scale and tail-weight of distributions. Communications in Statistics: Theory and Methods, 21 (2):539–545, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cartinhour:1990:ODM

[Car90] Jack Cartinhour. One-dimensional marginal density functions of a truncated multivariate normal density function. Communications in Statistics: Theory and Methods, 19(1):197–203, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Carriere:1993:CED

[Car93] Jacques F. Carriere. A continuous estimator of a distribution function that reproduces the empirical moments. Communications in Statistics: Theory and Methods, 22(7):1923–1931, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Carriere:1994:LST

[Car94] Jacques F. Carriere. A large sample test for one parameter families of copulas. Communications in Statistics: Theory and Methods, 23(5):1311–1317, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chaudhuri:1997:SDE

[CAS97] Arijit Chaudhuri, Arun Kumar Adhikary, and Arup Kumar Seal. Small domain estimation by empirical Bayes and Kalman filtering procedures — a case study. Communications in Statistics: Theory and Methods, 26(7):1613–1621, 1997. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chatterjee:1990:CMS

[CB90] Shoutir Kishore Chatterjee and Tathagata Banerjee. Combination of multiple scores for nonparametric testing in multivariate linear regression set-up. Communications in Statistics: Theory and Methods, 19(8):2967–2998, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cook:1994:BWA

[CB94] Peyton Cook and Lyle D. Broemeling. A Bayesian WLS approach to generalized linear models. Communications in Statistics: Theory and Methods, 23(12):3323–3347, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chan:1998:EPS

[CB98] Ivan S. F. Chan and Norman R. Bohidar. Exact power and sample size for vaccine efficacy studies. Communications in Statistics: Theory and Methods, 27(6):1305–1322, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cordeiro:1999:MML

[CBFCN99] Gauss M. Cordeiro, Denise A. Butter, Silvia L. P. Ferrari, and Francisco Cribari-Neto. Modified maximum likelihood estimation in one-parameter exponential family models. Communications in Statistics: Theory and Methods, 28(1):157–178, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cook:1990:BAM

[CBG90] Peyton Cook, Lyle D. Broemeling, and Mohammad Gharaff. A Bayesian analysis of the mixed linear model. Communications in Statistics: Theory and Methods, 19(3):987–1002, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chaudhuri:1991:BDB

[CBR91] G. Chaudhuri, J. D. Borwankar, and P. R. K. Rao. Bhattacharyya distance based linear discriminant function for stationary time series. Communications in Statistics: Theory and Methods, 20(7):2195–2205, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cyr:1990:CCE

[CBRS90] Louis Cyr, Harvey L. Bank, Philip F. Rust, and Marcia K. Schmehl. Comparative characteristics of estimators of the mean responding cell density in limiting dilution assays. Communications in Statistics: Theory and Methods, 19(7):2643–2666, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chandra:1990:ETW

[CC90a] Nimai Kumar Chandra and Arijit Chaudhuri. On the efficiency of a testimator for the Weibull shape parameter. Communications in Statistics: Theory and Methods, 19(4):1247–1259, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chang:1990:TFS

[CC90b] Chebg-Tao Chang and Joanna Cheng. Three-factor symmetrical balanced designs with full efficiency for main effects. Communications in Statistics: Theory and Methods, 19(12):4757–4772, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cox:1991:MSS

[CC91a] Trevor F. Cox and Michael A. A. Cox. Multidimensional scaling on a sphere. Communications in Statistics: Theory and Methods, 20(9):2943–2953, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cune:1991:JKQa

[CC91b] E. D. Mc Cune and Sandra Luna Mc Cune. Jackknifed kernel quantile estimators. Communications in Statistics: Theory and Methods, 20(8):2179–2725, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cune:1991:JKQb

[CC91c] E. D. Mc Cune and Sandra Luna Mc Cune. Jackknifed kernel quantile estimators. Communications in Statistics: Theory and Methods, 20(8):2719–2725, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cano:1993:RPM

[CC93] J. A. Cano and J. A. Cano. Robustness of the posterior mean in normal hierarchical models. Communications in Statistics: Theory and Methods, 22(7):1999–2014, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chai:1995:JEE

[CC95a] John J. Chai and Chen Chung. Joint effects of errors of measurement and grouping on the estimated product–moment correlation. Communications in Statistics: Theory and Methods, 24 (1):283–294, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chua:1995:AER

[CC95b] Tin Chiu Chua and Kok Leong Chiang. An alternative estimator for randomized response sampling with continuous distributions from a dichotomous population. Communications in Statistics: Theory and Methods, 24(11):2919–2928, 1995. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cabana:1996:BBT

[CC96] A. Cabana and E. M. Cabana. Bridge-to-bridge transformations and Kolmogorov–Smirnov tests. Communications in Statistics: Theory and Methods, 25(1):227–234, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chan:1996:DAM

[CCC96] H. T. Chan, R. Chinipardaz, and T. F. Cox. Discrimination of AR, MA and ARMA time series models. Communications in Statistics: Theory and Methods, 25(6):1247–1260, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chung:1998:BAE

[CD98] Younshik Chung and Dipak K. Dey. Bayesian approach to estimation of intraclass correlation using reference prior. Communications in Statistics: Theory and Methods, 27(9):2241–2255, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cerioli:1997:CTP

[Cer97] Andrea Cerioli. Comparing three partitions: an inferential approach based on multi-way contingency tables. Communications in Statistics: Theory and Methods, 26(10):2457–2471, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Consul:1992:GPR

[CF92] P. C. Consul and Felix Famoye. Generalized Poisson regression model. Communications in Statistics: Theory and Methods, 21 (1):89–109, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Consul:1995:GNB

[CF95] P. C. Consul and Felix Famoye. On the generalized negative binomial distribution. Communications in Statistics: Theory and Methods, 24(2):459–472, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Consul:1996:LKF

[CF96a] P. C. Consul and Felix Famoye. Lagrangean Katz family of distributions. Communications in Statistics: Theory and Methods, 25(2):415–434, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cordeiro:1996:BTC

[CF96b] Gauss M. Cordeiro and Silvia L. P. Ferrari. Bartlett-type corrections for some score tests in proper dispersion models. Communications in Statistics: Theory and Methods, 25(1):29–48, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cordeiro:1998:GBC

[CF98] Gauss M. Cordeiro and Silvia L. P. Ferrari. Generalized Bartlett correction. Communications in Statistics: Theory and Methods, 27(2):509–527, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cheah:1992:TOA

[CFRT92] P. K. Cheah, D. A. S. Fraser, N. Reid, and A. Tapia. Third order asymptotics: connections among test quantities. Communications in Statistics: Theory and Methods, 21(8):2127–2133, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Clemmens:1990:CLM

[CG90] Ann Clemmens and Z. Govindarajulu. A certain locally most powerful test for one-way random effects model: full distribution and power considerations. Communications in Statistics: Theory and Methods, 19(11):4135–4151, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chao:1991:THD

[CG91] C. C. Chao and A. K. Gupta. Testing of homogeneity of diagonal blocks with blockwise independence. Communications in Statistics: Theory and Methods, 20(7):1957–1969, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chakraborti:1992:NCT

[CG92a] S. Chakraborti and Jean D. Gibbons. Nonparametric comparison of treatments with a standard in the one-way layout: some design considerations. Communications in Statistics: Theory and Methods, 22(1):1–14, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chen:1992:DRM

[CG92b] James J. Chen and David W. Gaylor. Dose-response modeling of quantitative response data for risk assessment. Communications in Statistics: Theory and Methods, 21(8):2367–2381, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Clarke:1992:URE

[CG92c] B. R. Clarke and E. J. Godolphin. Uncorrelated residuals and an exact test for two variance components in experimental design. Communications in Statistics: Theory and Methods, 21(9):2501– 2526, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cretois:1999:PMT

[CGE99] Emmanuelle Crétois, Olivier Gaudoin, and Mhamed-Ali El Aroui. U-plot method for testing the goodness-of-fit of the power-law process. Communications in Statistics: Theory and Methods, 28(7):1731–1747, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Calabria:1990:BEP

[CGP90] R. Calabria, M. Guida, and G. Pulcini. Bayes estimation of prediction intervals for a power law process. Communications in Statistics: Theory and Methods, 19(8):3023–3035, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Calabria:1992:PBT

[CGP92] R. Calabria, M. Guida, and G. Pulcini. Power bounds for a test of equality of trends in k independent power law processes. Communications in Statistics: Theory and Methods, 21(11):3275– 3290, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chen:1991:SSS

[CH91a] Pinyucen Chen and Lifang Hsu. On a sequential subset selection procedure for the least probable multinomial cell. Communications in Statistics: Theory and Methods, 20(9):2845–2862, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hu:1991:LIL

[cH91b] Tien chung Hu. On the law of the iterated logarithm for arrays of random variables. Communications in Statistics: Theory and Methods, 20(7):1989–1994, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Conerly:1991:SEP

[CH91c] Michael D. Conerly and J. Michael Hardin. Shrunken estimators via a Pitman nearness criterion. Communications in Statistics: Theory and Methods, 20(11):3591–3604, 1991. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Carpenter:1994:EMM

[CH94] Mark Carpenter and Jaimie L. Hebert. Estimating the minimum and maximum location parameters for two gamma-exponential scale mixtures. Communications in Statistics: Theory and Methods, 23(8):2367–2377, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chen:1995:ADP

[CH95a] Sy-Mien Chen and Nai-Feng Hsu. The asymptotic distribution of the process capability index C_{pmk}. Communications in Statistics: Theory and Methods, 24(5):1279–1291, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chiou:1995:CIE

[CH95b] Paul Chiou and Chien-Pai Han. Conditional interval estimation of the exponential location parameter following rejection of a pre-test. Communications in Statistics: Theory and Methods, 24(6):1481–1492, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chakraborti:1996:MSI

[CH96] S. Chakraborti and T. P. Hettmansperger. Multi-sample inference for the simple-tree alternative based on one-sample confidence intervals. Communications in Statistics: Theory and Methods, 25(11):2819–2837, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Carpenter:1997:EGL

[CH97] Mark Carpenter and Jaimie L. Hebert. Estimating guaranteed lifetimes of systems in a random environment. Communications in Statistics: Theory and Methods, 26(2):309–316, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Charalambides:1990:ASD

[Cha90a] Ch. A. Charalambides. Abel series distributions with applications to fluctuations of sample functions of stochastic processes. Communications in Statistics: Theory and Methods, 19(1):317– 335, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chaudhry:1990:WJR

[Cha90b] Nazir Ahmed Chaudhry. The weighted jackknife for ratio estimation. Communications in Statistics: Theory and Methods, 19(9):3283–3313, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chan:1991:CML

[Cha91a] Shih-Huang Chan. On the calculations of the maximum likelihood estimates for the polynomial spline regression model with unknown knots and AR(1) errors. Communications in Statistics: Theory and Methods, 20(4):1199–1209, 1991. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chang:1991:EEK

[Cha91b] Myron N. Chang. Edgeworth expansion for the Kaplan–Meier estimator. Communications in Statistics: Theory and Methods, 20(8):2479–2494, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chaubey:1991:SRP

[Cha91c] Yogendra P. Chaubey. A study of ratio and product estimators under a super population model. Communications in Statistics: Theory and Methods, 20(5–6):1731–1746, 1991. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Champ:1992:SSR

[Cha92a] Charles W. Champ. Steady-state run length analysis of a Shewhart quality control chart with supplementary runs rules. Communications in Statistics: Theory and Methods, 21(3):765–777, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chan:1992:ODQ

[Cha92b] Ling-Yau Chan. D-optimal design for a quadratic log contrast model for experiments with mixtures. Communications in Statistics: Theory and Methods, 21(10):2909–2930, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chaturvedi:1993:RRE

[Cha93a] Ancop Chaturvedi. Ridge regression estimators in the linear regression models with non-spherical errors. Communications in Statistics: Theory and Methods, 22(8):2275–2284, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chaudhuri:1993:CVC

[Cha93b] Gopal Chaudhuri. Coefficient of variation for the L-class of life distributions. Communications in Statistics: Theory and Methods, 22(9):331–344, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chaturvedi:1995:BPA

[CHA95] Anoop Chaturvedi, Hikaru Hasegawa, and Sheel Asthana. Bayesian predictive analysis of the linear regression model with an Edgeworth series prior distribution. Communications in Statistics: Theory and Methods, 24(10):2469–2484, 1995. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chang:1996:ADI

[Cha96a] Myron N. Chang. On the asymptotic distribution of an isotonic window estimator for the generalized failure rate function. Communications in Statistics: Theory and Methods, 25(10):2239– 2249, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chang:1996:ODG

[Cha96b] Myron N. Chang. Optimal designs for group sequential clinical trials. Communications in Statistics: Theory and Methods, 25 (2):361–379, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chaturvedi:1997:BAL

[CHA97a] Anoop Chaturvedi, Hikaru Hasegawa, and Sheel Asthana. Bayesian analysis of the linear regression model with an Edgeworth series prior distribution. Communications in Statistics: Theory and Methods, 26(5):1145–1164, 1997. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chaudhuri:1997:PMS

[Cha97b] Arijit Chaudhuri. On a pragmatic modification of survey sampling in three stages. Communications in Statistics: Theory and Methods, 26(7):1805–1810, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chateau:1999:SDA

[Cha99] Frédéric Chateau. Structured discriminant analysis. Communications in Statistics: Theory and Methods, 28(2):255–266, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cheng:1998:GFT

[CHC98] K. F. Cheng, H. M. Hsueh, and T. H. Chien. Goodness of fit tests with misclassified data. Communications in Statistics: Theory and Methods, 27(6):1379–1393, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cheng:1990:AKR

[Che90] Philip E. Cheng. Applications of kernel regression estimation: survey. Communications in Statistics: Theory and Methods, 19(11):4103–4134, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chen:1992:BAB

[Che92a] Cathy W. S. Chen. Bayesian analysis of bilinear time series models: a Gibbs sampling approach. Communications in Statistics: Theory and Methods, 21(12):3407–3425, 1992. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chen:1992:BIF

[Che92b] Cathy W. S. Chen. Bayesian inferences and forecasting in bilinear time series models. Communications in Statistics: Theory and Methods, 21(6):1725–1743, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chen:1994:UFI

[Che94] Dafeng Chen. Unbiasedness of Fourier integral estimator under weaker assumptions. Communications in Statistics: Theory and Methods, 23(5):1275–1281, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cheng:1995:CSC

[Che95a] Chuen-Sheng Cheng. A cumulative score control scheme for detecting process shifts. Communications in Statistics: Theory and Methods, 24(3):755–774, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cheolyong:1995:SRC

[Che95b] Park Cheolyong. Some remarks on the chi-squared test with both margins fixed. Communications in Statistics: Theory and Methods, 24(3):653–631, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chen:1996:E

[Che96] Sy-Men Chen. Erratum. Communications in Statistics: Theory and Methods, 25(5):1151–1154, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chen:1997:RTW

[Che97] Chu-Ohih Chen. Rank transformations when covariables are measured with error or mismodelled. Communications in Statistics: Theory and Methods, 26(12):2967–2982, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chen:1998:PVA

[Che98a] James J. Chen. P-value adjustment for multiple binary endpoints. Communications in Statistics: Theory and Methods, 27(11):2791–2806, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chen:1998:TCP

[Che98b] Jie Chen. Testing for a change point in linear regression models. Communications in Statistics: Theory and Methods, 27(10): 2481–2493, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chen:1998:EPI

[Che98c] K. S. Chen. Estimation of the process incapability index. Communications in Statistics: Theory and Methods, 27(5):1263– 1274, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chen:1998:BAC

[Che98d] Ming-Hui Chen. Bayesian analysis of correlated mixed categorical data by incorporating historical prior information. Communications in Statistics: Theory and Methods, 27(6):1341–1361, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chechile:1999:VBG

[Che99] Richard A. Chechile. A vector-based goodness-of-fit metric for interval-scaled data. Communications in Statistics: Theory and Methods, 28(2):277–296, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chiou:1992:SER

[Chi92] Paul Chiou. Shrinkage estimation of reliability in the exponential distribution. Communications in Statistics: Theory and Methods, 21(6):1745–1758, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chiou:1993:EBS

[Chi93] Paul Chiou. Empirical Bayes shrinkage estimation of reliability in the exponential distribution. Communications in Statistics: Theory and Methods, 22(5):1483–1494, 1993. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chiou:1995:PPE

[Chi95] Paul Chiou. A pooling procedure for the exponential distribution. Communications in Statistics: Theory and Methods, 24 (2):445–458, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chikuse:1999:PAS

[Chi99a] Yasuko Chikuse. Procrustes analysis on some special manifolds. Communications in Statistics: Theory and Methods, 28(3–4): 885–903, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chiou:1999:CIE

[Chi99b] Paul Chiou. Conditional interval estimation of the extreme-value scale parameter following rejection of a pre-test. Communications in Statistics: Theory and Methods, 28(9):2065–2077, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chiu:1999:UES

[Chi99c] S. N. Chiu. An unbiased estimator for the survival function of censored data. Communications in Statistics: Theory and Methods, 28(9):2249–2260, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Choi:1990:DAD

[Cho90a] ByoungSeon Choi. A derivation of the asymptotic distribution of the partial autocorrelation function of an autoregressive process.

Communications in Statistics: Theory and Methods, 19(2):547– 553, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Choi:1990:BDB

[Cho90b] ByoungSeon Choi. On the block LU decomposition of block Toeplitz matrices using the vector Yule–Walker equations. Communications in Statistics: Theory and Methods, 19(8):2815– 2827, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Choi:1991:AJD

[Cho91a] ByoungSeon Choi. The asymptotic joint distribution of the correlation determinant estimates of an ARMA process. Communications in Statistics: Theory and Methods, 20(9):2823–2835, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Choi:1991:ADA

[Cho91b] ByoungSeon Choi. On the asymptotic distributions of the r and s arrays for identifying an ARMA process. Communications in Statistics: Theory and Methods, 20(10):3187–3206, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chou:1992:BND

[Cho92] Youn Min Chou. A bivariate noncentral T -distribution with applications. Communications in Statistics: Theory and Methods, 21(12):3427–3462, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Choi:1994:CME

[Cho94a] ByoungSeon Choi. On the covariance matrix estimators of the white noise process of a vector autoregressive model. Communications in Statistics: Theory and Methods, 23(1):249–256, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Choudhury:1994:UFO

[Cho94b] Askar H. Choudhury. Untransformed first observation problem in regression model with moving average process. Communications in Statistics: Theory and Methods, 23(10):2927–2937, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chow:1994:RET

[Cho94c] Hwee-Kwan Chow. Robust estimation in time series: an approximation to the Gaussian sum filter. Communications in Statistics: Theory and Methods, 23(12):3491–3505, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Christoffersson:1997:RAE

[Chr97] Jan Christoffersson. Resampling the autocovariance estimator in stationary Gaussian processes. Communications in Statistics: Theory and Methods, 26(7):1745–1766, 1997. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Castillo:1998:MEL

[CHS98a] Enrique Castillo, Ali S. Hadi, and José María Sarabia. A method for estimating Lorenz curves. Communications in Statistics: Theory and Methods, 27(8):2037–2063, 1998. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chang:1998:GSD

[CHS98b] Myron N. Chang, Irving K. Hwang, and Weichung J. Shin. Group sequential designs using both Type I and Type II error probability spending functions. Communications in Statistics: Theory and Methods, 27(6):1323–1339, 1998. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chu:1992:EDR

[Chu92] David P. T. Chu. Exact distribution of random volume from beta type-2 random points. Communications in Statistics: Theory and Methods, 21(3):589–597, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chu:1994:BMC

[Chu94a] C. K. Chu. Binned modified cross-validation with dependent errors. Communications in Statistics: Theory and Methods, 23(12):3515–3537, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chu:1994:ECP

[Chu94b] C. K. Chu. Estimation of change-points in a nonparametric regression function through kernel density estimation. Communications in Statistics: Theory and Methods, 23(11):3037–3062, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chung:1998:MML

[Chu98] Sai-Ho Chung. Modified maximum likelihood estimation. Communications in Statistics: Theory and Methods, 27(12):2925– 2942, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cipra:1998:RRE

[Cip98] T. Cipra. Robust recursive estimation in nonlinear time series. Communications in Statistics: Theory and Methods, 27(5):1071– 1082, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chakraborty:1998:ASS

[CJ98a] Sharmishtha Chakraborty and P. C. Joshi. Alternating sums of sub-diagonal product moments of order statistics. Communications in Statistics: Theory and Methods, 27(5):1049–1054, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cui:1998:TTF

[CJ98b] Xuewei Cui and Peter W. M. John. Time-trend free run orders with the minimum level changes. Communications in Statistics: Theory and Methods, 27(1):55–68, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Crouse:1995:URE

[CJH95] Robert H. Crouse, Chun Jin, and R. C. Hanumara. Unbiased ridge estimation with prior information and ridge trace. Communications in Statistics: Theory and Methods, 24(9):2341–2354, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Christopher:1995:EPM

[CK95] Susan Reade Christopher and Lawrence L. Kupper. On the effects of predictor misclassification in multiple linear regression analysis. Communications in Statistics: Theory and Methods, 24(1):13–37, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cing:1996:ESI

[CK96] Hülya Cing and Cem Kadilar. Estimation and smoothing from incomplete data for a class of lattice processes. Communications in Statistics: Theory and Methods, 25(3):683, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chan:1997:LHM

[CK97a] Nai N. Chan and Tsz-Kit Keung. Linear hypotheses made explicit. Communications in Statistics: Theory and Methods, 26 (2):387–400, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chung:1997:SEM

[CK97b] Younshik Chung and Chansoo Kim. Simultaneous estimation of the multivariate normal mean under balanced loss function. Communications in Statistics: Theory and Methods, 26(7):1599– 1611, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chan:1995:MIA

[CKH95] Nai N. Chan and Li Kim-Hung. A matrix identity and its applications to equivalent hypotheses in linear models. Communications in Statistics: Theory and Methods, 24(11):2769–2777, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chen:1990:URA

[CL90] Hanfeng Chen and Wei-Yin Loh. Uniform robustness against nonnormality of the t and F tests. Communications in Statistics: Theory and Methods, 19(10):3707–3723, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

chen:1991:CBT

[cL91] Hanfeng chen and Wei-Yin Loh. Consistency of the bootstrap for the transformed two-sample t-test. Communications in Statistics: Theory and Methods, 20(3):997–1014, 1991. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chow:1994:SCB

[CL94] Shein-Chung Chow and Jen-Pei Liu. Statistical considerations in bioequivalence trials. Communications in Statistics: Theory and Methods, 23(2):325–339, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Collins:1996:ARE

[CL96] John R. Collins and Shawn X. Liu. Asymptotic relative efficiency comparisions for some robust location estimators. Communications in Statistics: Theory and Methods, 25(1):63–82, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ciampi:1997:SMB

[CL97] Antonio Ciampi and Yves Lechevallier. Statistical models as building blocks of neural networks. Communications in Statistics: Theory and Methods, 26(4):991–1009, 1997. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Colbourn:1998:CPT

[CL98] Charles J. Colbourn and Alan C. H. Ling. A class of partial triple systems with applications in survey sampling. Communications in Statistics: Theory and Methods, 27(4):1009–1018, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chen:1997:TPP

[CLC97] Lin-An Chen, Tzong-Shi Lee, and Wenyaw Chan. Tensor product polynomial splines. Communications in Statistics: Theory and Methods, 26(9):2093–2111, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cleroux:1995:VCB

[CLL95] Robert Cléroux, Aziz Lazraq, and Yves Lepage. Vector correlation based on ranks and a nonparametric test of no association between vectors. Communications in Statistics: Theory and Methods, 24(3):713–733, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chaudhuri:1992:NTV

[CM92] Arijit Chaudhuri and Joydip Mitra. A note on two variance estimators for Rao–Hartley–Cochran estimator. Communications in Statistics: Theory and Methods, 21(12):3535–3543, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chaudhuri:1994:BSP

[CM94a] Arijit Chaudhuri and Tapabrata Maiti. Borrowing strength from past data in small domain prediction by Kalman filtering — a case study. Communications in Statistics: Theory and Methods, 23(12):3507–3514, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chaudhuri:1994:VEM

[CM94b] Arijit Chaudhuri and Tapabrata Maiti. Variance estimation in model assisted survey sampling. Communications in Statistics: Theory and Methods, 23(4):1203–1214, 1994. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chen:1994:IST

[CM94c] James J. Chen and Gary E. Moore. Impact of surviving time on tests for carcinogenicity. Communications in Statistics: Theory and Methods, 23(5):1375–1386, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chaudhuri:1995:ADM

[CM95a] Arijit Chaudhuri and Tapabrata Maiti. Asymptotic design-cum-model approach for convex weighting of direct and indirect components of small domain predictors. Communications in Statistics: Theory and Methods, 24(11):2803–2813, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chen:1995:ENO

[CM95b] Ling Chen and Jie Mi. Estimation of non-occurrence probability in periodic model. Communications in Statistics: Theory and Methods, 24(6):1493–1505, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cheng:1995:NFT

[CM95c] Peter H. Cheng and Cliff Y. K. Meng. A new formula for tail probabilities of Dunnett’s t with unequal sample sizes. Communications in Statistics: Theory and Methods, 24(2):523–532, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chaudhuri:1996:SCI

[CM96a] Arijit Chaudhuri and Joydip Mitra. Setting confidence intervals by ratio estimator. Communications in Statistics: Theory and Methods, 25(5):1135–1148, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cornell:1996:FMD

[CM96b] John A. Cornell and Douglas C. Montgomery. Fitting models to data: Interaction versus polynomial? Your choice. Communications in Statistics: Theory and Methods, 25(11):2531–2555, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chatterjee:1997:RRW

[CM97] Samprit Chatterjee and Martin Mächler. Robust regression: a weighted least squares approach. Communications in Statistics: Theory and Methods, 26(6):1381–1394, 1997. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chen:1998:SES

[CM98] Zhenmin Chen and Jie Mi. Statistical estimation for the scale parameter of the gamma distribution based on grouped data. Communications in Statistics: Theory and Methods, 27(12):3035– 3045, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Case:1994:TSD

[CMD94] L. Douglas Case, Timothy M. Morgan, and C. E. Davis. Threestage designs for monitoring clinical trials. Communications in Statistics: Theory and Methods, 23(7):1875–1893, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chen-Mok:1999:NEB

[CMS99] Mario Chen-Mok and Pranab K. Sen. Nondifferentiable errors in beta-compliance integrated logistic models. Communications in Statistics: Theory and Methods, 28(3–4):931–946, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cavanaugh:1999:GDS

[CN99a] Joseph E. Cavanaugh and Andrew A. Neath. Generalizing the derivation of the Schwarz information criterion. Communications in Statistics: Theory and Methods, 28(1):49–66, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Conti:1999:AEI

[CN99b] Pier Luigi Conti and Yakov Nikitin. Asymptotic efficiency of independence tests based on Gini’s rank association coefficient, Spearman’s footrule and their generalizations. Communications in Statistics: Theory and Methods, 28(2):453–465, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Choi:1990:SNP

[CO90] Byoung-Chul Choi and Donald B. Owen. A study of a new process capability index. Communications in Statistics: Theory and Methods, 19(4):1231–1245, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chou:1991:LRT

[CO91] Youn-Min Chou and D. B. Owen. A likelihood ratio test for the equality of proportions of two normal populations. Communications in Statistics: Theory and Methods, 20(8):2357–2374, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Coad:1994:STC

[Coa94] D. S. Coad. Sequential tests with covariates with an application to censored survival data. Communications in Statistics: Theory and Methods, 23(1):277–287, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Colombi:1998:MLM

[Col98] Roberto Colombi. A multivariate logit model with marginal canonical association. Communications in Statistics: Theory and Methods, 27(12):2953–2971, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Consul:1990:GDP

[Con90a] P. C. Consul. Geeta distribution and its properties. Communications in Statistics: Theory and Methods, 19(8):3051–3068, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Consul:1990:NCL

[Con90b] P. C. Consul. New class of location-parameter discrete probability distributions and their characterizations. Communications in Statistics: Theory and Methods, 19(12):4653–4666, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Consul:1990:SPA

[Con90c] P. C. Consul. On some properties and applications of quasibinomial distribution. Communications in Statistics: Theory and Methods, 19(2):477–504, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Consul:1990:TSM

[Con90d] P. C. Consul. Two stochastic models for the Geeta distribution. Communications in Statistics: Theory and Methods, 19 (10):3699–3706, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Consul:1994:SPD

[Con94a] P. C. Consul. On some probability distributions associated with random walks. Communications in Statistics: Theory and Methods, 23(11):3241–3255, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Consul:1994:SBF

[Con94b] P. C. Consul. Some bivariate families of Lagrangian probability distributions. Communications in Statistics: Theory and Methods, 23(10):2895–2906, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cornell:1990:EBU

[Cor90] Richard G. Cornell. The evaluation of bioequivalence using nonparametric procedures. Communications in Statistics: Theory and Methods, 19(11):4153–4165, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cornell:1991:MDP

[Cor91a] John A. Cornell. Mixture designs for product improvement studies. Communications in Statistics: Theory and Methods, 20(2): 391–416, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cornell:1991:NTD

[Cor91b] Richard G. Cornell. Nonparametric tests of dispersion for the two-period crossover design. Communications in Statistics: Theory and Methods, 20(3):1099–1106, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cordeiro:1992:BCB

[Cor92] Gauss M. Cordeiro. Bartlett corrections and bias correction for two heteroscedastic regression models. Communications in Statistics: Theory and Methods, 22(1):169–188, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Correa:1995:NEE

[Cor95] Juan C. Correa. A new estimator of entropy. Communications in Statistics: Theory and Methods, 24(10):2439–2449, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Costa:1998:VCS

[Cos98] Antonio F. B. Costa. VSSI X̄ charts with sampling at fixed times. Communications in Statistics: Theory and Methods, 27 (11):2853–2869, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cox:1992:TOS

[Cox92] James Robison Cox. Tables of order statistics of normal random variables under linear trend. Communications in Statistics: Theory and Methods, 21(12):3497–3520, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Calabria:1990:BEE

[CP90] R. Calabria and G. Pulcini. Bayes estimation in exponential censored samples with incomplete information. Communications in Statistics: Theory and Methods, 19(8):3037–3049, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Calabria:1992:BPI

[CP92] R. Calabria and G. Pulcini. Bayes probability intervals in a load-strength model. Communications in Statistics: Theory and Methods, 21(12):3393–3405, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Calabria:1994:BSP

[CP94] R. Calabria and G. Pulcini. Bayes 2-sample prediction for the inverse Weibull distribution. Communications in Statistics: Theory and Methods, 23(6):1811–1824, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Calabria:1995:BPN

[CP95] R. Calabria and G. Pulcini. Bayes prediction of number of failures in Weibull samples. Communications in Statistics: Theory and Methods, 24(2):487–499, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Calabria:1996:MLB

[CP96a] R. Calabria and G. Pulcini. Maximum likelihood and Bayes prediction of current system lifetime. Communications in Statistics: Theory and Methods, 25(10):2297–2309, 1996. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Calabria:1996:PEU

[CP96b] R. Calabria and G. Pulcini. Point estimation under asymmetric loss functions for left-truncated exponential samples. Communications in Statistics: Theory and Methods, 25(3):585–600, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chen:1996:TSR

[CP96c] Lin-An Chen and Stephen Portnoy. Two-stage regression quantiles and two-stage trimmed least squares estimators for structural equation models. Communications in Statistics: Theory and Methods, 25(5):1005–1032, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Calabria:1997:BIM

[CP97a] R. Calabria and G. Pulcini. Bayes inference for the modulated power law process. Communications in Statistics: Theory and Methods, 26(10):2421–2438, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chen:1997:ADE

[CP97b] Sy-Mien Chen and W. L. Pearn. The asymptotic distribution of the estimated process capability index C_{pk}. Communications in Statistics: Theory and Methods, 26(10):2489–2497, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Calabria:1999:TRE

[CP99a] R. Calabria and G. Pulcini. On testing for repair effect and time trend in repairable mechanical units. Communications in Statistics: Theory and Methods, 28(2):367–387, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chen:1999:UPT

[CP99b] Hongzi Chen and Steven Piantadosi. Uniqueness of permuted treatment estimator distributions in the proportional hazards model. Communications in Statistics: Theory and Methods, 28(12):2803–2819, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Carpenter:1992:ESL

[CPK92] Mark Carpenter, Nabendu Pal, and Debashis Kushary. Estimation of the smaller and larger of two exponential location parameters. Communications in Statistics: Theory and Methods, 21(4):883–895, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chen:1991:SRG

[CQ91] Jiahua Chen and Jing Qin. Some results on generalized regression quantiles. Communications in Statistics: Theory and Methods, 20(3):911–928, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chanda:1992:ANL

[CR92a] Kamal C. Chanda and Frits H. Ruymgaart. Asymptotic normality of linear combinations of functions of the concomitant order statistics. Communications in Statistics: Theory and Methods, 21(11):3247–3254, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Croux:1992:CHB

[CR92b] Christophe Croux and Peter J. Rousseeuw. A class of highbreakdown scale estimators based on subranges. Communications in Statistics: Theory and Methods, 21(7):1935–1951, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Campbell:1993:ALD

[CR93] Gregory Campbell and Makarand V. Ratnaparkhi. An application of Lomax distributions in receiver operating characteristic (ROC) curve analysis. Communications in Statistics: Theory and Methods, 22(6):1681–1687, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Crosier:1991:SME

[Cro91] Ronald B. Crosier. Symmetry in mixture experiments. Communications in Statistics: Theory and Methods, 20(5–6):1911–1935, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Crowell:1996:STP

[Cro96] John I. Crowell. On a score test for pure birth processes and a Poisson dispersion test. Communications in Statistics: Theory and Methods, 25(7):1525–1536, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cabuk:1990:DQN

[CS90a] Serafettin Cabuk and Melvin D. Springer. Distribution of the quotient of noncentral normal random variables. Communications in Statistics: Theory and Methods, 19(3):1157–1168, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Castillo:1990:BDS

[CS90b] Enrique Castillo and José María Sarabia. Bivariate distributions with second kind beta conditionals. Communications in Statistics: Theory and Methods, 19(9):3433–3445, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chaganty:1990:LDL

[CS90c] Narasinga Rao Chaganty and Sanjeev Sabnis. Large deviation local limit theorems for ratio statistics. Communications in Statistics: Theory and Methods, 19(11):4083–4101, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chatterjee:1993:CGN

[CS93] Samprit Chatterjee and Gary Simon. Confidentiality guaranteed: a noninvasive procedure for collecting sensitive information. Communications in Statistics: Theory and Methods, 22(6):

1629–1651, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chatterjee:1994:AAV

[CS94a] Shoutir Kishore Chatterjee and Sisir Kumar Samanta. An alternative approach to variable selection for prediction. Communications in Statistics: Theory and Methods, 23(8):2157–2174, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chen:1994:TMS

[CS94b] T. Timothy Chen and Richard Simon. Two multi-step selection procedures with possible selection of two treatments of equal prior preference. Communications in Statistics: Theory and Methods, 23(3):781–801, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chen:1995:SCB

[CS95] Hanfeng Chen and John Q. Su. Simultaneous confidence bands for comparing diagnostic markers. Communications in Statistics: Theory and Methods, 24(2):375–393, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chen:1996:ETS

[CS96a] T. Timothy Chen and Richard Simon. Extension of two-sided test to multiple treatment trials. Communications in Statistics: Theory and Methods, 25(5):947–965, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cohen:1996:TUA

[CS96b] Arthur Cohen and H. B. Sackrowitz. Tests for the umbrella alternative under normality. Communications in Statistics: Theory and Methods, 25(11):2807–2817, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chang:1997:TOD

[CS97a] Hsing Yi Chang and Chirayath M. Stichindran. Testing overdispersion in data with censoring using the mixture of exponential families. Communications in Statistics: Theory and Methods, 26(12):2945–2966, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cheng:1997:CM

[CS97b] Yuk W. Cheng and Deborah J. Street. The CGMANOVA model. Communications in Statistics: Theory and Methods, 26(5):1083– 1098, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cheng:1997:CON

[CS97c] Yuk W. Cheng and Deborah J. Street. Constructions for optimal non-strongly-balanced change-over designs. Communications in Statistics: Theory and Methods, 26(5):1073–1082, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chen:1998:IDF

[CSK^{+}98] T. Timothy Chen, Ph.D., Richard M. Simon, Ph.D., Edward L. Kom, Ph.D., Stewart J. Anderson, Ph.D., Anne S. Lindblad, Ph.D., Harry Sam Wieand, Ph.D., Harold O. Douglass, Jr., M.D., Bernard Fisher, M.D., J. Michael Hamilton, M.D., and Michael A. Friedman, M.D. Investigation of disease-free survival as a surrogate endpoint for survival in cancer clinical trials. Communications in Statistics: Theory and Methods, 27(6): 1363–1378, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chib:1991:RBA

[CT91] Siddhartha Chib and Ram C. Tiwari. Robust Bayes analysis in normal linear regression with an improper mixture prior. Communications in Statistics: Theory and Methods, 20(3):807–829, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chalton:1992:PGA

[CT92] Derek O. Chalton and Cas G. Troskie. Q plots, a graphical aid for regression analysis. Communications in Statistics: Theory and Methods, 21(3):625–636, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cameron:1993:RLS

[CT93a] Murray A. Cameron and T. Rolf Turner. Recursive location and scale estimators. Communications in Statistics: Theory and Methods, 22(9):2201–2226, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chalton:1993:CSP

[CT93b] Derek O. Chalton and Cas. G. Troskie. On the compatibility of sample and prior information in the mixed regression model. Communications in Statistics: Theory and Methods, 22(3):921– 928, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chang:1994:PBT

[CT94] Dong Shang Chang and Loon Ching Tang. Percentile bounds and tolerance limits for the Birnbaum–Saunders distribution. Communications in Statistics: Theory and Methods, 23(10): 2853–2863, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Craigmile:1997:PEF

[CT97] Peter F. Craigmile and D. M. Tirrerington. Parameter estimation for finite mixtures of uniform distributions. Communications in Statistics: Theory and Methods, 26(8):1981–1995, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chen:1998:TLS

[CT98] Lin-An Chen and Peter Thompson. Trimmed least squares estimator as best trimmed linear conditional estimator for linear regression model. Communications in Statistics: Theory and Methods, 27(7):1835–1849, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chalton:1999:PCM

[CT99] D. O. Chalton and C. G. Troskie. Parameter changes in the multiple regression model with autocorrelated errors: Bayesian analysis. Communications in Statistics: Theory and Methods, 28 (1):137–142, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chaturvedi:1993:MSE

[CTP93] Ajit Chaturvedi, Neeraj Tiwari, and S. K. Pandey. Multi-stage estimation of the common location parameter of several exponential distributions. Communications in Statistics: Theory and Methods, 22(5):1413–1423, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Castano-Tostado:1990:SCE

[CTT90] Eduardo Castaño-Tostado and Yutaka Tanaka. Some comments on Escoufier’s RV -coefficient as a sensitivity measure in principal component analysis. Communications in Statistics: Theory and Methods, 19(12):4619–4626, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Castano-Tostado:1991:SMI

[CTT91] Eduardo Castaño-Tostado and Yutaka Tanaka. Sensitivity measures of influence on the loading matrix in exploratory factor analysis. Communications in Statistics: Theory and Methods, 20(4):1329–1343, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cussens:1996:ESS

[Cus96] James Cussens. Effective sample size in a dichotomous process with noise. Communications in Statistics: Theory and Methods, 25(6):1233–1246, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cabrera:1990:SMR

[CW90] J. Cabrera and G. S. Watson. On a spherical median related distribution. Communications in Statistics: Theory and Methods, 19(6):1973–1986, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cheng:1992:EER

[CW92a] K. F. Cheng and J. W. Wu. Estimation of exponential regression parameters using binary data. Communications in Statistics: Theory and Methods, 21(8):2203–2214, 1992. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cuevas:1992:EGD

[CW92b] Antonio Cuevas and Gilbert G. Walter. On estimation of generalized densities. Communications in Statistics: Theory and Methods, 21(7):1807–1821, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chen:1993:NCC

[CW93] X. R. Chen and Y. Wu. On a necessary condition for the consistency of the L_{1} estimates in linear regression models. Communications in Statistics: Theory and Methods, 22(3):631–639, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chang:1994:DAS

[CW94a] Horng-Jinh Chang and Bor-Shyh Wen. The distribution of an actual sample size increment in stratified random sampling. Communications in Statistics: Theory and Methods, 23(6):1735– 1742, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chen:1994:CEC

[CW94b] X. R. Chen and Y. Wu. Consistency of L_{1} estimates in censored linear regression models. Communications in Statistics: Theory and Methods, 23(7):1847–1858, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cao:1997:CDO

[CW97] Guoliang Cao and Mike West. Computing distributions of order statistics. Communications in Statistics: Theory and Methods, 26(3):755–764, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Collins:1998:CAB

[CW98] John R. Collins and Boll Wu. Comparisons of asymptotic biases and variances of m-estimators of scale under asymmetric contamination. Communications in Statistics: Theory and Methods, 27(7):1791–1810, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Cheung:1993:SPI

[CWH93] Siu Hung Cheung, Ka Ho Wu, and Burt Holland. Simultaneous prediction intervals for all pairwise differences among several future sample means. Communications in Statistics: Theory and Methods, 22(11):3087–3096, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chan:1990:ADS

[CXZ90] Lai K. Chan, Z. Xiong, and D. Zhang. On the asymptotic distributions of some process capability indices. Communications in Statistics: Theory and Methods, 19(1):11–18, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chang-yu:1995:AIL

[Cy95] Lu Chang-yu. Admissibility of inhomogeneous linear estimators in linear models with respect to incomplete ellipsoidal restrictions. Communications in Statistics: Theory and Methods, 24(7):1737–1742, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chen:1997:NBD

[CZ97a] Jianbin Chen and Igor Zurbenko. Nonparametric boundary detection. Communications in Statistics: Theory and Methods, 26(12):2999–3014, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chen:1997:IFS

[CZ97b] Pinyuen Chen and Jun-Lue Zhang. An integrated formulation for selecting the best normal population: the common and unknown variance case. Communications in Statistics: Theory and Methods, 26(11):2701–2723, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Czado:1993:NRM

[Cza93] Claudia Czado. Norm restricted maximum likelihood estimators for binary regression models with parametric link. Communications in Statistics: Theory and Methods, 22(8):2259–2274, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dunsmore:1998:SPP

[DA98] Ian R. Dunsmore and Zeinab H. Amin. Some prediction problems concerning samples from the Pareto distribution. Communications in Statistics: Theory and Methods, 27(5):1221–1238, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dagne:1999:BAH

[Dag99] Getachew Asfaw Dagne. Bayesian analysis of hierarchical Poisson models with latent variables. Communications in Statistics: Theory and Methods, 28(1):119–136, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dahel:1995:MEP

[Dah95] S. Dahel. Mean estimation for a partitioned multinormal distribution. Communications in Statistics: Theory and Methods, 24 (3):601–606, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Damaraju:1993:CSC

[Dam93] C. V. Damaraju. Category-specific covariate effects in longitudinal study designs. Communications in Statistics: Theory and Methods, 22(5):1441–1469, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dasiou:1999:AEL

[DAM99] Despina Dasiou, Lefteris Angelis, and Chronis Moyssiadis. Alternative m-estimators of location and their linear convex combinations. Communications in Statistics: Theory and Methods, 28(1):19–33, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Datta:1991:PCH

[Dat91] Gauri Sankar Datta. Pitman closeness for hierarchical Bayes predictors in mixed linear models. Communications in Statistics: Theory and Methods, 20(11):3713–3727, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Davis:1990:CLR

[Dav90] Linda June Davis. Collapsibility of likelihood ratio tests in multidimensional contingency tables. Communications in Statistics: Theory and Methods, 19(2):465–476, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

David:1991:IEC

[Dav91] R. Fox David. Interval estimation in calibration models: an application of Fieller’s theorem. Communications in Statistics: Theory and Methods, 20(2):711–720, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Das:1995:PTE

[DB95a] G. Das and K. Bez. Preliminary test estimators in double sampling with two auxiliary variables. Communications in Statistics: Theory and Methods, 24(5):1211–1226, 1995. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

DeGooijer:1995:INL

[DB95b] Jan G. De Gooijer and Kurt Brännäs. Invertibility of non-linear time series models. Communications in Statistics: Theory and Methods, 24(11):2701–2714, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dey:1992:CPD

[DC92] Dipak K. Dey and Younshik Chung. Compound Poisson distributions: Properties and estimation*. Communications in Statistics: Theory and Methods, 21(11):3097–3121, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Draisma:1996:EEV

[DD96] G. Draisma and L. De Hann. An estimator for the extreme-value index. Communications in Statistics: Theory and Methods, 25 (4):685–694, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Delicado:1997:WNR

[DD97a] P. Delicado and M. Del Rio. Weighted nonparametric regression. Communications in Statistics: Theory and Methods, 26 (12):2983–2998, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dernidenko:1997:PMM

[DD97b] Eugene Dernidenko and Spiegelman Donna. A paradox: more measurement error can lead to more efficient estimates. Communications in Statistics: Theory and Methods, 26(7):1649–1675, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

DeOliveira:1992:IED

[De 92] Tiago J. De Oliveira. Intrinsic estimation of the dependence function in bivariate extremes: a new and simpler approach. Communications in Statistics: Theory and Methods, 21(3):599– 611, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Demetrios:1992:GSS

[Dem92] Papanastassiou Demetrios. A generalized state space model and its application to a set of extended Kalman filter equations. Communications in Statistics: Theory and Methods, 21(11):

3255–3262, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Demidenko:1996:ELS

[Dem96] Eugene Demidenko. On the existence of the least squares estimate in nonlinear growth curve models of exponential type. Communications in Statistics: Theory and Methods, 25(1):159– 182, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dennis:1991:HDT

[Den91] Samuel Y. Dennis III. On the hyper-Dirichlet Type 1 and hyper-Liouville distributions. Communications in Statistics: Theory and Methods, 20(12):4069–4081, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dennis:1994:DPI

[Den94] Samuel Y. Dennis III. On the distribution of products of independent beta variates. Communications in Statistics: Theory and Methods, 23(7):1895–1913, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dennis:1998:BER

[Den98] Samuel Y. Dennis III. Bayes estimators of ratios of proportions between tree-structured experiments. Communications in Statistics: Theory and Methods, 27(10):2495–2514, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dey:1991:C

[Dey91] Dipak K. Dey. Comments. Communications in Statistics: Theory and Methods, 20(11):3685–3688, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Drezner:1993:GBD

[DF93] Zvi Drezner and Nicholas Farnum. A generalized binomial distribution. Communications in Statistics: Theory and Methods, 22(11):3051–3063, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

DeLaHorra:1994:BRC

[DF94a] Julián De La Horra and Carmen Fernández. Bayesian robustness of credible regions in the presence of nuisance parameters. Communications in Statistics: Theory and Methods, 23(3):689–699, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Drezner:1994:CPD

[DF94b] Zvi Drezner and Nicholas Farnum. A correlated Poisson distribution for correlated events. Communications in Statistics: Theory and Methods, 23(3):841–857, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

DeLaHorra:1995:SPI

[DF95] Julian De La Horra and Carmen Fernandez. Sensitivity to prior independence via Farlie–Gumbel–Morgenstern model. Communications in Statistics: Theory and Methods, 24(4):987–996, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Doganaksoy:1991:ICQ

[DFT91] Necip Doganaksoy, Frederick W. Faltin, and William T. Tucker. Identification of out of control quality characteristics in a multivariate manufacturing environment. Communications in Statistics: Theory and Methods, 20(9):2775–2790, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dikta:1990:BAC

[DG90] G. Dikta and J. K. Ghorai. Bootstrap approximation with censored data under the proportional hazard model. Communications in Statistics: Theory and Methods, 19(2):573–581, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Daniel:1992:CAR

[DG92] J. Twedt Daniel and D. S. Gill. Comparison of algorithms for replacing missing data in discriminant analysis. Communications in Statistics: Theory and Methods, 21(6):1567–1578, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dahel:1994:SDR

[DG94] S. Dahel and N. Giri. Some distributions related to a noncentral Wishart distribution. Communications in Statistics: Theory and Methods, 23(1):229–237, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Das:1997:OBD

[DG97a] Ashish Das and Sudhir Gupta. Optimal block designs for triallel cross experiments. Communications in Statistics: Theory and Methods, 26(7):1767–1777, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dhawan:1997:SOS

[DG97b] Anil Kumar Dhawan and Amar Nath Gill. Simultaneous onesided confidence intervals for the ordered pairwise differences of exponential location parameters. Communications in Statistics: Theory and Methods, 26(1):247–262, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Das:1992:OBD

[DGD92] Ashish Das, V. K. Gupta, and Praggya Das. E-optimal block designs under heteroscedastic model. Communications in Statistics: Theory and Methods, 21(6):1651–1666, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

DelCastillo:1996:RSP

[DGMR96] Enrique Del Castillo, James M. Grayson, Douglas C. Montgomery, and George C. Runger. A review of statistical process control techniques for short run manufacturing systems. Communications in Statistics: Theory and Methods, 25(11):2723– 2737, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Davis:1990:UBT

[DH90] Barry R. Davis and Robert J. Hardy. Upper bounds for Type I and Type II error rates in conditional power calculations. Communications in Statistics: Theory and Methods, 19(10):3571– 3584, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Davis:1992:RCI

[DH92] Barry R. Davis and Robert J. Hardy. Repeated confidence intervals and prediction intervals using stochastic curtailment. Communications in Statistics: Theory and Methods, 21(2):351–368, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dorfman:1995:MST

[DH95] Jeffrey H. Dorfman and Arthur Havenner. Model specification tests for balanced representation state space models. Communications in Statistics: Theory and Methods, 24(1):97–119, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dietrich:1996:MDE

[DH96] D. Dietrich and J. Hüsler. Minimum distance estimators in extreme value distributions. Communications in Statistics: Theory and Methods, 25(4):695–703, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dumitrescu:1999:EFN

[DH99] Monica E. Dumitrescu and Norma Faris Hubele. An entropic framework for the normal distribution in capability analysis. Communications in Statistics: Theory and Methods, 28(6):1361– 1377, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dhar:1995:ENM

[Dha95] Sunil K. Dhar. Extension of a negative multinomial model. Communications in Statistics: Theory and Methods, 24(1):39– 57, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Davis:1990:PZC

[DHW90] Robert B. Davis, Anthony Homer, and William H. Woodall. Performance of the zone control chart. Communications in Statistics: Theory and Methods, 19(5):1581–1587, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Diaz:1990:BFM

[Dia90] Joaquin Diaz. Bayesian forecasting for AR(1) models with normal coefficients. Communications in Statistics: Theory and Methods, 19(6):2229–2246, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Diaz:1991:BSM

[Dia91] Joaquin Diaz. A Bayesian solution to the multiple outlier problem in time series. Communications in Statistics: Theory and Methods, 20(1):45–61, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dikta:1995:ANU

[Dik95] G. Dikta. Asymptotic normality under the Koziol–Green model. Communications in Statistics: Theory and Methods, 24(6):1537– 1549, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dixit:1991:EPS

[Dix91] Ulhas J. Dixit. On the estimation of the power of the scale parameter of the gamma distribution in the presence of outliers. Communications in Statistics: Theory and Methods, 20(4):1315– 1328, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Diaz:1998:TMR

[DJ98] Carroll J. Diaz, Jr. and William D. Johnson. An F -test for multivariate repeated measures data with the Wiener stochastic process pattern in the covariance matrix. Communications in Statistics: Theory and Methods, 27(2):275–289, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Danielsson:1996:MMR

[DJD96] Jon Danielsson, Dennis W. Jansen, and Casper G. De Vries. The method of moments ratio estimator for the tail shape parameter. Communications in Statistics: Theory and Methods, 25(4):711– 720, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Davis:1994:IPZ

[DJG94] Robert B. Davis, Chun Jin, and Ying-Yueh Guo. Improving the performance of the zone control chart. Communications in Statistics: Theory and Methods, 23(12):3557–3565, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Djojosugito:1995:CSS

[Djo95] Rianto A. Djojosugito. A cubic smoothing spline based lack of fit test for nonlinear regression models. Communications in Statistics: Theory and Methods, 24(9):2183–2197, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Deshaukh:1991:EPE

[DK91] S. R. Deshaukh and M. M. Kale. Estimation of probability of extinction of a multitype branching process. Communications in Statistics: Theory and Methods, 20(10):3405–3419, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Deshpande:1995:CBJ

[DK95] J. V. Deshpande and S. R. Karia. Confidence bounds for the joint survival function in the dependent competing risks setup. Communications in Statistics: Theory and Methods, 24(9):2315– 2327, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dey:1990:EPR

[DL90] Dipak K. Dey and Pei-San Liao Liu. Estimation of parameters and reliability from generalized life models. Communications in Statistics: Theory and Methods, 19(3):1073–1099, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Draper:1995:CPD

[DL95] Norman R. Draper and Dennis K. J. Lin. Characterizing projected designs: repeat and mirror-image runs. Communications in Statistics: Theory and Methods, 24(3):775–795, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

delaHorra:1991:BCE

[dlH91] Julián de la Horra. Bayesian confidence estimation: an alternative approach. Communications in Statistics: Theory and Methods, 20(7):2299–2306, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Damien:1997:BEU

[DLS97] P. Damien, P. W. Laud, and A. F. M. Smith. Bayesian estimation of unimodal distributions. Communications in Statistics: Theory and Methods, 26(2):429–440, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Deng:1996:MOD

[DLW96] Lih-Yuan Deng, Dennis K. J. Lin, and Jiannong Wang. Marginally oversaturated designs. Communications in Statistics: Theory and Methods, 25(11):2557–2573, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Douglas:1991:DFM

[DM91] Critchlow E. Douglas and Fligner A. Michael. On distributionfree multiple comparisons in the one-way analysis of variance. Communications in Statistics: Theory and Methods, 20(1):127– 139, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dhariyal:1994:SSE

[DM94] Ishwari D. Dhariyal and Neeraj Misra. Simultaneous selection of extreme populations: a Bayesian approach. Communications in Statistics: Theory and Methods, 23(7):1993–2027, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Deheuvels:1996:CMT

[DM96] Paul Deheuvels and Guennady V. Martynov. Cramér–von Misestype tests with applications to tests of independence for multivariate extreme-value distributions. Communications in Statistics: Theory and Methods, 25(4):871–908, 1996. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

deMoraes:1995:PCO

[dMD95] Aipore R. de Moraes and Ian R. Dunsmore. Predictive comparisons in ordinal models. Communications in Statistics: Theory and Methods, 24(8):2145–2164, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Das:1991:IEV

[DMG91] K. Das, Q. Meneghini, and N. Giri. Improved estimation of variance components in balanced hierarchical mixed models. Communications in Statistics: Theory and Methods, 20(5–6):1653– 1664, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dargahi-Noubary:1990:MPE

[DN90] G. R. Dargahi-Noubary. A method for parameter estimation of non-linear regression with autocorrelated errors. Communications in Statistics: Theory and Methods, 19(8):2907–2923, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dargahi-Noubary:1992:DBG

[DN92] G. R. Dargahi-Noubary. Discrimination between Gaussian time series based on their spectral differences. Communications in Statistics: Theory and Methods, 21(9):2439–2458, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dong:1993:APQ

[Don93] Fang Dong. Asymptotic properties of quantiles for truncated and contaminated data. Communications in Statistics: Theory and Methods, 22(11):3255–3261, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dourleijn:1993:CEV

[Dou93] C. Johan Dourleijn. Combining estimators from variety trials that are connected by control varieties only. Communications in Statistics: Theory and Methods, 22(7):1829–1847, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Douke:1997:GSS

[Dou97] Hideyuki Douke. Group sequential χ^{2} statistic for testing the difference of two mean vectors in clinical trials. Communications in Statistics: Theory and Methods, 26(12):3015–3029, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

delPino:1990:CFA

[dP90] Guido E. del Pino. A coordinate free approach to score tests. Communications in Statistics: Theory and Methods, 19(1):155– 167, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dyer:1993:CPD

[DP93] Danny Dyer and Rebecca L. Pierce. On the choice of the prior distribution in hypergeometric sampling. Communications in Statistics: Theory and Methods, 22(8):2125–2146, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dagne:1997:BPS

[DP97a] Getachew Asfaw Dagne and S. James Press. Bayesian prediction for small areas using SUR models. Communications in Statistics: Theory and Methods, 26(6):1355–1379, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dey:1997:BAC

[DP97b] Dipak K. Dey and Sumitra Purkayastha. Bayesian approach to change point problems. Communications in Statistics: Theory and Methods, 26(8):2035–2047, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

DeMasi:1998:FBF

[DQS98] Ralph A. DeMasi, Bahjat Qaqish, and Pranab Kumar Sen. A family of bivariate failure time distributions with proportional crude and conditional hazards. Communications in Statistics: Theory and Methods, 27(2):365–384, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

delRio:1994:PTN

[dR94] Alejandro Quintela del Río. A plug-in technique in nonparametric regression with dependence. Communications in Statistics: Theory and Methods, 23(9):2581–2603, 1994. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Davy:1996:MGD

[DR96a] P. J. Davy and J. C. W. Rayner. Multivariate geometric distributions. Communications in Statistics: Theory and Methods, 25(12):2971–2987, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Drees:1996:RLF

[DR96b] H. Drees and R. D. Reiss. Residual life functionals at great age. Communications in Statistics: Theory and Methods, 25(4): 823–835, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Draper:1990:ARA

[Dra90] Norman R. Draper. Applied regression analysis bibliography update 1988–89. Communications in Statistics: Theory and Methods, 19(4):1205–1229, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Draper:1992:ARA

[Dra92] Norman R. Draper. Applied regression analysis bibliography update 1990–91. Communications in Statistics: Theory and Methods, 21(9):2415–2437, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Draper:1994:ARA

[Dra94] Norman R. Draper. Applied regression analysis bibliography update 1992–93. Communications in Statistics: Theory and Methods, 23(9):2701–2731, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dragalin:1995:OSS

[Dra95] V. Dragalin. An optimal sequential selection procedure. Communications in Statistics: Theory and Methods, 24(2):429–443, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Draper:1998:ARA

[Dra98] Norman R. Draper. Applied regression analysis bibliography update 1994–97. Communications in Statistics: Theory and Methods, 27(10):2581–2623, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dittrich:1993:BLT

[DRB93] Regina Dittrich, Erhard Reschenhofer, and Immanuel M. Bomse. Behaviour of the length test for medium sample sizes. Communications in Statistics: Theory and Methods, 22(9):207–216, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

DeLaHorra:1997:ABP

[DRB97] Julian De La Horra and Maria Teresa Rodriguez-Bernal. Asymptotic behaviour of the posterior predictive p-value. Communications in Statistics: Theory and Methods, 26(11):2689–2699, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Drees:1996:RPE

[Dre96] Holger Drees. Refined Pickands estimators with bias correction. Communications in Statistics: Theory and Methods, 25(4):837– 851, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Driessen:1991:MCS

[Dri91] Stefan Driessen. Multiple comparisons with and selection of the best treatment in (incomplete) block designs. Communications in Statistics: Theory and Methods, 20(1):179–217, 1991.

CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Duchesne:1997:MCM

[DRL97] Thierry Duchesne, Jacques Rioux, and Andrew Luong. Minimum Cramér–von Mises distance methods for complete and grouped data. Communications in Statistics: Theory and Methods, 26(2):401–420, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Deutsch:1990:ESO

[DRS90] Stuart J. Deutsch, Jeery E. Richards, and James J. Swain. Effects of a single outlier on ARMA identification. Communications in Statistics: Theory and Methods, 19(6):2207–2227, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Drummer:1990:EPR

[Dru90] Tom Drummer. Estimation of proportions and ratios from line transect data. Communications in Statistics: Theory and Methods, 19(8):3069–3091, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

David:1991:TSC

[DS91] H. T. David and Shawki A. Salem. Three shrinkage constructions for Pitman-closeness in the one-dimensional location case. Communications in Statistics: Theory and Methods, 20(11):3605– 3627, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Djojosugito:1992:BBC

[DS92] Rianto A. Djojosugito and Paul L. Speckman. Boundary bias correction in nonparametric density estimation. Communications in Statistics: Theory and Methods, 21(1):69–88, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Doganaksoy:1993:OPC

[DS93] Necip Doganaksoy and Josef Schmee. Orthogonal parameters with censored data. Communications in Statistics: Theory and Methods, 22(3):669–685, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dasgupta:1997:CMC

[DS97] Nairanjana Dasgupta and John D. Spurrier. A class of multivariate chi-square distributions with applications to comparison with a control. Communications in Statistics: Theory and Methods, 26(7):1559–1573, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

dosSantos:1999:CWT

[dSC99] Sylvio José P. dos Santos and Gauss M. Cordeiro. Corrected Wald test statistics for one-parameter exponential family models. Communications in Statistics: Theory and Methods, 28(6): 1391–1414, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dube:1991:SRE

[DSTW91] M. Dube, V. K. Srivastava, H. Toutenburg, and P. Wijekoon. Stein-rule estimator under inclusion of superfluous variables in linear regression models. Communications in Statistics: Theory and Methods, 20(7):2009–2022, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dupuis:1998:HEG

[DT98] D. J. Dupuis and M. Tsao. A hybrid estimator for generalized Pareto and extreme-value distributions. Communications in Statistics: Theory and Methods, 27(4):925–941, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Donegani:1991:EVA

[DU91] M. Donegani and M. Unternährer. Estimation of the variance of an adaptive estimator by bootstrapping. Communications in Statistics: Theory and Methods, 20(12):3925–3931, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dunn:1995:CVP

[Dun95] Charles L. Dunn. Critical values and powers for tests of uniformity of directions under multivariate normality. Communications in Statistics: Theory and Methods, 24(10):2541–2560, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Duthie:1991:OCD

[Dut91] A. I. Duthie. E-optimality and cubic designs. Communications in Statistics: Theory and Methods, 20(7):1977–1981, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Duthie:1993:COH

[Dut93] A. I. Duthie. A class of E-optimal hypercubic designs. Communications in Statistics: Theory and Methods, 22(3):717–722, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dutta:1999:LST

[Dut99] Harinarayan Dutta. Large sample tests for a regression model with autoregressive conditional heteroscedastic errors. Communications in Statistics: Theory and Methods, 28(1):105–117, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Deutsch:1990:EIC

[DW90] Stuart Jay Deutsch and C. C. Wang. Estimation and identification considerations for the multiconsequence intervention model. Communications in Statistics: Theory and Methods, 19(3):857– 884, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dimaki:1990:IID

[DX90] Caterina Dimaki and Evdokia Xekalaki. Identifiability of income distributions in the context of damage and generating models. Communications in Statistics: Theory and Methods, 19(8):2757– 2766, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dey:1991:MET

[DY91] Dipak K. Dey and Chung Younshik. Multiparameter estimation in truncated power series distributions under the Stein’s loss. Communications in Statistics: Theory and Methods, 20(1):309– 326, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Desmond:1995:SPI

[DY95] Anthony F. Desmond and Zhenlin Yang. Shortest prediction intervals for the Birnbaum–Saunders distribution. Communications in Statistics: Theory and Methods, 24(6):1383–1401, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Dong:1996:MVK

[DY96] Jianping Dong and Qian Ye. A minimum variance kernel estimator and a discrete frequency polygon estimator for ordinal contingency tables. Communications in Statistics: Theory and Methods, 25(12):3217–3245, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Desmond:1998:CLB

[DY98] Anthony F. Desmond and Zhenlin Yang. A comparison of likelihood and Bayesian inference for the threshold parameter in the inverse Gaussian distribution. Communications in Statistics: Theory and Methods, 27(9):2173–2183, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Eric:1995:CNP

[ÉA95] Belzile Éric and Jean-François Angers. Choice of noninformative priors for the variance components of an unbalanced oneway random model. Communications in Statistics: Theory and Methods, 24(5):1325–1341, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Evans:1992:MRM

[EB92] Marc A. Evans and Douglas G. Bonett. Multiple recapture methods when sampling without replacement in the last occasion. Communications in Statistics: Theory and Methods, 21(9): 2603–2624, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

El-Bassiouni:1994:SCI

[EB94] M. Y. El-Bassiouni. Short confidence intervals for variance components. Communications in Statistics: Theory and Methods, 23(7):1915–1933, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Esa:1998:SSA

[EB98] Sahib Esa and Georgi Boshnakov. Sufficient statistics for ARMA models with some fixed parameters. Communications in Statistics: Theory and Methods, 27(5):1083–1099, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Esparza:1993:MCS

[ED93] S. O. Esparza and B. S. Duran. Multistate coherent systems and the three modules theorem. Communications in Statistics: Theory and Methods, 22(2):561–573, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Edgeman:1996:SCR

[Edg96] Rick L. Edgeman. Sprt and cusum results for inverse Gaussian processes. Communications in Statistics: Theory and Methods, 25(11):2797–2806, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Edwards:1994:AAE

[Edw94] Lloyd J. Edwards. An alternative approach to estimation in the functional measurement error problem. Communications in Statistics: Theory and Methods, 23(6):1651–1664, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ensor:1993:MSM

[EE93] Katherine B. Ensor and Joe E. Ensor. Model selection in magnetic resonance imaging. Communications in Statistics: Theory and Methods, 22(12):3419–3434, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ennis:1992:HOL

[EF92] M. Ennis and D. A. S. Fraser. Higher order local unbiasedness with computer algebra. Communications in Statistics: Theory and Methods, 21(11):3171–3176, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Escolano:1995:FNA

[EGM95] Sylvie Escolano, Jean-Louis Golmard, and Alain Mallet. A fully nonparametric approach in survival models with explanatory variables. Communications in Statistics: Theory and Methods, 24(12):3027–3054, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Eubank:1993:TGF

[EHL93] R. L. Eubank, J. D. Hart, and V. N. LaRiccia. Testing goodness of fit via nonparametric function estimation techniques. Communications in Statistics: Theory and Methods, 22(12):3327–3354, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Einmahl:1996:EHD

[Ein96] John H. J. Einmahl. Extension to higher dimensions of the Jaeschke–Eicker result on the standardized empirical process. Communications in Statistics: Theory and Methods, 25(4):813– 822, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ennis:1993:NCC

[EJ93] Daniel M. Ennis and Norman L. Johnson. Noncentral and central chi-square, F and beta distribution functions as special cases of the distribution function of an indefinite quadratic form. Communications in Statistics: Theory and Methods, 22(3):897–905, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

ElBarmi:1997:TAT

[El 97] Hammou El Barmi. Testing for or against a trend in odds ratios. Communications in Statistics: Theory and Methods, 26(8):1877– 1891, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Encarnacion:1992:NLR

[Enc92] Mark J. Encarnación. A note on linear regression functions. Communications in Statistics: Theory and Methods, 21(3):665– 666, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Entsuah:1990:RPA

[Ent90] A. Richard Entsuah. Randomization procedures for analyzing clinical trial data with treatment related withdrawals. Communications in Statistics: Theory and Methods, 19(10):3859–3880, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Elfessi:1991:LSP

[EP91] Abdulaziz Elfessi and Nabendu Pal. On location and scale parameters of exponential distributions with censored observations. Communications in Statistics: Theory and Methods, 20 (5–6):1579–1592, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Elfessi:1992:ESL

[EP92a] Abdulaziz Elfessi and Nabendu Pal. Estimation of the smaller and larger of two uniform scale parameters. Communications in Statistics: Theory and Methods, 21(10):2997–3015, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Elfessi:1992:NCM

[EP92b] Abdulaziz Elfessi and Nabendu Pal. A note on the common mean of two normal populations with order restricted variances. Communications in Statistics: Theory and Methods, 21(11):3177– 3184, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Estevez-Perez:1999:NEH

[EPdR99] Graciela Estévez-Pérez and Alejandro Quintela del Río. Nonparametric estimation of the hazard function under dependence conditions. Communications in Statistics: Theory and Methods, 28(10):2294–2331, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Epps:1995:TFL

[Epp95] T. W. Epps. A test of fit for lattice distributions. Communications in Statistics: Theory and Methods, 24(6):1455–1479, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Emond:1997:BGE

[ERO97] Mary Jane Emond, John Ritz, and David Oakes. Bias in GEE estimates from misspecified models for longitudinal data. Communications in Statistics: Theory and Methods, 26(1):15–32, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Epps:1992:MLE

[ES92] T. W. Epps and Eugene Seneta. Maximum-likelihood estimation for Galton–Watson processes. Communications in Statistics: Theory and Methods, 21(3):733–748, 1992. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Evans:1994:DTI

[ES94] M. Evans and T. Swartz. Distribution theory and inference for polynomial-normal densities. Communications in Statistics: Theory and Methods, 23(4):1123–1148, 1994. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

ElDin:1995:RRE

[ES95] M. M. Mohit El Din and K. S. Sultan. Recurrence relations for expectations of functions of order statistics for doubly truncated distributions and their applications. Communications in Statistics: Theory and Methods, 24(4):997–1010, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

El-Saidi:1996:SMP

[ESDC96] Mohammed A. El-Saidi, Boyan Dimitrov, and Stefanka Chukova. Some moment properties and limit theorems of the reversed generalized logistic distribution with applications. Communications in Statistics: Theory and Methods, 25(3):609–630, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Esteban:1996:THD

[Est96] M. D. Esteban. On testing hypotheses with divergence statistics. Communications in Statistics: Theory and Methods, 25(3):501– 518, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Eubank:1993:TSO

[ESY93] Randy Eubank, Edna Schechtman, and Shlomo Yitzhaki. A test for second order stochastic dominance. Communications in Statistics: Theory and Methods, 22(7):1893–1905, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Etemadi:1991:MIP

[Ete91] N. Etemadi. Maximal inequalities for partial sums of independent random vectors with multi dimensional time parameters.

Communications in Statistics: Theory and Methods, 20(12): 3909–3923, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Evans:1997:BIP

[Eva97] Michael Evans. Bayesian inference procedures derived via the concept of relative surprise. Communications in Statistics: Theory and Methods, 26(5):1125–1143, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Falk:1998:LAN

[Fal98] Michael Falk. Local asymptotic normality of intermediate and central order statistics. Communications in Statistics: Theory and Methods, 27(7):1729–1755, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Famoye:1993:RGP

[Fam93a] Felix Famoye. Restricted generalized Poisson regression model. Communications in Statistics: Theory and Methods, 22(5):1335– 1354, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Famoye:1993:THG

[Fam93b] Felix Famoye. Testing for homogeneity: the generalized Poisson distribution. Communications in Statistics: Theory and Methods, 22(3):705–715, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Fan:1991:BAS

[Fan91a] Da-Yin Fan. Bayesian acceptance sampling scheme for pass-fail components. Communications in Statistics: Theory and Methods, 20(8):2351–2355, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Fan:1991:DPI

[Fan91b] Da-Yin Fan. The distribution of the product of independent beta variables. Communications in Statistics: Theory and Methods, 20(12):4043–4052, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Fan:1991:POS

[Fan91c] Da-Yin Fan. On a property of the order statistics of the uniform distribution. Communications in Statistics: Theory and Methods, 20(5–6):1903–1909, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Fang:1998:PIO

[Fan98] B. Q. Fang. Prior information in optimal allocation for multinomial population. Communications in Statistics: Theory and Methods, 27(1):211–222, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Fay:1999:ACI

[Fay99] Michael P. Fay. Approximate confidence intervals for rate ratios from directly standardized rates with sparse data. Communications in Statistics: Theory and Methods, 28(9):2141–2160, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Fernandez:1997:LED

[FBD97] Arturo J. Fernández, José I. Bravo, and Íñigo De Fuentes. Lifetime estimation from doubly censored data and Dirichlet process prior knowledge on the observable random vector. Communications in Statistics: Theory and Methods, 26(6):1541–1558, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Fernandez:1999:PER

[FBD99] Arturo J. Fernández, José I. Bravo, and Íñigo De Fuentes. Parametric estimation of reliability from incomplete and double censored data. Communications in Statistics: Theory and Methods, 28(10):2395–2416, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Feiveson:1993:PP

[Fei93] Alan H. Feiveson. Proof of phenomenon B. Communications in Statistics: Theory and Methods, 22(12):3321–3326, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ferentinos:1990:MVU

[Fer90] K. K. Ferentinos. Minimum variance unbiased estimation in doubly type II censored samples from families of distributions involving two truncation parameters. Communications in Statistics: Theory and Methods, 19(3):847–856, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ferreri:1996:HP

[Fer96] Carlo Ferreri. On a hyperbinomial process. Communications in Statistics: Theory and Methods, 25(1):83–103, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Finner:1990:MMD

[Fin90] Helmut Finner. On the modified s-method and directional errors. Communications in Statistics: Theory and Methods, 19 (1):41–53, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Firinguetti:1999:GRR

[Fir99] Luis Firinguetti. A generalized ridge regression estimator and its finite sample properties. Communications in Statistics: Theory and Methods, 28(5):1217–1229, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Feng:1991:NEP

[FK91] Z. Feng and K. B. Kulasekera. Nonparametric estimation of the percentile residual life function. Communications in Statistics: Theory and Methods, 20(1):87–105, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Fei:1994:IEO

[FK94] Heliang Fei and Fanhui Kong. Internal estimations for one-and two-parameter exponential distributions under multiple Type-II censoring. Communications in Statistics: Theory and Methods, 23(6):1717–1733, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Fountain:1991:EAB

[FKR91] Robert L. Fountain, Jerome P. Keating, and C. Radhakrishna Rao. An example arising from Berkson’s conjecture. Communications in Statistics: Theory and Methods, 20(11):3457–3472, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Fei:1995:ETP

[FKT95] Heliang Fei, Fanhui Kong, and Yingcai Tang. Estimation for two-parameter Weibull distribution and extreme-value distribution under multiply Type-II censoring. Communications in Statistics: Theory and Methods, 24(8):2087–2104, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ferretti:1991:EBR

[FKY91] Nelida E. Ferretti, Diana M. Kelmansky, and Victor J. Yohai. Estimators based on ranks for ARMA models. Communications in Statistics: Theory and Methods, 20(12):3879–3907, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ferretti:1995:GFT

[FKY95] Nélida Ferretti, Diana Kelmansky, and Victor Yohai. A goodness-of-fit test based on ranks for ARMA models. Communications in Statistics: Theory and Methods, 24(2):295–318, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Frey:1992:AMN

[FM92] Carolin Malott Frey and Keith E. Muller. Analysis methods for nonlinear models with compound-symmetric covariance. Communications in Statistics: Theory and Methods, 21(5):1163– 1182, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Fernandez:1996:BTG

[FM96] J. M. Vilar Fernandez and W. Gonzalez Manteiga. Bootstrap test of goodness of fit to a linear model when errors are correlated. Communications in Statistics: Theory and Methods, 25(12):2925–2953, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Franses:1997:TNN

[FM97] Philip Hans Franses and Michael McAleer. Testing nested and non-nested periodically integrated autoregressive models. Communications in Statistics: Theory and Methods, 26(6):1461– 1475, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Frank:1998:ADW

[FMP98] O. Frank, M. L. Menendez, and L. Pardo. Asymptotic distributions of weighted divergence between discrete distributions.

Communications in Statistics: Theory and Methods, 27(4):867– 885, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Fukushima:1998:DQF

[FMV98] Júlia T. Fukushima, Regina C. C. P. Moran, and Ioannis G. Vlachonikolis. On the distribution of quadratic forms and applications in the analysis of repeated measurements. Communications in Statistics: Theory and Methods, 27(11):2625–2640, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Fearn:1991:MLR

[FN91] D. H. Fearn and E. Nebenzahl. On the maximum likelihood ratio method of deciding between the Weibull and gamma distributions. Communications in Statistics: Theory and Methods, 20 (2):579–593, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Fearn:1992:TEA

[FN92] Dean H. Fearn and Elliott Nebenzahl. Tests for exponentiality against increasing failure rate average alternatives with censored data. Communications in Statistics: Theory and Methods, 21(9): 2557–2567, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Fearn:1995:UPT

[FN95] Dean H. Fearn and Elliott Nebenzahl. Using power transformations when approximating quantiles. Communications in Statistics: Theory and Methods, 24(4):1073–1093, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Follmann:1992:GCM

[Fol92] Dean A. Follmann. Growth curve models with restrictions on random parameters. Communications in Statistics: Theory and Methods, 21(10):2775–2795, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Fountain:1991:PCC

[Fou91] Robert L. Fountain. Pitman closeness comparison of linear estimators: a canonical form. Communications in Statistics: Theory and Methods, 20(11):3535–3550, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Fountain:1990:SNS

[FP90] Robert L. Fountain and Pramod K. Pathak. Systematic and nonrandom sampling in the presence of polynomial trends. Communications in Statistics: Theory and Methods, 19(3):1061–1071, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Feuerverger:1992:SAP

[FR92] Andrey Feuerverger and Sheikh Rahman. Some aspects of probability forecasting. Communications in Statistics: Theory and Methods, 21(6):1615–1632, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Finner:1993:DFL

[FR93a] H. Finner and M. Roters. Distribution functions and logconcavity. Communications in Statistics: Theory and Methods, 22(8):2381–2396, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Fountain:1993:FIB

[FR93b] Robert L. Fountain and C. Radhakrishna Rao. Further investigations of Berkson’s example. Communications in Statistics: Theory and Methods, 22(3):613–629, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Frick:1998:PNS

[FR98] H. Frick and V. W. Rahlfs. On power and necessary sample sizes of the Wilcoxon–Mann–Whitney test. Communications in Statistics: Theory and Methods, 27(10):2445–2460, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Franck:1991:NMA

[Fra91] Wallace Franck. A new measure of association. Communications in Statistics: Theory and Methods, 20(9):3051–3070, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Freels:1992:ESD

[Fre92] Sally Freels, Ph.D. Eliminating a sparse data problem in nonparametric bivariate survival analysis with an application to women’s age at first marriage and age at first birth. Communications in Statistics: Theory and Methods, 21(2):295–318, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Frick:1990:PHL

[Fri90a] Hans Frick. On the procedure of Hodges and Lehmann for testing composite hypotheses of the mean of a normal distribution with unknown variance. Communications in Statistics: Theory and Methods, 19(2):665–687, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Frick:1990:UHL

[Fri90b] Hans Frick. The unbiasedness of Hodges and Lehman’s test for a composite hypothesis of the mean of a distribution with unknown variance. Communications in Statistics: Theory and Methods, 19(8):2799–2814, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Frick:1991:LFS

[Fri91] Hans Frick. On Lachin’s formulae for sample sizes of survival tests. Communications in Statistics: Theory and Methods, 20(7): 2267–2280, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Friedman:1992:NEM

[Fri92] Lea Friedman. A note on estimation in a minimum type model. Communications in Statistics: Theory and Methods, 21(3):749– 763, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Frick:1994:MLT

[Fri94] Hans Frick. A maxmin linear test of normal means and its application to Lachin’s data. Communications in Statistics: Theory and Methods, 23(4):1021–1029, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

From:1990:NEC

[Fro90] Steven G. From. Nonparametric estimation of a change point in positive shift models. Communications in Statistics: Theory and Methods, 19(9):3163–3173, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

From:1993:SRR

[Fro93] Steven G. From. Some recurrence relation-based estimators of the parameters of a generalized log-series distribution. Communications in Statistics: Theory and Methods, 22(6):1653–1664, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

From:1996:NGF

[Fro96] Steven G. From. A new goodness of fit test for the equality of multinomial cell probabilities versus trend alternatives. Communications in Statistics: Theory and Methods, 25(12):3167–3183, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Fernandez:1997:MLE

[FRS97] M. A. Fernández, C. Rueda, and B. Salvador. On the maximum likelihood estimator under order restrictions in uniform probability models. Communications in Statistics: Theory and Methods, 26(8):1971–1980, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Fawcett:1993:SAA

[FS93] Richard Fawcett and Penny Stewart. Sweeping, alignment and the analysis of unbalanced data. Communications in Statistics: Theory and Methods, 22(12):3453–3471, 1993. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Frangos:1995:IBC

[FS95a] Chris C. Frangos and William R. Schucany. Improved bootstrap confidence intervals in certain toxicological experiments. Communications in Statistics: Theory and Methods, 24(3):829–844, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Fu:1995:GBD

[FS95b] James Fu and Robert Sproule. A generalization of the binomial distribution. Communications in Statistics: Theory and Methods, 24(10):2645–2658, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Fujikoshi:1999:AEJ

[FS99] Yasunori Fujikoshi and Takashi Seo. Asymptotic expansions for the joint distribution of correlated Hotelling’s T^{2} statistics under normality. Communications in Statistics: Theory and Methods, 28(3–4):773–788, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Friedl:1995:VEL

[FT95] Herwig Friedl and Norbert Tilg. Variance estimates in logistic regression using the bootstrap. Communications in Statistics: Theory and Methods, 24(2):473–486, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Fung:1997:ALI

[FT97] Wing K. Fung and M. K. Tang. Assessment of local influence in multivariate regression analysis. Communications in Statistics: Theory and Methods, 26(4):821–837, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Fan:1999:BMD

[FT99] Tsai-Hung Fan and Chen-An Tsai. A Bayesian method in determining the order of a finite state Markov chain. Communications in Statistics: Theory and Methods, 28(7):1711–1730, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Fellingham:1994:LEM

[FTS94] G. W. Fellingham, H. D. Tolley, and D. T. Scott. Likelihood estimation of missing cell means in the fixed model analysis of variance. Communications in Statistics: Theory and Methods, 23(9):2429–2447, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Finkelstein:1999:CIE

[FTV99] Mark Finkelstein, Howard G. Tucker, and Jerry Alan Veeh. Confidence intervals for n in the exponential order statistics problem. Communications in Statistics: Theory and Methods, 28(6): 1415–1433, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Fujino:1990:EVC

[Fuj90] Yoritake Fujino. Estimating the value of control variable x with Pr{Y > c|x} = p in normal linear regression. Communications in Statistics: Theory and Methods, 19(1):1–9, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Fujii:1994:IUC

[Fuj94] Yoshinori Fujii. On the influence of unknown covariate on the binary data. Communications in Statistics: Theory and Methods, 23(10):2831–2840, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Fujisawa:1995:NML

[Fuj95] Hironori Fujisawa. A note on the maximum likelihood estimators for multivariate normal distribution with monotone data. Communications in Statistics: Theory and Methods, 24(6):1377– 1382, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Fujikoshi:1999:SCI

[Fuj99a] Yasunori Fujikoshi. Simultaneous confidence intervals in an extended Crow III curve model. Communications in Statistics: Theory and Methods, 28(3–4):671–682, 1999. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Fujisawa:1999:EUD

[Fuj99b] H. Fujisawa. Effects of unpaired data for estimating an interclass correlation. Communications in Statistics: Theory and Methods, 28(1):245–254, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Funo:1991:IRM

[Fun91] Eiichiro Funo. Inadmissibility results of the MLE for the multinomial problem when the parameter space is restricted or truncated. Communications in Statistics: Theory and Methods, 20(9):2863–2880, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Fung:1995:GSI

[Fun95] Wing K. Fung. Graphical summaries for influence of multiple observations. Communications in Statistics: Theory and Methods, 24(2):415–427, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Fung:1996:IOL

[Fun96a] Wing K. Fung. The influence of observations for local log-odd in linear discriminant analysis. Communications in Statistics: Theory and Methods, 25(2):257–268, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Fung:1996:IOM

[Fun96b] Wing K. Fung. The influence of observations on misclassification probability in multiple discriminant analysis. Communications in Statistics: Theory and Methods, 25(8):1917–1930, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Fung:1998:ETD

[Fun98] Wing K. Fung. On the equivalence of two diagnostic measures in discriminant analysis. Communications in Statistics: Theory and Methods, 27(8):1915–1922, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Follmann:1994:TTE

[FWG94] Dean Follmann, Margaret Wu, and Nancy L. Geller. Testing treatment efficacy in clinical trials with repeated binary measurements and missing observations. Communications in Statistics: Theory and Methods, 23(2):557–574, 1994. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Fong:1995:NIL

[FY95] Daniel Y. T. Fong and Paul S. F. Yip. A note on information loss in analyzing a mixture model of count data. Communications in Statistics: Theory and Methods, 24(12):3197–3209, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Fygenson:1991:OSE

[Fyg91] Mendel Fygenson. Optimal sequential estimation for semi-Markov and Markov renewal processes. Communications in Statistics: Theory and Methods, 20(4):1427–1444, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Fellag:1996:BLE

[FZ96] Hocine Fellag and Ryszard Zieliński. Bias of the LSE estimator of the first order autoregressive model under Tukey contamination. Communications in Statistics: Theory and Methods, 25(7):1537– 1551, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Feng:1997:SRM

[FZ97] Shiyong Feng and Guohua Zou. Sample rotation method with auxiliary variable. Communications in Statistics: Theory and Methods, 26(6):1497–1509, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gabr:1992:REB

[Gab92] M. M. Gabr. Recursive estimation of bilinear time series models. Communications in Statistics: Theory and Methods, 21(8):2261– 2277, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gabr:1998:REB

[Gab98] M. M. Gabr. Robust estimation of bilinear time series models. Communications in Statistics: Theory and Methods, 27(1):41– 53, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Galecki:1994:GCC

[Gal94] Andrzej T. Galecki. General class of covariance structures for two or more repeated factors in longitudinal data analysis. Communications in Statistics: Theory and Methods, 23(11):3105– 3119, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gharraf:1990:BGS

[GAN90] Mohamed K. Gharraf and Sumaia A. Al-Nasser. A Bayesian group sequential analysis procedure. Communications in Statistics: Theory and Methods, 19(3):977–985, 1990. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gao:1994:FSS

[Gao94] Feng Gao. Fitting smoothing splines to data from multiple sources. Communications in Statistics: Theory and Methods, 23(6):1665–1698, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gao:1995:APS

[Gao95a] Jiti Gao. Asymptotic properties of some estimators for partly linear stationary autoregressive models. Communications in Statistics: Theory and Methods, 24(8):2011–2026, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gao:1995:ATP

[Gao95b] Jiti Gao. Asymptotic theory for partly linear models. Communications in Statistics: Theory and Methods, 24(8):1985–2009, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gao:1997:APT

[Gao97] Jiti Gao. Adaptive parametric test in a semiparametric regression model. Communications in Statistics: Theory and Methods, 26(4):787–800, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Garthwaite:1992:PEL

[Gar92] Paul H. Garthwaite. Preposterior expected loss as a scoring rule for prior distributions. Communications in Statistics: Theory and Methods, 21(12):3601–3619, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Garfield:1996:ASL

[Gar96] Joan B. Garfield. Assessing student learning in the context of evaluating a chance course. Communications in Statistics: Theory and Methods, 25(11):2863–2873, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gastaldi:1992:TCH

[Gas92] Tommaso Gastaldi. Testing a censoring hypothesis through an identity test. Communications in Statistics: Theory and Methods, 21(5):1267–1272, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gaudoin:1998:CGF

[Gau98] O. Gaudoin. CPIT goodness-of-fit tests for the power-law process. Communications in Statistics: Theory and Methods, 27 (1):165–180, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gadag:1998:LRT

[GB98] Veeresh Gadag and Uditha Balasooriya. A likelihood ratio test for a shift in the lifelength distribution. Communications in Statistics: Theory and Methods, 27(8):1937–1945, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gimenez:1999:EWR

[GBC99] Patricia Gimenez, Heleno Bolfarine, and Enrico A. Colosimo. Estimation in Weibull regression model with measurement error. Communications in Statistics: Theory and Methods, 28(2):495– 510, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Guerrero-Cusumano:1998:MDM

[GC98a] José-Luis Guerrero-Cusumano. Measures of dependence for the multivariate t distribution with applications to the stock market. Communications in Statistics: Theory and Methods, 27(12): 2985–3006, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gupta:1998:ORC

[GC98b] Sudhir Gupta and Kuey Chung Choi. Optimal row-column designs for complete dial lei crosses. Communications in Statistics: Theory and Methods, 27(11):2827–2835, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Giraudeau:1999:LER

[GC99] Bruno Giraudeau and Claude Chastang. Letter to the editor: Remarks on Bhandary and Bansal’s test for detection of outliers in familial data. Communications in Statistics: Theory and Methods, 28(8):2021–2024, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Giron:1994:CFP

[GCD94] F. J. Girón, E. Caro, and J. I. Domínguez. A conjugate family for AR(1) processes with exponential errors. Communications in Statistics: Theory and Methods, 23(6):1771–1784, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gill:1999:OST

[GD99] Amar Nath Gill and Anil Kumar Dhawan. A one-sided test for testing homogeneity of scale parameters against simple ordered alternative. Communications in Statistics: Theory and Methods, 28(10):2417–2439, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gupta:1995:SRO

[GDK95] Sudhir Gupta, Ashish Das, and Sanpei Kageyama. Single replicate orthogonal block designs for circulant partial diallel crosses. Communications in Statistics: Theory and Methods, 24(10): 2601–2607, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Garisch:1991:UUE

[GdWG91] I. Garisch, D. J. de Waal, and P. C. N. Groenewald. The use of utilities and experts in decision-making. Communications in Statistics: Theory and Methods, 20(1):333–344, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Genshiro:1997:ICP

[Gen97] Kitagawa Genshiro. Information criteria for the predictive evaluation of Bayesian models. Communications in Statistics: Theory and Methods, 26(9):2223–2246, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Georg:1991:SLN

[Geo91] Neuhaus Georg. Some linear and nonlinear rank tests for competing risks models. Communications in Statistics: Theory and Methods, 20(2):667–701, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gross:1991:CLS

[GF91] Shulamith T. Gross and Martin R. Frankel. Confidence limits for small proportions in complex samples. Communications in Statistics: Theory and Methods, 20(3):951–975, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gastaldi:1992:KST

[GG92a] Tommaso Gastaldi and Tommaso Gastaldi. A Kolmogorov– Smirnov test procedure involving a possibly censored or truncated sample. Communications in Statistics: Theory and Methods, 22(1):31–39, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Godbole:1992:EAR

[GG92b] Anant P. Godbole and Michelle C. Gornowicz. Exact and approximate runs distributions. Communications in Statistics: Theory and Methods, 21(8):2151–2167, 1992. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Giles:1993:PTE

[GG93] Judith A. Giles and David E. A. Giles. Preliminary-test estimation of the regression scale parameter when the loss function is asymmetric. Communications in Statistics: Theory and Methods, 22(6):1709–1733, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gastaldi:1994:MTP

[GG94a] Tommaso Gastaldi and Shanti S. Gupta. Minimax type procedures for nonparametric selection of the “best” population with partially classified data. Communications in Statistics: Theory and Methods, 23(9):2503–2532, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gibbons:1994:BRC

[GG94b] Patrick C. Gibbons and Edward Greenberg. Bayesian reconstruction of contingency tables with partially categorized data. Communications in Statistics: Theory and Methods, 23(12): 3349–3359, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ghorai:1996:NEM

[GG96] J. K. Ghorai and Sanjoy Ghose. On the non-existence of ML estimates in the negative multinomial distributions derived from nonstationary Poisson processes. Communications in Statistics: Theory and Methods, 25(9):2117–2125, 1996. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gupta:1998:MFT

[GGG98] Ramesh C. Gupta, Pushpa L. Gupta, and Rameshwar D. Gupta. Modeling failure time data by Lehman alternatives. Communications in Statistics: Theory and Methods, 27(4):887–904, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gupta:1995:ERU

[GGHA95] Ramesh C. Gupta, Rameshwar D. Gupta, and Olcay H-Akman. Estimation of reliability under ordered restriction on the parameters. Communications in Statistics: Theory and Methods, 24(7): 1799–1812, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gupta:1996:AFT

[GGL96] Pushpa L. Gupta, Ramesh C. Gupta, and Sergey J. Lvin. Analysis of failure time data by Burr distribution. Communications in Statistics: Theory and Methods, 25(9):2013–2024, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Giles:1996:ERS

[GGO96] Judith A. Giles, David E. A. Giles, and Kazuhiro Ohtani. The exact risks of some pre-test and Stein-type regression estimators under balanced loss. Communications in Statistics: Theory and Methods, 25(12):2901–2924, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gupta:1995:IMP

[GGT95] Pushpa L. Gupta, Ramesh C. Gupta, and Ram C. Tripathi. Inflated modified power series distributions with applications. Communications in Statistics: Theory and Methods, 24(9):2355– 2374, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gomez:1998:MGP

[GGVM98] E. Gómez, M. A. Gomez-Viilegas, and J. M. Marín. A multivariate generalization of the power exponential family of distributions. Communications in Statistics: Theory and Methods, 27 (3):589–600, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Galambos:1992:CDF

[GH92] Janos Galambos and Charles Hagwood. The characterization of a distribution function by the second moment of the residual life. Communications in Statistics: Theory and Methods, 21(5): 1463–1468, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ghitany:1998:RGG

[Ghi98] M. E. Ghitany. On a recent generalization of gamma distribution. Communications in Statistics: Theory and Methods, 27 (1):223–233, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ghorai:1991:CMS

[Gho91] J. K. Ghorai. Cramér–von Mises statistic for testing goodness of fit under the proportional hazards model. Communications in Statistics: Theory and Methods, 20(3):1107–1126, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Giles:1991:PTM

[Gil91] Judith A. Giles. Pre-testing in a mis-specified regression model. Communications in Statistics: Theory and Methods, 20(10): 3221–3238, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gill:1992:MHG

[Gil92] Jeffrey I. Mc Gill. The multivariate hazard gradient and moments of the truncated multinormal distribution. Communications in Statistics: Theory and Methods, 21(11):3053–3060, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Giles:1993:ESP

[Gil93] Judith A. Giles. Estimation of the scale parameter after a pre-test for homogeneity in a mis-specified regression model. Communications in Statistics: Theory and Methods, 22(4):1007– 1029, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

George:1992:ERM

[GJ92a] Varghese T. George and William D. Johnson. Effect of regression to the mean in multivariate distributions. Communications in Statistics: Theory and Methods, 21(2):333–350, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ghosh:1992:ADE

[GJ92b] J. K. Ghosh and S. N. Joshi. On the asymptotic distribution of an estimate of the change point in a failure rate. Communications in Statistics: Theory and Methods, 21(12):3571–3588, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gonzalez:1997:OCP

[GJ97] Camino Gonzalez and Jesus Juan. Obtaining the confounding pattern for 2^{k−p} experiments. Communications in Statistics: Theory and Methods, 26(11):2769–2777, 1997. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ghosh:1996:ABA

[GJM96] Jayanta K. Ghosh, Shrikant N. Joshi, and Ohiranjit Mukhopadhyay. Asymptotics of a Bayesian approach to estimating changepoint in a hazard rate. Communications in Statistics: Theory and Methods, 25(12):3147–3166, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gaver:1990:RAH

[GJO90] D. P. Gaver, P. A. Jacobs, and I. G. O’Muircheartaigh. Regression analysis of hierarchical Poisson-like event rate data: superpopulation model effect on predictions. Communications in Statistics: Theory and Methods, 19(10):3779–3797, 1990. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gomez:1994:APL

[GJU94] Guadalupe Gómez, Olga Julià, and Frederic Utzet. Asymptotic properties of the left Kaplan–Meier estimator. Communications in Statistics: Theory and Methods, 23(1):123–135, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gupta:1990:RWD

[GK90] Ramesh C. Gupta and S. N. U. A. Kirmani. The role of weighted distributions in stochastic modeling. Communications in Statistics: Theory and Methods, 19(9):3147–3162, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ghosh:1991:BPC

[GK91] Malay Ghosh and Sen Pranab Kumar. Bayesian Pitman closeness. Communications in Statistics: Theory and Methods, 20 (11):3659–3678, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Govlndaraju:1992:NOC

[GK92] K. Govlndaraju and V. Kuralmani. A note on the operating characteristic curve of the known sigma single sampling variables plan. Communications in Statistics: Theory and Methods, 21(8): 2339–2347, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gupta:1998:HCS

[GK98] Rameshwar D. Gupta and Debasis Kundu. Hybrid censoring schemes with exponential failure distribution. Communications in Statistics: Theory and Methods, 27(12):3065–3083, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gupta:1990:SFS

[GL90] Shanti S. Gupta and Lii-Yuh Leu. Selecting the fairest of K(≥ 2) M-sided dice. Communications in Statistics: Theory and Methods, 19(6):2159–2177, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Giles:1993:BEH

[GL93] David Giles and Offer Lieberman. Bounds on the effect of heteroscedasticity on the chow test for structural change. Communications in Statistics: Theory and Methods, 22(3):687–703, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gupta:1994:EDR

[GL94] Rajan Gupta and Larry Lee. Estimating the discovery rate in a continuous time recapture model. Communications in Statistics: Theory and Methods, 23(3):871–881, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ganju:1997:BTS

[GL97] Jitendra Ganju and James M. Lucas. Bias in test statistics when restrictions in randomization are caused by factors. Communications in Statistics: Theory and Methods, 26(1):47–63, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Glendinning:1995:MSI

[Gle95] R. H. Glendinning. Model selection for infinite variance time series. Communications in Statistics: Theory and Methods, 24 (4):889–910, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Glendinning:1996:IIV

[Gle96] R. H. Glendinning. Identifying infinite variance ARMA models using a robust Pukkila–Koreisha–Kallinen strategy. Communications in Statistics: Theory and Methods, 25(12):3027–3047, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Glendinning:1998:DOA

[Gle98] R. H. Glendinning. Determining the order of an ARMA model from outlier contaminated data. Communications in Statistics: Theory and Methods, 27(1):13–40, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gupta:1990:LCB

[GLL90] Shanti S. Gupta, Lii-Yuh Leu, and TaChen Liang. On lower confidence bounds for PCS in truncated location parameter models. Communications in Statistics: Theory and Methods, 19(2): 527–546, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gijbels:1990:EDC

[GM90] Iène Gijbels and Jan Mielniczuk. Estimating the density of a copula function. Communications in Statistics: Theory and Methods, 19(2):445–464, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Genizi:1994:SMC

[GM94a] Abraham Genizi and Ruth Marcus. Some multiple comparison procedures for comparing treatments within several groups. Communications in Statistics: Theory and Methods, 23(3):763– 779, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Goel:1994:SAC

[GM94b] L. R. Goel and S. Z. Mumtaz. Stochastic analysis of a complex system with an auxiliary unit. Communications in Statistics:

Theory and Methods, 23(10):3003–3017, 1994. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gupta:1996:TEC

[GM96] Ramesh C. Gupta and Suchun Ma. Testing the equality of coefficients of variation in k normal populations. Communications in Statistics: Theory and Methods, 25(1):115–132, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Glueck:1998:TW

[GM98a] Deborah H. Glueck and Keith E. Muller. On the trace of a Wishart. Communications in Statistics: Theory and Methods, 27(9):2137–2141, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Guglielmi:1998:NII

[GM98b] Alessandra Guglielmi and Eugenio Melilli. Non-informative invariant priors yield peculiar marginals. Communications in Statistics: Theory and Methods, 27(9):2293–2306, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Geys:1997:PLI

[GMR97] Helena Geys, Geert Molenberghs, and Louise M. Ryan. Pseudolikelihood inference for clustered binary data. Communications in Statistics: Theory and Methods, 26(11):2743–2767, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Govindarajulu:1991:RCD

[GN91] Z. Govindarajulu and A. Nanthakumar. Rate of convergence for the distribution of the Spearman–Karber estimator. Communications in Statistics: Theory and Methods, 20(8):2577–2588, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gupta:1995:TSM

[GN95] A. K. Gupta and D. K. Nagar. On testing structure of mean vector of compound symmetric Gaussian model. Communications in Statistics: Theory and Methods, 24(12):3123–3136, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Guillamon:1998:KDE

[GNR98] A. Guillamón, J. Navarro, and J. M. Ruiz. Kernel density estimation using weighted data. Communications in Statistics: Theory and Methods, 27(9):2123–2135, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Godbole:1990:HRD

[God90] Anant P. Godbole. On hypergeometric and related distributions of order k. Communications in Statistics: Theory and Methods, 19(4):1291–1301, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Godbole:1992:EAD

[God92] Anant P. Godbole. The exact and asymptotic distribution of overlapping success runs. Communications in Statistics: Theory and Methods, 21(4):953–967, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Goldstein:1990:PAD

[Gol90] Larry Goldstein. Poisson approximation and DNA sequence matching. Communications in Statistics: Theory and Methods, 19(11):4167–4179, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Good:1991:SUD

[Goo91] I. J. Good. Serial unblocked designs balanced for correlation between neighbours, II. Communications in Statistics: Theory and Methods, 20(4):1225–1231, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Goudie:1995:PCA

[Gou95] I. B. J. Goudie. A plant-capture approach for achieving complete coverage of a population. Communications in Statistics: Theory and Methods, 24(5):1293–1305, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Govindaraju:1990:COL

[Gov90a] K. Govindaraju. Certain observations on lot sensitive sampling plan. Communications in Statistics: Theory and Methods, 19 (2):617–627, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Govindaraju:1990:SCC

[Gov90b] K. Govindaraju. Selection of ChSP-1 chain sampling plans for given acceptable quality level and limiting quality level. Communications in Statistics: Theory and Methods, 19(6):2179–2190, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Govindarajulu:1995:BSL

[Gov95] Z. Govindarajulu. A brief survey of LMP rank tests for random effects in one and two-way models. Communications in Statistics: Theory and Methods, 24(1):59–73, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Govindaraju:1997:CSS

[Gov97] K. Govindaraju. Conditional single sampling procedure. Communications in Statistics: Theory and Methods, 26(5):1215– 1225, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ghorai:1990:CKD

[GP90] J. K. Ghorai and L. M. Pattanaik. L_{1}-consistency of the kernel density estimators based on randomly right censored data. Communications in Statistics: Theory and Methods, 19(8):2853– 2870, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gosuami:1992:COF

[GP92] Kalyan Kumar Gosuami and Satyabrata Pal. On the construction of orthogonal factorial designs of resolution IV. Communications in Statistics: Theory and Methods, 21(12):3561–3570, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ghorai:1993:AOB

[GP93] J. K. Ghorai and L. M. Pattanaik. Asymptotically optimal bandwidth selection of the kernel density estimator under the proportional hazards model. Communications in Statistics: Theory and Methods, 22(5):1383–1401, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gulati:1994:SNE

[GP94] S. Gulati and W. J. Padgett. Smooth nonparametric estimation of the distribution and density functions from record-breaking data. Communications in Statistics: Theory and Methods, 23(5): 1259–1274, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gupta:1996:EBD

[GP96] V. K. Gupta and Rajender Parsad. A-efficient block designs for comparing two disjoint sets of treatments. Communications in Statistics: Theory and Methods, 25(5):967–983, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gather:1998:EGV

[GP98] Ursula Gather and Jörg Pawlitschko. On Efron’s and Gill’s version of the Kaplan–Meier integral. Communications in Statistics: Theory and Methods, 27(1):181–192, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Goudie:1998:PCA

[GPA98] I. B. J. Goudie, Kenneth H. Pollock, and J. Ashbridge. A plantcapture approach for population size estimation in continuous time. Communications in Statistics: Theory and Methods, 27 (2):433–451, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gao:1998:LLK

[GPW98] Jiti Gao, Anthony N. Pettitt, and Rodney C. L. Wolff. Local linear kernel estimation for discontinuous nonparametric regression functions. Communications in Statistics: Theory and Methods, 27(12):2871–2894, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ghorai:1990:RDK

[GR90] J. K. Ghorai and Lidia Rejtö. Relative deficiency of kernel type estimators of quantiles based on right censored data. Communications in Statistics: Theory and Methods, 19(5):1653–1670, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gray:1998:RAI

[Gra98] Robert J. Gray. On regression analysis with incomplete covariate data. Communications in Statistics: Theory and Methods, 27(11):2641–2662, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Grimshaw:1990:UTE

[Gri90] Scott D. Grimshaw. A unification of tail estimators. Communications in Statistics: Theory and Methods, 19(12):4841–4857, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Grimshaw:1992:ELP

[Gri92] Scott D. Grimshaw. Estimation of the linear-plateau segmented regression model in the presence of measurement error. Communications in Statistics: Theory and Methods, 21(8):2399–2413, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Griffiths:1997:BPE

[Gri97] William E. Griffiths. Bayesian predictors for an AR(1) error model. Communications in Statistics: Theory and Methods, 26 (2):441–448, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gross:1991:BCL

[Gro91] Shulamith T. Gross. Bayesian credibility limits for small finite population proportions: stratified and cluster sampling. Communications in Statistics: Theory and Methods, 20(3):1027– 1037, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gross:1994:TEP

[Gro94] Alan J. Gross. Three estimators of the probability of opiate use from incomplete data. Communications in Statistics: Theory and Methods, 23(2):575–588, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gross:1996:CEG

[Gro96] Jürgen Groß. On a class of estimators in the general Gauss– Markov model. Communications in Statistics: Theory and Methods, 25(2):381–388, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gross:1999:LIB

[Gro99] Jürgen Groß. Local improvement of best linear unbiased estimation and admissibility under the weakly singular Gauss–Markov model. Communications in Statistics: Theory and Methods, 28(8):1803–1812, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gruben:1993:MLE

[Gru93] M. S. David Gruben. Maximum likelihood estimation of the Poisson distribution of r × c independent samples in the presence of linear constraints on the vector of parameters. Communications in Statistics: Theory and Methods, 22(12):3473–3490, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ghosh:1990:CBM

[GS90] Malay Ghosh and Bikas Kumar Sinha. On the consistency between model-and design-based estimators in survey sampling. Communications in Statistics: Theory and Methods, 19(2):689– 702, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gelfand:1991:GSM

[GS91a] Alan E. Gelfand and Adrian F. M. Smith. Gibbs sampling for marginal posterior expectations. Communications in Statistics: Theory and Methods, 20(5–6):1747–1766, 1991. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gupta:1991:TDB

[GS91b] S. C. Gupta and O. P. Srivastava. A two-dimensional branching process with migration. Communications in Statistics: Theory and Methods, 20(8):2727–2736, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gould:1992:SSR

[GS92] A. Lawrence Gould and Weichung Joseph Shih. Sample size re-estimation without unblinding for normally distributed outcomes with unknown variance. Communications in Statistics: Theory and Methods, 21(10):2833–2853, 1992. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gillespie:1993:SOI

[GS93] John A. Gillespie and R. Srinivasan. Simultaneous outer and inner tolerance limits for screening and descriptive purposes. Communications in Statistics: Theory and Methods, 22(7):2057– 2066, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gillespie:1994:DFS

[GS94] Jhon A. Gillespie and R. Srinivasan. Distribution-free simultaneous outer and inner tolerance limits for screening and descriptive purposes. Communications in Statistics: Theory and Methods, 23(5):1357–1364, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gersch:1995:MAT

[GS95] Will Gersch and David Stone. Multivariate autoregressive time series modeling: one scalar autoregressive model at-a-time. Communications in Statistics: Theory and Methods, 24(11): 2715–2733, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Garcia-Soidan:1998:EET

[GS98a] Pilar H. García-Soidán. Edgeworth expansions for triangular arrays. Communications in Statistics: Theory and Methods, 27 (3):705–722, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gupta:1998:BED

[GS98b] R. P. Gupta and P. G. Sankaran. Bivariate equilibrium distribution and its applications to reliability. Communications in Statistics: Theory and Methods, 27(2):385–394, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gielens:1996:FTD

[GSD96] Geert Gielens, Stefan Straetmans, and Casper G. De Vries. Fat tail distributions and local thin tail alternatives. Communications in Statistics: Theory and Methods, 25(4):705–710, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gupta:1990:ELB

[GT90] Pushpa Lata Gupta and Ram C. Tripathi. Effect of length-biased sampling on the modeling error. Communications in Statistics: Theory and Methods, 19(4):1483–1491, 1990. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ghosh:1991:CNA

[GT91] D. K. Ghosh and Alex Thannippara. Construction on nonadaptive hypergeometric group testing designs for identifying at most two defectives. Communications in Statistics: Theory and Methods, 20(4):1257–1272, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gupta:1992:SIB

[GT92] Ramesh C. Gupta and Ram C. Tripathi. Statistical inference based on the length-biased data for the modified power series distributions. Communications in Statistics: Theory and Methods, 21(2):519–537, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gupta:1996:WBL

[GT96] Ramesh C. Gupta and Ram C. Tripathti. Weighted bivariate logarithmic series distributions. Communications in Statistics: Theory and Methods, 25(5):1099–1117, 1996. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gross:1997:EUA

[GT97] Jürgen Groß and Götz Trenkler. On the equality of usual and Amemiya’s partially generalized least squares estimator. Communications in Statistics: Theory and Methods, 26(9):2075– 2086, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Guerrero:1994:MMI

[Gue94] José-Luis Guerrero. Multivariate mutual information sampling distribution with applications. Communications in Statistics: Theory and Methods, 23(5):1319–1339, 1994. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Guo:1996:DFT

[Guo96] Jiin-Huarng Guo. Distribution-free test for slope in linear regression model with replications. Communications in Statistics:

Theory and Methods, 25(9):2069–2078, 1996. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gupta:1991:SRM

[Gup91] B. C. Gupta. Some results on main effect plus one plans for 2^{m} factorials. Communications in Statistics: Theory and Methods, 20(9):2955–2963, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gupta:1992:EME

[Gup92a] B. C. Gupta. On the existence of main effect plus k plans for 2^{m} factorials and tables for main effect plus 1 and 2 plans for 2^{7} factorials. Communications in Statistics: Theory and Methods, 21(4):1137–1143, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gupta:1992:ELM

[Gup92b] Sat N. Gupta. Estimation in long memory time series models. Communications in Statistics: Theory and Methods, 21(5):1327– 1338, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gupta:1999:GAP

[Gup99] Somesh Das Gupta. Gini association and pseudo Lorenz curve. Communications in Statistics: Theory and Methods, 28(9):2181– 2199, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Guven:1995:MLE

[Güv95] Bilgehan Güven. Maximum likelihood estimation in simple linear regression with one fold nested error. Communications in Statistics: Theory and Methods, 24(1):121–130, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gupta:1990:CCJ

[GV90] A. K. Gupta and T. Varga. Characterization of characterization of joint density by conditional densities. Communications in Statistics: Theory and Methods, 19(12):4643–4652, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gomez-Villegas:1992:BFT

[GVSM92] M. A. Gomez-Villegas and E. Gomez Sanchez-Manzano. Bayes factor in testing precise hypotheses. Communications in Statistics: Theory and Methods, 21(6):1707–1715, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gertsbakh:1991:PIE

[GW91] I. Gertsbakh and A. Winterbottom. Point and interval estimation of normal tail probabilities. Communications in Statistics: Theory and Methods, 20(4):1497–1514, 1991. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Geran:1992:MSS

[GW92] J. H. O’ Geran and H. P. Wynn. Multi-stage step-wise group screening. Communications in Statistics: Theory and Methods, 21(11):3291–3308, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gillespie:1994:NTS

[GW94] Brenda W. Gillespie and Douglas A. Wolfe. Nonparametric two-sample tests for dispersion differences based on placements. Communications in Statistics: Theory and Methods, 23(11): 3085–3103, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Grzegorzewski:1999:EBG

[GW99a] Przeniysław Grzegorzewski and Robert Wieczorkowski. Entropybased goodness-of-fit test for exponentiality. Communications in Statistics: Theory and Methods, 28(5):1183–1202, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Gupta:1999:STS

[GW99b] Arjun K. Gupta and Yining Wang. Some tests with specified size for the Behrens–Fisher problem. Communications in Statistics: Theory and Methods, 28(3–4):511–517, 1999. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wu:1997:EUM

[gWyJ97] Qi guang Wu and Zhen yu Jiang. The existence of the uniformly minimum risk equivariant estimator in sure model. Communications in Statistics: Theory and Methods, 26(1):113–128, 1997.

CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Garthwaite:1995:BAM

[GYH95] Paul H. Garthwaite, Keming Yu, and Peter B. Hope. Bayesian analysis of a multiple-recapture model. Communications in Statistics: Theory and Methods, 24(9):2229–2247, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ghitany:1997:ARE

[GZ97] M. E. Ghitany and S. Zhou. Asymptotic results for exponential failure models with imperfect signalling and covariates. Communications in Statistics: Theory and Methods, 26(4):801–820, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hadjicostas:1998:IPP

[Had98] Petors Hadjicostas. Improper and proper posteriors with improper priors in a hierarchical model with a beta-binomial likelihood. Communications in Statistics: Theory and Methods, 27(8):1905–1914, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Haines:1993:ODN

[Hai93] Linda M. Haines. Optimal design for nonlinear regression models. Communications in Statistics: Theory and Methods, 22(6): 1613–1627, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hajime:1990:PAS

[Haj90] Yamato Hajime. A probabilistic approach to Stirling numbers of the first kind. Communications in Statistics: Theory and Methods, 19(10):3915–3923, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Holiday:1993:CTT

[HAKM93] D. B. Holiday, H. R. Al-Khalidi, and J. W. McLarty. Comparing time trends in a two-arm longitudinal study with an ordinal categorical response. Communications in Statistics: Theory and Methods, 22(12):3569–3590, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hall:1996:MTP

[Hal96] Andreia O. Hall. Maximum term of a particular autoregressive sequence with discrete margins. Communications in Statistics: Theory and Methods, 25(4):721–736, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hall:1999:GBR

[Hal99] Daniel B. Hall. On Gee-based regression estimators under first moment misspecification. Communications in Statistics: Theory and Methods, 28(5):1021–1042, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hamilton:1991:MVA

[Ham91] Martin A. Hamilton. Model validation: an annotated bibliography. Communications in Statistics: Theory and Methods, 20(7): 2207–2266, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hamedani:1992:BMN

[Ham92] G. G. Hamedani. Bivariate and multivariate normal characterizations: a brief survey. Communications in Statistics: Theory and Methods, 21(9):2665–2688, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hanson:1990:SCI

[Han90] Morgan A. Hanson. Solutions for a class of invex programming problems. Communications in Statistics: Theory and Methods, 19(11):4181–4188, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Handcock:1991:CLH

[Han91] S. Mark Handcock. On cascading latin hypercube designs and additive models for experiments. Communications in Statistics: Theory and Methods, 20(2):417–439, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hanagal:1992:SIR

[Han92] David D. Hanagal. Some inference results in bivariate exponential distributions rased on censored samples. Communications in Statistics: Theory and Methods, 21(5):1273–1295, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hanagal:1993:SIR

[Han93] David D. Hanagal. Some inference results in several symmetric multivariate exponential models. Communications in Statistics: Theory and Methods, 22(9):276–288, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hanagal:1996:MPD

[Han96] David D. Hanagal. A multivariate Pareto distribution. Communications in Statistics: Theory and Methods, 25(7):1471–1488, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hanagal:1997:ERW

[Han97a] David D. Hanagal. Estimation of reliability when stress is censored at strength. Communications in Statistics: Theory and Methods, 26(4):911–919, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hanagal:1997:TBE

[Han97b] David D. Hanagal. Tests for bivariate exponentiality against BHNBUE alternatives. Communications in Statistics: Theory and Methods, 26(5):1239–1252, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hartigan:1990:PM

[Har90] J. A. Hartigan. Partition models. Communications in Statistics: Theory and Methods, 19(8):2745–2756, 1990. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hasegawa:1995:RCS

[Has95a] Hikaru Hasegawa. On risk comparisons of some estimators in a linear regression model under Inagaki’s loss function and nonnormal error terms. Communications in Statistics: Theory and Methods, 24(7):1655–1685, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hassler:1995:PRC

[Has95b] Uwe Hassler. The periodogram regression: correction and comments. Communications in Statistics: Theory and Methods, 24(12):3137–3146, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Haughton:1991:CCI

[Hau91] Dominique Haughton. Consistency of a class of information criteria for model selection in non-linear regression. Communications in Statistics: Theory and Methods, 20(5–6):1619–1629, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Herman:1990:BCL

[HB90] Herman and Neil Burstein. Binomial confidence limits adjusted for a confounding variable. Communications in Statistics: Theory and Methods, 19(1):337–346, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Huang:1991:RRM

[HB91] Mei Ling Huang and Percy Brill. Recurrence relations for the minimum variance unbiased estimator of the probability function of the R distribution. Communications in Statistics: Theory and Methods, 20(12):4005–4019, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Huang:1999:NCP

[HB99] Yeu-Shiang Huang and Vicki M. Bier. A natural conjugate prior for the nonhomogeneous Poisson process with an exponential intensity function. Communications in Statistics: Theory and Methods, 28(6):1479–1509, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hoa:1990:FRT

[HC90] Tran Van Hoa and A. Chatruvedi. Further results on the twostage hierarchical information (2SHI) estimators in the linear regression models. Communications in Statistics: Theory and Methods, 19(12):4697–4704, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hoa:1993:AAG

[HC93] Tran Van Hoa and A. Charurvedi. Asymptotic approximations to the gain of the 2SHI over Stein estimators in linear regression models when the disturbances are small. Communications in Statistics: Theory and Methods, 22(10):2777–2782, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hoekstra:1994:NBC

[HC94] Robert M. Hoekstra and Thomas C. Chenier. A note on the box correction factor for degrees of freedom in certain F tests. Communications in Statistics: Theory and Methods, 23(5):1483– 1505, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hogg:1996:PIM

[HC96] Robert V. Hogg and James A. Calvin. Preface: the Institute of Mathematical Statistics Central Regional Meeting: A meeting in honor of Robert V. Hogg. Communications in Statistics: Theory and Methods, 25(11):2459–2465, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hoa:1997:PEU

[HC97] Tran Van Hoa and Anoop Chaturvedi. Performance of the 2SHI estimator under the generalised Pitman nearness criterion. Communications in Statistics: Theory and Methods, 26(5): 1227–1238, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hung:1994:SSM

[HCL94] H. M. James Hung, George Y. H. Chi, and Raymond J. Lipicky. On some statistical methods for analysis of combination drug studies. Communications in Statistics: Theory and Methods, 23 (2):361–376, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Haber:1991:ARC

[HCW91] Michael Haber, Catherine C. H. Chen, and C. David Williamson. Analysis of repeated categorical responses from fully and partially cross-classified data. Communications in Statistics: Theory and Methods, 20(10):3293–3313, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hossain:1992:EPY

[HD92] Syed A. Hossain and Ram C. Dahiya. Estimation of parameters of Yamada et al. model and modification of the model. Communications in Statistics: Theory and Methods, 21(10):2881–2898, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Huque:1994:MAM

[HD94] Mohammad F. Huque and Satya D. Dubey. A meta-analysis methodology for utilizing study-level covariate information from clinical trials. Communications in Statistics: Theory and Methods, 23(2):377–394, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hannon:1999:EPT

[HD99a] Patrick M. Hannon and Ram C. Dahiya. Estimation of parameters for the truncated exponential distribution. Communications in Statistics: Theory and Methods, 28(11):2591–2612, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hayakawa:1999:WSH

[HD99b] Takesi Hayakawa and Masataka Doi. Wald statistic for homogeneity of multiple parameters. Communications in Statistics: Theory and Methods, 28(3–4):755–771, 1999. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

He:1990:LRT

[He90] Xuming He. Likelihood ratio test for discordancy with slippage alternatives. Communications in Statistics: Theory and Methods, 19(10):3585–3594, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hamdy:1992:FWI

[HEB92] H. I. Hamdy and M. Y. Ei-Bassiouni. Fixed width interval estimation in a random one way model. Communications in Statistics: Theory and Methods, 21(1):169–177, 1992. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hebert:1994:GME

[Heb94a] Jaimie L. Hebert. Generating moments of exponential scale mixtures. Communications in Statistics: Theory and Methods, 23(4):1181–1189, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hebert:1994:MIS

[Heb94b] Jaimie L. Hebert. Moments of inverted scale mixtures. Communications in Statistics: Theory and Methods, 23(4):1173–1180, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Heine:1993:NEV

[Hei93] Barbara Heine. Nonnegative estimation of variance components in an unbalanced one way random effects model. Communications in Statistics: Theory and Methods, 22(8):2351–2372, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Henze:1992:NFC

[Hen92] Norbert Henze. A new flexible class of omnibus tests for exponentiality. Communications in Statistics: Theory and Methods, 22(1):115–133, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Henze:1994:ABV

[Hen94a] Norbert Henze. The asymptotic behavior of a variant of multivariate kurtosis. Communications in Statistics: Theory and Methods, 23(4):1047–1061, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Henze:1994:MKT

[Hen94b] Norbert Henze. On Mardia’s kurtosis test for multivariate normality. Communications in Statistics: Theory and Methods, 23(4):1031–1045, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Herbst:1992:EMU

[Her92] Tomáŝ Herbst. Estimation of moments under Koziol–Green model for random censorship. Communications in Statistics: Theory and Methods, 21(3):613–624, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Herbst:1993:ERM

[Her93] Tomas Herbst. On estimation of residual moments under Koziol– Green model of random censorship. Communications in Statistics: Theory and Methods, 22(9):1769–1772, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Herbert:1996:ATS

[Her96] Büning Herbert. Adaptive tests for the c-sample location problem — the case of two-sided alternatives. Communications in Statistics: Theory and Methods, 25(7):1569–1582, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Huang:1994:EBE

[HFP94] Wen-Tao Huang, Hung-Lin Fu, and Hsiao-Tien Pao. Empirical Bayes estimations for some distribution families. Communications in Statistics: Theory and Methods, 23(5):1429–1445, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hedayat:1994:BSD

[HGM94] A. S. Hedayat, Chi Y. H. George, and Kooros Mahjoob. Blends of statistical designs for dose-response studies. Communications in Statistics: Theory and Methods, 23(2):341–360, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Herndon:1990:RCS

[HH90] James E. Herndon II and Frank E. Harrell, Jr. The restricted cubic spline hazard model. Communications in Statistics: Theory and Methods, 19(2):639–663, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Han:1994:SAE

[HH94] Chien-Pai Han and D. L. Hawkins. A smooth adaptive estimator of the mean of a symmetric or asymmetric distribution. Communications in Statistics: Theory and Methods, 23(1):1–10, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Howlader:1995:BEP

[HH95] H. A. Howlader and A. Hossain. On Bayesian estimation and prediction from Rayleigh based on type II censored data. Communications in Statistics: Theory and Methods, 24(9):2251– 2259, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hidekazu:1991:DIO

[Hid91] Takeuchi Hidekazu. Detecting influential observations by using a new expression of Cook’s distance. Communications in Statistics: Theory and Methods, 20(1):261–274, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hillis:1993:ETB

[Hil93] Stephen L. Hillis. The equivalence of two Buckley–James variance estimators. Communications in Statistics: Theory and Methods, 22(2):479–484, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hadi:1990:HFC

[HJ90] Ali S. Hadi and Brian Jersky. How fair can employers be? Communications in Statistics: Theory and Methods, 19(12):4545– 4558, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Handa:1990:NMD

[HK90a] B. R. Handa and N. S. Kambo. A note on a modified double stage shrinkage estimator. Communications in Statistics: Theory and Methods, 19(12):4833–4840, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Haq:1990:PDL

[HK90b] M. Safiul Haq and Shahjahan Khan. Prediction distribution for a linear regression model with multivariate Student-t error distribution. Communications in Statistics: Theory and Methods, 19(12):4705–4712, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Haq:1990:SIP

[HK90c] M. Safiul Haq and Shahjahan Khan. Structural inference of the parameters of the heteroscedastic simultaneous equation model. Communications in Statistics: Theory and Methods, 19(12): 4713–4729, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hseih:1990:IUM

[HK90d] H. K. Hseih and R. M. Korwar. Inadmissibility of the uniformly minimum variance unbiased estimator of the inverse Gaussian variance. Communications in Statistics: Theory and Methods, 19(7):2509–2516, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hanagal:1991:LST

[HK91] David D. Hanagal and B. K. Kale. Large sample tests of independence for absolutely continuous bivariate exponential distribution. Communications in Statistics: Theory and Methods, 20(4):1301–1313, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hanagal:1992:LST

[HK92] David D. Hanagal and B. K. Kale. Large sample tests for testing symmetry and independence in some bivariate exponential models. Communications in Statistics: Theory and Methods, 21(9): 2625–2643, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hossain:1994:MLE

[HK94] Syed Shahadat Hossain and Azmeri Khan. Maximum likelihood estimation of variance components of heteroscedastic random ANOVA model. Communications in Statistics: Theory and Methods, 23(9):2449–2467, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Henze:1995:BBG

[HK95] Norbert Henze and Bernhard Klar. Bootstrap based goodness of fit tests for the generalized Poisson model. Communications in Statistics: Theory and Methods, 24(7):1875–1896, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Haq:1997:PIL

[HK97a] M. Safiul Haq and B. M. Golam Kibria. Predictive inference for linear and multivariate linear models with MA(1) error processes. Communications in Statistics: Theory and Methods, 26 (2):331–353, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kim:1997:IPF

[hK97b] Sung ho Kim. Iterative proportional fitting for nonhierarchical log-linear models. Communications in Statistics: Theory and Methods, 26(6):1443–1460, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hughes:1999:IAV

[HK99] Anthony W. Hughes and Maxwell L. King. An iterative approach to variable selection based on the Kullback–Leibler information. Communications in Statistics: Theory and Methods, 28(5):1043– 1057, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hayter:1990:PAT

[HL90] A. J. Hayter and W. Liu. Power assessment for tests of the equality of several proportions. Communications in Statistics: Theory and Methods, 19(1):19–30, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hayter:1992:MPA

[HL92] A. J. Hayter and W. Liu. A method of power assessment for tests comparing several treatments with a control. Communications in Statistics: Theory and Methods, 21(7):1871–1889, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ho:1993:ASE

[HL93a] Wai-Tung Ho and Sik-Yum Lee. Analysis of structural equation models with exact and stochastic constraints using a Bayesian approach. Communications in Statistics: Theory and Methods, 22(8):2171–2183, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hung:1993:CPP

[HL93b] Wen Jang Hung and Shun-Hwa Li. Characterizations of the Poisson process using, the variance. Communications in Statistics: Theory and Methods, 22(5):1371–1382, 1993. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hawkins:1994:AEL

[HL94] D. L. Hawkins and C. K. Lakshminarayan. An adaptive estimator of location based on the t-family. Communications in Statistics: Theory and Methods, 23(3):747–761, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hydorn:1999:PEE

[HM99] Debra L. Hydorn and Robb J. Muirhead. Polynomial estimation of eigenvalues. Communications in Statistics: Theory and Methods, 28(3–4):581–596, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hossain:1997:EPP

[HN97] Anwar M. Hossain and Shyamal K. Nath. Estimation of parameters in the presence of outliers for a Burr XII distribution. Communications in Statistics: Theory and Methods, 26(3):637–652, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Holland:1990:BTT

[HO90] Burt Holland and Nasser K. Ordoukhani. Balancing Type I and Type II error probabilities: further comments on proof of safety vs. proof of hazard. Communications in Statistics: Theory and Methods, 19(10):3557–3570, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ho:1998:RST

[Ho98] Chih-Hsiang Ho. Repeated significance tests on accumulating data of repairable systems. Communications in Statistics: Theory and Methods, 27(5):1181–1200, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hobart:1991:CES

[Hob91] E. J. G. Pitman Hobart. The “closest” estimates of statistical parameters. Communications in Statistics: Theory and Methods, 20(11):3423–3437, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hocking:1990:NAV

[Hoc90] R. R. Hocking. A new approach to variance component estimation with diagnostic implications. Communications in Statistics: Theory and Methods, 19(12):4591–4617, 1990. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hodoshima:1990:GCS

[Hod90] Jiro Hodoshima. Gains by the common structural variance. Communications in Statistics: Theory and Methods, 19(7):2473– 2484, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hogg:1996:I

[Hog96a] Robert V. Hogg. Introduction. Communications in Statistics: Theory and Methods, 25(11):2483–2484, 1996. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hogg:1996:TM

[Hog96b] Robert V. Hogg. Thanks for the memories. Communications in Statistics: Theory and Methods, 25(11):2485–2491, 1996.

CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Holiday:1991:CON

[Hol91] D. B. Holiday. Considerations for optimal nonparametric regression under a generalized error structure. Communications in Statistics: Theory and Methods, 20(8):2387–2406, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Holm:1995:CSF

[Hol95] Sture Holm. Confidence sets of fixed size with predetermined confidence level. Communications in Statistics: Theory and Methods, 24(6):1521–1536, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Honsi:1991:OSP

[Hon91] A. K. Honsi. Optimal smoothing parameter of Fourier series density estimates under an autoregressive dependence model. Communications in Statistics: Theory and Methods, 20(5–6): 1839–1851, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Honda:1994:ICI

[Hon94] Yuzo Honda. Information criteria in identifying regression models. Communications in Statistics: Theory and Methods, 23(10): 2815–2829, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Honda:1998:TGF

[Hon98] Toshio Honda. Testing the goodness of fit of a linear model by kernel regression. Communications in Statistics: Theory and Methods, 27(3):529–546, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Honda:1999:RCS

[Hon99] Toshio Honda. Root-n-consistent semiparametric estimation of partially linear models for weakly dependent observations. Communications in Statistics: Theory and Methods, 28(8):2001– 2020, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hoover:1990:CIB

[Hoo90] Donald R. Hoover. Cohpasisons of improved Bonferroni and Sidak/Slepian bounds with applications to normal Markov processes. Communications in Statistics: Theory and Methods, 19(5):1623–1637, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Horn:1992:QNU

[Hor92] Paul S. Horn. Quasi-nonparametric upper tolerance regions based on the bootstrap. Communications in Statistics: Theory and Methods, 21(12):3351–3367, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hossain:1990:DIO

[Hos90] Anwar Hossain. Detection of influential observations in regression model with autocorrelated errors. Communications in Statistics: Theory and Methods, 19(3):1047–1060, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Harris:1992:LQL

[HP92] Ian R. Harris and Donna K. Pauler. Locally quadratic log likelihood and data-based transformations. Communications in Statistics: Theory and Methods, 21(3):637–646, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Huh:1998:MPL

[HPS98] Keun Huh, Anthony K. Postert, and Robin C. Sickles. Maximum penalized likelihood estimation of mixed proportional hazard models. Communications in Statistics: Theory and Methods, 27(9):2143–2164, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hanagal:1998:TBE

[HR98] David D. Hanagal and T. V. Ramanathan. Tests for bivariate exponentiality against BIFRA alternatives based on censored samples. Communications in Statistics: Theory and Methods, 27(8):1947–1960, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hodges:1993:APS

[HRS93] J. L. Hodges, Jr., Philip H. Ramsey, and Juliet Popper Shaffer. Accurate probabilities for the sign test. Communications in Statistics: Theory and Methods, 22(5):1235–1255, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hadi:1990:SPR

[HS90a] Ali S. Hadi and Mun S. Son. Some properties of, and relationships among, several uncorrelated and homoscedastic residual vectors. Communications in Statistics: Theory and Methods, 19(7):2625–2642, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hasofer:1990:TMN

[HS90b] A. M. Hasofer and G. Z. Stein. Testing for multivariate normality after coordinate transformation. Communications in Statistics: Theory and Methods, 19(4):1403–1418, 1990. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hooda:1990:IIG

[HS90c] D. S. Hooda and Umed Singh. An information improvement generating function. Communications in Statistics: Theory and Methods, 19(3):1037–1046, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Huang:1993:DEU

[HS93] Su-Yun Huang and W. J. Studden. Density estimation using spline projection kernels. Communications in Statistics: Theory and Methods, 22(11):3263–3285, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hinkle:1996:ECS

[HS96] John E. Hinkle and Arnold J. Stromberg. Efficient computation of statistical procedures based on all subsets of a specified size. Communications in Statistics: Theory and Methods, 25(3):489– 500, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hwang:1999:MLE

[HS99] Jungsywan Hwang and Steven J. Sepanski. Maximum likelihood estimation of a random-effects model. Communications in Statistics: Theory and Methods, 28(1):143–156, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hsieh:1990:ICV

[Hsi90] H. K. Hsieh. Inferences on the coefficient of variation of an inverse Gaussian distribution. Communications in Statistics: Theory and Methods, 19(5):1589–1605, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hooda:1993:EFD

[HSS93] B. K. Hooda, N. P. Singh, and Umed Singh. On estimation in Fréchet distribution of m-th maxima. Communications in Statistics: Theory and Methods, 22(3):907–919, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hirji:1998:CTT

[HT98] Karim F. Hirji and Man-Lai Tang. A comparison of tests for trend. Communications in Statistics: Theory and Methods, 27 (4):943–963, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hu:1994:MPL

[Hu94] Chin-Yuan Hu. The most powerful location and scale invariant test under the assumption of symmetry. Communications in Statistics: Theory and Methods, 23(1):11–26, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Huang:1990:RRD

[Hua90] Mei Ling Huang. Recurrence relations for the r-distribution. Communications in Statistics: Theory and Methods, 19(1):145– 154, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Huda:1991:SDO

[Hud91] S. Huda. On some D_{S}-optimal designs in spherical regions. Communications in Statistics: Theory and Methods, 20(9):2965– 2985, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hughey:1991:SCM

[Hug91] Richard L. Hughey. A survey and comparison of methods for estimating extreme right tail-area quantiles. Communications in Statistics: Theory and Methods, 20(4):1463–1496, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hung:1990:NRA

[Hun90] Hsien-Ming Hung. Nonlinear regression analysis for complex surveys. Communications in Statistics: Theory and Methods, 19(9):3447–3470, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hurlimann:1998:CML

[Hür98] Werner Hürlimann. On the characterization of maximum likelihood estimators for location-scale families. Communications in Statistics: Theory and Methods, 27(2):495–508, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Husler:1996:MOP

[Hüs96] Jürg Hüsler. Multivariate option price models and extremes. Communications in Statistics: Theory and Methods, 25(4):853– 869, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Huwang:1993:CSS

[Huw93] Longcheen Huwang. A confidence set for the slope in the Dolby’s ultrastructural model. Communications in Statistics: Theory and Methods, 22(5):1403–1412, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Haagen:1991:RCD

[HV91] Klaus Haagen and Giorgio Vittadini. Regression component decomposition in structural analysis. Communications in Statistics: Theory and Methods, 20(4):1153–1161, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Horn:1996:SSB

[HV96] Manfred Horn and Rüdiger Vollandt. Subset selection of the best treatment via many-one tests. Communications in Statistics: Theory and Methods, 25(6):1335–1349, 1996. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hamedani:1990:EBE

[HW90] G. G. Hamedani and G. G. Walter. Empirical Bayes estimation of the binomial parameter n. Communications in Statistics: Theory and Methods, 19(6):2065–2084, 1990. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hinkley:1991:ERS

[HW91] D. V. Hinkley and S. Wang. Efficiency of robust standard errors for regression coefficients. Communications in Statistics: Theory and Methods, 20(1):1–11, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hwang:1992:EBA

[Hwa92] Leng-Cheng Hwang. Empirical Bayes approach to Bayes sequential estimation: the Poisson and Bernoulli cases. Communications in Statistics: Theory and Methods, 21(8):2293–2308, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hwang:1994:EDD

[Hwa94] Chang-Ha Hwang. On estimating the dimensionality in discriminant analysis. Communications in Statistics: Theory and Methods, 23(8):2197–2215, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hoshino:1997:NTB

[HY97] Naoto Hoshino and Seki Yoichi. Nonparametric tests based on contrasts for d-star, bipartite and broom tree alternatives. Communications in Statistics: Theory and Methods, 26(1):129–138, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Hyodo:1994:CIM

[Hyo94] Yoshifumi Hyodo. Characterization of information matrices for balanced two-level fractional factorial designs of odd resolution derived from two-symbol simple arrays. Communications in Statistics: Theory and Methods, 23(7):1859–1874, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Henze:1990:CIC

[HZ90] N. Henze and B. Zirkler. A class of invariant consistent tests for multivariate normality. Communications in Statistics: Theory and Methods, 19(10):3595–3617, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zhou:1993:MLE

[hZ93] Xiao hua Zhou. Maximum likelihood estimators of sensitivity and specificity corrected for verification bias. Communications in Statistics: Theory and Methods, 22(11):3177–3198, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ibrahim:1991:EPM

[Ibr91] Ibrahim Hassan Ibrahim. Evaluating the power of the Mann– Whitney test using the bootstrap method. Communications in Statistics: Theory and Methods, 20(9):2919–2931, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Il:1998:MCM

[ID98] Hyo Park Il and M. M. Desu. Multivariate control median test for right censored data. Communications in Statistics: Theory and Methods, 27(8):1923–1935, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Islam:1994:TST

[IH94] Mohamed Zahorul Islam and John E. Hewett. Two stage tests for location when covariates are present. Communications in Statistics: Theory and Methods, 23(4):1239–1246, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ihaka:1993:SAE

[Iha93] Ross Ihaka. Statistical aspects of earthquake source parameter estimation in the presence of signal generated noise. Communications in Statistics: Theory and Methods, 22(5):1425–1440, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Inada:1997:SIE

[IK97] Koichi Inada and Hea-Jung Kim. On shrinkage to interval estimators of the binomial p. Communications in Statistics: Theory and Methods, 26(1):95–112, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Koike:1999:LBB

[iK99] Ken ichi Koike. A lower bound for the Bayes risk in the sequential case. Communications in Statistics: Theory and Methods, 28(3–4):857–871, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Inoue:1999:AIG

[Ino99] Kiyoshi Inoue. Asymptotic improvement of the Graybill–Deal estimator. Communications in Statistics: Theory and Methods, 28 (2):389–407, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Isaki:1992:MBE

[Isa92] Cary T. Isaki. Model bias effects in small area coverage error estimation. Communications in Statistics: Theory and Methods, 21(5):1213–1231, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Isogawa:1992:CML

[Iso92a] Yoshiko Isogawa. Comparing multivariate linear functional relationships. Communications in Statistics: Theory and Methods, 21(2):319–331, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Isohgawa:1992:ADS

[Iso92b] Yoshiko Isohgawa. Asymptotic distributions of some test statistics for a linear functional relationship. Communications in Statistics: Theory and Methods, 21(12):3521–3534, 1992. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Isogawa:1993:STM

[Iso93] Yoshiko Isogawa. On some tests for a multivariate linear functional relationship with replication. Communications in Statistics: Theory and Methods, 22(4):1173–1190, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ihorst:1993:MMS

[IT93] G. Ihorst and G. Trenkler. MSE-matrix superiority of the mixed over the least squares estimator in the presence of outliers. Communications in Statistics: Theory and Methods, 22(7):1865– 1877, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Isogai:1993:NMR

[IU93] Eiichi Isogai and Chikara Uno. A note on minimum risk point estimation of the variance of a normal population. Communications in Statistics: Theory and Methods, 22(8):2309–2319, 1993.

CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Iwata:1994:LBB

[Iwa94] Shigeru Iwata. Lower bounds on Bayes factors for a linear regression model. Communications in Statistics: Theory and Methods, 23(6):1825–1834, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Irotsu:1999:EOR

[IY99] C. Irotsu and Y. Yamada. Estimating odds ratios through the connected comparative experiments. Communications in Statistics: Theory and Methods, 28(3–4):905–929, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Joarder:1992:DCM

[JA92] Anwarul H. Joarder and Mir M. Ali. Distribution of the correlation matrix for a class of elliptical models. Communications in Statistics: Theory and Methods, 21(7):1953–1964, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Jahn:1991:CSO

[Jah91] W. Jahn. Criterion and selection of optimal subsets in the linear regression, two stage procedures. Part I: Selection in the univariate multiple model. Communications in Statistics: Theory and Methods, 20(5–6):1631–1652, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Jain:1991:EMQ

[Jai91] Sudha Jain. Estimation in M/E_{r} /1 queueing systems. Communications in Statistics: Theory and Methods, 20(5–6):1871–1879, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Jain:1994:DCT

[Jai94] Ram B. Jain. On the design of clinical trials for treatment of opiate dependence. Communications in Statistics: Theory and Methods, 23(2):589–613, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Janardan:1995:ERA

[Jan95] K. G. Janardan. Environmental risk assessment model for low dose extrapolation. Communications in Statistics: Theory and Methods, 24(10):2621–2633, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Janardan:1997:WCL

[Jan97] K. G. Janardan. A wider class of Lagrange distributions of the second kind. Communications in Statistics: Theory and Methods, 26(9):2087–2091, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Janardan:1998:GPE

[Jan98a] K. G. Janardan. Generalized Pólya–Eggenberger family of distributions and its relation to Lagrangian Katz family. Communications in Statistics: Theory and Methods, 27(10):2423–2442, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Janardan:1998:SMN

[Jan98b] K. G. Janardan. A stochastic model for the number of accidents and a class of distributions for the number of fatalities. Communications in Statistics: Theory and Methods, 27(7):1823–1833, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Janardan:1999:EPG

[Jan99] K. G. Janardan. Estimation of parameters of the GPED. Communications in Statistics: Theory and Methods, 28(9):2167– 2179, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Jarpe:1999:SIP

[Jär99] E. Järpe. Surveillance of the interaction parameter of the Ising model. Communications in Statistics: Theory and Methods, 28(12):3009–3027, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Jin:1993:EEN

[JB93] Chun Jin and Calvin J. Berry. Equivariant estimation of a normal mean vector using a normal concomitant vector for covariance adjustment. Communications in Statistics: Theory and Methods, 22(2):335–346, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Judge:1990:SSI

[JBY90] George Judge, Robert Bohrer, and Thomas Yancey. Some statistical implications of multivariate inequality constrained testing. Communications in Statistics: Theory and Methods, 19(2): 413–430, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Jin:1998:IED

[JC98a] Chun Jin and Robert H. Crouse. An identity for exponential distributions with the common location parameter and its applications. Communications in Statistics: Theory and Methods, 27 (2):409–422, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Jin:1998:NCL

[JC98b] Chun Jin and Robert H. Crouse. A note on the common location parameter of several exponential populations. Communications in Statistics: Theory and Methods, 27(11):2777–2789, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Jelenkowska:1995:BAM

[Jel95] T. H. Jelenkowska. A Bayesian analysis of the multivariate mixed-linear model. Communications in Statistics: Theory and Methods, 24(12):3183–3196, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Jelenkowska:1999:BEI

[Jel99] T. H. Jelenkowska. Bayesian estimation of the intraclass correlation coefficients in multivariate mixed linear model. Communications in Statistics: Theory and Methods, 28(5):1203–1215, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Johaadien:1995:AML

[JG95] Hassan Johaadien and Constantinos Goutis. Approximate maximum likelihood estimation in the presence of extra-Poisson variation. Communications in Statistics: Theory and Methods, 24(6):1625–1636, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Jimenez-Gamero:1998:BSM

[JGMGMR98] M. D. Jiménez-Gamero, J. Muñoz-García, and A. Muñoz-Reyes. Bootstrapping the sample median. Communications in Statistics: Theory and Methods, 27(8):1979–1990, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Jackson:1999:MAC

[JH99] Michael J. Jackson and John E. Hewett. Multivariate analysis of covariance based on residuals. Communications in Statistics: Theory and Methods, 28(11):2687–2715, 1999. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Jose:1997:EJP

[JI97] C. T. Jose and B. Ismail. Estimation of jump points in nonparametric regression through residual analysis. Communications in Statistics: Theory and Methods, 26(11):2583–2607, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Jose:1999:CPN

[JI99] C. T. Jose and B. Ismall. Change points in nonparametric regression functions. Communications in Statistics: Theory and Methods, 28(8):1883–1902, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Jiang:1999:MHL

[Jia99] Jiming Jiang. On maximum hierarchical likelihood estimators. Communications in Statistics: Theory and Methods, 28(8):1769– 1775, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Jie:1990:ARK

[Jie90] Mi Jie. Asymptotic results for the Koziol–Green model. Communications in Statistics: Theory and Methods, 19(8):2767–2779, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Jin:1995:SEM

[Jin95] Chun Jin. Shrinkage estimation of a mean vector in a two-sample problem. Communications in Statistics: Theory and Methods, 24(11):2881–2888, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Jing:1997:RPB

[Jin97] Bing-Yi Jing. On the relative performance of the block bootstrap for dependent data. Communications in Statistics: Theory and Methods, 26(6):1313–1328, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Jensen:1991:SAE

[JK91] J. L. Jensen and L. B. Kristensen. Saddlepoint approximations to exact tests and improved likelihood ratio tests for the gamma distribution. Communications in Statistics: Theory and Methods, 20(4):1515–1532, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Joshi:1991:TSP

[JN91] S. M. Joshi and S. P. Nabar. Testing the scale parameter of the exponential distribution with known coefficient of variation. Communications in Statistics: Theory and Methods, 20(2):747– 756, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Jeevanand:1994:EES

[JN94] E. S. Jeevanand and N. Unnikrishnan Nair. Estimating P [X > Y] from exponential samples containing spurious observations. Communications in Statistics: Theory and Methods, 23(9):2629– 2642, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Jain:1995:MWD

[JN95] Kanchan Jain and Asok K. Nanda. On multivariate weighted distributions. Communications in Statistics: Theory and Methods, 24(10):2517–2539, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Jang:1996:GME

[JN96] Dae-Heung Jang and Ho-Jun Na. A graphical method for evaluating mixture designs with respect to the slope. Communications in Statistics: Theory and Methods, 25(5):1043–1058, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

John:1995:OFD

[Joh95] Peter W. M. John. Optimal foldover designs. Communications in Statistics: Theory and Methods, 24(7):1821–1827, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Johnson:1996:LEA

[Joh96] Norman L. Johnson. Letter to the editor: the asymptotic distribution of the process capability index c_{pmk}. Communications in Statistics: Theory and Methods, 25(6):1405, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Jones:1994:KDD

[Jon94] M. C. Jones. On kernel density derivative estimation. Communications in Statistics: Theory and Methods, 23(8):2133–2139, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Jorgensen:1999:MRP

[Jor99] Murray Jorgensen. Model-robust parameter dispersions for iteratively re-weighted least squares. Communications in Statistics: Theory and Methods, 28(8):1903–1919, 1999. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Jayade:1990:EPM

[JP90] V. D. Jayade and M. S. Prasad. Estimation of parameters of mixed failure time distribution. Communications in Statistics: Theory and Methods, 19(12):4667–4678, 1990. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Jin:1991:ELP

[JP91a] Chun Jin and Nabendu Pal. Estimation of location parameters of several exponential distributions. Communications in Statistics: Theory and Methods, 20(12):3995–4003, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Jin:1991:NLP

[JP91b] Chun Jin and Nabendu Pal. A note on the location parameters of two exponential distributions under order restrictions. Communications in Statistics: Theory and Methods, 20(10):3147–3158, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Juan:1992:SMI

[JP92] Jesús Juan and Daniel Pefña. A simple method to identify significant effects in unreplicated two-level factorial designs. Communications in Statistics: Theory and Methods, 21(5):1383–1403, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Jang:1993:MGM

[JP93] Dae H. Jang and Sung H. Park. A measure and a graphical method for evaluating slope rotatability in response surface designs. Communications in Statistics: Theory and Methods, 22(7):1849–1863, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Jani:1999:BES

[JP99] P. N. Jani and Mayuri Pandya. Bayes estimation of shift point in left truncated exponential sequence. Communications in Statistics: Theory and Methods, 28(11):2623–2639, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Jensen:1990:DDT

[JR90] D. R. Jensen and D. E. Ramirez. Dispersion-diminishing transformations. Communications in Statistics: Theory and Methods, 19(9):3259–3266, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Jensen:1991:MTL

[JR91] D. R. Jensen and D. E. Ramiez. Misspecified T^{2} tests. I. Location and scale. Communications in Statistics: Theory and Methods, 20(1):249–259, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Joshi:1990:SAA

[JS90] S. Joshi and M. Shah. Sequential analysis applied to testing the mean of an inverse Gaussian distribution with known coefficient of variation. Communications in Statistics: Theory and Methods, 19(4):1457–1466, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Joshi:1991:SIA

[JS91a] P. C. Joshi and Shubha. Some identities among moments of order statistics. Communications in Statistics: Theory and Methods, 20(9):2837–2843, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Joshi:1991:EMI

[JS91b] S. Joshi and M. Shah. Estimating the mean of an inverse Gaussian distribution with known coefficient of variation. Communications in Statistics: Theory and Methods, 20(9):2907–2912, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Janner:1993:CML

[JS93] Michael Jänner and Peter Stahlecker. A comparison of maximum likelihood and quasi minimax approach to missing values. Communications in Statistics: Theory and Methods, 22(2):319–333, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Johnson:1995:NTO

[JS95] Richard A. Johnson and Songyong Sim. Nonparametric tests for ordered c sample contaminated distributions. Communications in Statistics: Theory and Methods, 24(4):861–879, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Jani:1997:STP

[JSM97] P. N. Jani, A. Shanubhogue, and K. Muralidharan. Some tests for a Poisson process and their powers with respect to trend in the intensity. Communications in Statistics: Theory and Methods, 26(11):2671–2688, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Jani:1998:TGD

[JSM98] P. N. Jani, Ashok Shanubhogue, and K. Muralidharan. A test for gamma DFR alternative. Communications in Statistics: Theory and Methods, 27(10):2535–2546, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Jinadasa:1990:MLU

[JT90] K. G. Jinadasa and D. S. Tracy. Multivariate linear ultrastructural relationships model. Communications in Statistics: Theory and Methods, 19(3):805–814, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Jinadasa:1992:MLE

[JT92] K. G. Jinadasa and D. S. Tracy. Maximum likelihood estimation for multivariate normal distribution with monotone sample. Communications in Statistics: Theory and Methods, 21(1): 41–50, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Johnson:1993:NRM

[JU93] Alan M. Johnson and Ida S. Umphers. A nonlinear regression model for describing averaged growth curves of mice. Communications in Statistics: Theory and Methods, 22(12):3405–3418, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Janssen:1992:BSE

[JV92] Paul Janssen and Noël Veraverbeke. Bootstrapping U-statistics with estimated parameters. Communications in Statistics: Theory and Methods, 21(6):1585–1603, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Janardan:1995:SSB

[JVM95] K. G. Janardan, R. Vasudeva, and N. R. Mohan. On some special birth processes in biology. Communications in Statistics: Theory and Methods, 24(4):1033–1038, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Joseph:1992:EMP

[JW92] Lawrence Joseph and David B. Wolfson. Estimation in multipath change-point problems. Communications in Statistics: Theory and Methods, 21(4):897–913, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Jing:1996:SBK

[JZ96] Ping Jing and Li-Xing Zhu. Some Blum–Kiefer–Rosenblatt type tests for the joint independence of variables. Communications in Statistics: Theory and Methods, 25(9):2127–2139, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Krehbiel:1991:OFF

[KA91] Timothy C. Krehbiel and Donald A. Anderson. Optimal fractional factorial designs for estimating interactions of one factor with all others: 3^{m} series. Communications in Statistics: Theory and Methods, 20(3):1055–1072, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Krehbiel:1992:UML

[KA92a] Timothy C. Krehbiel and Donald A. Anderson. The use of a monetary loss function to determine the optimal fractional replicate of factorial experiments. Communications in Statistics: Theory and Methods, 21(8):2383–2398, 1992. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kulperger:1992:BTP

[KA92b] R. J. Kulperger and S. E. Ahmed. A bootstrap theorem for a preliminary test estimator. Communications in Statistics: Theory and Methods, 21(7):2071–2082, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Krehbiel:1997:SAS

[KA97a] Timothy C. Krehbiel and Donald A. Anderson. Sequential analysis of 3^{m} supersaturated factorial experiments. Communications in Statistics: Theory and Methods, 26(3):699–713, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kumar:1997:AEP

[KA97b] Pranesh Kumar and S. K. Agarwal. Alternative estimators for the population totals in multiple characteristic surveys. Communications in Statistics: Theory and Methods, 26(10):2527–2537, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kabe:1990:END

[Kab90] D. G. Kabe. The exact noncentral distribution of the generalized multiple correlation matrix. Communications in Statistics: Theory and Methods, 19(9):3497–3504, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kabe:1991:MNL

[Kab91a] D. G. Kabe. On a multivariate normal linear regression hypothesis testing problem with additional observations. Communications in Statistics: Theory and Methods, 20(3):889–898, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kabe:1991:WFN

[Kab91b] D. G. Kabe. On Wilks’ A factorization noncentral distribution theory. Communications in Statistics: Theory and Methods, 20(7):1995–2007, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kabaila:1997:AVS

[Kab97] Paul Kabaila. Admissible variable-selection procedures when fitting misspecified regression models by least squares. Communications in Statistics: Theory and Methods, 26(10):2299–2306, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kershner:1994:CEP

[KAHK94] Ronald Kershner, P. H. Ingrid Amara, M. S. Suzanne Hackett, and Gary Koch. Constructing and evaluating a partially complete crossover design for a multi-dose study. Communications in Statistics: Theory and Methods, 23(2):395–402, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kaigh:1996:RPI

[Kai96] W. D. Kaigh. Random partition indicator and spacing statistics. Communications in Statistics: Theory and Methods, 25(11): 2699–2722, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Keselman:1999:ARM

[KAKW99] H. J. Keselman, James Algina, Rhonda K. Kowalchuk, and Russell D. Wolfinger. The analysis of repeated measurements: a comparison of mixed-model Satterthwaite F tests and a nonpooled adjusted degrees of freedom multivariate test. Communications in Statistics: Theory and Methods, 28(12):2967–2999, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kamps:1994:RPR

[Kam94] Udo Kamps. Reliability properties of record values from nonidentically distributed random variables. Communications in Statistics: Theory and Methods, 23(7):2101–2112, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kamstra:1999:ORB

[Kam99] Mark Kamstra. An observation on regression-based specification tests. Communications in Statistics: Theory and Methods, 28(6): 1435–1446, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kano:1990:CSN

[Kan90] Yutaka Kano. Comparative studies of non-iterative estimators based on Ihara and Kano’s method in exploratory factor analysis. Communications in Statistics: Theory and Methods, 19 (2):431–444, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kanjo:1993:ETN

[Kan93] Anis I. Kanjo. An exact test for NBUE class of survival functions. Communications in Statistics: Theory and Methods, 22 (3):787–795, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kappenman:1992:RER

[Kap92] Russell F. Kappenman. Robust estimation of the ratio of scale parameters for positive random variables. Communications in Statistics: Theory and Methods, 21(10):2983–2996, 1992. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Karmous:1994:ESL

[Kar94] Azza R. Karmous. M-estimation of k-sample location parameters; a two-sided preliminary test approach. Communications in Statistics: Theory and Methods, 23(11):3063–3073, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Karagrigoriou:1995:AEM

[Kar95] Alexandros Karagrigoriou. Asymptotic efficiency of model selection criteria: the nonzero mean Gaussian AR(∞) case. Communications in Statistics: Theory and Methods, 24(4):911–930, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Karlsson:1999:TNC

[Kar99] Niklas Karlsson. Testing for normally in censored regressions. Communications in Statistics: Theory and Methods, 28(7):1635– 1651, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Katkovnik:1995:NLP

[Kat95] Vladimir Katkovnik. Nonparametric local polynomial approximation of the time-varying frequency and amplitude. Communications in Statistics: Theory and Methods, 24(12):3001–3025, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kim:1990:APF

[KB90] Won Kyung Kim and L. Billard. Asymptotic properties for the first-order bilinear time series model. Communications in Statistics: Theory and Methods, 19(4):1171–1183, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Koti:1996:STR

[KB96] Kallapa M. Koti and Gutti Jogesh Babu. Sign test for ranked-set sampling. Communications in Statistics: Theory and Methods, 25(7):1617–1630, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kish:1990:DIIa

[KC90a] Charles W. Kish, Jr. and Vernon M. Chinchilli. Diagnostics for identifying influential cases in MANOVA. Communications in Statistics: Theory and Methods, 19(7):2667–2681, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kish:1990:DIIb

[KC90b] Charles W. Kish, Jr. and Vernon M. Chinchilli. Diagnostics for identifying influential cases in MANOVA. Communications in Statistics: Theory and Methods, 19(7):2683–2704, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kaigh:1991:SQEb

[KC91a] W. D. Kaigh and Cheng Cheng. Subsampling quantile estimator standard errors with applications. Communications in Statistics: Theory and Methods, 20(3):977–995, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kaigh:1991:SQEa

[Kc91b] W. D. Kaigh and Cheng cheng. Subsampling quantile estimators and uniformity criteria. Communications in Statistics: Theory and Methods, 20(2):539–560, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kao:1998:NPC

[KC98] L. H. Kao and S. Chakraborti. Nonparametric procedures for comparing treatments with a control in a randomized complete block design. Communications in Statistics: Theory and Methods, 27(3):687–704, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kodell:1994:CDT

[KCM94] Ralph L. Kodell, James J. Chen, and Gary E. Moore. Comparing distributions of time to onset of disease in animal tumorigenicity experiments. Communications in Statistics: Theory and Methods, 23(4):959–980, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kale:1995:MLE

[KD95] Mandar Kale and S. R. Deshmukh. Maximum likelihood estimation in birth-death process allowing catastrophes. Communications in Statistics: Theory and Methods, 24(7):1863–1873, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Karia:1999:RFC

[KD99] S. R. Karia and J. V. Deshpande. The response functions for competing risks. Communications in Statistics: Theory and Methods, 28(3–4):947–963, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kale:1993:EEP

[KE93] Mohan M. Kale and S. R. Eshmokh. Estimation of extinction probability of a multitype Markov branching process. Communications in Statistics: Theory and Methods, 22(4):1195–1206, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Keating:1991:KCT

[Kea91] Jerome P. Keating. Karlin’s corollary: a topological approach to Pitman’s measure. Communications in Statistics: Theory and Methods, 20(11):3729–3750, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Keating:1996:SMS

[Kea96] Jerome P. Keating. A statistical method for the serial comparison of vector cardiograms. Communications in Statistics: Theory and Methods, 25(9):1971–1983, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kejian:1993:NCB

[Kej93] Liu Kejian. A new class of biased estimate in linear regression. Communications in Statistics: Theory and Methods, 22(2):393– 402, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kelleher:1996:BME

[Kel96] Thomas Kelleher. A Bayes method for estimating the common mean of two normal populations. Communications in Statistics: Theory and Methods, 25(9):2141–2157, 1996. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kemp:1992:HEE

[Kem92] Adrienne W. Kemp. Heine–Euler extensions of the Poisson distribution. Communications in Statistics: Theory and Methods, 21(3):571–588, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kerstin:1995:DEM

[Ker95] Vännman Kerstin. On the distribution of the estimated mean from nonstandard mixtures of distributions. Communications in Statistics: Theory and Methods, 24(6):1569–1584, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Koons:1990:EMA

[KF90] Bruce K. Koons and Robert V. Foutz. Estimating moving average parameters in the presence of measurement error. Communications in Statistics: Theory and Methods, 19(9):3179–3187, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kim:1997:KTT

[KF97a] Kangkyun Kim and Robert V. Foutz. Kolmogorov-type tests of goodness-of-fit against stochastic ordering. Communications in Statistics: Theory and Methods, 26(4):885–897, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kwan:1997:FIM

[KF97b] C. W. Kwan and W. K. Fung. Further influence measures in factor analysis. Communications in Statistics: Theory and Methods, 26(6):1337–1353, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kambo:1992:HTS

[KFAH92] N. S. Kambo, B. R. Fanda, and Z. A. Al-Hemyari. On Huntsberger type shrinkage estimator. Communications in Statistics: Theory and Methods, 21(3):823–841, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kshirsagar:1990:DCF

[KG90] A. M. Kshirsagar and R. P. Gupta. Direction and collinearity factors of Wilks’ A — a review. Communications in Statistics: Theory and Methods, 19(4):1359–1402, 1990. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Khattree:1992:SPD

[KG92a] Ravindra Khattree and Rameshwar D. Gupta. Some probability distributions connected with beta and gamma matrices. Communications in Statistics: Theory and Methods, 21(2):369–390, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kuralmani:1992:SMD

[KG92b] V. Kuralmani and K. Govlndaraju. Selection of multiple deferred (dependent) state sampling plans. Communications in Statistics: Theory and Methods, 21(5):1339–1366, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kuralmani:1995:MTS

[KG95] V. Kuralmani and K. Govindaraju. Modified tables for the selection of double sampling attribute plan indexed by AQL and LQL. Communications in Statistics: Theory and Methods, 24(7):

1897–1927, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kumar:1994:DFT

[KGM94] Narinder Kumar, Amar Nath Gill, and Gobind P. Mehta. Distribution-free test for homogeneity against ordered alternatives. Communications in Statistics: Theory and Methods, 23(4): 1247–1256, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Katti:1990:SDA

[KH90] S. K. Katti and Ferrin Harrison. Sampling distribution after splitting experimental units. Communications in Statistics: Theory and Methods, 19(11):4189–4198, 1990. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kim:1997:GFT

[KHJY97] Choongrak Kim, Changkon Hong, Meeseon Jeong, and Meeyoung Yang. Goodness-of-fit test for density estimation. Communications in Statistics: Theory and Methods, 26(11):2725–2741, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kambo:1997:ESL

[KHR97] N. S. Kambo, B. R. Handa, and V. Ravindranath. Estimation of the smaller and larger of truncation parameters of two nonregular distributions. Communications in Statistics: Theory and Methods, 26(10):2439–2456, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ki:1992:MSE

[Ki92] Yuen-Ching Fanny Ki. Multiple shrinkage estimators in multiple linear regression. Communications in Statistics: Theory and Methods, 21(1):111–136, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kim:1994:DSM

[KI94] Hea-Jung Kim and Koichi Inada. On double stage minimum discrimination information estimators of the interval constrained normal mean. Communications in Statistics: Theory and Methods, 23(10):2733–2753, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kibria:1996:PTR

[Kib96] B. M. Golam Kibria. On preliminary test ridge regression estimators for linear restrictions in a regression model with nonnormal disturbances. Communications in Statistics: Theory and Methods, 25(10):2349–2369, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kibria:1999:PDH

[Kib99] B. M. Golam Kibria. The prediction distribution for the heteroscedastic multivariate linear models. Communications in Statistics: Theory and Methods, 28(10):2477–2486, 1999. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kim:1993:TPR

[Kim93] Hyune-Ju Kim. Two-phase regression with nonhomogeneous errors. Communications in Statistics: Theory and Methods, 22 (3):647–657, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kim:1994:DSO

[Kim94a] Choongrak Kim. Dispersion statistics in overdispersed mixture models. Communications in Statistics: Theory and Methods, 23 (1):27–45, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kim:1994:ICC

[Kim94b] Myung Geun Kim. Influence curve for the Cholesky root of a covariance matrix. Communications in Statistics: Theory and Methods, 23(5):1399–1412, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kim:1994:GPA

[Kim94c] Sung-Ho Kim. A general property among nested, pruned subtrees of a decision-support tree. Communications in Statistics: Theory and Methods, 23(4):1227–1238, 1994. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kim:1995:BQC

[Kim95a] Hea-Jung Kim. On a balanced quadratic classification rule. Communications in Statistics: Theory and Methods, 24(3):607– 623, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kim:1995:LIM

[Kim95b] Myung Genn Kim. Local influence in multivariate regression. Communications in Statistics: Theory and Methods, 24(5):1271– 1278, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kim:1995:ICC

[Kim95c] Myung Geun Kim. The influence in comparing covariance matrices. Communications in Statistics: Theory and Methods, 24(6): 1431–1441, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kim:1996:LIR

[Kim96a] Choongrak Kim. Local influence and replacement measure. Communications in Statistics: Theory and Methods, 25(1):49– 61, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kim:1996:LIB

[Kim96b] Myung Genu Kim. Local influence in Box–Cox transformation for multivariate regression data. Communications in Statistics: Theory and Methods, 25(12):2955–2969, 1996. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kotz:1990:SCP

[KJ90] Samuel Kotz and Norman L. Johnson. Some comments on a paper by Aiala Barr. Communications in Statistics: Theory and Methods, 19(3):815–819, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). See [Bar89].

Kung-Jong:1991:DLB

[KJ91] Lui Kung-Jong. A discussion on the limitations of B. Prasad’s unbiased estimator in finite population sampling. Communications in Statistics: Theory and Methods, 20(1):293–298, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kocherlakota:1990:BLS

[KK90] Subrahmaniam Kocherlakota and Kathleen Kocherlakota. The bivariate logarithmic series distribution. Communications in Statistics: Theory and Methods, 19(9):3387–3432, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kano:1991:RRP

[KK91a] Kazuhiko Kano and Kazutomo Kawamura. On recurrence relations for the probability function of multivariate generalized Poisson distribution. Communications in Statistics: Theory and Methods, 20(1):165–178, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kocherlakota:1991:PCI

[KK91b] Subrahmaniam Kocherlakota and Kathleen Kocherlakota. Process capability index: bivariate normal distribution. Communications in Statistics: Theory and Methods, 20(8):2529–2547, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kulkarni:1991:ECT

[KK91c] P. M. Kulkarni and Debashis Kushary. β expectation and βcontent tolerance intervals for dependent observations. Communications in Statistics: Theory and Methods, 20(3):1043–1054, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kemp:1992:GDM

[KK92] Adrienne W. Kemp and David C. Kemp. A group-dynamic model and the lost-games distribution. Communications in Statistics: Theory and Methods, 21(3):791–798, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kocherlakota:1994:CIP

[KK94] Subrahmaniam Kocherlakota and Kathleen Kocherlakota. Confidence intervals for the process capability ratio based on robust estimators. Communications in Statistics: Theory and Methods, 23(1):257–276, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kang:1998:LCD

[KK98a] Seung-Ho Kang and Jerome Klotz. Limiting conditional distribution for tests of independence in the two way table. Communications in Statistics: Theory and Methods, 27(8):2075–2082, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kang:1998:MLE

[KK98b] Shin-Soo Kang and Kenneth J. Koehler. Maximum likelihood estimation of bivariate survival probabilities and its existence and uniqueness. Communications in Statistics: Theory and Methods, 27(8):1961–1977, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kim:1998:SDR

[KK98c] Choongrak Kim and Woochul Kim. Some diagnostic results in nonparametric density estimation. Communications in Statistics: Theory and Methods, 27(2):291–303, 1998. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kshirsagar:1990:CPU

[KKK90] A. M. Kshirsagar, S. Kocherlakota, and K. Kocherlakota. Classification procedures using principal component analysis and stepwise discriminant function. Communications in Statistics: Theory and Methods, 19(1):91–109, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kirmani:1991:EPC

[KKK91] S. N. U. A. Kirmani, Kathleen Kocherlakota, and Subrahmaniam Kocherlakota. Estimation of σ and the process capability index based on subsamples. Communications in Statistics: Theory and Methods, 20(1):275–291, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kim:1999:STB

[KKP^{+}99] Choongrak Kim, Woochul Kim, Byeong U. Park, Changkon Hong, and Meeseon Jeong. Smoothing techniques via the Bézier curve. Communications in Statistics: Theory and Methods, 28(7):1577–1597, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kariya:1999:CSE

[KKS99] Takeaki Kariya, Yoshihiko Konno, and William E. Strawderman. Construction of shrinkage estimators for the regression coefficient matrix in the GMANOVA model. Communications in Statistics: Theory and Methods, 28(3–4):597–611, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kumar:1998:SMC

[KKT98] A. Vijay Kumar, B. Krishna Kumar, and B. Thilaka. On semi-Markov compartmental models with branching particles under the influence of disasters. Communications in Statistics: Theory and Methods, 27(5):1101–1122, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Koolaard:1996:LED

[KL96] J. P. Koolaard and C. R. O. Lawoko. The linear and Euclidean discriminant functions: a comparison via asymptotic expansions and simulation study. Communications in Statistics: Theory and Methods, 25(12):2989–3011, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Keyes:1999:CKL

[KL99a] Tim K. Keyes and Martin S. Levy. On calibration of Kullback– Leibler divergence via prediction. Communications in Statistics: Theory and Methods, 28(1):67–85, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kim:1999:CTD

[KL99b] Chul-Ki Kim and Tze Leung Lai. For censored and truncated data. Communications in Statistics: Theory and Methods, 28 (11):2717–2747, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Klar:1996:SAC

[Kla96] Neil Klar. Stratified analysis of correlated binary outcome data: a comparison of model dependent and robust tests of significance. Communications in Statistics: Theory and Methods, 25(10):2431–2458, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Klefsjo:1994:LE

[Kle94] Bengt Klefsjö. Letter to the editor. Communications in Statistics: Theory and Methods, 23(8):2417–2421, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kacker:1991:EES

[KLH^{+}91] Raghu N. Kacker, Eric S. Laferfren, Michael D. Hill, Winnie Wong-Ng, Chawn K. Chiang, and Edwin R. Fuller. An efficient experiment to study superconducting ceramics. Communications in Statistics: Theory and Methods, 20(2):441–456, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Klotz:1990:MLE

[Klo90] Jerome Klotz. Maximum likelihood estimators for the two way mixed model. Communications in Statistics: Theory and Methods, 19(10):3741–3749, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kageyama:1990:RBD

[KM90] Sanpei Kageyama and Dibyen Majumdar. Resistant BTIB designs. Communications in Statistics: Theory and Methods, 19(6):2145–2158, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kanwar:1991:TVR

[KM91a] Sen Kanwar and Bala Jain Madhu. A test for the variance residual life. Communications in Statistics: Theory and Methods, 20 (1):327–331, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Krug:1991:ILM

[KM91b] T. Krug and R. J. Martin. Information loss on the mean for spatial processes when some values are missing. Communications in Statistics: Theory and Methods, 20(7):2169–2194, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kunchur:1993:ERM

[KM93] S. H. Kunchur and Surekha B. Munoli. Estimation of reliability for a multicomponent survival stress–strength model based on exponential distributions. Communications in Statistics: Theory and Methods, 22(3):769–779, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kim:1994:AEO

[KM94a] Byung Hwee Kim and Glen Meeden. Admissible estimation in an one parameter nonregular family of absolutely continuous distributions. Communications in Statistics: Theory and Methods, 23(10):2993–3001, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kuk:1994:FAE

[KM94b] Anthony Y. C. Kuk and T. K. Mak. A functional approach to estimating finite population distribution functions. Communications in Statistics: Theory and Methods, 23(3):883–896, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kunchur:1994:ERF

[KM94c] S. H. Kunchur and Surekha B. Munoli. Estimation of reliability in Freund model for two component system. Communications in Statistics: Theory and Methods, 23(11):3273–3283, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Krug:1995:ILD

[KM95] T. Krug and R. J. Martin. Information loss on the dependence parameter due to incomplete data from a one-parameter spatial Gaussian process. Communications in Statistics: Theory and Methods, 24(12):3055–3072, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kundu:1996:ATL

[KM96] Debasis Kundu and Amit Mitra. Asymptotic theory of least squares estimator of a nonlinear time series regression model. Communications in Statistics: Theory and Methods, 25(1):133– 141, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kolev:1997:DDR

[KM97] Nikolai Kolev and Leda Minkova. Discrete distributions related to success runs of length k in multi-state Markov chain. Communications in Statistics: Theory and Methods, 26(4):1031–1049, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kolev:1999:QRS

[KM99a] Nikolai Kolev and Leda Minkova. Quotas on runs of successes and failures in a multi-state Markov chain. Communications in Statistics: Theory and Methods, 28(9):2235–2248, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kolev:1999:RFQ

[KM99b] Nikolai Kolev and Leda Minkova. Run and frequency quotas in a multi-state Markov chain. Communications in Statistics: Theory and Methods, 28(9):2223–2233, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kowalski:1999:DIP

[KMBTG99] Jeanne Kowalski, Jose R. Mendoza-Blanco, Xim M. Tu, and Leon J. Gleser. On the difference in inference and prediction between the joint and independent F -error models for seemingly unrelated regressions. Communications in Statistics: Theory and Methods, 28(9):2119–2140, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kmietowicz:1995:SSL

[Kmi95] Z. W. Kmietowicz. Significance of the sample lead: some theoretical and empirical results. Communications in Statistics: Theory and Methods, 24(6):1507–1519, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kushner:1991:BCF

[KML91a] H. B. Kushner, Morris Meisner, and Eugene M. Laska. A Bartlett correction factor for a likelihood ratio statistic in multivariate bioassay. Communications in Statistics: Theory and Methods, 20(5–6):1549–1555, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kushner:1991:CRM

[KML91b] H. B. Kushner, Morris Meisner, and Eugene M. Laska. A confidence region with multiple intervals in multivariate bioassay. Communications in Statistics: Theory and Methods, 20(5–6): 1533–1547, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kala:1990:ACI

[KMM90] R. Kala, A. Molińska, and K. Moliński. Approximate confidence intervals for a ratio of two variance components. Communications in Statistics: Theory and Methods, 19(5):1889–1898, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kadiyala:1994:CSL

[KO94a] K. R. Kadiyala and Dennis Oberhelman. A comparison of Steinlike procedures for estimating linear regression models with multicollinear data. Communications in Statistics: Theory and Methods, 23(5):1447–1469, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kozumi:1994:GEM

[KO94b] Hideo Kozumi and Kazuhiro Ohtani. The general expressions for the moments of Lawless and Wang’s ordinary ridge regression estimator. Communications in Statistics: Theory and Methods, 23(10):2755–2774, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kochar:1990:SPO

[Koc90] Subhash C. Kochar. Some partial ordering results on record values. Communications in Statistics: Theory and Methods, 19 (1):299–306, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kocherlakota:1995:DBW

[Koc95] Subrahmaniam Kocherlakota. Discrete bivariate weighted distributions under multiplicative weight function. Communications in Statistics: Theory and Methods, 24(2):533–551, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Koop:1990:SST

[Koo90] J. C. Koop. Systematic sampling of two-dimensional surfaces and related problems. Communications in Statistics: Theory and Methods, 19(5):1701–1750, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Koop:1993:CIA

[Koo93a] J. C. Koop. Conditional inference about an existing universe from independent interpenetrating samples. Communications in Statistics: Theory and Methods, 22(4):985–1006, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kooreman:1993:BRC

[Koo93b] Peter Kooreman. Bounds on the regression coefficients when a covariate is categorized. Communications in Statistics: Theory and Methods, 22(8):2373–2380, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Koti:1990:SIE

[Kot90] Kallappa M. Koti. Some inequalities in elementary special functions with applications to nonparametric statistical inference. Communications in Statistics: Theory and Methods, 19(10): 3671–3676, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Koti:1991:NSQ

[Kot91] Kallappa M. Koti. A note on sum–quota sampling. Communications in Statistics: Theory and Methods, 20(3):1039–1042, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kozek:1990:NTF

[Koz90] Andrzej S. Kozek. A nonparametric test of fit of a linear model. Communications in Statistics: Theory and Methods, 19(1):169– 179, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kozumi:1994:TEM

[Koz94] Hideo Kozumi. Testing equality of the means in two independent multivariate t distributions. Communications in Statistics: Theory and Methods, 23(1):215–227, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Khatri:1992:AMT

[KP92a] C. G. Khatri and H. I. Patel. Analysis of a multicenter trial using a multivariate approach to a mixed linear model. Communications in Statistics: Theory and Methods, 21(1):21–39, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kim:1992:SRT

[KP92b] Byung Soo Kim and Cheolyong Park. Some remarks on testing goodness of fit for the Poisson assumption. Communications in Statistics: Theory and Methods, 21(4):979–995, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Klein:1993:IRA

[KP93a] Ruben Klein and S. James Press. Image reconstruction by adaptive Bayesian classification with a locally proportional prior. Communications in Statistics: Theory and Methods, 22(10): 2925–2940, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kumar:1993:TST

[KP93b] B. Krishna Kumar and P. R. Parthasarathy. On the total sojourn time of particles in semi-Markov compartmental system with branching particles. Communications in Statistics: Theory and Methods, 22(4):1207–1220, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kherwa:1995:NGB

[KP95] G. Ram Kherwa and Jagdish Prasad. A note on generalized binary proper efficiency-balanced block designs. Communications in Statistics: Theory and Methods, 24(11):2797–2801, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Krishnamoorthy:1990:UEC

[KR90] K. Krishnamoorthy and Vijay K. Rohatgi. Unbiased estimation of the common mean of a multivariate normal distribution. Communications in Statistics: Theory and Methods, 19(5):1803– 1810, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Krauth:1992:CID

[Kra92] J. Krauth. Confidence intervals for discrete distributions based on a single observation. Communications in Statistics: Theory and Methods, 21(4):1103–1114, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kreienbrock:1992:GMD

[Kre92] Lothar Kreienbrock. Generalized measures of dispersion to solve the allocation problem in multivariate stratified random sampling. Communications in Statistics: Theory and Methods, 22 (1):219–239, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Krishnamoorthy:1991:ENC

[Kri91] K. Krishnamoorthy. Estimation of normal covariance and precision matrices with incomplete data. Communications in Statistics: Theory and Methods, 20(3):757–770, 1991. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Krouse:1994:PMM

[Kro94a] Donal P. Krouse. Patterned matrices and the misspecification robustness of 2^{k−p} designs. Communications in Statistics: Theory and Methods, 23(11):3285–3301, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Krouse:1994:PDT

[Kro94b] Donal P. Krouse. The power of d’Agostino’s d test of normality against a normal mixture alternative. Communications in Statistics: Theory and Methods, 23(1):47–57, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Krouse:1997:AMB

[Kro97] Donal P. Krouse. Admissible minimum bias estimation of response surfaces. Communications in Statistics: Theory and Methods, 26(8):1937–1949, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kaur:1990:EDF

[KS90a] Amarjot Kaur and Harshinder Singh. Estimation of a distribution function bounded by two known distribution functions. Communications in Statistics: Theory and Methods, 19(9):3343– 3360, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kianifard:1990:MCC

[KS90b] Farid Kianifard and William H. Swallow. A Monte Carlo comparison of five procedures for identifying outliers in linear regression. Communications in Statistics: Theory and Methods, 19(5): 1913–1938, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kwiatkowski:1990:DFT

[KS90c] Denis Kwiatkowski and Peter Schmidt. Dickey-fuller tests with trend. Communications in Statistics: Theory and Methods, 19 (10):3645–3656, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Konishi:1994:MLE

[KS94a] Sadanori Konishi and Kunio Shimizu. Maximum likelihood estimation of an intraclass correlation in a bivariate normal distribution with missing observations. Communications in Statistics: Theory and Methods, 23(6):1593–1604, 1994. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kumar:1994:GMS

[KS94b] E. Vasantha Kumar and T. Srivenkataramana. A generalization of Midzuno–Sen sampling scheme for finite populations. Communications in Statistics: Theory and Methods, 23(9):2541–2559, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kim:1995:NSN

[KS95] Tae-Sung Kim and Hye-Young Seo. A note on some negative dependence notions. Communications in Statistics: Theory and Methods, 24(3):845–858, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Khare:1997:TRT

[KS97] B. B. Khare and S. Srivastava. Transformed ratio type estimators for the population mean in the presence of nonresponse. Communications in Statistics: Theory and Methods, 26(7):1779– 1791, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Khan:1998:CEM

[KS98] Shahjahan Khan and A. K. Md. Ehsanes Saleh. Comparison of estimators of means based on p-samples from multivariate Student-t population. Communications in Statistics: Theory and Methods, 27(1):193–210, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kaciranlar:1998:MSE

[KSA98] Selahattin Kaçiranlar, Sadullah Sakallioğlu, and Fikri Akdeniz. Mean squared error comparisons of the modified ridge regression estimator and the restricted ridge regression estimator. Communications in Statistics: Theory and Methods, 27(1):131–138, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kshirsagar:1998:NTL

[Ksh98] Anant M. Kshirsagar. A non-testable linear hypothesis and its effective equivalent. Communications in Statistics: Theory and Methods, 27(9):2117–2122, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kumar:1996:URP

[KSm96] E. Vasantha Kumar, T. Srivenkataramana, and K. P. Sri math. Use of ranks in probability proportional to size sampling. Communications in Statistics: Theory and Methods, 25(12):3195– 3215, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kooda:1990:EPT

[KSmS90] B. K. Kooda, N. P. Singh, and U. med Singh. Estimates of the parameters of Type-II extreme value distribution from censored samples. Communications in Statistics: Theory and Methods, 19(8):3093–3110, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Khan:1991:FEM

[KT91] Khushnood A. Khan and Derrick S. Tracy. Fourth exact moment results for MRBP tests with two or three treatments. Communications in Statistics: Theory and Methods, 20(12):3863–3877, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kariya:1999:TMA

[KTTL99] Takeaki Kariya, Ruey S. Tsay, Nobuhike Terni, and Hong Li. Tests for multinormality with applications to time series. Communications in Statistics: Theory and Methods, 28(3–4):519– 536, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kubokawa:1991:EEU

[Kub91] Tatsuya Kubokawa. Equivariant estimation under the Pitman closeness criterion. Communications in Statistics: Theory and Methods, 20(11):3499–3523, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kubokawa:1999:SMT

[Kub99] Tatsuya Kubokawa. Shrinkage and modification techniques in estimation of variance and the related problems: A review. Communications in Statistics: Theory and Methods, 28(3–4):613– 650, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kulkarni:1990:RDP

[Kul90] P. M. Kulkarni. Robust Durbin’s and portmanteau goodnessof-fit test statistics and their limit distributions. Communications in Statistics: Theory and Methods, 19(8):2829–2845, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kulinskaya:1994:LSR

[Kul94] Elena Kulinskaya. Large sample results for permutation tests of association. Communications in Statistics: Theory and Methods, 23(10):2939–2963, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kulinskaya:1995:LSR

[Kul95] Elena Kulinskaya. Large sample results for tests of association II. Multinomial and stratified sampling. Communications in Statistics: Theory and Methods, 24(5):1121–1150, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kulldorff:1997:SSS

[Kul97] Martin Kulldorff. A spatial scan statistic. Communications in Statistics: Theory and Methods, 26(6):1481–1496, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kulasekera:1999:CPF

[Kul99] K. B. Kulasekera. Crossing points of failure rates. Communications in Statistics: Theory and Methods, 28(1):87–104, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kumar:1992:IAM

[Kum92] Kuldeep Kumar. Identification of ARMA models with non-Gaussian innovations. Communications in Statistics: Theory and Methods, 21(4):1145–1161, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kumar:1997:CTS

[Kum97] Narinder Kumar. A class of two-sample tests for location based on sub-sample medians. Communications in Statistics: Theory and Methods, 26(4):943–951, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kumar:1998:LE

[Kum98] Narinder Kumar. Letter to the editor. Communications in Statistics: Theory and Methods, 27(8):2115–2116, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kundu:1991:APC

[Kun91] Debasis Kundu. Asymptotic properties of the complex valued non-linear regression model. Communications in Statistics: Theory and Methods, 20(12):3793–3803, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kunst:1993:SCB

[Kun93] Robert M. Kunst. Stability conditions for a bivariate arch system which is cointegrated in mean. Communications in Statistics: Theory and Methods, 22(10):2941–2953, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kunbu:1995:CUE

[Kun95] Debasis Kunbu. Consistency of the undamped exponential signal model on a restricted parameter space. Communications in Statistics: Theory and Methods, 24(1):241–251, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kunst:1997:FOM

[Kun97] Robert M. Kunst. Fourth-order moments of augmented arch processes. Communications in Statistics: Theory and Methods, 26(6):1425–1441, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Knautz:1993:RLM

[KW93] Henning Knautz and Joachim Wilde. R^{2} in restricted linear models. Communications in Statistics: Theory and Methods, 22 (3):641–645, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Krijnen:1996:BLP

[KWtB96] Wim P. Krijnen, Tom Wansbeek, and Jos M. F. ten Berge. Best linear predictors for factor scores. Communications in Statistics: Theory and Methods, 25(12):3013–3025, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kikuchi:1991:IHC

[KY91a] Yasuki Kikuchi and Takashi Yanagawa. Incorporating historical controls using a random-effects model with a normal prior. Communications in Statistics: Theory and Methods, 20(4):1273– 1291, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ko:1991:CSA

[KY91b] Jang-Hwan Ko and Bong-Jin Yum. The chi-square approximation approach to ranking and selection procedures under type II censoring. Communications in Statistics: Theory and Methods, 20(3):1127–1147, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Kshirsagar:1993:ERP

[KY93] Anant M. Kshirsagar and Weiying Yuan. Estimation of relative potency in multivariate parallel line bioassays. Communications in Statistics: Theory and Methods, 22(12):3355–3361, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lee:1990:MLE

[LAC90] Tao-Lin Lee, Rohan Attele, and Keh-Wei Chen. Maximum likelihood estimation in convex hull models. Communications in Statistics: Theory and Methods, 19(10):3759–3777, 1990. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Laitila:1998:AHO

[Lai98] Thomas Laitila. Adaptation of Honda’s one-sided test of random effects to repeated measurements experiments. Communications in Statistics: Theory and Methods, 27(12):3015–3033, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Laib:1999:ETI

[Lai99] Naâmane Laib. Exponential-type inequalities for martingale difference sequences. Application to nonparametric regression estimation. Communications in Statistics: Theory and Methods, 28(7):1565–1576, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lavine:1992:NBM

[Lav92] Michael Lavine. A note on bounding Monte Carlo variances. Communications in Statistics: Theory and Methods, 21(10): 2855–2860, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lawoko:1990:MLE

[Law90] Charles R. O. Lawoko. On the maximum likelihood estimation of parameters of a spatial discrimination model. Communications in Statistics: Theory and Methods, 19(12):4627–4641, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lawson:1993:CIS

[Law93] Andrew Lawson. On composite intensity score tests. Communications in Statistics: Theory and Methods, 22(11):3223–3235, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lawson:1997:ARD

[Law97] Andrew B. Lawson. An angular-radial dependence model for point process data. Communications in Statistics: Theory and Methods, 26(1):237–245, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lopez-Blazques:1998:NUE

[LB98] F. López-Blázques. A note on UMVU estimation in the transformed chi-square family. Communications in Statistics: Theory and Methods, 27(8):2065–2073, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lien:1992:MOS

[LBB92] Da-Hsiang Donald Lien, N. Balakrishnan, and K. Balasubramanian. Moments of order statistics from a non-overlapping mixture model with applications to truncated Laplace distribution. Communications in Statistics: Theory and Methods, 21(7): 1909–1928, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lopez-Blazquez:1997:UEC

[LBBCMR97] F. López-Blázquez, I. Barranco-Chamorro, and J. L. Moreno-Rebollo. UMVU estimation for certain family of exponential distributions. Communications in Statistics: Theory and Methods, 26(2):469–482, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lachenbruch:1991:AKW

[LC91] Peter A. Lachenbruch and Philip J. Clements. ANOVA, Kruskal–Wallis, normal scores and unequal variance. Communications in Statistics: Theory and Methods, 20(1):107–126, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Liu:1992:AVB

[LC92] Jen-Pei Liu and Shein-Chung Chow. On the assessment of variability in bioavailability/bioequivalence studies. Communications in Statistics: Theory and Methods, 21(9):2591–2607, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lai:1996:NAD

[LC96a] Dejian Lai and Wenyaw Chan. A note on asymptotic distribution of permutation statistics from mixing processes. Communications in Statistics: Theory and Methods, 25(10):2395–2401, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lee:1996:MVC

[LC96b] Hyan Suk Lee and Yogendra P. Chaubey. MINQUE of variance components in generalized linear model with random effects. Communications in Statistics: Theory and Methods, 25(6): 1375–1382, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Li:1999:NDS

[LC99] Shaoyi Li and Hubert J. Chen. A note on the determination of sample sizes for hypergeometric distributions. Communications in Statistics: Theory and Methods, 28(7):1749–1757, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Luceno:1995:DEB

[LD95] Alberto Luceño and Federico De Ceballos. Describing extrabinomial variation with partially correlated models. Communications in Statistics: Theory and Methods, 24(6):1637–1653, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Li:1999:FAR

[LD99a] Baibing Li and Bart De Moor. A family of adaptive robust estimates based on the symmetrical generalized logistic distribution. Communications in Statistics: Theory and Methods, 28(6): 1293–1310, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Li:1999:MLI

[LD99b] Baibing Li and Bart De Moor. More on local influence in principal components analysis. Communications in Statistics: Theory and Methods, 28(10):2487–2495, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Leite:1992:WTE

[LdBPR92] José Galvao Leite, Carlos Alberto de Bragança Pereira, and Flávio Wagner Rodrigues. Waiting time to exhaust lottery numbers. Communications in Statistics: Theory and Methods, 22 (1):301–310, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Le:1994:STL

[Le94] Chap T. Le. Some tests for linear trend of variances. Communications in Statistics: Theory and Methods, 23(8):2269–2282, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lee:1990:BLR

[Lee90] Kee-Won Lee. Bootstrapping logistic regression models with random regressors. Communications in Statistics: Theory and Methods, 19(7):2527–2539, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lee:1994:EMD

[Lee94a] H. S. Lee. Estimates for mean and dispersion effects in unreplicated factorial designs. Communications in Statistics: Theory and Methods, 23(12):3593–3608, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lee:1994:TSM

[Lee94b] Mei-Ling Ting Lee. A two-state Markov model. Communications in Statistics: Theory and Methods, 23(2):615–623, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lee:1995:CLG

[Lee95a] Austin F. S. Lee. Coefficients of Lee–Gurland two-sample test on normal means. Communications in Statistics: Theory and Methods, 24(7):1743–1768, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lee:1995:FID

[Lee95b] C. P. Lee. Forecasting with incomplete data. Communications in Statistics: Theory and Methods, 24(5):1255–1269, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lee:1996:PAS

[Lee96] Mei-Ling Ting Lee. Properties and applications of the Sarmanov family of bivariate distributions. Communications in Statistics: Theory and Methods, 25(6):1207–1222, 1996. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lee:1997:REM

[Lee97] Hyun Suk Lee. Random effects models for count data. Communications in Statistics: Theory and Methods, 26(8):1893–1904, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lee:1998:DDP

[Lee98] Sunho Lee. Detecting departures from a Poisson model. Communications in Statistics: Theory and Methods, 27(5):1201–1210, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lee:1999:SSP

[Lee99] Oesook Lee. Strict stationarity of AR(p) processes generated by nonlinear random functions with additive perturbations. Communications in Statistics: Theory and Methods, 28(11):2527– 2537, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Leipnik:1990:MRE

[Lei90] Roy B. Leipnik. A maximum relative entropy principle for distribution of personal income with derivations of several known inc.ome distributions. Communications in Statistics: Theory and Methods, 19(3):1003–1036, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Letellier:1997:RCE

[Let97] Julie A. Letellier. Rates of convergence for an estimator of a density function based on Jacobi polynomials. Communications in Statistics: Theory and Methods, 26(1):197–220, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Leung:1992:PEM

[Leu92a] Chi-Ying Leung. Proportion estimation in mixtures with covariates. Communications in Statistics: Theory and Methods, 21(11):3123–3141, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Leung:1992:EES

[Leu92b] Pui Lam Leung. Estimation of eigenvalues of the scale matrix of the multivariate F distribution. Communications in Statistics: Theory and Methods, 21(7):1845–1856, 1992. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Leung:1994:CDC

[Leu94a] Chi-Ying Leung. Classification of dichotomous and continuous variables with incomplete samples. Communications in Statistics: Theory and Methods, 23(6):1581–1592, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Leung:1994:IEF

[Leu94b] Pui Lam Leung. Improved estimation of functions of the canonical correlation coefficients. Communications in Statistics: Theory and Methods, 23(3):831–840, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Levin:1992:TOR

[Lev92] Bruce Levin. Tests of odds ratio homogeneity with improved power in sparse fourfold tables. Communications in Statistics: Theory and Methods, 21(6):1469–1500, 1992. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Leysieffer:1990:SFY

[Ley90] Frederick W. Leysieffer. The second fifteen years: a continuation of the chronicle of the Department of Statistics at the Florida State University. Communications in Statistics: Theory and Methods, 19(11):3961–3983, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lee:1991:FTB

[LF91] Austin F. S. Lee and Naomi S. Fineberg. A fitted test for the Behrens–Fisher problem. Communications in Statistics: Theory and Methods, 20(2):653–666, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Landon:1993:ECB

[LF93] R. Landon and L. S. Freedman. Estimating correlations between questionnaires and food records/true dietary intake. Communications in Statistics: Theory and Methods, 22(8):2327–2340, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Logan:1993:BDU

[LG93] T. P. Logan and A. K. Gupta. Bayesian discrimination using multiple observations. Communications in Statistics: Theory and Methods, 22(6):1735–1754, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Le:1994:TAB

[LG94] Chap T. Le and Patricia M. Grambsch. Tests of association between survival time and a continuous covariate. Communications in Statistics: Theory and Methods, 23(4):1009–1019, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Long:1998:ALC

[LG98] Tianyong Long and Rameshwar D. Gupta. Alternative linear classification rules under order restrictions. Communications in Statistics: Theory and Methods, 27(3):559–575, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lin:1993:OEF

[LH93] Yan-Xia Lin and C. C. Heyde. Optimal estimating functions and Wedderburn’s quasi-likelihood. Communications in Statistics: Theory and Methods, 22(8):2341–2350, 1993. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Liang:1999:LST

[LH99] Hua Liang and Wolfgang Hardle. Large sample theory of the estimation of the error distribution for a semiparametric model. Communications in Statistics: Theory and Methods, 28(9):2025– 2036, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Li:1994:PIT

[Li94] W. K. Li. Parameter inference for time series with regular and seasonal unit roots. Communications in Statistics: Theory and Methods, 23(3):721–733, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Li:1995:IFD

[Li95] Guibin Li. Identification of fractional differencing autoregressive models. Communications in Statistics: Theory and Methods, 24(10):2635–2643, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Li:1998:SIS

[Li98] Lung-An Li. Some issues in stochastic curtailing. Communications in Statistics: Theory and Methods, 27(6):1409–1418, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Li:1999:GSC

[Li99] Zhengqing Li. Group sequential checking of the proportional hazards assumption based on a time-dependent covariate. Communications in Statistics: Theory and Methods, 28(8):1931– 1943, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Liang:1990:EBE

[Lia90] TaChen Liang. Empirical Bayes estimation of binomial parameter with symmetric priors. Communications in Statistics: Theory and Methods, 19(5):1671–1683, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Liang:1994:BEB

[Lia94] Hua Liang. The Berry–Esséen bounds of error variance estimation in semiparametric regression models. Communications in Statistics: Theory and Methods, 23(12):3439–3451, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Liao:1999:OTL

[Lia99] C. T. Liao. Orthogonal three-level parallel flats designs for user-specified resolution. Communications in Statistics: Theory and Methods, 28(8):1945–1960, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lim:1993:SET

[Lim93] K. S. Lim. On the stability and ergodicity of a threshold AR(l) without intercepts. Communications in Statistics: Theory and Methods, 22(2):369–373, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Link:1993:EOA

[Lin93] William A. Link. Efficiency and optimal allocation in the staggered entry design. Communications in Statistics: Theory and Methods, 22(2):485–503, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lin:1995:ACB

[Lin95] Yachen Lin. Asymptotic confidence bands for normal distributions. Communications in Statistics: Theory and Methods, 24 (3):641–652, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Linxiong:1997:LSN

[Lin97] Li Linxiong. Large sample nonparametric estimation of the mean residual life. Communications in Statistics: Theory and Methods, 26(5):1165–1182, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lindkvist:1999:CEI

[Lin99] Marie Lindkvist. Case and event influence on the proportional hazards assumption. Communications in Statistics: Theory and Methods, 28(10):2347–2366, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Liseo:1992:NCA

[Lis92] Brunero Liseo. A note on a counterexample against the likelihood principle. Communications in Statistics: Theory and Methods, 21(2):547–556, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Liu:1990:MPE

[Liu90] Kung-Jong Liu. Modified product estimators of finite population mean in finite sampling. Communications in Statistics: Theory and Methods, 19(10):3799–3807, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Liu:1991:BOW

[Liu91] Zhijun Liu. Bootstrapping one way analysis of Rao’s quadratic entropy. Communications in Statistics: Theory and Methods, 20(5–6):1683–1703, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Liu:1992:PPT

[Liu92a] Jingyu Liu. Precedence probabilities and their applications. Communications in Statistics: Theory and Methods, 21(6):1667– 1682, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Liu:1992:PS

[Liu92b] W. Liu. Predictive screening. Communications in Statistics: Theory and Methods, 21(8):2349–2366, 1992. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Liu:1992:NEN

[Liu92c] Z. J. Liu. Nonparametric estimates of the nuisance parameter in the LAD tests. Communications in Statistics: Theory and Methods, 21(4):861–881, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Liu:1993:CPC

[Liu93a] W. Liu. Conditional probability of correct selection under the continuum partition with applications. Communications in Statistics: Theory and Methods, 22(4):1085–1111, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Liu:1993:SPI

[Liu93b] W. Liu. Simultaneous prediction intervals for all distances from the “best”. Communications in Statistics: Theory and Methods, 22(2):505–518, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Liu:1994:SBS

[Liu94a] Han-Chu Liu. Selecting the best and the second best multinomial cells. Communications in Statistics: Theory and Methods, 23(4): 1191–1201, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Liu:1994:NSP

[Liu94b] W. Liu. A new sequential procedure with elimination for selecting the best of several normal populations. Communications in Statistics: Theory and Methods, 23(11):3157–3170, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Liu:1995:CFA

[Liu95a] Shu-Ing Liu. Comparison of forecasts for ARMA models between a random coefficient approach and a Bayesian approach. Communications in Statistics: Theory and Methods, 24(2):319–333, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Liu:1995:CEL

[Liu95b] Wei Liu. On the construction of exact lower confidence bounds on the distance between any two ranked location parameters of K populations. Communications in Statistics: Theory and Methods, 24(8):2073–2085, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Liu:1996:NPF

[Liu96] Wei Liu. A note on a power function of Tukey’ test for comparing three normal means with unequal sample sizes. Communications in Statistics: Theory and Methods, 25(7):1611–1616, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Liu:1997:MLE

[Liu97a] Aiyi Liu. On the maximum likelihood estimate for the drift of Brownian motion following a symmetric sequential probability ratio test. Communications in Statistics: Theory and Methods, 26(4):977–989, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Liu:1997:ROD

[Liu97b] Shawn X. Liu. Restricted optimal designs for linear and quadratic polynomial regressions. Communications in Statistics: Theory and Methods, 26(1):33–46, 1997. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Long:1996:GCC

[LK96] Dou Long and Roman Krzysztofowicz. Geometry of a correlation coefficient under a copula. Communications in Statistics: Theory and Methods, 25(6):1397–1404, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lyles:1998:UEP

[LK98] Robert H. Lyles and Lawrence L. Kupper. Umvu estimators for the population mean and variance based on random effects models for lognormal data. Communications in Statistics: Theory and Methods, 27(4):795–818, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Leu:1990:SBP

[LL90] Lii-Yuh Leu and TaChen Liang. Selection of the best with a preliminary test for two-parameter exponential distributions. Communications in Statistics: Theory and Methods, 19(4):1443– 1455, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Liu:1991:SSC

[LL91] Han-Chu Liu and Jer-Yan Lin. On selecting a subset containing the most probable multinomial event. Communications in Statistics: Theory and Methods, 20(10):3381–3387, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ling:1993:SLW

[LL93] K. D. Ling and T. Y. Low. On the soonest and latest waiting time distributions: succession quotas. Communications in Statistics: Theory and Methods, 22(8):2207–2221, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lin:1994:EBR

[LL94] I-Chin Lin and Lii-Yuh Leu. Empirical Bayes rules for selecting the best exponential population with location parameter. Communications in Statistics: Theory and Methods, 23(6):1797– 1809, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lo:1996:DTE

[LL96] Milton Lo and Pui Lam Leung. Decision theoretic estimation of functions of the canonical correlation coefficients. Communications in Statistics: Theory and Methods, 25(9):1985–1995, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lou:1998:NGW

[LL98] W. Y. Wendy Lou and K. K. Gordon Lan. A note on the gehan–wilcoxon statistic. Communications in Statistics: Theory and Methods, 27(6):1453–1459, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lipsitz:1990:UJE

[LLH90] Stuart R. Lipsitz, Nan M. Laird, and David P. Harrington. Using the jackknife to estimate the variance of regression estimators from repeated measures studies. Communications in Statistics: Theory and Methods, 19(3):821–845, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Levine:1992:PPC

[LLL92] A. Levine, J. Liukkonen, and D. W. Levine. Predicting power changes under transformations in ANOVA tests. Communications in Statistics: Theory and Methods, 21(3):679–692, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Levine:1993:DCT

[LLL93] A. Levine, J. Liukkonen, and D. W. Levine. Decision and choice in transforming data. Communications in Statistics: Theory and Methods, 22(4):1221–1233, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Levine:1996:EIU

[LLL96] A. Levine, J. Liukkonen, and D. W. Levine. Equivalent inferences using transformations. Communications in Statistics: Theory and Methods, 25(5):1059–1072, 1996. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lloyd:1995:ACI

[Llo95] Christopher J. Lloyd. Accurate confidence intervals from recapture experiments. Communications in Statistics: Theory and Methods, 24(5):1235–1253, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lloyd:1999:EFD

[Llo99] Chris J. Lloyd. Estimating a frequency distribution when the sampling is biased. Communications in Statistics: Theory and Methods, 28(5):1115–1132, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lim:1998:BIR

[LLP98] Jae-Hak Lim, Kun-Liang Lu, and Dong Ho Park. Bayesian imperfect repair model. Communications in Statistics: Theory and Methods, 27(4):965–984, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lesaffre:1993:CGL

[LM93] Emmanuel Lesaffre and Brian D. Marx. Collinearity in generalized linear regression. Communications in Statistics: Theory and Methods, 22(7):1933–1952, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lu:1994:EGL

[LM94] Wang-Shu Lu and Carl N. Morris. Estimation in the generalized linear empirical Bayes model using the extended quasilikelihood. Communications in Statistics: Theory and Methods, 23(3):661–687, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lalitha:1996:MML

[LM96] S. Lalitha and Anand Mishra. Modified maximum likelihood estimation for Rayleigh distribution. Communications in Statistics: Theory and Methods, 25(2):389–401, 1996. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Legault:1997:AME

[LM97] Claudine Legault and Timothy M. Morgan. Analyzing multiple end points with a two-stage design in clinical trials. Communications in Statistics: Theory and Methods, 26(8):1951–1969, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Levy:1990:TLF

[LN90] Martin S. Levy and James W. Neill. Testing for lack of fit in linear multiresponse models based on exact or near replicates. Communications in Statistics: Theory and Methods, 19(6):1987– 2002, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Locke:1990:UMG

[Loc90] Charles S. Locke. Use of a more general model for bioavailability studies. Communications in Statistics: Theory and Methods, 19(9):3361–3373, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lombard:1999:RTC

[Lom99] F. Lombard. A rank test for constancy of a location oil scale parameter. Communications in Statistics: Theory and Methods, 28(3–4):537–547, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lopes:1993:AEM

[Lop93] Silvia Lopes. Amplitude estimation in multiple frequency spectrum. Communications in Statistics: Theory and Methods, 22 (10):2955–2967, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Louani:1998:ANK

[Lou98] Djamal Louani. On the asymptotic normality of the kernel estimators of the density function and its derivatives under censoring. Communications in Statistics: Theory and Methods, 27(12):

2909–2924, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Loynes:1990:VRD

[Loy90] R. M. Loynes. The variance and the range of I.I.D. random variables. Communications in Statistics: Theory and Methods, 19(4):1419–1432, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Liao:1994:MMS

[LP94] Pen-Yuan Liao and Elinor S. Pape. Mixture model for sampling efficiency of an alternating Poisson process. Communications in Statistics: Theory and Methods, 23(6):1699–1703, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lee:1998:RBF

[LP98a] Sang Eun Lee and S. James Press. Robustness of Bayesian factor analysis estimates. Communications in Statistics: Theory and Methods, 27(8):1871–1893, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lim:1998:FTT

[LP98b] Jae-Hak Lim and Dong Ho Park. A family of tests for trend change in mean residual life. Communications in Statistics: Theory and Methods, 27(5):1163–1179, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lee:1995:TDO

[LPK95] Sunho Lee, Cheolyong Park, and Bynng Soo Kim. Tests for detecting overdispersion in Poisson models. Communications in Statistics: Theory and Methods, 24(9):2405–2420, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lakshminarayana:1999:BPD

[LPR99] J. Lakshminarayana, S. N. N. Pandit, and K. Srinivasa Rao. On a bivariate Poisson distribution. Communications in Statistics: Theory and Methods, 28(2):267–276, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lien:1992:ICP

[LR92] Da-Hsiang Donald Lien and David Rearden. On the information content of partial observations. Communications in Statistics:

Theory and Methods, 21(10):2869–2880, 1992. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Loganathan:1993:BEM

[LR93] A. Loganathan and M. Rajagopalan. Bayes estimates of mixing parameters in finite mixture distributions. Communications in Statistics: Theory and Methods, 22(11):3013–3029, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lien:1996:PCV

[LR96] Donald Lien and David Rearden. A property of coefficients of variation for ordered bivariate log-normal variables. Communications in Statistics: Theory and Methods, 25(8):1903–1916, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lien:1998:CME

[LR98] Donald Lien and David Rearden. Cross moments of extreme observations from a multivariate lognormal distribution. Communications in Statistics: Theory and Methods, 27(3):601–607, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lan:1994:IIF

[LRD94] K. K. Gordon Lan, David M. Reboussin, and David L. DeMets. Information and information fractions for design and sequential monitoring of clinical trials. Communications in Statistics: Theory and Methods, 23(2):403–420, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lin:1992:CPT

[LS92] C. H. Lin and Shashikala Sukhatme. On the choice of precedence tests. Communications in Statistics: Theory and Methods, 21(10):2949–2968, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lee:1994:CRP

[LS94] Chu-In Charles Lee and Shi-Yun Shen. Convergence rates and powers of six power-divergence statistics for testing independence in 2 by 2 contingency table. Communications in Statistics: Theory and Methods, 23(7):2113–2126, 1994. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lee:1997:MLE

[LS97a] Jong Hyup Lee and Dong Wan Shin. Maximum likelihood estimation for ARMA models in the presence of ARMA errors. Communications in Statistics: Theory and Methods, 26(5):1057– 1072, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Li:1997:OFD

[LS97b] Junfang Li and Jagdish N. Srivastava. Optimal 2^{n} factorial designs of parallel flats type. Communications in Statistics: Theory and Methods, 26(10):2473–2488, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Li:1994:BCU

[LSR94] Yuanzhang Li, Li Shi, and H. Daniel Roth. The bias of the commonly-used estimate of variance in meta-analysis. Communications in Statistics: Theory and Methods, 23(4):1063–1085, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lin:1990:PCC

[LSU90] Fu-Jen Lin, Eila Seppänen, and Esa Uusipaikka. On pairwise comparisons of the components of the mean vector in multivariate normal distribution. Communications in Statistics: Theory and Methods, 19(2):395–411, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lee:1992:ASE

[LT92] Sik-Yum Lee and Man-Lai Tang. Analysis of structural equation models with incomplete polytomous data. Communications in Statistics: Theory and Methods, 21(1):213–232, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Luo:1994:RCS

[LTCC94] Xiaolong Luo, Bruce W. Turnbull, Haiyan Cai, and Larry C. Clark. Regression for censored survival data with lag effects. Communications in Statistics: Theory and Methods, 23(12): 3417–3438, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lauer:1997:OWB

[LTvC^{+}97] S. Lauer, J. Timmer, D. van Calker, D. Maier, and J. Honerkamp. Optimal weighted Bayesian design applied to doseresponse-curve analysis. Communications in Statistics: Theory and Methods, 26(12):2879–2903, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lu:1990:LSE

[Lu90] Jye-Chyi Lu. Least squares estimation for the multivariate Weibull model of Bougaard based on accelerated life test of system and component. Communications in Statistics: Theory and Methods, 19(10):3725–3739, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lu:1993:ENR

[Lu93] Wang-Shu Lu. Estimations in the normal regression empirical Bayes model. Communications in Statistics: Theory and Methods, 22(6):1773–1794, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Luceno:1992:NFP

[Luc92a] Alberto Luceño. A new family of probability distributions with applications to data analysis. Communications in Statistics: Theory and Methods, 21(2):391–409, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Luceno:1992:PEV

[Luc92b] Alberto Luceño. Performance of EWMA versus last observation for feedback control. Communications in Statistics: Theory and Methods, 22(1):241–255, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lucas:1997:ARL

[Luc97a] André Lucas. Asymptotic robustness of least median of squares for autoregressions with additive outliers. Communications in Statistics: Theory and Methods, 26(10):2363–2380, 1997. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lucas:1997:RSB

[Luc97b] André Lucas. Robustness of the Student t based M-estimator. Communications in Statistics: Theory and Methods, 26(5):1165– 1182, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Luong:1994:MQD

[Luo94] Andrew Luong. Minimum quadratic distances methods for censored data. Communications in Statistics: Theory and Methods, 23(6):1605–1615, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lei:1998:LMC

[LV98] Ding-Hwa Lei and Stephen B. Vardeman. The LRT method of constructing a two-sided “variables” acceptance region and its comparison with other methods. Communications in Statistics: Theory and Methods, 27(2):329–351, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lahiff:1990:EIF

[LW90] Maureen Lahiff and Kathleen M. Whitcomb. Empirical influence functions for misclassification rates in discriminant analysis. Communications in Statistics: Theory and Methods, 19(8): 2999–3009, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Liu:1994:ELT

[LW94] Jen-Pei Liu and Chung-Sing Weng. Evaluation of logtransformation in assessing bioequivalence. Communications in Statistics: Theory and Methods, 23(2):421–434, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lee:1995:CLN

[LW95] Seungyeoun Lee and Robert A. Wolfe. A class of linear nonparametric rank tests for dependent censoring. Communications in Statistics: Theory and Methods, 24(3):687–712, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lin:1998:ESE

[LWH98] Stan C. Lin, David M. Whipple, and Charles S. Ho. Evaluation of statistical equivalence using limits of agreement and associated sample size calculation. Communications in Statistics: Theory and Methods, 27(6):1419–1432, 1998. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lam:1994:PAS

[LY94] K. Lam and P. L. H. Yu. A predictive approach for the selection of a fixed number of good treatments. Communications in Statistics: Theory and Methods, 23(9):2469–2492, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Linxiong:1997:SCE

[LY97] Li Linxiong and Qiqing Yu. Self-consistent estimators of survival functions with doubly-censored data. Communications in Statistics: Theory and Methods, 26(11):2609–2621, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lye:1991:IUI

[Lye91] J. N. Lye. Implications of uncorrelatedness and independence in linear regression models. Communications in Statistics: Theory and Methods, 20(8):2407–2430, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lee:1996:CRT

[LYS96] Chu-In Charles Lee, Xiaosong Yan, and Ning-Zhong Shi. On the consistency of rank transformed regression. Communications in Statistics: Theory and Methods, 25(9):2159–2168, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lee:1996:SPG

[LZ96a] Andy H. Lee and Yuejen Zhao. Sensitivity of Pearson’s goodnessof-fit statistic in generalized linear models. Communications in Statistics: Theory and Methods, 25(1):143–157, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Liang:1996:ALE

[LZ96b] Hua Liang and Guohua Zou. Asymptotically linear estimation in a generalized linear model. Communications in Statistics: Theory and Methods, 25(5):937–945, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Liang:1998:ANS

[LZ98a] Hua Liang and Yong Zhou. Asymptotic normality in a semiparametric partially linear model with right-censored data. Communications in Statistics: Theory and Methods, 27(12):2895–2907, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Liang:1998:MEE

[LZ98b] Hua Liang and Yong Zhou. A modified estimator of error variance in a partly linear model. Communications in Statistics: Theory and Methods, 27(4):819–825, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Liang:1999:AEE

[LZZ99] Hua Liang, Lixing Zhu, and Yong Zhou. Asymptotically efficient estimation based on wavelet of expectation value in a partial linear model. Communications in Statistics: Theory and Methods, 28(9):2045–2055, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mukhopadhyay:1992:SSM

[MA92] N. Mukhopadhyay and A. D. Abid. Sequential sampling methodologies for comparing strata means. Communications in Statistics: Theory and Methods, 21(2):485–500, 1992. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Macchiavelli:1994:VOA

[MA94] Raul E. Macchiavelli and Steven F. Arnold. Variable order ante-dependence models. Communications in Statistics: Theory and Methods, 23(9):2683–2699, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Macchiavelli:1995:DUP

[MA95] Raül E. Macchiavelli and Steven F. Arnold. Difficulties with the use of penalized likelihood criteria in antedependence and polynomial models. Communications in Statistics: Theory and Methods, 24(2):501–522, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

MacGillivray:1992:SPJ

[Mac92] H. L. MacGillivray. Shape properties of the g-and-h and Johnson families. Communications in Statistics: Theory and Methods, 21(5):1233–1250, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mack:1998:TSC

[Mac98] Y. P. Mack. Testing for the shoulder condition in transect sampling. Communications in Statistics: Theory and Methods, 27 (2):423–432, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Madhusudan:1992:DNO

[Mad92] Bhandary Madhusudan. Detection of the numbers of outliers present in a data set using an information theoretic criterion. Communications in Statistics: Theory and Methods, 21(11): 3263–3274, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Madi:1993:MSS

[Mad93a] Mohamed T. Madi. Multiple step-stress accelerated life test: the tampered failure rate model. Communications in Statistics: Theory and Methods, 22(9):295–306, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Madsen:1993:GBD

[Mad93b] Richard W. Madsen. Generalized binomial distributions. Communications in Statistics: Theory and Methods, 22(11):3065– 3086, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Maeshiro:1990:PBP

[Mae90] Asatoshi Maeshiro. Peculiar bias properties of the OLS estimator when applied to a dynamic model with autocorrelated disturbances. Communications in Statistics: Theory and Methods, 19(4):1185–1204, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Maesono:1996:NBJ

[Mae96] Yoshihiko Maesono. A note on biases of jackknife estimators of third central moments. Communications in Statistics: Theory and Methods, 25(8):1931–1942, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Maesono:1998:ARR

[Mae98] Yoshihiko Maesono. An asymptotic representation of a ratio of two statistics and its applications. Communications in Statistics: Theory and Methods, 27(2):305–327, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Makov:1991:BAB

[Mak91] Udi E. Makov. Bayesian and approximate Bayesian solutions to simultaneous estimation of multiple dynamic processes. Communications in Statistics: Theory and Methods, 20(3):851–871, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mak:1994:MLE

[Mak94] T. K. Mak. Maximum likelihood estimation of familial correlations. Communications in Statistics: Theory and Methods, 23(8):2351–2365, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mansouri:1991:DFT

[Man91] Hossein Mansouri. Distribution-free tests in randomized blocks. Communications in Statistics: Theory and Methods, 20(9):2753– 2773, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Madigan:1996:BMA

[MAPV96] David Madigan, Steen A. Andersson, Michael D. Perlman, and Chris T. Volinsky. Bayesian model averaging and model selection for Markov equivalence classes of acyclic digraphs. Communications in Statistics: Theory and Methods, 25(11):2493–2519, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mariotto:1990:SBE

[Mar90a] Angela B. Mariotto. Sensitivity of Bayes and empirical Bayes estimates. Communications in Statistics: Theory and Methods, 19(12):4471–4484, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Martin:1990:UTS

[Mar90b] R. J. Martin. The use of time-series models and methods in the analysis of agricultural field trials. Communications in Statistics: Theory and Methods, 19(1):55–81, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Martin:1992:ADT

[Mar92a] R. J. Martin. Approximations to the determinant term in Gaussian maximum likelihood estimation of some spatial models.

Communications in Statistics: Theory and Methods, 22(1):189– 205, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Martin:1992:LIR

[Mar92b] R. J. Martin. Leverage, influence and residuals in regression models when observations are correlated. Communications in Statistics: Theory and Methods, 21(5):1183–1212, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Martin:1995:SRR

[Mar95] R. J. Martin. Some results on randomization when observations are spatially dependent and the analysis uses generalized leastsquares. Communications in Statistics: Theory and Methods, 24(9):2199–2210, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Markatou:1996:RSI

[Mar96] Marianthi Markatou. Robust statistical inference: weighted likelihoods or usual m-estimation? Communications in Statistics: Theory and Methods, 25(11):2597–2613, 1996. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Martin:1997:LOS

[Mar97] R. J. Martin. Low-order spatially-distinct latin squares. Communications in Statistics: Theory and Methods, 26(6):1511–1539, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mason:1991:C

[Mas91] Robert L. Mason. Comments. Communications in Statistics: Theory and Methods, 20(11):3453–3456, 1991. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Misra:1994:ELL

[MAS94] Neeraj Misra, Rajesh Anand, and Harshinder Singh. Estimation of the largest location parameter of exponential distributions. Communications in Statistics: Theory and Methods, 23 (10):2865–2880, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mathai:1991:FMQ

[Mat91] A. M. Mathai. On fractional moments of quadratic expressions in normal variables. Communications in Statistics: Theory and Methods, 20(10):3159–3174, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Matsuda:1998:DSF

[Mat98] Yasumasa Matsuda. A diagnostic statistic for functionalcoefficient autoregressive models. Communications in Statistics: Theory and Methods, 27(9):2257–2273, 1998. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Matsuda:1999:TLA

[Mat99] Yasumasa Matsuda. A test of linearity against functional coefficient autoregressive models. Communications in Statistics: Theory and Methods, 28(11):2539–2551, 1999. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mason:1991:RRE

[MB91a] Robert L. Mason and Neil W. Blaylock, Jr. Ridge regression estimator comparisons using Pitman’s measure of closeness. Communications in Statistics: Theory and Methods, 20(11):3629– 3641, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

McDonald:1991:OPG

[MB91b] James B. McDonald and Richard M. Bookstaber. Option pricing for generalized distributions. Communications in Statistics: Theory and Methods, 20(12):4053–4068, 1991. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Maine:1991:DNC

[MBR91] Mary Maine, Thomas Boullion, and Gaspard T. Rizzuto. Detecting the number of components in a finite mixture having normal components. Communications in Statistics: Theory and Methods, 20(2):611–620, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mayer:1991:MCR

[MC91] Lawrence S. Mayer and Steven S. Carroll. Modeling the contemporaneous relationship in a continuous variable panel model. Communications in Statistics: Theory and Methods, 20(5–6):

1593–1617, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mukherjee:1992:SDH

[MC92] S. P. Mukherjee and A. Chatterjee. Stochastic dominance of higher orders and its implications. Communications in Statistics: Theory and Methods, 21(7):1977–1986, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Chen:1996:RLE

[mC96] Sy mien Chen. Response to the “Letter to the editor”. Communications in Statistics: Theory and Methods, 25(6):1407–1408, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Muller:1997:DCS

[McM97] Keith E. Muller and Mario chen Mok. The distribution of Cook’s d statistic. Communications in Statistics: Theory and Methods, 26(3):525–546, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Muttlak:1990:RSS

[MD90] Hassen A. Muttlak and Lynan L. Mc Donald. Ranked set sampling with respect to concomitant variables and with size biased probability of selection. Communications in Statistics: Theory and Methods, 19(1):205–219, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

MimosoPaulino:1992:BAC

[MD92] Carlos Daniel Mimoso Paulino and Carlos Alberto De Bragança Pereira. Bayesian analysis of categorical data informatively censored. Communications in Statistics: Theory and Methods, 21(9):2689–2705, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Miazaki:1993:EPT

[MD93a] E. S. Miazaki and C. C. Y. Dorea. Estimation of the parameters of a time series subject to the error of rotation sampling. Communications in Statistics: Theory and Methods, 22(3):805–825, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Misra:1993:SCI

[MD93b] Neeraj Misra and Ishwari D. Dimriya. Simultaneous confidence intervals for all distances from the worst and best populations. Communications in Statistics: Theory and Methods, 22(11): 3097–3116, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Martin:1994:ACP

[MD94a] R. J. Martin and J. N. Dwyer. Approximations to the covariance properties of processes averaged over irregular spatial regions. Communications in Statistics: Theory and Methods, 23(3):913– 945, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Misra:1994:SSE

[MD94b] Neeraj Misra and Ishwari D. Dhariyal. Simultaneous selection of extreme populations: Bayes-p^{∗} and minimax rules. Communications in Statistics: Theory and Methods, 23(7):1963–1992, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Misra:1995:SIR

[MD95] Neeraj Misra and Ishwari D. Dhariyal. Some inadmissibility results for estimating ordered uniform scale parameters. Communications in Statistics: Theory and Methods, 24(3):675–685, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mensah:1993:CEV

[MEC93] Robert D. Mensah, R. K. Elswick, Jr., and Vernon M. Chinchilli. Consistent estimators of the variance-covariance matrix of the GMANOVA model with missing data. Communications in Statistics: Theory and Methods, 22(6):1495–1514, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Meeter:1990:RSB

[Mee90] Duane Meeter. Row similarity in binary-valued matrices. Communications in Statistics: Theory and Methods, 19(11):4213– 4223, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mehring:1993:OTG

[Meh93] Günther H. Mehring. On optimal tests for general intervalhypotheses. Communications in Statistics: Theory and Methods, 22(5):1257–1297, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Menzefricke:1991:BAP

[Men91] Ulrich Menzefricke. A Bayesian analysis of the product of medians or means from a lognormal distribution. Communications in Statistics: Theory and Methods, 20(1):37–44, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Menzefricke:1992:DAL

[Men92a] Ulrich Menzefricke. Designing accelerated life tests when there is type II censoring. Communications in Statistics: Theory and Methods, 21(9):2569–2590, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Menzefricke:1992:VTW

[Men92b] Ulrich Menzefricke. On the variance of total warranty claims. Communications in Statistics: Theory and Methods, 21(3):779– 790, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Menzefricke:1995:PGS

[Men95] Ulrich Menzefricke. On the performance of the Gibbs sampler for the multivariate normal distribution. Communications in Statistics: Theory and Methods, 24(1):191–213, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Menzefricke:1997:ECA

[Men97] Ulrich Menzefricke. Estimation of the correlation among destructively measured variables. Communications in Statistics: Theory and Methods, 26(1):181–195, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Murtaugh:1990:BBM

[MF90] Paul A. Murtaugh and Lloyd D. Fisher. Bivariate binary models of efficacy and toxicity in dose-ranging trials. Communications in Statistics: Theory and Methods, 19(6):2003–2020, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Miller:1997:AIC

[MF97] G. Edward Miller and Carol J. Feltz. Asymptotic inference for coefficients of variation. Communications in Statistics: Theory and Methods, 26(3):715–726, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Marcus:1991:SOS

[MG91] Ruth Marcus and Abraham Genizi. Simultaneous one-sided tolerance intervals for sets of contrasts from normal populations. Communications in Statistics: Theory and Methods, 20(5–6): 1861–1870, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Marchand:1993:JSE

[MG93] E. Marchand and N. C. Giri. James-stein estimation with constraints on the norm. Communications in Statistics: Theory and Methods, 22(10):2903–2924, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mandowara:1994:OPS

[MG94] V. L. Mandowara and P. C. Gupta. Optimum points of stratification for two or more stage designs. Communications in Statistics: Theory and Methods, 23(3):947–958, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Meeden:1998:AML

[MGLF98] Glen Meeden, Charles Geyer, Joseph Long, and Eiichiro Funa. The admissibility of the maximum likelihood estimator for decomposable log-linear interaction models for contingency tables. Communications in Statistics: Theory and Methods, 27(2):473– 493, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Molina:1998:BIB

[MGM98] M. Molina, M. González, and M. Mota. Bayesian inference for bisexual Galton–Watson processes. Communications in Statistics: Theory and Methods, 27(5):1055–1070, 1998. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Madigan:1995:EPI

[MGR95] David Madigan, Jonathan Gavrin, and Adrian E. Raftery. Eliciting prior information to enhance the predictive performance of Bayesian graphical models. Communications in Statistics: Theory and Methods, 24(9):2271–2292, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mark:1995:EMM

[MH95] Carpenter Mark and Jaimie L. Hebert. Estimating the minimum and maximum location parameters for two ig-exponential scale mixtures. Communications in Statistics: Theory and Methods, 24(5):1227–1233, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mudholkar:1996:EWF

[MH96] Govind S. Mudholkar and Alan D. Hutson. The exponentiated Weibull family: some properties and a flood data application. Communications in Statistics: Theory and Methods, 25(12):3059–3083, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ma:1998:ADS

[MH98] Jun Ma and H. Malcolm Hudson. An augmented data scoring algorithm for maximum likelihood. Communications in Statistics: Theory and Methods, 27(11):2761–2776, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mi:1990:NMC

[Mi90] Jie Mi. Notes on the MLE of correlation coefficient in meta analysis. Communications in Statistics: Theory and Methods, 19(6):2035–2052, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mielniczuk:1991:GEC

[Mie91] Jan Mielniczuk. Grade estimation of chi-square divergence. Communications in Statistics: Theory and Methods, 20(12): 4021–4041, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Migliorati:1998:OSM

[Mig98] Sonia Migliorati. Order statistics as minimal sufficient statistics. Communications in Statistics: Theory and Methods, 27(12): 2943–2951, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Miller:1991:ATS

[Mil91] G. Edward Miller. Asymptotic test statistics for coefficients of variation. Communications in Statistics: Theory and Methods, 20(10):3351–3363, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Miller:1996:FSE

[Mil96] Robert B. Miller. A future for statistics education — we will do more with less. Communications in Statistics: Theory and Methods, 25(11):2839–2851, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Minami:1998:MLD

[Min98] Mihoko Minami. Multivariate Lagrange distributions and a subfamiliy. Communications in Statistics: Theory and Methods, 27(9):2185–2197, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Minami:1999:IRL

[Min99] Mihoko Minami. Inverse relationship and limiting forms of Lagrange distributions. Communications in Statistics: Theory and Methods, 28(2):409–429, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mishra:1996:GBB

[Mis96] A. Mishra. A generalised bivariate binomial distribution applicable in four-fold sampling. Communications in Statistics: Theory and Methods, 25(8):1943–1956, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Modarres:1992:TEC

[MJ92a] Reza Modarres and Robert W. Jernigan. Testing the equality of correlation matrices. Communications in Statistics: Theory and Methods, 21(8):2107–2125, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Morrell:1992:EIR

[MJ92b] Christopher H. Morrell and Richard A. Johnson. Estimation from k independent randomly truncated samples: application to neutrino energies. Communications in Statistics: Theory and Methods, 21(12):3335–3350, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Meintanis:1990:AEC

[MK90a] Simos Meintanis and Ioannis A. Koutrouvelis. Adaptive estimation of the center of symmetry based on the empirical characteristic function. Communications in Statistics: Theory and Methods, 19(1):381–394, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mukerjee:1990:RGD

[MK90b] Rahul Mukerjee and Sanpei Kageyama. Robustness group divisible designs. Communications in Statistics: Theory and Methods, 19(9):3189–3203, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mudholkar:1994:GWF

[MK94] Govind S. Mudholkar and Georgia D. Kollia. Generalized Weibull family: a structural analysis. Communications in Statistics: Theory and Methods, 23(4):1149–1171, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Miller:1991:EAA

[ML91] Michael E. Miller and J. Richard Landis. Evaluation of an analysis approach used to account for extra-variation in clustered categorical responses. Communications in Statistics: Theory and Methods, 20(8):2645–2661, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Malani:1993:ACE

[ML93] Hina Mehta Malani and Ying Lu. Animal carcinogenicity experiments with and without serial sacrifice. Communications in Statistics: Theory and Methods, 22(6):1557–1584, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mountassir:1996:RTB

[ML96] M. Mountassir and Y. Lepage. A rank test for bivariate location and scale problem for elliptically symmetric populations. Communications in Statistics: Theory and Methods, 25(5):1073– 1088, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Matzner-Lofber:1998:NFC

[MLGD98] Eric Matzner-Løfber, Ali Gannoun, and Jan G. De Gooijer. Nonparametric forecasting: a comparison of three kernel-based methods. Communications in Statistics: Theory and Methods, 27(7):1593–1617, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

McCormick:1992:DPJ

[MLH92] W. P. McCormick, N. I. Lyons, and K. Hutcheson. Distributional properties of Jaccard’s index of similarity. Communications in Statistics: Theory and Methods, 21(1):51–68, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Marzec:1993:TDC

[MM93] Leszek Marzec and Pawel Marzec. Tests for dispersive comparison of two distributions. Communications in Statistics: Theory and Methods, 22(3):843–851, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mays:1996:BAD

[MM96a] Darcy P. Mays and Raymond H. Myers. Bayesian approach for the design and analysis of a two level factorial experiment in the presence of dispersion effects. Communications in Statistics: Theory and Methods, 25(7):1409–1428, 1996. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Moser:1996:MLR

[MM96b] Barry Kurt Moser and Melinda H. McCann. Maximum likelihood and restricted maximum likelihood estimators as functions of ordinary least squares and analysis of variance estimators. Communications in Statistics: Theory and Methods, 25(3):631– 646, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Michiels:1997:PEL

[MM97] Bart Michiels and Geert Molenberghs. Protective estimation of longitudinal categorical data with nonrandom dropout. Communications in Statistics: Theory and Methods, 26(1):65–94, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Martinez-Morales:1993:TRB

[MMD93] Manuel Martinez-Morales and B. S. Duran. A test for randomness based on a complexity measure. Communications in Statistics: Theory and Methods, 22(3):879–895, 1993. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Michiels:1999:PMO

[MML99] Bart Michiels, Geert Molenberghs, and Stuart R. Lipsitz. A pattern-mixture odds ratio model for incomplete categorical data. Communications in Statistics: Theory and Methods, 28 (12):2843–2869, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Menendez:1996:ADB

[MMPV96] M. L. Menéndez, D. Morales, L. Pardo, and I. Vajda. About divergence-based goodness-of-fit tests in the Dirichletmultinomial model. Communications in Statistics: Theory and Methods, 25(5):1119–1133, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Menendez:1998:TAG

[MMPV98] M. Menendez, D. Morales, L. Pardo, and I. Vajda. Two approaches to grouping of data and related disparity statistics. Communications in Statistics: Theory and Methods, 27(3):609– 633, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mudholkar:1991:CAP

[MMS91] Anil Mudholkar, Govind S. Mudholkar, and Deo Kumar Srivastava. A construction and appraisal of pooled trimmed-t statistics. Communications in Statistics: Theory and Methods, 20(4): 1345–1359, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mentz:1997:RVE

[MMT97] Raul P. Mentz, Pedro A. Morettin, and Clélia M. C. Toloi. Residual variance estimation in moving average models. Communications in Statistics: Theory and Methods, 26(8):1905–1923, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

McDonald:1993:BEM

[MN93] James B. McDonald and Ray D. Nelson. Beta estimation in the market model: skewness and leptokurtosis. Communications in Statistics: Theory and Methods, 22(10):2843–2862, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Morais:1998:IUO

[MN98] Manuel Morais and Isabel Natário. Improving an upper onesided c-chart. Communications in Statistics: Theory and Methods, 27(2):353–364, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

McKean:1996:EHB

[MNS96] Joseph W. McKean, Joshua D. Naranjo, and Simon J. Sheather. An efficient and high breakdown procedure for model criticism. Communications in Statistics: Theory and Methods, 25(11): 2575–2595, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Marsaglia:1990:DBR

[MNZ90] George Marsaglia, B. Narasimhan, and Arif Zaman. The distance between random points in rectangles. Communications in Statistics: Theory and Methods, 19(11):4199–4212, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Massiuma:1991:FST

[MO91] D. K. Massiuma and N. Ordoukhani. Forecasting of some transformed series. Communications in Statistics: Theory and Methods, 20(10):3207–3220, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Moffatt:1995:GPR

[Mof95] Peter G. Moffatt. Grouped Poisson regression models: theory and an application to public house visit frequency. Communications in Statistics: Theory and Methods, 24(11):2779–2796, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mohammed:1990:PEV

[Moh90] Ohidul Haque Mohammed. Preliminary evaluation of the Victorian zero blood alcohol content legislation at alcohol times: an Australian case study in intervention analysis. Communications in Statistics: Theory and Methods, 19(10):3881–3899, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Monlezun:1999:FPM

[Mon99] Charles J. Monlezun. Fitting polynomials to main effects in multi factor experiments: unequal spacing and unbalanced data with no missing cells. Communications in Statistics: Theory and Methods, 28(12):2899–2916, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Moore:1996:USP

[Moo96] Thomas L. Moore. Using student projects in an introductory course for liberal arts students. Communications in Statistics: Theory and Methods, 25(11):2647–2661, 1996. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Morton:1990:BCI

[Mor90] Sally C. Morton. Bootstrap confidence intervals in a complex situation: a sequential paired clinical trial. Communications in Statistics: Theory and Methods, 19(1):181–195, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Morgenthaler:1991:CBB

[Mor91] S. Morgenthaler. Confidence bands for the bias of linear estimators of a function of one variable. Communications in Statistics: Theory and Methods, 20(3):899–909, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

McEwen:1991:MES

[MP91a] Robert P. McEwen and Bernard R. Parresol. Moment expressions and summary statistics for the complete and truncated Weibull distribution. Communications in Statistics: Theory and Methods, 20(4):1361–1372, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Moothathu:1991:UCE

[MP91b] T. S. K. Moothathu and M. S. Prasad. On uniformly consistent estimation of parametric Lorenz curves. Communications in Statistics: Theory and Methods, 20(5–6):1893–1902, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Moreno:1993:CPQ

[MP93] Elias Moreno and Luis Raul Pericchi. On ϵ-contaminated priors with quantile and piece-wise unimodality constraints. Communications in Statistics: Theory and Methods, 22(7):1963–1978, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Muller:1997:BLM

[MP97] Keith E. Muller and Virginia B. Pasour. Bias in linear model power and sample size due to estimating variance. Communications in Statistics: Theory and Methods, 26(4):839–851, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

McAleer:1990:AAT

[MPB90] Michael McAleer, M. Hashem Pesaran, and Anil K. Bera. Alternative approaches to testing non-nested models with autocorrelated disturbances. Communications in Statistics: Theory and Methods, 19(10):3619–3644, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mudholkar:1997:ESC

[MPH97] Govind S. Mudholkar, Marshall Preimer, and Alan D. Hutson. On the efficiencies of some common quick estimators. Communications in Statistics: Theory and Methods, 26(7):1623–1647, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Morales:1991:BSE

[MPQ91] D. Morales, L. Pardo, and V. Quesada. Bayesian survival estimation for incomplete data when the life distribution is proportionally related to the censoring time. Communications in Statistics: Theory and Methods, 20(3):831–850, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mittnik:1998:TEI

[MPR98] Stefan Mittnik, Marc S. Paolella, and Svetlozar T. Rachev. A tail estimator for the index of the stable Paretian distribution. Communications in Statistics: Theory and Methods, 27(5):1239– 1262, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mack:1999:KET

[MQZ99] V. P. Mack, Pham X. Qyang, and Shunpy Zhang. Kernel estimation in transect sampling without the shoulder condition. Communications in Statistics: Theory and Methods, 28(10):2277– 2296, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mehra:1990:CNT

[MR90] K. L. Mehra and K. S. Madhava Rao. On a class of nonparametric tests for the treatment vs control problem. Communications in Statistics: Theory and Methods, 19(6):2323–2336, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

McGilchrist:1995:GCV

[MR95a] C. A. McGilchrist and T. E. Ragunathan. Generalised components of variance models. Communications in Statistics: Theory and Methods, 24(2):349–361, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mishra:1995:CKT

[MR95b] M. N. Mishra and B. L. S. Prakasa Rao. On L_{1}-consistency of kernel-type density estimator for stationary Markov processes. Communications in Statistics: Theory and Methods, 24(3):581– 591, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Menendez:1992:TNO

[MRS92] J. A. Menéndez, C. Rueda, and B. Salvador. Testing non-oblique hypotheses. Communications in Statistics: Theory and Methods, 21(2):471–484, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Meaux:1990:PIQ

[MS90a] Laurie M. Meaux and John W. Seanam, Jr. Probability inequalities for quadratic forms. Communications in Statistics: Theory and Methods, 19(5):1863–1874, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mohebbi:1990:EMT

[MS90b] Cyrus Mohebbi and Lewis H. Shoemaker. An extension of the median test to analysis of variance. Communications in Statistics: Theory and Methods, 19(3):1101–1117, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mortera:1990:CPA

[MS90c] Julia Mortera and Aristide San Martini. Comparison of proportions with asymmetric prior information. Communications in Statistics: Theory and Methods, 19(9):3205–3222, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Maravalle:1995:STH

[MS95] Maurizio Maravalle and Bruno Simeone. A spanning tree heuristic for regional clustering. Communications in Statistics: Theory and Methods, 24(3):625–639, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Murat:1998:NZI

[MS98] M. Murat and D. Szynal. Non-zero inflated modified power series distributions. Communications in Statistics: Theory and Methods, 27(12):3047–3064, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mudholkar:1999:PHM

[MS99] Govind S. Mudholkar and Ila C. Sarkar. A proportional hazards modeling of multisample reliability data. Communications in Statistics: Theory and Methods, 28(9):2079–2101, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mudholkar:1995:SVA

[MSL95] Govind S. Mudholkar, Deo Kumar Srivastava, and C. Thomas Lin. Some p-variate adaptations of the Shapiro–Wilk test of normality. Communications in Statistics: Theory and Methods, 24 (4):953–985, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Menendez:1997:DMB

[MSMP97] M. Menéndez, M. Salicrú, D. Morales, and L. Pardo. Divergence measures between populations: applications in the exponential family. Communications in Statistics: Theory and Methods, 26(5):1099–1117, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mangat:1997:VRP

[MSS97] Naurang Singh Mangat, Ravindra Singh, and Sarjinder Singh. Violation of respondent’s privacy in Moors’ model — its rectification through a random group strategy. Communications in Statistics: Theory and Methods, 26(3):743–754, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Madi:1990:ECS

[MT90] Mohamed Madi and Kam-Wah Tsui. Estimation of the common scale of several shifted exponential distributions with unknown locations. Communications in Statistics: Theory and Methods, 19(6):2295–2313, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Marcus:1992:FRT

[MT92] Ruth Marcus and Hovav Talpaz. Further results on testing homogeneity of normal means against simple tree alternatives. Communications in Statistics: Theory and Methods, 21(8):2135– 2149, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Militino:1997:GEL

[MU97] Ana F. Militino and M. Dolores Ugarfe. A GM estimation of the location parameters in a spatial linear model. Communications in Statistics: Theory and Methods, 26(7):1701–1725, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mukhopadhyay:1994:BMP

[Muk94] P. Mukhopadhyay. Bayes and minimax procedures for finite population sampling under measurement error models. Communications in Statistics: Theory and Methods, 23(7):1953–1961, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mukhopadhyay:1995:BME

[Muk95] P. Mukhopadhyay. Bayes and minimax estimators for two-stage sampling from a finite population under measurement error models. Communications in Statistics: Theory and Methods, 24(3): 663–674, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mukhopadhyay:1997:NUE

[Muk97] Parimal Mukhopadhyay. Note on UMVSU-estimation under randomized response model. Communications in Statistics: Theory and Methods, 26(10):2415–2420, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Martinez:1990:OPC

[MVM90] José Raúl Martinez, Av Valparaísoy, and R. Martínez. An optimal property of the “(r, k) class estimators”. Communications in Statistics: Theory and Methods, 19(4):1281–1289, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

MacLean:1996:EMR

[MW96] Leonard C. MacLean and K. Laurence Weldon. Estimating multivariate random effects without replication. Communications in Statistics: Theory and Methods, 25(7):1447–1469, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Meyer:1998:BVA

[MW98] R. Daniel Meyer and Robert G. Wilkinson. Bayesian variable assessment. Communications in Statistics: Theory and Methods, 27(11):2675–2705, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Mulekar:1991:AFS

[MY91] Madhuri S. Mulekar and Linda J. Young. Approximations for a fixed sample size selection procedure for negative binomial populations. Communications in Statistics: Theory and Methods, 20(5–6):1767–1776, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

McGilchrist:1995:DBM

[MY95] C. A. McGilchrist and K. K. W. Yau. The derivation of BLUP, ML, REML estimation methods for generalised linear mixed models. Communications in Statistics: Theory and Methods, 24(12):2963–2980, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Myers:1991:RSM

[Mye91] Raymond H. Myers. Response surface methodology in quality improvement. Communications in Statistics: Theory and Methods, 20(2):457–476, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Myongsik:1995:MLE

[Myo95] Oh Myongsik. On maximum likelihood estimation of cell probabilities in 2 × k contingency tables under negative dependence restrictions with various sampling schemes. Communications in Statistics: Theory and Methods, 24(8):2127–2143, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ncube:1990:TPC

[NA90] Matoteng M. Ncube and Raid W. Amin. Two parameter cuscore quality control procedures. Communications in Statistics: Theory and Methods, 19(6):2191–2205, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Nandini:1994:CSE

[NA94] Kannan Nandini and Raychaudhuri Aparna. Composite sampling: effectiveness under a two stage scheme. Communications in Statistics: Theory and Methods, 23(9):2533–2540, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Nabeya:1993:CTF

[Nab93] Seiji Nabeya. Coincidence of two failure rate models. Communications in Statistics: Theory and Methods, 22(3):781–785, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Nagaraja:1990:SRP

[Nag90] H. N. Nagaraja. Some reliability properties of order statistics. Communications in Statistics: Theory and Methods, 19(1):307– 316, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Nagata:1995:RBI

[Nag95] Yasushi Nagata. The relationship between the improvement on the point estimation and the improvement on the interval estimation for the disturbance variance in a linear regression model. Communications in Statistics: Theory and Methods, 24(7):1687– 1704, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Nagata:1996:NAW

[Nag96] Yasushi Nagata. The Neyman accuracy and the Wolfowitz accuracy of the Stein type confidence interval for the disturbance variance. Communications in Statistics: Theory and Methods, 25 (5):985–1004, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Nagata:1997:STC

[Nag97] Yasushi Nagata. Stein type confidence interval of the disturbance variance in a linear regression model with multivariate Studentt distributed errors. Communications in Statistics: Theory and Methods, 26(2):503–523, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Nagao:1999:RUS

[Nag99] Hisao Nagao. The risks for usual sequential estimates and stopping times of multivariate normal mean for conjugate distribution. Communications in Statistics: Theory and Methods, 28(3– 4):873–883, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Naik:1990:DOS

[Nai90] Dayanand N. Naik. On detection of outliers in symmetric normal models. Communications in Statistics: Theory and Methods, 19(6):2315–2321, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Naito:1996:MSB

[Nai96] Kanta Naito. Modification of statistics based on the empirical characteristic function to yield asymptotic normality. Communications in Statistics: Theory and Methods, 25(1):105–114, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Nanayakkaraa:1992:RJM

[Nan92] Nuwan Nanayakkaraa. On the robustness and Johnson’s modification of the one sample T -statistic. Communications in Statistics: Theory and Methods, 21(11):3079–3096, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Nassiuma:1991:NNA

[Nas91] Dankit Nassiuma. A note on nonstationary ARMA processes with infinite variance. Communications in Statistics: Theory and Methods, 20(5–6):1821–1826, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Nigm:1996:BPR

[NAW96] A. M. Nigm and N. Y. Abd Al-Wahab. Bayesian prediction with a random sample size for the Burr lifetime distribution.

Communications in Statistics: Theory and Methods, 25(6):1289– 1303, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Nayak:1991:C

[Nay91] Tapan K. Nayak. Comments. Communications in Statistics: Theory and Methods, 20(11):3679–3684, 1991. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Nayak:1994:RRS

[Nay94] Tapan K. Nayak. On randomized response surveys for estimating a proportion. Communications in Statistics: Theory and Methods, 23(11):3303–3321, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Nayak:1996:ECP

[Nay96] Tapan K. Nayak. On estimating the conditional probability of discovering a new species. Communications in Statistics: Theory and Methods, 25(9):2039–2056, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Neath:1997:RTS

[NC97] Andrew A. Neath and Joseph E. Cavanaugh. Regression and time series model selection using variants of the Schwarz information criterion. Communications in Statistics: Theory and Methods, 26(3):559–580, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Newby:1999:MGF

[New99] Martin Newby. Moments and generating functions for the absorption distribution and its negative binomial analogue. Communications in Statistics: Theory and Methods, 28(12):2935– 2945, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ng:1990:BAL

[Ng90] Vee Ming Ng. Bayesian analysis of linear models exhibiting changes in mean and precision at an unknown time point. Communications in Statistics: Theory and Methods, 19(1):111–120, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ng:1993:RBP

[Ng93] Meei Pyng Ng. A relation between product-limit estimate and joint-risk estimate. Communications in Statistics: Theory and Methods, 22(5):1355–1369, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ng:1994:IPT

[Ng94] Tie-Hua Ng. The impact of a preliminary test for interaction in a 2 × 2 factorial trial. Communications in Statistics: Theory and Methods, 23(2):435–450, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Naoto:1995:LPU

[NHY95] Hoshino Naoto, Miyazaki Haruo, and Seki Yoichi. On the level probabilities for useful partially ordered alternatives in the analysis of variance. Communications in Statistics: Theory and Methods, 24(8):2059–2071, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Nieto:1997:NIA

[Nie97] Fabio H. Nieto. A note on interpolation of ARIMA processes. Communications in Statistics: Theory and Methods, 26(10): 2381–2392, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Nishii:1994:CGC

[Nis94] Ryuei Nishii. Convergence of the Gram–Charier expansion of the signed log likelihood ratio. Communications in Statistics: Theory and Methods, 23(9):2605–2614, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Na:1999:IMR

[NK99] Myung Hwan Na and Jae Joo Kim. On inference for mean residual life. Communications in Statistics: Theory and Methods, 28(12):2917–2933, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Nishu:1999:SVN

[NKN99] Ryuei Nishu, Saeko Kusanobu, and Noriyuki Nakaoka. Selection of variables and neighborhoods for spatial enhancement of thermal infrared images. Communications in Statistics: Theory and Methods, 28(3–4):965–976, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Norris:1992:RBL

[NM92] James L. Norris III and Duane A. Meeter. A robust Bayesian lower limit for the expected number of unobserved classes. Communications in Statistics: Theory and Methods, 22(1):81–95, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Norris:1994:NEN

[NM94] James L. Norris III and Duane A. Meeter. A nonparametric estimator of the number of classes based on a stratified random sample. Communications in Statistics: Theory and Methods, 23(9):2561–2580, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Nieto:1996:RAE

[NM96] Fabio H. Nieto and Jorge Martínez. A recursive approach for estimating missing observations in an univariate time series. Communications in Statistics: Theory and Methods, 25(9):2101– 2116, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Nirel:1998:ORS

[NMB98] Ronit Nirel, Moira A. Mugglestone, and Vic Barnett. Outlierrobust spectral estimation for spatial lattice processes. Communications in Statistics: Theory and Methods, 27(12):3095–3111, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Nagarsenker:1991:MTE

[NN91] P. B. Nagarsenker and B. N. Nagarsenker. On a modified test of equality of scale parameters of exponential distributions. Communications in Statistics: Theory and Methods, 20(2):735–745, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Norton:1993:STD

[Nor93] Robert M. Norton. Some tests for differences between proportions where the expected number of exposures to an inferior treatment is kept small. Communications in Statistics: Theory and Methods, 22(9):156–162, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Norris:1995:SDO

[Nor95] James L. Norris III. Sample design for observing classes of objects. Communications in Statistics: Theory and Methods, 24(12):3147–3162, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Nurhonen:1992:PPG

[NP92] Markku Nurhonen and Simo Puntanen. A property of partitioned generalized regression. Communications in Statistics: Theory and Methods, 21(6):1579–1583, 1992. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Nakamura:1993:EDA

[NPA93] Miguel Nakamura and Victor Pérez-Abreu. Exploratory data analysis for counts using the empirical probability generating function. Communications in Statistics: Theory and Methods, 22 (3):827–842, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Nasri-Roudsari:1999:LDG

[NR99] Dirk Nasri-Roudsari. Limit distributions of generalized order statistics under power normalization. Communications in Statistics: Theory and Methods, 28(6):1379–1389, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Nebebe:1990:RPI

[NS90] Fassil Nebebe and Ah Boon Sim. The relative performances of improved ridge estimators and an empirical Bayes estimator: some Monte Carlo results. Communications in Statistics: Theory and Methods, 19(9):3469–3495, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Nomakuchi:1992:NOP

[NS92] Kentaro Nomakuchi and Ning-Zhong Shi. A note on the orthant probability of the elliptically contoured distribution. Communications in Statistics: Theory and Methods, 21(1):161–167, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Neudecker:1993:BQP

[NS93] Heinz Neudecker and Liu Shuangzhe. Best quadratic and positive semidefinite unbiased estimation of the variance matrix of the multivariate normal distribution. Communications in Statistics: Theory and Methods, 22(10):2723–2732, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ngai:1994:HBS

[NS94] Hung Man Ngai and T. W. F. Stroud. Hierarchical Bayes simultaneous estimation of Poisson means. Communications in Statistics: Theory and Methods, 23(10):2965–2991, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Nelson:1996:RTG

[NS96] Peter R. Nelson and Paul L. Stephenson. Runs tests for group control charts. Communications in Statistics: Theory and Methods, 25(11):2739–2765, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Nishimura:1997:ESF

[NS97] Kazuo Nishimura and Masaaki Sibuya. Extended Stirling family of discrete probability distributions. Communications in Statistics: Theory and Methods, 26(7):1727–1744, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Nandi:1999:SIT

[NS99] S. N. Nandi and M. Sreehari. Some improvements in Taguchi’s economic method allowing continued quality deterioration in production process. Communications in Statistics: Theory and Methods, 28(5):1169–1181, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Niki:1990:CRS

[NSI90] Naota Niki, Nakagawa Shigekazu, and Hideyuki Inoue. Cumulants of random sum distributions. Communications in Statistics: Theory and Methods, 19(5):1857–1861, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Newman:1999:PPR

[NSS99] R. M. Newman, W. B. Smith, and F. M. Speed. Profile parallelism in repeated measures designs. Communications in Statistics: Theory and Methods, 28(11):2749–2781, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Nassiuma:1992:SEN

[NT92] D. K. Nassiuma and A. Thavaneswaran. Smoothed estimates for nonlinear time series models with irregular data. Communications in Statistics: Theory and Methods, 21(8):2247–2259, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Nasser:1991:FVD

[NTS91] Ordoukhani Nasser, A. Thavaneswarn, and M. Samanta. Functional version of the delta method and its application. Communications in Statistics: Theory and Methods, 20(1):373–388, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Nel:1990:EDU

[NvdMM90] D. G. Nel, Cay A. van der Merwe, and B. K. Moser. The exact distributions of the univariate and multivariate Behrens–Fisher statistics with a comparison of several solutions in the univariate case. Communications in Statistics: Theory and Methods, 19 (1):279–298, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ogden:1999:DMD

[OC99] R. Todd Ogden and Geoffrey L. Collier. On detecting and modeling deterministic drift in long run sequences of tapping data. Communications in Statistics: Theory and Methods, 28(3–4): 977–987, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Oh:1998:BTS

[Oh98] Man-Suk Oh. A Bayes test for simple versus one-sided hypothesis on the mean vector of a multivariate normal distribution. Communications in Statistics: Theory and Methods, 27(10):2371–2389, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ohtani:1991:EVN

[Oht91] Kazuhiro Ohtani. Estimation of the variance in a normal population after the one-sided pre-test for the mean. Communications in Statistics: Theory and Methods, 20(1):219–234, 1991.

CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ohtani:1993:DDF

[Oht93a] Kazuhiro Ohtani. Distribution and density functions of the feasible generalized ridge regression estimator. Communications in Statistics: Theory and Methods, 22(10):2733–2746, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ohtani:1993:EDD

[Oht93b] Kazuhiro Ohtani. The exact distribution and density functions of the Stein-type estimator for normal variance. Communications in Statistics: Theory and Methods, 22(10):2863–2876, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ohtani:1996:ESS

[Oht96a] Kazuhiro Ohtani. Exact small sample properties of an operational variant of the minimum mean squared error estimator. Communications in Statistics: Theory and Methods, 25(6):1223– 1231, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ohtani:1996:ADF

[Oht96b] Kazuhiro Ohtani. On an adjustment of degrees of freedom in the minimum mean squared error estimator. Communications in Statistics: Theory and Methods, 25(12):3049–3058, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Oluyede:1993:MCS

[Olu93] Broderick O. Oluyede. A modified chi-square test for testing equality of two multinomial populations against an order restricted alternative. Communications in Statistics: Theory and Methods, 22(4):1133–1155, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Oluyede:1994:FBB

[Olu94] Broderick O. Oluyede. A family of bivariate binomial distributions generated by extreme Bernoulli distributions. Communications in Statistics: Theory and Methods, 23(5):1531–1547, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ong:1990:MFB

[Ong90] S. H. Ong. Mixture formulations of a bivariate negative binomial distribution with applications. Communications in Statistics: Theory and Methods, 19(4):1303–1322, 1990. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ong:1993:CEC

[Ong93] S. H. Ong. Canonical expansions, correlation structure, and conditional distributions of bivariate distributions generated by mixtures. Communications in Statistics: Theory and Methods, 22 (9):163–168, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ong:1995:CPG

[Ong95] S. H. Ong. Computation of probabilities of a generalized logseries and related distributions. Communications in Statistics: Theory and Methods, 24(1):253–271, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

OReilly:1998:NFL

[OR98] F. J. O’Reilly and R. Rueda. A note on the fit for the Lévy distribution. Communications in Statistics: Theory and Methods, 27(7):1811–1821, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Oberg:1993:OCS

[OS93a] Karen M. Oberg and David L. Snyder. One-cluster sampling designs for an autocorrelated population. Communications in Statistics: Theory and Methods, 22(7):1979–1998, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Overton:1993:PDS

[OS93b] W. Scott Overton and Stephen V. Stehman. Properties of designs for sampling continuous spatial resources from a triangular grid. Communications in Statistics: Theory and Methods, 22 (9):251–264, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ouyang:1991:REU

[OSL91] Zhao Ouyang, Hans T. Schreuder, and Junfang Li. Regression estimation under sampling with one unit per stratum. Communications in Statistics: Theory and Methods, 20(8):2431–2449, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ouyang:1992:RSS

[OSL92] Z. Ouyang, H. T. Schreuder, and H. G. Li. Robust sampling strategies for regression estimation. Communications in Statistics: Theory and Methods, 21(11):3309–3326, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ordoukhani:1991:CAP

[OT91] N. Ordoukhani and A. Thavaneswarn. Convergence analysis of parametric identification methods for jump processes. Communications in Statistics: Theory and Methods, 20(5–6):1827–1837, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

OGorman:1990:SSM

[OW90] Thomas W. O’Gorman and Robert F. Woolson. The stability of several measures of association in small contingency tables. Communications in Statistics: Theory and Methods, 19(3):1141– 1155, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

OGorman:1993:ECC

[OW93] Thomas W. O’Gorman and Robert F. Woolson. The effect of category choice on the odds ratio and several measures of association in case-control studies. Communications in Statistics: Theory and Methods, 22(4):1157–1171, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ozturk:1991:GAU

[Ozt91] Aydin Ozturk. A general algorithm for univariate and multivariate goodness of fit tests based on graphical representation. Communications in Statistics: Theory and Methods, 20(10):3111– 3137, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ozturk:1999:OTS

[Özt99] Ömer Öztürk. One-and two-sample sign tests for ranked set sample selective designs. Communications in Statistics: Theory and Methods, 28(5):1231–1245, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Paik:1995:EAQ

[Pai95] Myunghee Cho Paik. The em algorithm for the quasi-likelihood regression model. Communications in Statistics: Theory and Methods, 24(6):1403–1430, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Pan:1996:DFT

[Pan96] Guohua Pan. Distribution-free tests for umbrella alternatives. Communications in Statistics: Theory and Methods, 25(12): 3185–3194, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Pandey:1997:TSP

[Pan97] B. N. Pandey. Testimator of the scale parameter of the exponential distribution using LINEX loss function. Communications in Statistics: Theory and Methods, 26(9):2191–2202, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Pape:1990:SEA

[Pap90] Elinor S. Pape. Sampling efficiency for an alternating Poisson process. Communications in Statistics: Theory and Methods, 19(9):3175–3178, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Papp:1992:TST

[Pap92] Zoltan Papp. Two- sided tolerance limits and confidence bounds for normal populations. Communications in Statistics: Theory and Methods, 21(5):1309–1318, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Park:1990:DED

[Par90] Joong-Yang Park. Designs for estimating the difference between two responses. Communications in Statistics: Theory and Methods, 19(12):4773–4787, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Park:1993:NRO

[Par93a] Hyo-Ii Park. Nonparametric rank-order tests for the right censored and grouped data in linear model. Communications in Statistics: Theory and Methods, 22(11):3143–3158, 1993. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Park:1993:EML

[Par93b] Teasung Park. Equivalence of maximum likelihood estimation and iterative two-stage estimation for seemingly unrelated regression models. Communications in Statistics: Theory and Methods, 22(8):2285–2296, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Parzen:1993:CPC

[Par93c] Emanuel Parzen. Change P P plot and continuous sample quantile function. Communications in Statistics: Theory and Methods, 22(12):3287–3304, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Park:1994:MRM

[Par94] Taesung Park. Multivariate regression models for discrete and continuous repeated measurements. Communications in Statistics: Theory and Methods, 23(6):1547–1564, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Park:1998:DEC

[Par98] Changsoon Park. Design of and EWMA charts in a variance components model. Communications in Statistics: Theory and Methods, 27(3):659–672, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Passi:1994:LME

[Pas94] Ranjit M. Passi. Linear model estimation errors with estimated design parameters. Communications in Statistics: Theory and Methods, 23(11):3189–3209, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Patel:1994:DRP

[Pat94a] H. I. Patel. Dose-response in pharmacokinetics. Communications in Statistics: Theory and Methods, 23(2):451–465, 1994.

CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Patel:1994:GSA

[Pat94b] H. I. Patel. Group sequential analysis of a clinical trial with repeated measurements. Communications in Statistics: Theory and Methods, 23(4):981–995, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Palmer:1990:CBM

[PB90a] Judy L. Palmer, Ph.D. and Lyle D. Broemeling, Ph.D. A comparison of Bayes and maximum likelihood estimation of the intraclass correlation coefficient. Communications in Statistics: Theory and Methods, 19(3):953–975, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Patel:1990:UET

[PB90b] S. R. Patel and M. B. Bhatt. Unbiased estimation in type II censored samples from two-truncation parameter family. Communications in Statistics: Theory and Methods, 19(10):3809–3831, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Patel:1991:UED

[PB91] S. R. Patel and M. B. Bhatt. Unbiased estimation in doubly Type II censored samples from two two-truncation parameter families. Communications in Statistics: Theory and Methods, 20(7):2119–2148, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Patel:1992:UEP

[PB92] S. R. Patel and M. B. Bhatt. Unbiastld estimation of parametric functions in sampling from two two-truncation parameter families. Communications in Statistics: Theory and Methods, 21 (3):843–860, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Park:1993:CIA

[PB93] Dong Joon Park and Richard K. Burdick. Confidence intervals on the among group variance component in a simple linear regression model with a balanced one-fold nested error structure. Communications in Statistics: Theory and Methods, 22 (12):3435–3452, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Paark:1998:CIM

[PB98] Kern Paark and Richard K. Burdick. Confidence intervals for the mean in a balanced two factor random effects model. Communications in Statistics: Theory and Methods, 27(11):2807–2825, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Paula:1995:BCI

[PC95] Gilberto A. Paula and Gauss M. Cordeiro. Bias correction and improved residuals for non-exponential family nonlinear models. Communications in Statistics: Theory and Methods, 24(5):1193– 1210, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Pan:1998:ESC

[PC98] Wei Pan and Rick Chappell. Estimating survival curves with left truncated and interval censored data via the EMS algorithm. Communications in Statistics: Theory and Methods, 27(4):777– 793, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Priebe:1999:GWM

[PC99] Carey E. Priebe and Lenore J. Cowen. A generalized Wilcoxon– Mann–Whitney statistic. Communications in Statistics: Theory and Methods, 28(12):2871–2878, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Park:1992:SIR

[PD92] D. K. Park and A. M. Dean. Structurally incomplete row-column designs with adjusted orthogonality. Communications in Statistics: Theory and Methods, 21(12):3545–3560, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Pankratz:1992:IRD

[PE92] Alan Pankratz and Brooks Elliott. Identifying rational distributed lag models using a joint test applied to the corner table. Communications in Statistics: Theory and Methods, 21(8): 2279–2292, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Peiris:1990:AMA

[Pei90] M. Shelton Peiris. Analysis of multivariate ARMA processes with non-stationary innovations. Communications in Statistics: Theory and Methods, 19(8):2847–2852, 1990. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Peixoto:1992:FED

[Pei92] Julio L. Peixoto. Four equivalent definitions of reparametrizations and restrictions in linear models. Communications in Statistics: Theory and Methods, 22(1):283–299, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Pepple:1990:HBE

[Pep90] Patricia A. Pepple. Hierarchical Bayes estimation of normal variances with application to a random effect model. Communications in Statistics: Theory and Methods, 19(6):2085–2108, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Peristiani:1991:SDA

[Per91] Steve Peristiani. The F -system distribution as an alternative to multivariate normality: An application in multivariate models with qualitative dependent variables. Communications in Statistics: Theory and Methods, 20(1):147–163, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Pettit:1994:BAD

[Pet94] L. I. Pettit. Bayesian approaches to the detection of outliers in Poisson samples. Communications in Statistics: Theory and Methods, 23(6):1785–1795, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Parsian:1993:AIE

[PF93] A. Parsian and N. Sanjari Farsipour. On the admissibility and inadmissibility of estimators of scale parameters using an asymmetric loss function. Communications in Statistics: Theory and Methods, 22(10):2877–2901, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Parsian:1992:MPT

[PFN92] A. Parsian, N. Sanjari Farsipour, and N. Nematollahi. On the minimaxity of Pitman type estimator under a LINEX loss function. Communications in Statistics: Theory and Methods, 22 (1):97–113, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Pandey:1991:ENS

[PFU91a] M. Pandey, J. Ferdous, and Md. Borhan Uddin. Effect of a new selection procedure on adaptive estimation. Communications in Statistics: Theory and Methods, 20(10):3365–3379, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Pandey:1991:SPD

[PFU91b] M. Pandey, Jannatual Ferdous, and Md. Borhan Uddin. Selection of probability distribution for life testing data. Communications in Statistics: Theory and Methods, 20(4):1373–1388, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Patel:1992:MLE

[PG92] M. N. Patel and A. V. Gajjar. Maximum likelihood estimation in compound exponential failure model with changing failure rates from Type I progressively censored and group censored samples. Communications in Statistics: Theory and Methods, 21(10):2899–2908, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Pham-Gia:1994:VBP

[PG94] T. Pham-Gia. Value of the beta prior information. Communications in Statistics: Theory and Methods, 23(8):2175–2195, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Patel:1995:SRM

[PG95] M. N. Patel and A. V. Gajjar. Some results on maximum likelihood estimators of parameters of exponential distribution under Type I progressive censoring with changing failure rates. Communications in Statistics: Theory and Methods, 24(9):2421– 2435, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Pleguezuelo:1997:NCC

[PGC97] Rafael Herrerías Pleguezuelo, Federico Palacios González, and José Caliejón Céspedes. A new condition to check the independence of random variables. Communications in Statistics: Theory and Methods, 26(2):421–428, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Pleguezuelo:1998:UGF

[PGC98] Rafael Herrerías Pleguezuelo, Federico Palacios González, and José Callejón Céspedes. Using generator function system to check independence of random variables. Communications in Statistics: Theory and Methods, 27(5):1211–1219, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Pham-Gia:1992:ILN

[PGT92] T. Pham-Gia and N. Turkkan. Information loss in a noisy dichotomous sampling process. Communications in Statistics: Theory and Methods, 21(7):2001–2018, 1992. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Pham-Gia:1998:DLC

[PGT98] T. Pham-Gia and N. Turkkan. Distribution of the linear combination of two general beta variables and applications. Communications in Statistics: Theory and Methods, 27(7):1851–1869, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Pincus:1992:AES

[PH92] Steven M. Pincus and Wei-Min Huang. Approximate entropy: Statistical properties and applications. Communications in Statistics: Theory and Methods, 21(11):3061–3077, 1992. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Phillips:1996:NPA

[Phi96] Robert F. Phillips. A note on partially adaptive estimation via a normal mixture. Communications in Statistics: Theory and Methods, 25(6):1199–1206, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Park:1999:MEE

[PJ99] Sang-Hyun Park and Dae-Heung Jang. Measures for evaluating the effect of blocking in response surface designs. Communications in Statistics: Theory and Methods, 28(7):1599–1616, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Parthasarathy:1991:BDP

[PK91a] P. R. Parthasarathy and B. Krishna Kumar. A birth and death process with logistic mean population. Communications in Statistics: Theory and Methods, 20(2):621–629, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Peddada:1991:CEL

[PK91b] Shyamal Das Peddada and Ravindra Khattree. Comparison of estimators of the location parameter using Pitman’s closeness criterion. Communications in Statistics: Theory and Methods, 20(11):3525–3534, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Park:1994:DBN

[PK94] Pil S. Park and Anant M. Kshirsagar. Distances between normal populations when covariance matrices are unequal. Communications in Statistics: Theory and Methods, 23(12):3549–3556, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Park:1998:SRD

[PK98] Sung H. Park and Hyo T. Kwon. Slope-rotatable designs with equal maximum directional variance for second order response surface models. Communications in Statistics: Theory and Methods, 27(11):2837–2851, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Peps:1996:OCD

[PL96] Tim K. Peps and Martin S. Levy. Optimal control designs using predicting densities for the multivariate linear model. Communications in Statistics: Theory and Methods, 25(6):1351–1374, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Perez:1997:SSD

[PL97] Adriana Perez and John J. Lefante. Sample size determination and the effect of censoring when estimating the arithmetic mean of a lognormal distribution. Communications in Statistics: Theory and Methods, 26(11):2779–2801, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Liu:1998:SEI

[pL98] Jen pei Liu. Statistical evaluation of individual bioequivalence. Communications in Statistics: Theory and Methods, 27(6):1433– 1451, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Pu:1999:RMA

[PL99] Zongwei Pu and Linxiong Li. Regression models with arbitrarily interval-censored observations. Communications in Statistics: Theory and Methods, 28(7):1547–1563, 1999. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Pearn:1998:DIP

[PLC98] W. L. Pearn, G. H. Lin, and K. S. Chen. Distributional and inferential properties of the process accuracy and process precision indices. Communications in Statistics: Theory and Methods, 27 (4):985–1000, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Park:1993:TMD

[PLW93] Taesung Park, Seungyeoun Lee, and Robert F. Woolson. A test of the missing data mechanism for repeated measures data. Communications in Statistics: Theory and Methods, 22(10):2813– 2829, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Pandey:1990:TMI

[PM90] B. N. Pandey and H. J. Malik. Testing for the mean of the inverse Gaussian distribution and adaptive estimation of parameters. Communications in Statistics: Theory and Methods, 19(2):629– 637, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Pandey:1991:WEN

[PM91a] B. N. Pandey and G. C. Mishra. Weighted estimators for a normal population variance. Communications in Statistics: Theory and Methods, 20(1):235–247, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Pina:1991:IOM

[PM91b] Daniel Piña and Agustin Maravall. Interpolation, outliers mid inverse autocorrelations. Communications in Statistics: Theory and Methods, 20(10):3175–3186, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Patel:1992:SWS

[PM92] M. S. Patel and M. M. Manene. Step-wise screening of defective factors. Communications in Statistics: Theory and Methods, 21(9):2527–2539, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Pino:1993:CNE

[PM93] Francisco A. Pino and Pedro A. Morettin. The consistency of the L_{1} norm estimates in ARMA models. Communications in Statistics: Theory and Methods, 22(8):2185–2206, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Pandey:1994:SIE

[PM94] B. N. Pandey and H. J. Malik. Some improved estimation for common variance of two populations. Communications in Statistics: Theory and Methods, 23(10):3019–3035, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Pandey:1995:BSE

[PMD95] B. N. Pandey, H. J. Malik, and P. K. Dubey. Bayesian shrinkage estimators for a measure of dispersion of an inverse Gaussian distribution. Communications in Statistics: Theory and Methods, 24(9):2261–2270, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Pardo:1997:LSB

[PMSM97] L. Pardo, D. Morales, M. Salicrú, and M. L. Menéndez. Large sample behavior of entropy measures when parameters are estimated. Communications in Statistics: Theory and Methods, 26 (2):483–501, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Pardo:1993:CEB

[PMTP93] J. A. Pardo, M. L. Menendez, I. J. Taneja, and L. Pardo. Comparison of experiments based on generalized entropy measures. Communications in Statistics: Theory and Methods, 22(4):1113– 1132, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Parsian:1995:AOP

[PN95] A. Parsian and N. Nematollahi. On the admissibility of ordered Poisson parameter under the entropy loss function. Communications in Statistics: Theory and Methods, 24(10):2451–2467, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Perez-Ocon:1995:CAP

[POGP95] Rafael Pérez-Ocón and M. Luz Gámiz-Pérez. Conditions on the arrival process to obtain HNBUE survival using a shock model. Communications in Statistics: Theory and Methods, 24(4):931– 944, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Perez-Ocon:1996:HPS

[POGP96] Rafael Pérez-Ocón and M. Luz Gámiz-Pérez. HNBUE property in a shock model with cumulative damage threshold. Communications in Statistics: Theory and Methods, 25(2):345–360, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Posse:1992:PPD

[Pos92a] C. G. Posse. Projection pursuit discriminant analysis for two groups. Communications in Statistics: Theory and Methods, 21 (1):1–19, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Posten:1992:RTS

[Pos92b] Harry O. Posten. Robustness of the two-sample t-test under violations of the homogeneity of variance assumption, Part II. Communications in Statistics: Theory and Methods, 21(8):2169– 2184, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Psarakis:1990:FD

[PP90] S. Psarakis and J. Panaretoes. The folded t distribution. Communications in Statistics: Theory and Methods, 19(7):2717– 2734, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Patel:1992:MEE

[PP92] S. R. Patel and N. P. Patel. Mixed effect exponential linear model. Communications in Statistics: Theory and Methods, 21(9):2721–2740, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Park:1997:AOC

[PPS97] Chul Gyu Park, Taesung Park, and Dong Wan Shin. Analysis of ordered covariate effects among groups with repeated measurements. Communications in Statistics: Theory and Methods, 26(10):2291–2301, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Pandey:1992:BEM

[PR92] B. N. Pandey and Omkar Rai. Bayesian estimation of mean and square of mean of normal distribution using LINEX loss function. Communications in Statistics: Theory and Methods, 21(12):3369–3391, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Penev:1993:CES

[PR93] Spiridon Penev and Tenko Raykov. On choosing estimators in a simple linear errors-in-variables model. Communications in Statistics: Theory and Methods, 22(9):101–115, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Prasad:1989:SIR

[Pra89] B. Prasad. Some improved ratio type estimators of population mean and ratio in finite population sample surveys. Communications in Statistics: Theory and Methods, 18(1):379–392, 1989. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). See erratum [Pra91].

Prasad:1991:ESI

[Pra91] B. Prasad, Research Officer. Erratum to “Some improved ratio type estimators of population mean and ratio in finite population sample survey”. Communications in Statistics: Theory and Methods, 20(7):2331–2332, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic). See [Pra89].

Prescott:1998:NOO

[Pre98] Philip Prescott. Nearly optimal orthogonally blocked designs for a quadratic mixture model with q components. Communications in Statistics: Theory and Methods, 27(10):2559–2580, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Provost:1990:ATS

[Pro90] Serge B. Provost. Alternate test statistics for a functional errorsin-variables model. Communications in Statistics: Theory and Methods, 19(6):2247–2259, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Phatak:1992:GBP

[PRP92] A. Phatak, P. M. Reilly, and A. Penlidis. The geometry of 2block partial least squares regression. Communications in Statistics: Theory and Methods, 21(6):1517–1553, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Prasad:1990:SIR

[PS90] B. Prasad and H. P. Singh. Some improved ratio-type estimators of finite population variance in sample surveys. Communications in Statistics: Theory and Methods, 19(3):1127–1139, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ponnuswamy:1991:CPT

[PS91] K. N. Ponnuswamy and M. R. Srinivasan. Construction of partial triallel crosses (PTC) using a class of balanced incomplete block designs (BIBD). Communications in Statistics: Theory and Methods, 20(10):3315–3323, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Prasad:1992:UEF

[PS92] B. Prasad and H. P. Singh. Unbiased estimators of finite population variance using auxiliary information in sample surveys. Communications in Statistics: Theory and Methods, 21(5):1367– 1376, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Parsa:1993:SUO

[PS93a] A. Rahulji Parsa and William B. Smith. Scoring under ordered constraints in contingency tables. Communications in Statistics: Theory and Methods, 22(12):3537–3551, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Poskitt:1993:SSF

[PS93b] D. S. Poskitt and M. O. Salau. Stable spectral factorization with applications to the estimation of time series models. Communications in Statistics: Theory and Methods, 22(2):347–367, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Park:1995:CRC

[PS95a] D. K. Park and K. R. Shah. On connectedness of row-column designs. Communications in Statistics: Theory and Methods, 24 (1):87–96, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Park:1995:PEG

[PS95b] Taesung Park and Min-Woong Shin. A practical extension of the generalized estimating equation approach for longitudinal data. Communications in Statistics: Theory and Methods, 24 (10):2561–2579, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Pape:1996:ESA

[PS96] Elinor S. Pape and Ming-Hung Shu. Erlang sampling of an alternating Poisson process. Communications in Statistics: Theory and Methods, 25(6):1283–1288, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Pawlas:1999:RRS

[PS99a] Plotr Pawlas and Dominik Szynal. Recurrence relations for single and product moments of k-th record values from Pareto, generalized Pareto and Burr distributions. Communications in Statistics: Theory and Methods, 28(7):1699–1709, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Press:1999:NCN

[PS99b] S. James Press and Kazuo Shigemasu. A note on choosing the number of factors. Communications in Statistics: Theory and Methods, 28(7):1653–1670, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Pandey:1996:BES

[PSM96] B. N. Pandey, B. P. Singh, and C. S. Mishra. Bayes estimation of shape parameter of classical Pareto distribution under LIVEX loss function. Communications in Statistics: Theory and Methods, 25(12):3125–3145, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Paul:1992:MST

[PT92] S. R. Paul and K. Thiagarajah. Multi-sample test of equal gamma distribution scale parameters in presence of unknown common shape parameter. Communications in Statistics: Theory and Methods, 21(6):1633–1650, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Pham-Gia:1993:BAD

[PT93] T. Pham-Gia, Ph.D., FIS and N. Turkkan, P.Eng. Bayesian analysis of the difference of two proportions. Communications in Statistics: Theory and Methods, 22(6):1755–1771, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Puntanen:1996:SMR

[Pun96] Simo Puntanen. Some matrix results related to a partitioned singular linear model. Communications in Statistics: Theory and Methods, 25(2):269–279, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Puntanen:1997:SFR

[Pun97] Simo Puntanen. Some further results related to reduced singular linear models. Communications in Statistics: Theory and Methods, 26(2):375–385, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

PaezBorrallo:1999:NFC

[PZ99] Jose-M. Paez Borrallo and Santiago Zazo. A new formula for the computation of multivariate factorized moments by using the joint cumulative distribution function. Communications in Statistics: Theory and Methods, 28(11):2613–2622, 1999. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Qin:1998:CRT

[QCL98] Huaizhen Qin, Hengjian Cui, and Yong Li. The N.S. conditions for the ratio of two quadratic forms to have an F -distribution and its applications. Communications in Statistics: Theory and Methods, 27(2):453–471, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Qiu:1994:ENJ

[Qiu94] Peihua Qiu. Estimation of the number of jumps of the jump regression functions. Communications in Statistics: Theory and Methods, 23(8):2141–2155, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Qu:1992:ABC

[QL92] Peng Qu and Wei-Yin Loh. Application of Box–Cox transformations to discrimination for the two-class problem. Communications in Statistics: Theory and Methods, 21(10):2757–2774, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Quan:1994:TSS

[QSC94] Hui Quan, Weichung J. Shih, and Thomas Capizzi. A twostage sequential design and analysis for comparing binomial rates with possibly random dropouts. Communications in Statistics: Theory and Methods, 23(2):467–484, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Qu:1998:ACO

[QT98] Yinsheng Qu and Ming Tan. Analysis of clustered ordinal data with subclusters via a Bayesian hierarchical model. Communications in Statistics: Theory and Methods, 27(6):1461–1475, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Qumsiyeh:1994:BEE

[Qum94] Maher B. Qumsiyeh. Bootstrapping and empirical Edgeworth expansions in multiple linear regression models. Communications in Statistics: Theory and Methods, 23(11):3227–3239, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Qumsiyeh:1997:AEW

[Qum97] Maher B. Qumsiyeh. Asymptotic expansion for weighted least squares in linear regression models. Communications in Statistics: Theory and Methods, 26(3):547–558, 1997. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ranger:1996:CPC

[RA96] George C. Ranger and Frank B. Alt. Choosing principal components for multivariate statistical process control. Communications in Statistics: Theory and Methods, 25(5):909–922, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rabinowitz:1998:CSL

[Rab98] Daniel Rabinowitz. Calculating significance levels when LOD scores are maximized over nuisance parameters. Communications in Statistics: Theory and Methods, 27(6):1477–1485, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Radson:1997:MRC

[Rad97] Darrell Radson. The moving-range chart and autocorrelated processes. Communications in Statistics: Theory and Methods, 26(9):2113–2127, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Raghunath:1990:CDM

[Rag90] Arnab Raghunath. On commutativity of design and model expectations in randomized response surveys. Communications in Statistics: Theory and Methods, 19(10):3751–3757, 1990. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rai:1996:STP

[Rai96] Omkar Rai. A some times pool estimator of mean life under LINEX loss function. Communications in Statistics: Theory and Methods, 25(9):2057–2067, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Raju:1991:TSC

[Raj91] C. Raju. Three-stage chain sampling plans. Communications in Statistics: Theory and Methods, 20(5–6):1777–1801, 1991.

CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rakab:1992:NOTa

[Rak92a] Kamel Rakab. A nearly optimal two stage procedure. Communications in Statistics: Theory and Methods, 21(1):197–201, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rakab:1992:NOTb

[Rak92b] Kamel Rakab. A nearly optimal two stage procedure. Communications in Statistics: Theory and Methods, 21(1):197–201, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Runger:1996:CMS

[RAM96] George C. Runger, Frank B. Alt, and Douglas C. Montgomery. Contributors to a multivariate statistical process control chart signal. Communications in Statistics: Theory and Methods, 25(10):2203–2213, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rao:1990:STT

[Rao90a] B. Raja Rao. A simple time to tumor distribution and the ‘setting the clock back to zero’ property. Communications in Statistics: Theory and Methods, 19(2):583–598, 1990. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rao:1990:OST

[Rao90b] K. S. Madhava Rao. A one- sample test based on subsamples medians. Communications in Statistics: Theory and Methods, 19(12):4559–4568, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rao:1991:CMI

[Rao91] T. J. Rao. On certain methods of improving ratio and regression estimators. Communications in Statistics: Theory and Methods, 20(10):3325–3340, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rao:1992:ETR

[Rao92] B. Raja Rao. Equivalence of the tampered random variable and the tampered failure rate models in accelerated lifetesting for a class of life distributions having the ‘setting the clock back to zero property’. Communications in Statistics: Theory and Methods, 21(3):647–664, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rai:1995:AST

[RAP95] S. N. Rai, B. Abraham, and S. Peiris. Analysis of short time series with an over-dispersion model. Communications in Statistics: Theory and Methods, 24(2):335–348, 1995. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Raqab:1996:LPF

[Raq96] Mohammad Z. Raqab. Linear predictors of future order statistics. Communications in Statistics: Theory and Methods, 25(7): 1553–1568, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rasch:1990:OED

[Ras90] Dieter Rasch. Optimum experimental design in nonlinear regression. Communications in Statistics: Theory and Methods, 19(12):4786–4806, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rodrigues:1997:NQB

[RBC97] Josemar Rodrigues, Heleno Bolfarine, and Gauss M. Cordeiro. Nonlinear quasi-Bayesian theory and inverse linear regression. Communications in Statistics: Theory and Methods, 26(10): 2347–2361, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rodrigues:1993:CSA

[RBLN93] Josemar Rodrigues, Heleno Bolfarine, and Francisco Louzada-Neto. Comparing several accelerated life models. Communications in Statistics: Theory and Methods, 22(8):2297–2308, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rayner:1995:ISC

[RBM95] J. C. W. Rayner, D. J. Best, and K. L. Mathews. Interpreting the skewness coefficient. Communications in Statistics: Theory and Methods, 24(3):593–600, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rom:1994:GFM

[RC94] Dror M. Rom and Laira Connell. A generalized family of multiple test procedures. Communications in Statistics: Theory and Methods, 23(11):3171–3187, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rossetti:1995:EES

[RC95] Manuel D. Rossetti and Gordon M. Clark. Estimating expected sojourn times for a Markov renewal process. Communications in Statistics: Theory and Methods, 24(2):553–579, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Randies:1996:PCR

[RC96] Ronald H. Randies and James A. Calvin. The professional contributions of Robert V. Hogg. Communications in Statistics: Theory and Methods, 25(11):2467–2481, 1996. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Roman:1994:ULT

[RD94] Mureika A. Roman and Dupuis J. Debbie. Using Laplace transform theory to estimate knots. Communications in Statistics: Theory and Methods, 23(10):2775–2786, 1994. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Reboussin:1996:EPI

[RD96] David M. Reboussin and David L. DeMets. Exact permutation inference for two sample repeated measures data. Communications in Statistics: Theory and Methods, 25(10):2223–2238, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rao:1992:CEL

[RDA92] B. Raja Rao, C. V. Damaraju, and Jasem M. Alhumoud. Covariate effect on the life expectancy and percentile residual life functions under the proportional hazards and the accelerated life models. Communications in Statistics: Theory and Methods, 22 (1):257–281, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rao:1993:SCB

[RDA93] B. Raja Rao, C. V. Damaraju, and Jasem M. Alhumoud. Setting the clock back to zero property of a class of bivariate life distributions. Communications in Statistics: Theory and Methods, 22(7):2067–2080, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Robinson:1998:ECA

[RDLT98] Laurence D. Robinson, J. Robert Dorroh, Donald Lien, and Moti L. Tiku. The effects of covariate adjustment in generalized linear models. Communications in Statistics: Theory and Methods, 27(7):1653–1675, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rao:1992:ERR

[RDM92] B. Raja Rao, Richard D. Day, and Gary M. Marsh. Estimation of relative risks from individual and ecological correlation studies. Communications in Statistics: Theory and Methods, 21(1):241– 268, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Reeve:1996:RCC

[Ree96] Russell Reeve. The randomized concentration-controlled trial: mathematical definitions, a dose-adjusting algorithm, and sample size efficiency. Communications in Statistics: Theory and Methods, 25(9):2169–2188, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Reiland:1990:IUS

[Rei90] Thomas W. Reiland. Invexity: an updated survey and new results for nonsmooth functions. Communications in Statistics: Theory and Methods, 19(11):4239–4268, 1990. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Reschenhofer:1996:PVP

[Res96] Erhard Reschenhofer. Prediction with vague prior knowledge. Communications in Statistics: Theory and Methods, 25(3):601– 608, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Reiser:1992:CSS

[RFG92] B. Reiser, D. Faraggi, and I. Guttman. Choice of sample size for testing the P (X > Y). Communications in Statistics: Theory and Methods, 21(3):559–569, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rahman:1992:FTC

[RG92] M. S. Rahman and R. P. Gupta. Family of transformed chisquare distributions. Communications in Statistics: Theory and Methods, 22(1):135–146, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Redner:1994:FEU

[RG94] Richard A. Redner and Kevin Gehringer. Function estimation using partitions of unity. Communications in Statistics: Theory and Methods, 23(7):2059–2078, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ruiz:1996:NRE

[RG96] JoséM Ruiz and Antonio Guillamón. Nonparametrie recursive estimator for mean residual life and vitality function under dependence conditions. Communications in Statistics: Theory and Methods, 25(9):1997–2011, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rahman:1997:TPD

[RGU97] Mezbahur Rahman, D. V. Gokhale, and Aman Ullah. On testimation of a probability density: the normal case. Communications in Statistics: Theory and Methods, 26(2):263–278, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Riedel:1993:SSG

[RI93] Kurt S. Riedel and Kaya Imre. Smoothing spline growth curves with covariates. Communications in Statistics: Theory and Methods, 22(7):1795–1818, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Richardson:1990:TSP

[Ric90] Dejuran Richardson. A two-sample parametric group sequential procedure for monitoring clinical trials. Communications in Statistics: Theory and Methods, 19(8):3111–3126, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rieck:1995:PEI

[Rie95] James R. Rieck. Parameter estimation for ihe Birnbaum– Saunders distribution based on symmetrically censored samples.

Communications in Statistics: Theory and Methods, 24(7):1721– 1736, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rieck:1999:MGF

[Rie99] James R. Rieck. A moment-generating function with application to the Birnbaum–Saunders distribution. Communications in Statistics: Theory and Methods, 28(9):2213–2222, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Raychaudhuri:1996:ISC

[RK96] Aparna Raychaudhuri and Nandini Kannan. Inverse sampling: a composite approach. Communications in Statistics: Theory and Methods, 25(5):1033–1042, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rao:1991:LEL

[RKN91] A. Vasudeva Rao, R. R. L. Kantam, and V. L. Narasimham. Linear estimation of location and scale parameters in the generalized gamma distribution. Communications in Statistics: Theory and Methods, 20(12):3823–3848, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rajagopalan:1991:BEM

[RL91] M. Rajagopalan and A. Loganathan. Bayes estimates of mixing proportions in finite mixture distributions. Communications in Statistics: Theory and Methods, 20(8):2337–2349, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ren:1994:ELM

[RL94] Jian-Jian Ren and Kun-Liang Lu. On L-estimation in linear models. Communications in Statistics: Theory and Methods, 23 (1):137–151, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Russell:1997:ECD

[RL97] K. G. Russell and S. M. Lewis. Efficient cross-over designs allowing a check on the assumption that direct and carry-over effects do not interact. Communications in Statistics: Theory and Methods, 26(12):2835–2854, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ratnaparkhi:1990:NTB

[RM90] Makarand V. Ratnaparkhi and James E. Mosimann. On the normality of transformed beta and unit- gamma random variables. Communications in Statistics: Theory and Methods, 19 (10):3833–3854, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rousseeuw:1993:TNP

[RM93] Peter J. Rousseeuw and Geert Molenberghs. Transformation of non positive semidefinite correlation matrices. Communications in Statistics: Theory and Methods, 22(4):965–984, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Robins:1994:CNC

[Rob94] James M. Robins. Correcting for non-compliance in randomized trials using structural nested mean models. Communications in Statistics: Theory and Methods, 23(8):2379–2412, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rodrigues:1992:KLA

[Rod92] Josemar Rodrigues. The Kullback–Leibler approximation of the marginal posterior density: An application to the linear functional model. Communications in Statistics: Theory and Methods, 21(10):2861–2868, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rodrigues:1998:ISR

[Rod98] Josemar Rodrigues. Inference for the software reliability using asymmetric loss functions: a hierarchical Bayes approach. Communications in Statistics: Theory and Methods, 27(9):2165– 2171, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rogatko:1992:BAM

[Rog92] André Rogatko. Bayesian approach for meta-analysis of controlled clinical trials. Communications in Statistics: Theory and Methods, 21(5):1441–1462, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rosenbaum:1996:SAT

[Ros96] Paul R. Rosenbaum. Sensitivity analysis for two control groups. Communications in Statistics: Theory and Methods, 25(11): 2687–2698, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Randles:1990:MRT

[RP90] Ronald H. Randles and Dawn Peters. Multivariate rank tests for the two-sample location problem. Communications in Statistics: Theory and Methods, 19(11):4225–4238, 1990. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rudy:1991:KSS

[RR91] Gideon A. Rudy and Sister Adele Mareie Rothen. The Kolmogorov–Smirnov statistics in variation testing. Communications in Statistics: Theory and Methods, 20(1):73–86, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ramanathan:1997:THT

[RR97] T. V. Ramanathan and M. B. Rajarshi. Testing homogeneity over time of a parameter of a Markov sequence. Communications in Statistics: Theory and Methods, 26(2):317–330, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Radhakrishna:1992:EPT

[RRA92] C. Radhakrishna, A. V. Dattatreya Rao, and G. V. S. R. Anjaneyulu. Estimation of parameters in a two-component mixture generalized gamma distribution. Communications in Statistics: Theory and Methods, 21(6):1799–1805, 1992. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rashid:1993:RBA

[RRAB93] M. Mushfiqur Rashid, M. Mushfiqur Rashid, James C. Aubuchon, and Ansuman Bagchi. Rank-based analysis of repeated measures block designs. Communications in Statistics: Theory and Methods, 22(10):2783–2811, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rueda:1991:GFP

[RRrA91] R. Rueda, F. O’ Reilly, and V. Pé rez Abreu. Goodness of fit for the Poisson distribution based on the probability generating function. Communications in Statistics: Theory and Methods, 20(10):3093–3110, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rohatgi:1990:MOS

[RS90] Vijay K. Rohatgi and Kandasamy Selvavel. Moments of order statistics from one truncation parameter densities. Communications in Statistics: Theory and Methods, 19(6):2053–2063, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rattihalli:1992:ICR

[RS92] R. N. Rattihalli and D. T. Shirke. An improved confidence region for the common mean vector of two multivariate homoscedastic normal distributions. Communications in Statistics: Theory and Methods, 21(5):1319–1325, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rueda:1995:RRU

[RS95] Cristina Rueda and Bomfacio Salvador. Reduction of risk using restricted estimators. Communications in Statistics: Theory and Methods, 24(4):1011–1022, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Raveh:1996:SDT

[RS96] Adi Raveh and Gideon Schwarz. Simple diagnostic tools for investigating linear trends in time series. Communications in Statistics: Theory and Methods, 25(7):1505–1524, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rai:1998:ERC

[RS98] Anil Rai and A. K. Srivastava. Estimation of regression coefficient from survey data based on test of significance. Communications in Statistics: Theory and Methods, 27(3):761–773, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rancel:1999:MPI

[RS99] M. Mercedes Suárez Rancel and Miguel A. Gonzalez Sierra. Measures and procedures for the identification of locally influential observations in linear regression. Communications in Statistics: Theory and Methods, 28(2):343–366, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rohatgi:1990:LMP

[RSAJ90] Vijay K. Rohatgi, A. K. Md Ehsanes Sleh, Rashpal Ahluwalia, and Ping Ji. Locally most powerful tests for the two sample problem when the combined sample is Type II censored. Communications in Statistics: Theory and Methods, 19(6):2337–2355, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Robins:1991:CNC

[RT91] James M. Robins and Anastasios A. Tsiatis. Correcting for non-compliance in randomized trials using rank preserving structural failure time models. Communications in Statistics: Theory and Methods, 20(8):2609–2631, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rekab:1996:EBP

[RT96a] Kamel Rekab and Mohamed Tahir. An empirical Bayes procedure for testing a truncation parameter. Communications in Statistics: Theory and Methods, 25(6):1383–1390, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Roberts:1996:MCC

[RT96b] Harry V. Roberts and Ruey S. Tsay. Making control charts more effective by time series analysis: three illustrative applications. Communications in Statistics: Theory and Methods, 25(11):2767–2796, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ratnasabapathi:1992:OEP

[RTS92] D. Ratnasabapathi, A. Thavaneswaran, and Jagbir Singh. Optimal estimation of polynomial hazard functions. Communications in Statistics: Theory and Methods, 21(8):2309–2323, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ruggeri:1990:PRF

[Rug90] Fabrizio Ruggeri. Posterior ranges of functions of parameters under priors with specified quantiles. Communications in Statistics: Theory and Methods, 19(1):127–144, 1990. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ruggeri:1991:RBA

[Rug91] Fabrizio Ruggeri. Robust Bayesian analysis given a lower bound on the probability of a set. Communications in Statistics: Theory and Methods, 20(5–6):1881–1891, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ruggeri:1993:RBA

[Rug93] Fabrizio Ruggeri. Robust Bayesian analysis given bounds on the probability of a set. Communications in Statistics: Theory and Methods, 22(11):2983–2998, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rodriguez:1995:CMT

[RV95] Juan C. Rodriguez and Aldo J. Viollaz. A Cramér-von Mises type goodness of fit test with asymmetric weight function. Communications in Statistics: Theory and Methods, 24(4):1095– 1120, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rodriguez:1999:WCM

[RV99] Juan C. Rodríguez and Aldo J. Viollaz. A weighted Cramér–von Mises statistic derived from a decomposition of the Anderson– Darling statistic. Communications in Statistics: Theory and Methods, 28(10):2333–2346, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Raghavarao:1994:EDA

[RW94] Damaraju Raghavarao and James B. Wiley. Experimental designs for availability effects and cross effects with one attribute. Communications in Statistics: Theory and Methods, 23(6):1835– 1846, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rosen:1996:CEM

[RW96] Ori Rosen and Ishay Weissman. Comparison of estimation methods in extreme value theory. Communications in Statistics:

Theory and Methods, 25(4):759–773, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Runger:1995:ARL

[RWP95] George C. Runger, Thomas R. Willemain, and Sharad Prabhu. Average run lengths for cusum control charts applied to residuals. Communications in Statistics: Theory and Methods, 24 (1):273–282, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rychlik:1993:SBE

[Ryc93] Tomasz Rychlik. Sharp bounds on L-estimates and their expectations for dependent samples. Communications in Statistics: Theory and Methods, 22(4):1053–1068, 1993. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rychlik:1994:C

[Ryc94] T. Rychlik. Corrigendum. Communications in Statistics: Theory and Methods, 23(1):305–306, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Rodrigues:1994:WBL

[RZ94] Josemar Rodrigues and Arnold Zellner. Weighted balanced loss function and estimation of the mean time to failure. Communications in Statistics: Theory and Methods, 23(12):3609–3616, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Raghavarao:1998:UOU

[RZ98] Damaraju Raghavarao and Bei Zhou. Universal optimality of UE 3-designs for a competing effects model. Communications in Statistics: Theory and Methods, 27(1):153–164, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

SanchezdeNaranjo:1999:EMD

[S99] M. V. Sánchez de Naranjo. Estimating mixing density from a system of observations. Communications in Statistics: Theory and Methods, 28(5):1005–1020, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sadooghi-Alvandi:1990:SPI

[SA90a] S. M. Sadooghi-Alvandi. Simultaneous prediction intervals for regression models with an intercept. Communications in Statistics: Theory and Methods, 19(4):1433–1441, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Saweris:1990:DSS

[SA90b] Nashat Saweris and Masoom M. Ali. The distributions of some statistics associated with two multivariate hypotheses under violations in the complex case. Communications in Statistics: Theory and Methods, 19(4):1343–1357, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Senyo:1990:PIU

[SA90c] B. Senyo and S. K. Adjibolosoo. A procedure for improving upon the performance of the traditional heteroskedasticity pretest estimator. Communications in Statistics: Theory and Methods, 19(5):1899–1912, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Soundararajan:1992:CDS

[SA92a] V. Soundararajan and S. Devaraj Arumainayagam. Conditional double sampling system. Communications in Statistics: Theory and Methods, 21(7):2019–2044, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Soundararajan:1992:LSD

[SA92b] V. Soundararajan and S. Devaraj Arumainayagam. Lot sensitive double sampling plan (LSDSP). Communications in Statistics: Theory and Methods, 21(10):2931–2948, 1992. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Soundararajan:1993:IID

[SA93] V. Soundararajan and S. Devaraj Arumainayagam. IQL indexed double sampling plans that are compatible with MIL-STD-105 D in structure. Communications in Statistics: Theory and Methods, 22(9):2569–2597, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sharma:1996:MRS

[SA96] Subhash C. Sharma and Mukhtar M. Ali. Moments of regression F -statistics under non-normality. Communications in Statistics: Theory and Methods, 25(7):1631–1658, 1996. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sung:1998:PRI

[SA98] Chang Sup Sung and Sung Jin Ahn. A proportional-reduction-in-impurity measure of association for categorical variables. Communications in Statistics: Theory and Methods, 27(8):2083– 2110, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Samawi:1999:PET

[SA99] Hani M. Samawi and Raed R. K. Abu Awwad. Power estimation for two sample tests using balanced resampling. Communications in Statistics: Theory and Methods, 28(5):1073–1092, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shen:1997:ABC

[SAK97] Larry Z. Shen and Hussein R. Al-Khalidi. Application of boundary calculation methodologies to group sequential testing in general parametric models. Communications in Statistics: Theory and Methods, 26(9):2173–2190, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Salicru:1990:EMA

[Sal90] Miquel Salicrú. Entropy measures associated with j-divergences. Communications in Statistics: Theory and Methods, 19(10): 3687–3698, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sankoh:1992:BOS

[San92] Abdul J. Sankoh. Bayesian optimal stratified sampling design. Communications in Statistics: Theory and Methods, 21(11): 3185–3196, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Alvandi:1992:EBP

[SAP92] S. M. Sadooghi-Alvandi and A. Parsian. Estimation of the binomial parameter_{n} using a LINEX loss function. Communications in Statistics: Theory and Methods, 21(5):1427–1439, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sapatinas:1994:LER

[Sap94] Theofanis Sapatinas. Letter to the editor: A remark on Consul’s (1990) counter-example. Communications in Statistics: Theory and Methods, 23(7):2127–2128, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sarkar:1991:ECM

[Sar91] Sanat K. Sarkar. On estimating the common mean of several normal populations under the Pitman closeness criterion. Communications in Statistics: Theory and Methods, 20(11):3487–3498, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Saran:1992:CRG

[Sar92a] Jagdish Saran. Correction to a result of S. G. Mohanty and B. R. Handa on a generalized random walk. Communications in Statistics: Theory and Methods, 21(4):1115–1123, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sarkar:1992:NEC

[Sar92b] Nityananda Sarkar. A new estimator combining the ridge regression and the restricted least squares methods of estimation. Communications in Statistics: Theory and Methods, 21(7):1987– 2000, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sarabia:1995:CNC

[Sar95] José-María Sarabia. The centered normal conditionals distribution. Communications in Statistics: Theory and Methods, 24(11):2889–2900, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Saracco:1997:ATS

[Sar97] JÉRÔ Me Saracco. An asymptotic theory for sliced inverse regression. Communications in Statistics: Theory and Methods, 26(9):2141–2171, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Saracco:1999:SIR

[Sar99] Jerome Saracco. Sliced inverse regression under linear constraints. Communications in Statistics: Theory and Methods, 28(10):2367–2393, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sartawi:1991:BPB

[SAS91a] Hana Ahmad Sartawi and Muhammad S. Abu-Salih. Bayesian prediction bounds for the Burr type X model. Communications in Statistics: Theory and Methods, 20(7):2307–2330, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sastry:1991:SGD

[Sas91b] D. V. S. Sastry. A series of group divisible designs. Communications in Statistics: Theory and Methods, 20(5–6):1677–1682, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sathe:1990:NMO

[Sat90] Y. S. Sathe. A note on the mean of the order statistics based on independent exponential random variables. Communications in Statistics: Theory and Methods, 19(10):3855–3858, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Satam:1991:CNP

[Sat91a] Meena R. Satam. A class of new PBIB designs. Communications in Statistics: Theory and Methods, 20(7):1971–1976, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sathe:1991:FNT

[Sat91b] Y. S. Sathe. A further note on testing for multisource contamination in location/scale families. Communications in Statistics: Theory and Methods, 20(9):2913–2917, 1991. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Stewart:1990:SAO

[SB90] Frances P. Stewart and Ralph A. Bradley. Some adjusted orthogonal, variance balanced, multidimensional block designs. Communications in Statistics: Theory and Methods, 19(6):2109– 2144, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Stewart:1991:SPV

[SB91] Frances P. Stewart and Ralph A. Bradley. Some partially variance balanced structurally incomplete row-column designs. Communications in Statistics: Theory and Methods, 20(8):2451– 2465, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Suresh:1993:DTN

[SB93] K. K. Suresh and S. Balamurali. Designing of tightened-normal-tightened (TNT) plans indexed by MAPD. Communications in Statistics: Theory and Methods, 22(7):2043–2056, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shitan:1995:ATS

[SB95] Mahendran Shitan and Peter J. Brockwell. An asymptotic test for separability of a spatial autoregressive model. Communications in Statistics: Theory and Methods, 24(8):2027–2040, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Schaalje:1997:RIP

[SB97] G. Bruce Schaalje and Brian D. Beus. A reliability index for presence-absence data. Communications in Statistics: Theory and Methods, 26(2):355–374, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shin:1996:SSC

[SBS96] Dong Wan Shin, Ayanendranath Basu, and Sahadeb Sarkar. Small sample comparisons for the blended weight chi-square goodness-of-fit test statistics. Communications in Statistics: Theory and Methods, 25(1):211–226, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Steinberg:1992:ELR

[SC92] Dan Steinberg and N. Scott Cardell. Estimating logistic regression models when the dependent variable has no variance. Communications in Statistics: Theory and Methods, 21(2):423–450, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shinozaki:1993:MEB

[SC93] Nobuo Shinozaki and Yuan-Tsung Chang. Minimaxity of empirical Bayes estimators of the means of independent normal variables with unequal variances. Communications in Statistics: Theory and Methods, 22(8):2147–2169, 1993. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shinozaki:1996:MEB

[SC96] Nobuo Shinozaki and Yuan-Tsung Chang. Minimaxity of empirical Bayes estimators shrinking toward the grand mean when variances are unequal. Communications in Statistics: Theory and Methods, 25(1):183–199, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Samanta:1997:VSD

[SC97] Sisir Kumar Samanta and Shoutir Kishore Chatterjee. Variable selection for discrimination among several populations. Communications in Statistics: Theory and Methods, 26(11):2565–2582, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Scarf:1992:EFP

[Sca92] P. A. Scarf. Estimation for a four parameter generalized extreme value distribution. Communications in Statistics: Theory and Methods, 21(8):2185–2201, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Schechtman:1990:TC

[Sch90] Kenneth B. Schechtman, Ph.D. Triserial correlation. Communications in Statistics: Theory and Methods, 19(8):3011–3022, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Schmoyer:1991:ORG

[Sch91] Richard L. Schmoyer. Order-restricted goodness-of-fit tests based on spacings. Communications in Statistics: Theory and Methods, 20(4):1409–1425, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Schmid:1997:PET

[Sch97] Uwe Schmid. Percentile estimators for the three-parameter Weibull distribution for use when all parameters are unknown.

Communications in Statistics: Theory and Methods, 26(3):765– 785, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Scollnik:1995:BAI

[Sco95a] David P. M. Scollnik. Bayesian analysis of an intervened Poisson distribution. Communications in Statistics: Theory and Methods, 24(3):735–754, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Scollnik:1995:BAT

[Sco95b] David P. M. Scollnik. Bayesian analysis of two overdispersed Poisson regression models. Communications in Statistics: Theory and Methods, 24(11):2901–2918, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Scollnik:1997:ICS

[Sco97] David P. M. Scollnik. Inference concerning the size of the zero class from an incomplete Poisson sample. Communications in Statistics: Theory and Methods, 26(1):221–236, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Scott:1999:WCM

[Sco99] W. F. Scott. A weighted Cramér–von Mises statistic, with some applications to clinical trials. Communications in Statistics: Theory and Methods, 28(12):3001–3008, 1999. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shein-Chung:1994:EVC

[SCSG94] Chow Shein-Chung and Wang Song-Gui. On the estimation of variance components in stability analysis. Communications in Statistics: Theory and Methods, 23(1):289–303, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Simpson:1996:WLR

[SCXG96] Douglas G. Simpson, Raymond J. Carroll, Minge Xie, and Daniel J. Guth. Weighted logistic regression and robust analysis of diverse toxicology data. Communications in Statistics: Theory and Methods, 25(11):2615–2632, 1996. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Swanepoel:1990:PTU

[SD90] J. W. Swanepoel II and C. E. De Beer. A powerful test for uniformity to detect highly peaked alternative densities. Communications in Statistics: Theory and Methods, 19(8):2781–2798, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shah:1991:MPT

[SD91] K. R. Shah and Sheela S. Deo. Missing plot technique in linear models. Communications in Statistics: Theory and Methods, 20(10):3239–3252, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Swanepoel:1992:NNE

[SD92] J. W. H. Swanepoel and C. F. De Beer. A new nonparametric estimator for the mean of the selected population. Communications in Statistics: Theory and Methods, 21(11):3229–3246, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Steyn:1993:ODT

[SD93] P. W. Steyn and D. J. De Waal. Occupancy distributions for testing marginal homogeneity. Communications in Statistics: Theory and Methods, 22(5):1299–1314, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sutradhar:1995:AEF

[SD95] Brajendra C. Sutradhar and Kalyan Das. Analysing exponential family based messy data. Communications in Statistics: Theory and Methods, 24(10):2683–2699, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sutradhar:1997:GLM

[SD97] Brajendra C. Sutradhar and Kalyan Das. Generalized linear models for beta correlated binary longitudinal data. Communications in Statistics: Theory and Methods, 26(3):617–635, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Streit:1997:EGM

[SE97] R. L. Streit and Luginbuhl T. E. Estimation of Gaussian mixtures with rotationally invariant covariance matrices. Communications in Statistics: Theory and Methods, 26(12):2927–2944, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Segerstedt:1992:ORR

[Seg92] Bo Segerstedt. On ordinary ridge regression in generalized linear models. Communications in Statistics: Theory and Methods, 21(8):2227–2246, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Seidel:1998:MMM

[Sei98] Wilfried Seidel. Moments in mixture models with non-identically distributed observations. Communications in Statistics: Theory and Methods, 27(8):2013–2024, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sen:1991:C

[Sen91] P. K. Sen. Comments. Communications in Statistics: Theory and Methods, 20(11):3439–3443, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sethuraman:1990:ADR

[Set90] Jayaram Sethuraman. The asymptotic distribution of the Rényi maximal correlation. Communications in Statistics: Theory and Methods, 19(11):4291–4298, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sambamoorthi:1994:SPI

[SET94] N. Sambamoorthi, V. J. Ervin, and G. Thomas. Simultaneous prediction intervals for multinomial logistic regression models. Communications in Statistics: Theory and Methods, 23(3):815– 829, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Spurrier:1996:ESL

[SET96] John D. Spurrier, Don Edwards, and Lori A. Thombs. Elementary statistics laboratory. Communications in Statistics: Theory and Methods, 25(11):2663–2673, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sugiyama:1999:RRC

[SFT99] Takakazu Sugiyama, Masafumi Fukuda, and Yuichi Takeda. Recurrence relations of coefficients of the generalized hypergeometric function in multivariate analysis. Communications in Statistics: Theory and Methods, 28(3–4):825–837, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Singh:1990:EST

[SG90] Murari Singh and Sudhir Gupta. Estimating the sum of two ratios of parameters and an application. Communications in Statistics: Theory and Methods, 19(9):3267–3282, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Srinivasan:1991:CIP

[SG91a] Syamala Srinivasan and Franklin A. Graybill. Confidence intervals for proportions of total variations in unbalanced two-way components of variance models using unweighted means. Communications in Statistics: Theory and Methods, 20(2):511–526, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Srivastava:1991:UEM

[SG91b] V. K. Srivastava and D. E. A. Giles. Unbiased estimation of the mean squared error of the feasible generalised ridge regression estimator. Communications in Statistics: Theory and Methods, 20(8):2375–2386, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shi:1994:RCM

[SG94] Peide Shi and Li Guoying. On the rates of convergence of “minimum L_{1}-norm” estimates in a partly linear model. Communications in Statistics: Theory and Methods, 23(1):175–196, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Salsas:1999:PVO

[SGA99] Pau Salsas, Montserrat Guillen, and Ramon Alemany. Perfect value and outlier detection in logistic binary choice models. Communications in Statistics: Theory and Methods, 28(6): 1447–1460, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shetty:1998:CDF

[SGB98] I. D. Shetty, Z. Govindarajulu, and Sharada V. Bhat. A class of distribution-free tests for testing homogeneity against ordered location alternatives. Communications in Statistics: Theory and Methods, 27(3):747–760, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Srivastava:1990:RSD

[SGD90] R. Srivastava, V. K. Gupta, and A. Dey. Robustness of some designs against missing observations. Communications in Statistics: Theory and Methods, 19(1):121–126, 1990. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shahane:1995:EBR

[SGJ95] Aditi Shahane, Varghese George, and William D. Johnson. Effect of bivariate regression toward the mean in uncontrolled clinical trials. Communications in Statistics: Theory and Methods, 24(8):2165–2181, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Singh:1993:EPE

[SGN93] Harshinder Singh, Rameshwar D. Gupta, and Neeraj Nisra. Estimation of parameters of an exponential distribution when the parameter space is restricted with an application to two-sample problem. Communications in Statistics: Theory and Methods, 22 (2):461–477, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sharyn:1994:DBS

[SGS94] Campbell A. Sharyn, Anant P. Godbole, and Stephanie Schaller. Discriminating between sequences of Bernoulli and Markov– Bernoulli trials. Communications in Statistics: Theory and Methods, 23(10):2787–2814, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Son:1990:ASP

[SH90] M. S. Son and H. I. Hamdy. On accelerating sequential procedures for estimating exponential populations. Communications in Statistics: Theory and Methods, 19(9):3315–3324, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Salahuddin:1991:CVS

[SH91] Salahuddin and A. G. Hawkes. Cross-validation in stepwise regression. Communications in Statistics: Theory and Methods, 20(4):1163–1182, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shahjahan:1994:ETR

[SH94] Khan Shahjahan and M. Safiul Haq. β-expectation tolerance region for the multilinear model with matrix-t error distribution. Communications in Statistics: Theory and Methods, 23(7):1935– 1951, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shanubhogue:1990:CDF

[Sha90] Ashok Shanubhogue. A class of distribution-free tests for twoway layout. Communications in Statistics: Theory and Methods, 19(7):2517–2526, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shalaby:1992:BIT

[Sha92a] O. A. Shalaby. Bayesian inference in truncated and censored exponential distribution and reliability estimation. Communications in Statistics: Theory and Methods, 22(1):57–79, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shao:1992:SCD

[Sha92b] Jun Shao. Some conditions on the design of a regression model. Communications in Statistics: Theory and Methods, 21(4):943– 952, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shanmugam:1998:IMD

[Sha98] Ramalingam Shanmugam. Index method of determining the sample size. Communications in Statistics: Theory and Methods, 27(4):1001–1008, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shanmugam:1999:KLI

[Sha99] Ramalingam Shanmugam. Kullback–Leibler information and interval estimation. Communications in Statistics: Theory and Methods, 28(9):2057–2063, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shen:1994:ORI

[She94] Larry Zaiqian Shen. On optimal B-robust influence functions in multidimensional parametric models. Communications in Statistics: Theory and Methods, 23(4):1103–1122, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shetty:1998:LE

[She98] I. D. Shetty. Letter to the editor. Communications in Statistics: Theory and Methods, 27(8):2113, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shi:1995:MIR

[Shi95] Ning-Zhong Shi. The minimal L_{1} isotonic regression. Communications in Statistics: Theory and Methods, 24(1):175–189, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shin:1999:WST

[Shi99] Dong Wan Shin. Weighted symmetric tests for cointegration based on residual. Communications in Statistics: Theory and Methods, 28(1):179–195, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shyr:1998:CFA

[Shy98] Yu Shyr. A canonical form of Anderson’s classification statistic. Communications in Statistics: Theory and Methods, 27(3):577– 587, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sugiura:1997:RPB

[SI97] Nariaki Sugiura and Hiroki Ishibayashi. Reference prior Bayes estimator for bivariate normal covariance matrix with risk comparison. Communications in Statistics: Theory and Methods, 26(9):2203–2221, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Silvapulle:1994:TAD

[Sil94] Paramsothy Silvapulle. Testing for AR(1) against IMA(1, 1) disturbances in the linear regression model. Communications in Statistics: Theory and Methods, 23(3):701–720, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Singh:1990:SRI

[Sin90] Randhir Singh. A study on the role of interviewer in surveys. Communications in Statistics: Theory and Methods, 19(5):1751– 1761, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Singh:1991:DVR

[Sin91] B. Singh. Distribution of variance ratio in unbalanced random model under heterogeneity of error variances. Communications in Statistics: Theory and Methods, 20(9):3021–3028, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sinha:1992:CBT

[Sin92] Kishore Sinha. Construction of balanced treatment incomplete block designs. Communications in Statistics: Theory and Methods, 21(5):1377–1382, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Singh:1997:GSM

[Sin97] V. K. Singh. A generalised stochastic model for the most recent inter-arrival time. Communications in Statistics: Theory and Methods, 26(6):1411–1424, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sen:1990:TBE

[SJ90] Kanwar Sen and Madhu Bala Jain. A test for bivariate exponentiality against BHNBUE alternative. Communications in Statistics: Theory and Methods, 19(5):1827–1835, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sen:1991:BEA

[SJ91a] Kanwar Sen and Madhu Bala Jain. Bayes estimators of the average failure rates. Communications in Statistics: Theory and Methods, 20(7):1983–1988, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sen:1991:NTB

[SJ91b] Kanwar Sen and Madhu Bala Jain. A new test for bivariate distributions: exponential vs new-better-than-used alternative.

Communications in Statistics: Theory and Methods, 20(3):881– 887, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sen:1991:TBE

[SJ91c] Kanwar Sen and Madhu Bala Jain. A test for bivariate exponentiality against BIFR alternative. Communications in Statistics: Theory and Methods, 20(10):3139–3145, 1991. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sen:1991:TBM

[SJ91d] Kanwar Sen and Madhu Bala Jain. Tests for bivariate mean residual life. Communications in Statistics: Theory and Methods, 20(8):2549–2558, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shanubhogue:1993:CDI

[SJ93] A. Shanubhogue and P. N. Jani. Consequences of departure from independence in two component systems when component life times depend through FGM model. Communications in Statistics: Theory and Methods, 22(10):2969–2981, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sen:1997:MGP

[SJ97] Kanwar Sen and Ritu Jain. A multivariate generalized Pólya– Eggenberger probability model — first passage approach. Communications in Statistics: Theory and Methods, 26(4):871–884, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sinha:1991:ESE

[SK91a] Kishore Sinha and Sanpei Kageyama. Equiblock-sized efficiencybalanced designs. Communications in Statistics: Theory and Methods, 20(9):2933–2941, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sungur:1991:LDM

[SK91b] Engin A. Sungur and Milorad S. Kovacevic. Lower dimensional marginal density functions of absolutely continuous truncated multivariate distributions. Communications in Statistics: Theory and Methods, 20(5–6):1569–1578, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Saleh:1993:PSN

[SK93] A. K. Md. Ehsanes Saleh and B. M. Golam Kibria. Performance of some new preliminary test ridge regression estimators and their properties. Communications in Statistics: Theory and Methods, 22(10):2747–2764, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sarkar:1995:MIB

[SK95] A. Sarkar and P. P. Kanjilal. On a method of identification of best subset model from full AR-model. Communications in Statistics: Theory and Methods, 24(6):1551–1567, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sohn:1996:BEP

[SK96] Joong Kweon Sohn and Sang Gil Kang. Bayesian estimation procedure in multiprocess non-linear dynamic generalized model. Communications in Statistics: Theory and Methods, 25(10): 2281–2296, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shyr:1997:FMP

[SK97a] Yu Shyr and Anant M. Kshirsagar. A formula for a missing plot in a general incomplete block design, when recovery of interblock information is used. Communications in Statistics: Theory and Methods, 26(12):2855–2860, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shyr:1997:SCA

[SK97b] Yu Shyr and Anant M. Kshirsagar. Stepwise canonical analysis in categorical data. Communications in Statistics: Theory and Methods, 26(7):1575–1583, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sugiura:1999:DMA

[SK99] Nariaki Sugiura and Yutaka Kano. Delta method approach in a certain irregular condition. Communications in Statistics: Theory and Methods, 28(3–4):789–807, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Seo:1994:EDS

[SKF94] Takashi Seo, Takashi Kanda, and Yasunori Fujikoshi. The effects on the distributions of sample canonical correlations under nonnormality. Communications in Statistics: Theory and Methods, 23(9):2615–2628, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sproule:1993:ARF

[SKS93] Robert A. Sproule, M. Kendall, and A. Stuart. An alternative reduction formula for the central moments of the general Bernoulli distribution. Communications in Statistics: Theory and Methods, 22(4):1191–1194, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Seely:1994:NSC

[SL94a] Justus F. Seely and Youngjo Lee. A note on the Satterthwaite confidence interval for a variance. Communications in Statistics: Theory and Methods, 23(3):859–869, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sukhatme:1994:DRT

[SL94b] Shashikala Sukhatme and C. H. Lin. Distribution of ranks in a two-sample problem with right censored data. Communications in Statistics: Theory and Methods, 23(6):1753–1769, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shih:1998:BSS

[SL98a] Weichung J. Shih and Jianmin Long. Blinded sample size reestimation with unequal variances and center effects in clinical trials. Communications in Statistics: Theory and Methods, 27 (2):395–408, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Lau:1998:MMH

[sL98b] Tai shing Lau. Mixture models in hazard rates estimation. Communications in Statistics: Theory and Methods, 27(6):1395– 1407, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shi:1999:REL

[SL99] Jian Shi and Tai-Shing Lau. Robust empirical likelihood for linear models under median constraints. Communications in Statistics: Theory and Methods, 28(10):2456–2476, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Stromberg:1997:ISC

[SLL97] Ulf Strömberg, Nedjad Losic, and Jan Lanke. Incorporation of a stopping criterion for futility into the design of a clinical trial with one interim analysis. Communications in Statistics: Theory and Methods, 26(4):1011–1030, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Saleh:1996:RIB

[SLS96] Amina A. Saleh, Youngio Lee, and Justus Seely. Recovery of inter-block information: extensions in a two variance component model. Communications in Statistics: Theory and Methods, 25(9):2189–2200, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sengupta:1991:OTF

[SM91a] S. Sengupta and Rahul Mukerjee. On optimal trend free sampling designs. Communications in Statistics: Theory and Methods, 20(1):299–307, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sohn:1991:MCS

[SM91b] So Young Sohn and Mainak Mazumdar. A Monte Carlo study of variable selection and inferences in a two stage random coefficient linear regression model with unbalanced data. Communications in Statistics: Theory and Methods, 20(12):3761–3791, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sengupta:1992:SOS

[SM92] S. Sengupta and Rahul Mukerjee. Some optimal sampling designs in the presence of polynomial trends. Communications in Statistics: Theory and Methods, 22(1):207–218, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Schwertman:1994:APC

[SM94] Neil C. Schwertman and Ricardo A. Martinez. Approximate Poisson confidence limits. Communications in Statistics: Theory and Methods, 23(5):1507–1529, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Schuster:1996:BDQ

[SM96] Christof Schuster and Reinhard Meisier. Benchmarks derived from quantitative endpoints — an approach accounting for litter effects. Communications in Statistics: Theory and Methods, 25(12):3085–3099, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sen:1997:BSR

[SM97] Kanwar Sen and Sunil Kumar Mathur. A bivariate signed rank test for two sample location problem. Communications in Statistics: Theory and Methods, 26(12):3031–3050, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Small:1993:LPP

[Sma93] John P. Small. The limiting power of point optimal autocorrelation tests. Communications in Statistics: Theory and Methods, 22(9):3907–3916, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Smith:1992:CED

[Smi92a] Murray D. Smith. Comparing the exact distribution of the tstatistic to Student’s distribution when its constituent normal and chi-squared variables are dependent. Communications in Statistics: Theory and Methods, 21(12):3589–3600, 1992. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Smith:1992:E

[Smi92b] William B. Smith. Editorial. Communications in Statistics: Theory and Methods, 21(1):v, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Smith:1994:E

[Smi94] William B. Smith. Editorial. Communications in Statistics: Theory and Methods, 23(1):v, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Smith:1995:E

[Smi95] William B. Smith. Editorial. Communications in Statistics: Theory and Methods, 24(1):v, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Smith:1996:E

[Smi96] William B. Smith. Editorial. Communications in Statistics: Theory and Methods, 25(1):v, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Smith:1997:E

[Smi97] William B. Smith. Editorial. Communications in Statistics: Theory and Methods, 26(1):v, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Smith:1998:EC

[Smi98] William B. Smith. Editorial conundrum. Communications in Statistics: Theory and Methods, 27(8):2111, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Salicru:1993:ADE

[SMMP93] M. Salicrú, M. L. Menendez, D. Morales, and L. Pardo. Asymptotic distribution of (h, ϕ)-entropies. Communications in Statistics: Theory and Methods, 22(7):2015–2031, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Singh:1993:MSD

[SMS93] Ravindra Singh, N. S. Mangat, and Sarjinder Singh. A mail survey design for sensitive character without using randomization device. Communications in Statistics: Theory and Methods, 22 (9):63–69, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Singh:1997:ESM

[SMS97] Sarjinder Singh, N. S. Mangat, and Ravindra Singh. Estimation of size and mean of a sensitive quantitative variable for a subgroup of a population. Communications in Statistics: Theory and Methods, 26(7):1793–1804, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Snapinn:1994:URD

[Sna94] Steven M. Snapinn. Use of the randomization distribution in the analysis of sequential clinical trials. Communications in Statistics: Theory and Methods, 23(2):485–498, 1994. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Snell:1996:TCC

[Sne96] J. Laurie Snell. Teaching a chance course. Communications in Statistics: Theory and Methods, 25(11):2853–2862, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sen:1991:CED

[SNK91] Pranab Kumar Sen, Tapan K. Nayak, and Ravindra Khattree. Comparison of estimators of a dispersion matrix under generalized Pitman nearness criterion. Communications in Statistics: Theory and Methods, 20(11):3473–3486, 1991. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shimizu:1997:MIT

[SNM97] Kunio Shimizu, Nobuharu Nishii, and Mihoko Minami. The multivariate inverse trinomial distribution as a Lagrangian probability model. Communications in Statistics: Theory and Methods, 26(7):1585–1598, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Srivastava:1995:CICa

[SO95] Viren K. Srivastava and Kazuhiro Ohtani. A comparison of interval constrained least squares and mixed regression estimators. Communications in Statistics: Theory and Methods, 24 (2):395–413, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Stanek:1998:ERR

[SO98] Edward J. Stanek III and James R. Ohearn. Estimating realized random effects. Communications in Statistics: Theory and Methods, 27(5):1021–1048, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sobel:1990:CRP

[Sob90] Marc J. Sobel. Complete ranking procedures with appropriate loss functions. Communications in Statistics: Theory and Methods, 19(12):4525–4544, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sobel:1992:CVA

[Sob92a] Marc Sobel. Cross-validated assessments of test scoring models. Communications in Statistics: Theory and Methods, 21(11):

3037–3051, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sobel:1992:SPO

[Sob92b] Marc Sobel. On selecting the Pareto-optimal subset of a class of populations. Communications in Statistics: Theory and Methods, 21(4):1085–1102, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sobel:1995:CES

[Sob95] Marc Sobel. Comparisons of estimators of a slippage probability. Communications in Statistics: Theory and Methods, 24(4):1039– 1055, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sobel:1996:OTP

[Sob96] Mare J. Sobel. Optimal transitive procedures for comparing large numbers of parameters. Communications in Statistics: Theory and Methods, 25(6):1155–1172, 1996. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Song:1992:EMS

[Son92] Ruiguang Song. On estimating the mean of the selected uniform population. Communications in Statistics: Theory and Methods, 21(9):2707–2719, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Song:1993:MVS

[Son93] Ruiguang Song. Moments of variables summing to gamma order statistics. Communications in Statistics: Theory and Methods, 22(3):797–803, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sinha:1994:SAS

[SOSW94] Bimal K. Sinha, Robert F. O’Brien, William P. Smith, and John Warren. Some aspects of simultaneous inference. Communications in Statistics: Theory and Methods, 23(1):59–87, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Smith:1990:CTS

[SP90] William B. Smith and Amba R. Parsa. Contingency tables with structural zeros. Communications in Statistics: Theory and Methods, 19(12):4485–4496, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sanso:1994:LCP

[SP94] Bruno Sansó and Luis Raúl Pericchi. Large classes of proper priors for linear models. Communications in Statistics: Theory and Methods, 23(9):2493–2501, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Smith:1999:AVP

[SP99] Peter J. Smith and Lalith W. Peiris. Added variable plots for linear regression with censored data. Communications in Statistics: Theory and Methods, 28(8):1987–2000, 1999. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shao:1998:LSE

[SPQ98] Jun Shao, Mari Palta, and Roger P. Qu. Least squares estimation of regression parameters in mixed effects models with unmeasured covariates. Communications in Statistics: Theory and Methods, 27(6):1487–1501, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Singh:1994:GCS

[SPS94] R. Karan Singh, S. K. Pandey, and V. K. Srivastava. A generalized class of shrinkage estimators in linear regression when disturbances are not normal. Communications in Statistics: Theory and Methods, 23(7):2029–2064, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Spurrier:1991:IBM

[Spu91] John D. Spurrier. Improved bounds for moments of some rank statistics. Communications in Statistics: Theory and Methods, 20(8):2603–2608, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Spurrier:1995:MBM

[Spu95] John D. Spurrier. More bounds for moments of some rank statistics. Communications in Statistics: Theory and Methods, 24(10): 2679–2682, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Saran:1991:DBU

[SR91] Jagdish Saran and Sarita Rani. Distributions based on upcrossings, upward crossings, positive reflections and their runs in a generalized random walk. Communications in Statistics: Theory and Methods, 20(10):3389–3404, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Srivastava:1995:CICb

[Sri95] M. S. Srivastava. Comparison of the inverse and classical estimators in multi-univariate linear calibration. Communications in Statistics: Theory and Methods, 24(11):2753–2767, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Srivastava:1996:EPA

[Sri96] M. S. Srivastava. Economical process adjustment with sampling interval. Communications in Statistics: Theory and Methods, 25(10):2403–2430, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Srivastava:1997:CPM

[Sri97] M. S. Srivastava. Cusum procedure for monitoring variability. Communications in Statistics: Theory and Methods, 26(12): 2905–2926, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Srivastava:1999:LCP

[Sri99] M. S. Srivastava. On-line control procedures for integrated moving average process of order one. Communications in Statistics: Theory and Methods, 28(8):1857–1882, 1999. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sampson:1990:RAA

[SS90a] Allan R. Sampson and David S. Stoffer. Regression from another angle. Communications in Statistics: Theory and Methods, 19(11):4269–4290, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Schick:1990:IPD

[SS90b] A. Schick and V. Susarla. Inference with paired data and partial control. Communications in Statistics: Theory and Methods, 19(10):3901–3913, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Saleh:1991:PCS

[SS91a] A. K. Md. Ehsanes Saleh and P. K. Sen. Pitman-closeness of some preliminary test and shrinkage estimators. Communications in Statistics: Theory and Methods, 20(11):3643–3657, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Schwertman:1991:DAV

[SS91b] Neil C. Schwertman and Kathryn L. Schenk. Determining analysis of variance sample size for detecting specified trends or interactions. Communications in Statistics: Theory and Methods, 20 (2):527–538, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sen:1991:CPC

[SS91c] Pranab K. Sen and Debapriya Sengupta. On characterizations of Pitman closeness of some shrinkage estimators. Communications in Statistics: Theory and Methods, 20(11):3551–3580, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shirahata:1992:EVS

[SS92] S. Shirahata and Y. Sakamoto. Estimate of variance of Ustatistics. Communications in Statistics: Theory and Methods, 21(10):2969–2981, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sarkar:1993:NNR

[SS93a] Sahadeb Sarkar and Dong Wan Shin. A note on nonlinear regression for the autoregressive moving average with non-iid errors. Communications in Statistics: Theory and Methods, 22 (9):501–513, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Schork:1993:SSE

[SS93b] Nicholas J. Schork and M. Anthony Schork. Some sampling effects of pairwise correlated observations on likelihood ratio tests for the difference between two means. Communications in Statistics: Theory and Methods, 22(9):123–129, 1993. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Singh:1993:GFM

[SS93c] Sarjinder Singh and Ravindra Singh. Generalized Franklin’s model for randomized response sampling. Communications in Statistics: Theory and Methods, 22(3):741–755, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shin:1994:LRT

[SS94a] Dong Wah Shin and Sahadeb Sarkar. Likelihood ratio type unit root tests for AR(1) models with nonconsecutive observations. Communications in Statistics: Theory and Methods, 23(5):1387– 1397, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shin:1994:PER

[SS94b] Dong Wan Shin and Sahadeb Sarkar. Parameter estimation in regression models with autocorrelated errors using irregular data. Communications in Statistics: Theory and Methods, 23(12):3567–3580, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shin:1994:TUR

[SS94c] Dong Wan Shin and Sahadeb Sarkar. Testing for a unit root in an ARIMA(p, 1, 0) signal observed with MA(q) noise. Communications in Statistics: Theory and Methods, 23(9):2643–2670, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Speed:1994:AHD

[SS94d] Fred Michael Speed and Fred Michael Speed, Jr. An ad-hoc diagnostic tool for checking for interaction in a nonreplicated experiment. Communications in Statistics: Theory and Methods, 23(5):1365–1374, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Suzukawa:1994:CPE

[SS94e] Akio Suzukawa and Yoshiharu Sato. Comparison of powers and exact Bahadur slopes of some tests for dimensionality in MANOVA models. Communications in Statistics: Theory and Methods, 23(12):3581–3591, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

See:1998:TIT

[SS98] Kyoungah See and Eric P. Smith. Testing interaction in threeway ANOVA tables with a single observation per cell. Communications in Statistics: Theory and Methods, 27(4):839–866, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sakamoto:1999:LBC

[SS99a] Wataru Sakamoto and Shingo Shirahata. Likelihood-based crossvalidation score for selecting the smoothing parameter in maximum penalized likelihood estimation. Communications in Statistics: Theory and Methods, 28(7):1671–1698, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sree:1999:SBG

[SS99b] P. Udaya Sree and R. J. R. Swamy. Selection of the best generalized inverse-Gaussian Markov sequence. Communications in Statistics: Theory and Methods, 28(9):2161–2166, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Stidley:1999:AMR

[SS99c] Christine A. Stidley and Ronald M. Scharder. The analysis of mixed and random effects models for equireplicate variancebalanced block designs. Communications in Statistics: Theory and Methods, 28(10):2509–2526, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sun:1999:ECI

[SS99d] Yijun Sun and Bimal K. Sinha. Exact confidence intervals for a common location of several truncated exponentials. Communications in Statistics: Theory and Methods, 28(3–4):683–692, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sinclair:1990:MAD

[SSA90] C. D. Sinclair, B. D. Spurr, and M. I. Ahmad. Modified Anderson–Darling test. Communications in Statistics: Theory and Methods, 19(10):3677–3686, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Speed:1999:ATD

[SSCS99] F. Michael Speed, William B. Smith, Rong Chen, and F. Michael Speed, Jr. Analysis of tidal data and datums: accessible examples of harmonic modeling with autocorrelation and imputation. Communications in Statistics: Theory and Methods, 28(12):2947–2965, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sinha:1990:NAL

[SSM90] Bikas K. Sinha, Rita Saharay, and Anis C. Mukhopadhyay. Nonadditive linear models: estimability and efficient estimations of interaction. Communications in Statistics: Theory and Methods, 19(2):739–764, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

See:1997:CPB

[SSS97] Kyoungah See, Sung Y. Song, and John Stufken. On a class of partially balanced incomplete block designs with applications in survey sampling. Communications in Statistics: Theory and Methods, 26(1):1–13, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Schwertman:1990:MCC

[SSSS90] Neil C. Schwertman, Todd Stewart, Katherine L. Schenk, and Kathy S. Stone. A Monte Carlo comparison of tests of parallelism of the responses over time of two populations vith non-uniform observation times. Communications in Statistics: Theory and Methods, 19(10):3539–3555, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Singh:1994:GCC

[SSSS94] Vijay K. Singh, Hari P. Singh, Housila P. Singh, and D. Shukla. A general class of chain estimators for ratio and product of two means of a finite population. Communications in Statistics: Theory and Methods, 23(5):1341–1355, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Singh:1993:PFL

[SSW93] Bahadur Singh, Michael J. Schell, and F. T. Wright. The power functions of the likelihood ratio tests for a simple tree ordering in normal means: unequal weights. Communications in Statistics:

Theory and Methods, 22(2):425–449, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Samanta:1990:NPE

[ST90a] M. Samanta and A. Thavaneswaran. Non-parametric estimation of the conditional mode. Communications in Statistics: Theory and Methods, 19(12):4515–4524, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shaked:1990:PDR

[ST90b] Moshe Shaked and Y. L. Tong. Positive dependence of random variables with a common marginal distribution. Communications in Statistics: Theory and Methods, 19(11):4299–4313, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Stemann:1993:LCO

[ST93a] D. Stemann and G. Trenkler. Leverage and Cochrane–Orcutt estimation in linear regression. Communications in Statistics: Theory and Methods, 22(5):1315–1333, 1993. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Subbaiah:1993:ICI

[ST93b] Perla Subbaiah and Winson Taam. Inference on the capability index: Cpm. Communications in Statistics: Theory and Methods, 22(2):537–560, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sugiura:1996:DJS

[ST96] Nariaki Sugiura and Yoshiharu Takagi. Dominating James–Stein positive-part estimator for normal mean with unknown covariance matrix. Communications in Statistics: Theory and Methods, 25(12):2875–2900, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sugiura:1999:HVN

[ST99] Nariaki Sugiura and Naoko Takaoka. Homogeneity of variances in normal linear regression with a change point. Communications in Statistics: Theory and Methods, 28(12):2785–2801, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Steele:1990:PSS

[Ste90] J. Michael Steele. Probability and statistics in the service of computer science: illustrations using the assignment problem. Communications in Statistics: Theory and Methods, 19(11):4315– 4329, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Stewart:1991:GAF

[Ste91] Paul W. Stewart. The graphical advantages of finite interval confidence band procedures. Communications in Statistics: Theory and Methods, 20(12):3975–3993, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Steffey:1992:HBM

[Ste92] Duane Steffey. Hierarchical Bayesian modeling with elicited prior information. Communications in Statistics: Theory and Methods, 21(3):799–821, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sadaghiani:1991:MTW

[STL91] M. K. Sadaghiani, J. L. Teugels, and K. U. Leuven. Models for two-way contingency tables. Communications in Statistics: Theory and Methods, 20(7):2059–2072, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Strickert:1990:ECA

[Str90] Donald P. Strickert. Estimating consumer acceptance limits. Communications in Statistics: Theory and Methods, 19(7):2365– 2472, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Strawderman:1991:C

[Str91a] William E. Strawderman. Comments. Communications in Statistics: Theory and Methods, 20(11):3689–3691, 1991. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Stroud:1991:HBP

[Str91b] T. W. F. Stroud. Hierarchical Bayes predictive means and variances with application to sample survey inference. Communications in Statistics: Theory and Methods, 20(1):13–36, 1991.

CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Stromberg:1995:CLM

[Str95] Arnold J. Stromberg. Consistency of the least median of squares estimator in nonlinear regression. Communications in Statistics: Theory and Methods, 24(8):1971–1984, 1995. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sahli:1997:OSA

[STR97] A. Sahli, P. Trecourt, and S. Robin. One sided acceptance sampling by variables: the case of the Laplace distribution. Communications in Statistics: Theory and Methods, 26(11):2817–2834, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Srivastava:1995:CEC

[STS95] Anil K. Srivastava, Derrick S. Tracy, and Praveen Shukla. Concentration of estimators of coefficients in mixed regression model. Communications in Statistics: Theory and Methods, 24(3):797– 811, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Stufken:1991:BOE

[Stu91] John Stufken. Bayes A-optimal and efficient block designs for comparing test treatments with a standard treatment. Communications in Statistics: Theory and Methods, 20(12):3849–3862, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Stute:1994:IEU

[Stu94] Winfried Stute. Improved estimation under random censorship. Communications in Statistics: Theory and Methods, 23(9):2671– 2682, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Srivastava:1995:SRE

[SU95] Virendra K. Srivastava and Aman Ullah. Stein-rule estimation in models with a lagged-dependent variable. Communications in Statistics: Theory and Methods, 24(5):1343–1353, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Subramani:1991:IQU

[Sub91] J. Subramani. On invariant quadratic unbiased estimation of variance components. Communications in Statistics: Theory and Methods, 20(5–6):1705–1730, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Subramani:1992:AME

[Sub92] J. Subramani. Analysis of means for experimental designs with missing observations. Communications in Statistics: Theory and Methods, 21(7):2045–2057, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sugiura:1999:GBM

[Sug99] Nariaki Sugiura. Generalized Bayes and multiple contrasts tests for simple loop-ordered normal means. Communications in Statistics: Theory and Methods, 28(3–4):549–580, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sundaraiyer:1996:EPC

[Sun96] Vani H. Sundaraiyer. Estimation of a process capability index for inverse Gaussian distribution. Communications in Statistics: Theory and Methods, 25(10):2381–2393, 1996. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sun:1997:WCG

[Sun97] Yanqing Sun. Weak convergence of the generalized parametric empirical processes and goodness-of-fit tests for parametric models. Communications in Statistics: Theory and Methods, 26(10): 2393–2413, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sun:1998:IEG

[Sun98] Xiaoqian Sun. Improved estimation of the generalized precision under the squared loss. Communications in Statistics: Theory and Methods, 27(11):2725–2742, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sungur:1999:TID

[Sun99] Engin A. Sungur. Truncation invariant dependence structures. Communications in Statistics: Theory and Methods, 28(11): 2553–2568, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Singh:1995:BET

[SUS95] Rajesh Singh, S. K. Upadhyay, and Umesh Singh. Bayes estimators for two parameter exponential distribution. Communications in Statistics: Theory and Methods, 24(1):227–240, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Smith:1997:LRL

[SV97] Tom Smith and Penelope Vounatsou. Logistic regression and latent class models for estimating positivities in diagnostic assays with poor resolution. Communications in Statistics: Theory and Methods, 26(7):1677–1700, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Swartz:1995:SPB

[SVM95] T. B. Swartz, C. Villegas, and C. J. Martinez. On semiparametric pivotal Bayesian inference for quantiles. Communications in Statistics: Theory and Methods, 24(10):2499–2515, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Srivastava:1991:SOA

[SW91] M. S. Srivastava and Yanhong Wu. A second order approximation to Taguchi’s online control procedure. Communications in Statistics: Theory and Methods, 20(7):2149–2168, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Scott:1996:MAA

[SW96] David W. Scott and Gerald Whittaker. Multivariate applications of the ASH in regression. Communications in Statistics: Theory and Methods, 25(11):2521–2530, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Srivastava:1997:EOW

[SW97] M. S. Srivastava and Yanhong Wu. Evaluation of optimum weights and average run lengths in EWMA control schemes. Communications in Statistics: Theory and Methods, 26(5):1253– 1267, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Song:1998:OTP

[SW98] Dale Song and Weng Kee Wong. Optimal two-point designs for the Michaelis–Menten model with heteroscedastic errors. Communications in Statistics: Theory and Methods, 27(6):1503– 1516, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Srivastava:1999:EPA

[SW99] M. S. Srivastava and Yanhong Wu. Economical process adjustment with measurement error. Communications in Statistics: Theory and Methods, 28(3–4):989–1003, 1999. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Swartz:1993:SPD

[Swa93] T. Swartz. Subjective priors for the Dirichlet process. Communications in Statistics: Theory and Methods, 22(11):2999–3011, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Swartz:1999:NGF

[Swa99] Tim Swartz. Nonparametric goodness-of-fit. Communications in Statistics: Theory and Methods, 28(12):2821–2841, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Scott:1993:TCC

[SWD93] Douglas M. Scott, Janet L. Waryanka, and Micheal R. Domagala. Teaching control and capability at Alcoa. Communications in Statistics: Theory and Methods, 22(12):3393–3405, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Swift:1998:DLF

[Swi98] Louise Swift. The distribution of the low frequency periodogram ordinates of fractionally differenced series and their inclusion in two estimators of the differencing parameter. Communications in Statistics: Theory and Methods, 27(3):723–745, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Shayib:1991:ERG

[SY91] Mohammed A. Shayib and David H. Young. The extreme residuals in gamma regression. Communications in Statistics: Theory and Methods, 20(2):561–577, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Srivastava:1993:LEM

[SY93] M. S. Srivastava and Wu Yanhong. Local efficiency of moment estimators in beta-binomial model. Communications in Statistics: Theory and Methods, 22(9):257–261, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Siepman:1994:GLS

[SY94] Howard R. Siepman and Shie-Shien Yang. Generalized least squares estimation of multivariate nonlinear models with missing data. Communications in Statistics: Theory and Methods, 23(6): 1565–1579, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Sungur:1996:DCA

[SY96] Engin A. Sungur and Yimin Yang. Diagonal copulas of Archimedean class. Communications in Statistics: Theory and Methods, 25(7):1659–1676, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Szroeter:1995:EPF

[Szr95] Jerzy Szroeter. The exact power function of an exact test of a regression model against multiple separate alternatives. Communications in Statistics: Theory and Methods, 24(9):2329–2339, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Takada:1996:TSP

[TA96] Yoshikazu Takada and Makoto Aoshiina. Two-stage procedures for the difference of two multinormal means with covariance matrices different only by unknown scalar multipliers. Communications in Statistics: Theory and Methods, 25(10):2371–2379, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tahir:1991:BSE

[Tah91] Mohamed Tahir. Bayesian sequential estimation of the ratio of failure rates of two exponential distributions censored data. Communications in Statistics: Theory and Methods, 20(1):63– 71, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tahir:1992:FWS

[Tah92] Mohamed Tahir. Fixed-width sequential confidence interval for the correlation coefficient of a bivariate normal distribution. Communications in Statistics: Theory and Methods, 21(2):501– 506, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Takada:1994:CPP

[Tak94] Yoshikazu Takada. Conditional properties of prediction intervals. Communications in Statistics: Theory and Methods, 23 (1):89–98, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Takada:1997:ECD

[Tak97a] Yoshikazu Takada. Effects of correlated disturbances of some regression problems. Communications in Statistics: Theory and Methods, 26(8):1925–1935, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Takeuchi:1997:GIR

[Tak97b] Hidekazu Takeuchi. A generalized inverse regression estimator in multi-univariate linear calibration. Communications in Statistics: Theory and Methods, 26(11):2645–2669, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Takada:1999:LBB

[Tak99] Yoshikazu Takada. Lower bounds on the Bayes risk EOR statistical prediction problems. Communications in Statistics: Theory and Methods, 28(3–4):693–703, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tang:1994:TSO

[Tan94] Boxin Tang. A theorem for selecting OA-based latin hypercubes using a distance criterion. Communications in Statistics: Theory and Methods, 23(7):2047–2058, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tang:1998:THN

[Tan98] Dei-In Tang. Testing the hypothesis of a normal mean lying outside a convex cone. Communications in Statistics: Theory and Methods, 27(6):1517–1534, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tappin:1992:UEP

[Tap92] Linda Tappin. Unbiased estimation of the parameter of a selected binomial population. Communications in Statistics: Theory and Methods, 21(4):1067–1083, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Taplin:1998:RKF

[Tap98] Ross H. Taplin. The reverse Kalman filter. Communications in Statistics: Theory and Methods, 27(10):2547–2558, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tardif:1990:AOR

[Tar90] Serge Tardif. Asymptotic optimality of rank tests for randomized blocks. Communications in Statistics: Theory and Methods, 19(12):4497–4513, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Targhetta:1991:LBB

[Tar91] M. Luisa Targhetta. Lower bounds for the Bayes risk in estimating a mixing parameter. Communications in Statistics: Theory and Methods, 20(3):873–880, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tripathi:1991:EMU

[TB91] T. P. Tripathi and Shashi Bahl. Estimation of mean using double sampling for stratification and multivariate auxiliary information. Communications in Statistics: Theory and Methods, 20(8):2589–2602, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tanaka:1990:SAM

[TC90] Yutaka Tanaka and Eduardo Tostado Castaño. Sensitivity analysis in multivariate methods: Decomposition of an arbitrary influence into a finite number of components. Communications in Statistics: Theory and Methods, 19(4):1323–1341, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tong:1991:SHT

[TC91] Lee-Ing Tong and Paul L. Cornelius. Studies on the hypothesis testing of the slope parameter in the simple linear regression model with one-fold nested error structure. Communications in Statistics: Theory and Methods, 20(7):2023–2043, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tripathi:1992:IEF

[TC92] T. P. Tripathi and Y. P. Chaubey. Improved estimation of a finite population mean based on paired observations. Communications in Statistics: Theory and Methods, 21(11):3327–3333, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Troskie:1994:DOI

[TCSJ94] C. G. Troskie, D. O. Chalton, T. J. Stewart, and M. Jacobs. Detection of outliers and influential observations in regression analysis using stochastic prior information. Communications in Statistics: Theory and Methods, 23(12):3453–3476, 1994. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tanaka:1990:QPE

[TCT90] Yutaka Tanaka and Eduardo Castaño-Tostado. Quadratic perturbation expansions of certain functions of eigenvalues and eigenvectors and their application to sensitivity analysis in multivariate methods. Communications in Statistics: Theory and Methods, 19(8):2943–2965, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Temple:1994:SSD

[Tem94] Robert J. Temple, M.D. Special study designs: early escape, enrichment, studies in non-responders. Communications in Statistics: Theory and Methods, 23(2):499–531, 1994. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Terui:1990:TSS

[Ter90] Nobuhiko Terui. An F -type small sample simultaneous test for nested linear regression models. Communications in Statistics:

Theory and Methods, 19(2):703–722, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tu:1994:BRJ

[TG94] D. Tu and A. J. Gross. Bias reduction for jackknife skewness. Communications in Statistics: Theory and Methods, 23(8):2323– 2341, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Thail:1992:STT

[Tha92] Peter F. Thail. Score tests in the two-way layout of counts. Communications in Statistics: Theory and Methods, 21(10):3017– 3036, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Thomas:1993:SII

[Tho93] Yageen P. Thomas. On some identities involving the moments of extremes from gamma distribution. Communications in Statistics: Theory and Methods, 22(8):2321–2326, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Thompson:1996:BNE

[Tho96a] Peter Thompson. A bivariate normal example where the locally most powerful and distantly most powerful test statistics are beaten everywhere from a median p-value standpoint. Communications in Statistics: Theory and Methods, 25(10):2343–2347, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Thompson:1996:IAS

[Tho96b] Peter Thompson. Improving the admissibility screen: evaluating test statistics on the basis of p-values rather than power. Communications in Statistics: Theory and Methods, 25(3):537–553, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Timm:1996:NMM

[Tim96] Neil H. Timm. A note on the MANOVA model with double linear restrictions. Communications in Statistics: Theory and Methods, 25(6):1391–1395, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tiwari:1994:NBA

[Tiw94] Ram C. Tiwari. A nonparametric Bayesian approach to analysis of drug dependence data. Communications in Statistics: Theory and Methods, 23(2):625–642, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tracy:1991:FMR

[TK91a] Derrick S. Tracy and Khushnood A. Khan. Fourth moment results for MRBP and related power performance. Communications in Statistics: Theory and Methods, 20(8):2701–2718, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tripathi:1991:CSS

[TK91b] R. C. Tripathi and B. K. Kale. Combined sample scores versus two-sample scores under censoring. Communications in Statistics: Theory and Methods, 20(7):2095–2118, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tiku:1992:EHT

[TK92] M. L. Tiku and N. S. Kambo. Estimation and hypothesis testing for a new family of bivariate nonnormal distributions. Communications in Statistics: Theory and Methods, 21(6):1683–1705, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tsai:1993:SWT

[TK93] Kao-Tai Tsai and James A. Koziol. Score and Wald tests for the multivariate growth curve model with missing data and a patterned covariance matrix. Communications in Statistics: Theory and Methods, 22(2):311–317, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tsai:1994:AME

[TK94] Kao-Tai Tsai and James A. Kozial. Assessing multiple endpoints with ordered alternatives. Communications in Statistics: Theory and Methods, 23(2):533–544, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tripathi:1997:EMV

[TK97a] T. P. Tripathi and B. B. Khare. Estimation of mean vector in presence of non-response. Communications in Statistics: Theory and Methods, 26(9):2255–2269, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tsuchiya:1997:GSA

[TK97b] Takahiro Tsuchiya and Sadanori Konishi. General saddlepoint approximations and normalizing transformations for multivariate statistics. Communications in Statistics: Theory and Methods, 26(11):2541–2563, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Takemura:1999:SSN

[TK99a] Akimichi Takemura and Satoshi Kuriki. Shrinkage to smooth non-convex cone: Principal component analysis as Stein estimation. Communications in Statistics: Theory and Methods, 28(3– 4):651–669, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tsuchiya:1999:GSA

[TK99b] Takahiro Tsuchiya and Sadanori Konishi. General saddlepoint approximations to distributions under an elliptical population. Communications in Statistics: Theory and Methods, 28(3–4): 727–754, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tang:1991:EPA

[TL91] Dei-In Tang and Shang P. Lin. Extension of the pool-adjacent-violators algorithm. Communications in Statistics: Theory and Methods, 20(8):2633–2643, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tsimikas:1994:RBL

[TL94] John Tsimikas and Johnnes Ledolter. Reml and best linear unbiased prediction in state space models. Communications in Statistics: Theory and Methods, 23(8):2253–2268, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Thoresen:1999:IVE

[TL99] Magne Thoresen and Petter Laake. Instrumental variable estimation in logistic measurement error models by means of factor scores. Communications in Statistics: Theory and Methods, 28 (2):297–313, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Thomas:1991:RRM

[TM91] P. Yageen Thomas and T. S. K. Moothathu. Recurrence relations for moments of different orders of extremes from gamma distribution. Communications in Statistics: Theory and Methods, 20(3):945–950, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tudor:1995:ASD

[TM95] Gail E. Tudor and Timothy M. Morgan. Analyzing survival data in conjunction with time-dependent surrogate endpoints in clinical trials. Communications in Statistics: Theory and Methods, 24(9):2293–2313, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tanizaki:1996:NFB

[TM96a] Hisashi Tanizaki and Roberto S. Mariano. Nonlinear filters based on Taylor series expansions. Communications in Statistics: Theory and Methods, 25(6):1261–1282, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Taylor:1996:BLM

[TM96b] Douglas J. Taylor and Keith E. Muller. Bias in linear model power and sample size calculation due to estimating noncentrality. Communications in Statistics: Theory and Methods, 25(7): 1595–1610, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tomizawa:1992:MMF

[Tom92] Sadao Tomizawa. Multiplicative models with further restrictions on the usual symmetry model. Communications in Statistics: Theory and Methods, 21(3):693–710, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tong:1996:CREb

[Ton96a] Hengqing Tong. Convergence rates for empirical Bayes estimators of parameters in linear regression models. Communications in Statistics: Theory and Methods, 25(6):1325–1334, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tong:1996:CREa

[Ton96b] Hengqing Tong. Convergence rates for empirical Bayes estimators of parameters in multi-parameter exponential families. Communications in Statistics: Theory and Methods, 25(5):1089– 1098, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tripsiannis:1997:NMI

[TP97] Gregory A. Tripsiannis and Andreas N. Philippou. A new multivariate inverse Pólya distribution of order k. Communications in Statistics: Theory and Methods, 26(1):149–158, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Trejo:1994:TET

[Tre94] Stephen J. Trejo. A transformation for estimating the trinomial logit model with grouped data. Communications in Statistics: Theory and Methods, 23(12):3539–3548, 1994. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Trosset:1993:OSK

[Tro93] Michael W. Trosset. Optimal shapes for kernel density estimation. Communications in Statistics: Theory and Methods, 22 (2):375–391, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tracy:1990:CSL

[TS90] D. S. Tracy and V. K. Srivastava. Comparison of some linear calibration estimators. Communications in Statistics: Theory and Methods, 19(6):2281–2293, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tripathi:1992:CUP

[TS92] T. P. Tripathi and H. P. Singh. A class of unbiased producttype estimators for the mean suitable for positive and negative correlation situations. Communications in Statistics: Theory and Methods, 21(2):507–518, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tracy:1994:COV

[TS94] Derrick S. Tracy and Anil K. Srivastava. Comparison of operational variants of best homogeneous and heterogeneous estimators in linear regression. Communications in Statistics: Theory and Methods, 23(8):2313–2322, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Trosset:1995:CDF

[TS95] Michael W. Trosset and Barbara N. Sands. On the choice of a discrepancy functional for model selection. Communications in Statistics: Theory and Methods, 24(11):2841–2863, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tsai:1990:PTDb

[Tsa90a] Hsien-Tang Tsai. Probabilistic tolerance design for a subsystem under Burr distribution using Taguchi’s loss functions. Communications in Statistics: Theory and Methods, 19(12):4679–4696, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tsai:1990:PTDa

[Tsa90b] Hsien-Tang Tsai. The probabilistic tolerance design for a subsystem using Taguchi’s quadratic loss function. Communications in Statistics: Theory and Methods, 19(9):3243–3258, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Taneichi:1998:PAT

[TSI98] Nobuhiro Taneichi, Yuri Sekiya, and Hideyuki Imai. A power approximation of the test of homogeneity for multinomial populations based on a normalizing transformation. Communications in Statistics: Theory and Methods, 27(7):1677–1691, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tsionas:1999:BAM

[Tsi99] Efthymios G. Tsionas. Bayesian analysis of the multivariate Poisson distribution. Communications in Statistics: Theory and Methods, 28(2):431–451, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Thavaneswaran:1991:NRS

[TSQ91] A. Thavaneswaran, M. Samanta, and N. Qrdoukhani. A note on recursive smoothers for semimartingales. Communications in Statistics: Theory and Methods, 20(7):2281–2289, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tsukibayashi:1998:JDM

[Tsu98] Shomei Tsukibayashi. The joint distribution and moments of an extreme of the dependent variable and the concomitant of an extreme of the independent variable. Communications in Statistics: Theory and Methods, 27(7):1639–1651, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tsuda:1999:AUT

[Tsu99] Yoshiyuki Tsuda. The approximation of an unbiased test with a small level for the bioequivalence problem. Communications in Statistics: Theory and Methods, 28(5):1105–1114, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tiwari:1992:SWD

[TTG92] Ram C. Tiwari, Ram C. Tripathi, and Rainesh C. Gupta. Some weighted distributions of order k. Communications in Statistics: Theory and Methods, 21(2):411–422, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tu:1999:CBC

[Tu99] Renjin Tu. Comparison between confidence intervals of linear regression models with and without restriction. Communications in Statistics: Theory and Methods, 28(12):2879–2898, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Turner:1990:EWT

[Tur90] David L. Turner. An easy way to tell what you are testing in analysis of variance. Communications in Statistics: Theory and Methods, 19(12):4807–4832, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Turrero:1995:REC

[Tur95] Agustin Turrero. Relative efficiency of a censored experiment in terms of Fisher information matrix. Communications in Statistics: Theory and Methods, 24(5):1169–1191, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tiku:1991:TEL

[TV91] M. L. Tiku and D. C. Vaughan. Testing equality of location parameters of two exponential distributions from censored samples. Communications in Statistics: Theory and Methods, 20 (3):929–944, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tanaka:1992:SAC

[TW92] Yutaka Tanaka and Shingo Watadani. Sensitivity analysis in covariance structure analysis with equality constraints. Communications in Statistics: Theory and Methods, 21(6):1501–1515, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tiku:1999:EPA

[TWB99a] Moti L. Tiku, Wing-Keung Wong, and Guorui Bian. Estimating parameters in autoregressive models in non-normal situations: symmetric innovations. Communications in Statistics: Theory and Methods, 28(2):315–341, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tiku:1999:TSM

[TWB99b] Moti L. Tiku, Wing Keung Wong, and Guorui Bian. Time series models with asymmetric innovations. Communications in Statistics: Theory and Methods, 28(6):1331–1360, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tanaka:1991:ICS

[TWM91] Yutaka Tanaka, Shingo Watadani, and Sung Ho Moon. Influence in covariance structure analysis: with an application to confirmatory factor analysis. Communications in Statistics: Theory and Methods, 20(12):3805–3821, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Troendle:1999:CEF

[TY99] James F. Troendle and Kai F. Yu. Conditional estimation following a group sequential clinical trial. Communications in Statistics: Theory and Methods, 28(7):1617–1634, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Taylor:1990:SCN

[TZ90a] R. L. Taylor and H. M. Zhang. On a strongly consistent nonparametric density estimator for the deconvolution problem. Communications in Statistics: Theory and Methods, 19(9):3325– 3342, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tiwari:1990:NBE

[TZ90b] Ram C. Tiwari and Jyoti N. Zalkikar. Nonparametric Bayes estimation of the survival function from a record of failures and follow-ups. Communications in Statistics: Theory and Methods, 19(11):4331–4354, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Tsairidis:1997:FIM

[TZF97] C. H. Tsairidis, K. Zografos, and K. Ferentinos. Fisher’s information matrix and ϕ-divergence for finite and optimal partitions of the sample space. Communications in Statistics: Theory and Methods, 26(9):2271–2289, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Uddin:1998:OTP

[Udd98] Nizam Uddin. On optimal two-period repeated measurements designs. Communications in Statistics: Theory and Methods, 27 (4):827–837, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Upton:1995:PCD

[UG95] Graham J. G. Upton and Montserrat Guillen. Perfect cells, direct models and contingency table outliers. Communications in Statistics: Theory and Methods, 24(7):1843–1862, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Uno:1999:SPE

[UI99] C. Uno and E. Isogai. A sequential procedure for estimating ratio of normal parameters. Communications in Statistics: Theory and Methods, 28(1):233–244, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ullah:1992:EDN

[Ull92] Aman Ullah. The exact density of nonparametric regression estimators: fixed design case. Communications in Statistics: Theory and Methods, 21(5):1251–1254, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ukita:1992:TLH

[UNM92] Yoshimasa Ukita, Kazuo Noda, and Etsuo Miyaoka. On F -tests and linear hypothesis in a general linear model. Communications in Statistics: Theory and Methods, 21(7):1857–1870, 1992.

CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Ushizawa:1992:EPC

[US92] K. Ushizawa and T. Sugiyama. Exact power comparison of three criteria to assess the independence between two sets of variates. Communications in Statistics: Theory and Methods, 21(10):2741–2756, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

VanAarde:1993:STS

[Van93] I. M. R. Van Aarde. Semirange tests for shortfall. Communications in Statistics: Theory and Methods, 22(2):519–536, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Vannman:1997:DMS

[Vän97a] Kerstin Vännman. Distribution and moments in simplified form for a general class of capability indices. Communications in Statistics: Theory and Methods, 26(1):159–179, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Vannman:1997:GCC

[Vän97b] Kerstin Vännman. A general class of capability indices in the case of asymmetric tolerances. Communications in Statistics: Theory and Methods, 26(8):2049–2072, 1997. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Vardeman:1996:ISP

[Var96] Stephen B. Vardeman. Independent student projects in undergraduate engineering statistics and quality control courses. Communications in Statistics: Theory and Methods, 25(11):2633– 2646, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Varbanov:1998:BAO

[Var98] Alexandre Varbanov. Bayesian approach to outlier detection in multivariate normal samples and linear models. Communications in Statistics: Theory and Methods, 27(3):547–557, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Vasdekis:1996:NDB

[Vas96] V. G. S. Vasdekis. A note on the difference between two likelihood based methods in growth curves. Communications in Statistics: Theory and Methods, 25(9):2093–2100, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Vaughan:1992:TSM

[Vau92a] David C. Vaughan. On the Tiku–Suresh method of estimation. Communications in Statistics: Theory and Methods, 21(2):451– 469, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Vaughan:1992:VSI

[Vau92b] Timothy S. Vaughan. Variable sampling interval NP process control chart. Communications in Statistics: Theory and Methods, 22(1):147–167, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Vaughan:1993:EDR

[Vau93] Timothy S. Vaughan. Experimental design for response surface gradient estimation. Communications in Statistics: Theory and Methods, 22(6):1535–1555, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Verheijen:1997:CTC

[VCvdL97] J. H. M. Verheijen, F. P. A. Coolent, and P. van der Laan. Combining two classical approaches for statistical selection. Communications in Statistics: Theory and Methods, 26(6):1291–1312, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

vanderMerwe:1990:EAB

[vdMvdM90] A. J. van der Merwe and C. A. van der Merwe. Exact-and approximate Bayes credibility intervals for the one-way analysis of variance model. Communications in Statistics: Theory and Methods, 19(7):2541–2565, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Velilla:1992:IOT

[Vel92] S. Velilla. Influential observations in transformations: A Bayesian approach. Communications in Statistics: Theory and Methods, 21(11):3197–3209, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Vellaisamy:1993:UEF

[Vel93] P. Vellaisamy. On UMVU estimation following selection. Communications in Statistics: Theory and Methods, 22(4):1031– 1043, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Veraverbeke:1994:BQP

[Ver94] Noël Veraverbeke. Bootstrapping quantiles in the proportional hazards model of random censorship. Communications in Statistics: Theory and Methods, 23(4):997–1007, 1994. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Versluis:1996:CTB

[Ver96] Cokki Versluis. Comparison of tests for bivariate normality with unknown parameters by transformation to an univariate statistic. Communications in Statistics: Theory and Methods, 25(3): 647–665, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

vanHui:1992:AIT

[vHL92] Yer van Hui and Andy H. Lee. Assessing influence in time series. Communications in Statistics: Theory and Methods, 21(12): 3463–3478, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Vecchia:1991:MQA

[VI91] Dominic F. Vecchia and Hari K. Iyer. Moments of the quartic assignment statistic with an application to multiple regression. Communications in Statistics: Theory and Methods, 20 (10):3253–3269, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Viollaz:1995:CWS

[Vio95] Aldo J. Viollaz. Convergence in weighted supremum norms of the Skorokhod representation of the estimated empirical process. Communications in Statistics: Theory and Methods, 24 (11):2829–2839, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Velilla:1992:ROD

[VL92] S. Velilla and R. Llosa. A result in optimal design with application in particle physics. Communications in Statistics: Theory and Methods, 21(1):203–212, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

VonRosen:1991:GCM

[Von91] Dietrich Von Rosen. The growth curve model: a review. Communications in Statistics: Theory and Methods, 20(9):2791–2822, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

VonTress:1993:LMP

[Von93] Mark Von Tress. Longitudinal models for polytomous responses. Communications in Statistics: Theory and Methods, 22(12): 3523–3536, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Vos:1994:LBM

[Vos94] Paul W. Vos. Likelihood-based measures of influence for generalized linear models. Communications in Statistics: Theory and Methods, 23(12):3477–3490, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Venkateswarlu:1998:EVC

[VP98] K. Venkateswarlu and K. N. Ponnuswamy. Estimation of variance components based on bio-model of dial lei crosses (balanced). Communications in Statistics: Theory and Methods, 27 (1):139–152, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Victor:1990:MBR

[VR90] Yohai J. Victor and Maronna A. Ricardo. The maximum bias of robust covariances. Communications in Statistics: Theory and Methods, 19(10):3925–3933, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Viollaz:1996:CMT

[VR96] Aldo J. Viollaz and Juan C. Rodriguez. A Cramér–von Mises type goodness-of-fit test with asymmetric weight function. The Gaussian and exponential cases. Communications in Statistics:

Theory and Methods, 25(1):235–256, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Viana:1993:BAP

[VRL93] M. A. G. Viana, V. Ramakrishnan, and P. S. Levy. Bayesian analysis of prevalence from the results of small screening samples. Communications in Statistics: Theory and Methods, 22 (2):575–585, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Vijayasree:1991:SET

[VS91] G. Vijayasree and Harshinder Singh. Simultaneous estimation of two ordered exponential parameters. Communications in Statistics: Theory and Methods, 20(8):2559–2576, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Viollaz:1995:AUE

[VSS95] Aldo J. Viollaz, Ana M. Sfer, and Stella M. Salvatierra. An approximation of the unconditional error rates of the sample linear discriminant function. Communications in Statistics: Theory and Methods, 24(8):1941–1969, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Vaughan:1993:TEL

[VT93] D. C. Vaughan and M. L. Tiku. Testing the equality of location parameters of exponential populations from censored samples. Communications in Statistics: Theory and Methods, 22(9):2567– 2581, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Vlachonikolis:1994:CCP

[VV94] I. G. Vlachonikolis and V. G. S. Vasdekis. On a class of changepoint models in covariance structures for growth curves and repeated measurements. Communications in Statistics: Theory and Methods, 23(4):1087–1102, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

VanKeilegom:1996:USC

[VV96] Ingrid Van Keilegom and Noël Veraverbeke. Uniform strong convergence results for the conditional Kaplan–Meier estimator and its quantiles. Communications in Statistics: Theory and Methods, 25(10):2251–2265, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

VanKeilegom:1997:WCB

[VV97] Ingrid Van Keilegom and Noël Veraverbeke. Weak convergence of the bootstrapped conditional Kaplan–Meier process and its quantile process. Communications in Statistics: Theory and Methods, 26(4):853–869, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wesolowski:1995:CSM

[WA95] Jacek Wesolowski and Mohammad Ahsanullah. Conditional specifications of multivariate Pareto and Student distributions. Communications in Statistics: Theory and Methods, 24(4):1023– 1031, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wirjanto:1996:NRM

[WA96] Tony S. Wirjanto and Robert A. Amano. Nonstationary regression models with a lagged dependent variable. Communications in Statistics: Theory and Methods, 25(7):1489–1503, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wang:1991:SAC

[Wan91] You-Gan Wang. Sequential allocation in clinical trials. Communications in Statistics: Theory and Methods, 20(3):791–805, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wan:1993:AEC

[Wan93a] Jim Y. Wan. Asymptotic efficiency of a competing risks model: A two-sample case. Communications in Statistics: Theory and Methods, 22(11):3159–3176, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wang:1993:ACE

[Wan93b] C. Y. Wang. Alternative covariance estimates in a replicated measurement error model with correlated, heteroscedastic errors. Communications in Statistics: Theory and Methods, 22(7): 1819–1828, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wan:1994:NOI

[Wan94a] Alan T. K. Wan. The non-optimality of interval restricted and pre-test estimators under squared error loss. Communications in Statistics: Theory and Methods, 23(8):2231–2252, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wan:1994:ILC

[Wan94b] Jim Y. Wan. Information loss in a competing risks model. Communications in Statistics: Theory and Methods, 23(11):3211– 3226, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wang:1994:ASR

[Wan94c] Gang Wang. Almost sure representation for multivariate mestimators. Communications in Statistics: Theory and Methods, 23(8):2217–2229, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wan:1996:BMS

[Wan96] Alan T. K. Wan. On the bias and mean square error of the least square estimator in a regression model with two inequality constraints and multivariate t error terms. Communications in Statistics: Theory and Methods, 25(9):2079–2091, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wang:1998:CPB

[Wan98a] Sue-Jane Wang. A closed procedure based on Follmann’s test for the analysis of multiple endpoints. Communications in Statistics: Theory and Methods, 27(10):2461–2480, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wang:1998:BFD

[Wan98b] Xiaodong Wang. Bayesian forecasting and detecting structural changepoints in dynamic models. Communications in Statistics: Theory and Methods, 27(6):1535–1546, 1998. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wang:1999:PSC

[Wan99] Peijie Wang. Phase-shifting common cycles and common trends. Communications in Statistics: Theory and Methods, 28(6):1311– 1329, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wassmer:1994:TEC

[Was94] Gernot Wassmer. Testing equivalence in clinical trials using a new principle for constructing statistical tests. Communications in Statistics: Theory and Methods, 23(5):1413–1427, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Washburn:1999:DKS

[Was99] Alan Washburn. A dithered K–S statistic. Communications in Statistics: Theory and Methods, 28(7):1535–1546, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wun:1993:ADR

[WB93] Lap-Ming Wun and Charles C. Brown. An alternative decision rule for quality control in sequential sampling plans. Communications in Statistics: Theory and Methods, 22(7):2033–2042, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wang:1997:MMB

[WB97] Weiren Wang and Douglas C. Bice. A model with mixed binary responses and censored observations. Communications in Statistics: Theory and Methods, 26(4):921–941, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wang:1996:RIE

[WC96a] Yazhen Wang and K. S. Chen. The L_{2} risk of an isotonic estimate. Communications in Statistics: Theory and Methods, 25 (2):281–294, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Weissman:1996:EPU

[WC96b] Ishay Weissman and Ayala Cohen. The estimation of the parameter of uniformly distributed failure-times under censoring. Communications in Statistics: Theory and Methods, 25(4):775– 788, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wang:1997:ECP

[WD97] Ching-Cheng Wang and Stuart Jay Deutsch. Existence conditions for purely spatial processes. Communications in Statistics:

Theory and Methods, 26(8):1997–2007, 1997. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wolfe:1990:NRB

[WDH90] Douglas A. Wolfe, Angela M. Dean, and Bradley A. Hartlaub. Nonparametric rank-based test procedures for nonadditive models in the two-way layout I. No replications. Communications in Statistics: Theory and Methods, 19(11):4355– 4382, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Weerakkody:1992:NRI

[Wee92] Govinda J. Weerakkody. A note on the recovery of inter-block information in balanced incomplete block designs. Communications in Statistics: Theory and Methods, 21(4):1125–1136, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Weng:1994:TDA

[Wen94] T. S. Weng. Toward a dynamic analysis of disease-state transition monitored by serial clinical laboratory tests. Communications in Statistics: Theory and Methods, 23(2):643–660, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wessman:1998:SPS

[Wes98] Peter Wessman. Some principles for surveillance adopted for multivariate processes with a common change point. Communications in Statistics: Theory and Methods, 27(5):1143–1161, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Woldie:1995:CIG

[WF95] Mammo Woldie and J. Leroy Folks. Calibration for inverse Gaussian regression. Communications in Statistics: Theory and Methods, 24(10):2609–2620, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wax:1990:STM

[WG90] Yohanan Wax and Zvi Gilula. A score test for monotonic trend in hazard rates for grouped survival data in stratified populations.

Communications in Statistics: Theory and Methods, 19(9):3375– 3386, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Weerakkody:1995:ECC

[WG95] Govinda J. Weerakkody and Sumalee Givaruangsawat. Estimating the correlation coefficient in the presence of correlated observations from a bivariate normal population. Communications in Statistics: Theory and Methods, 24(7):1705–1719, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Westfall:1996:APA

[WG96] Peter H. Westfall and Mithat Gönen. Asymptotic properties of ANOVA Bayes factors. Communications in Statistics: Theory and Methods, 25(12):3101–3123, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wellmann:1999:NCM

[WG99] Järgen Wellmann and Ursula Gather. A note on contamination models and outliers. Communications in Statistics: Theory and Methods, 28(8):1793–1802, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wang:1999:APV

[WH99] Tonghui Wang and Ruichan He. An alternative proof for a version of Cochran’s theorem under elliptical settings. Communications in Statistics: Theory and Methods, 28(5):1133–1144, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wijesinha:1993:ODI

[Wij93] Manel Cooray Wijesinha. Optimal designs to improve the power of multiresponse hypothesis tests. Communications in Statistics: Theory and Methods, 22(2):587–602, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wiklander:1998:CTE

[Wik98] Kerstin Wiklander. A comparison of two estimators of dispersion effects. Communications in Statistics: Theory and Methods, 27 (4):905–923, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wilcox:1991:SHM

[Wil91] Rand R. Wilcox. A step-down heteroscedastic multiple comparison procedure. Communications in Statistics: Theory and Methods, 20(3):1087–1097, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wilcox:1992:CRR

[Wil92] Rand R. Wilcox. Comparing robust regression lines corresponding to two independent groups. Communications in Statistics: Theory and Methods, 21(5):1255–1266, 1992. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wilson:1994:DAS

[Wil94] Berry Wilson. Discriminant analysis with stratified prior probabilities. Communications in Statistics: Theory and Methods, 23(5):1283–1295, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wilcox:1995:SSS

[Wil95] Rand R. Wilcox. Some small-sample results on a bounded influence rank regression method. Communications in Statistics: Theory and Methods, 24(4):881–888, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wilcox:1996:ESL

[Wil96] Rand R. Wilcox. Estimation in the simple linear regression model when there is heteroscedasticity of unknown form. Communications in Statistics: Theory and Methods, 25(6):1305– 1324, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wilson:1998:LPR

[Wil98] P. David Wilson. Longitudinal Poisson regression with disturbed random intercept. Communications in Statistics: Theory and Methods, 27(9):2275–2292, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Withers:1990:SEB

[Wit90] C. S. Withers. Shrinkage estimates based on orthogonal decomposition of the sample space. Communications in Statistics: Theory and Methods, 19(2):505–526, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Withers:1991:MEM

[Wit91a] C. S. Withers. Moment estimates for mixtures with common scale. Communications in Statistics: Theory and Methods, 20(4): 1445–1461, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Withers:1991:TCI

[Wit91b] C. S. Withers. Tests and confidence intervals for the shape parameter of a gamma distribution. Communications in Statistics: Theory and Methods, 20(8):2663–2685, 1991. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Weerakkody:1990:VCE

[WJ90] Govinda J. Weerakkody and Dallas E. Johnson. Variance component estimation in multiple regression models having a nested error structure. Communications in Statistics: Theory and Methods, 19(8):2879–2905, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wijesinha:1991:RDF

[WK91a] Manel Cooray Wijesinha and Andre I. Khuri. Robust designs for first-order multiple design multivariate models. Communications in Statistics: Theory and Methods, 20(9):2987–2999, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Woollcott:1991:CRM

[WK91b] Smith Woollcott and Ghosh Krishnendu. A continuous repeated measurements model: the experimental design problem with fixed time constraints. Communications in Statistics: Theory and Methods, 20(1):361–372, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

West:1993:SMR

[WK93] Ronnie W. West and Ralph L. Kodell. Statistical methods of risk assessment for continuous variables. Communications in Statistics: Theory and Methods, 22(12):3363–3376, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wang:1997:OSS

[WK97] Yue Wang and Lawrence L. Kupper. Optimal sample sizes for estimating the difference in means between two normal populations treating confidence interval length as a random variable. Communications in Statistics: Theory and Methods, 26(3):727– 741, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Whitcomb:1992:ISI

[WL92] Kathleen M. Whitcomb and Maureen Lahiff. An investigation of sample influence functions for the DS estimate of the optimum error rate in discriminant analysis. Communications in Statistics: Theory and Methods, 22(1):41–56, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wang:1993:BAC

[WL93] China-Yuan Wang and Chung-Bow Lee. Bayesian analysis for a change in the intercept of simple linear regression. Communications in Statistics: Theory and Methods, 22(11):3031–3050, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wang:1996:SPO

[WL96] C. Ming Wang and Hari K. Lyer. Sampling plans for obtaining tolerance intervals in a balanced one-way random-effects model. Communications in Statistics: Theory and Methods, 25(2):313– 324, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wang:1999:BCR

[WLG99] Naisyin Wang, Xihong Lin, and Roberto G. Guttierrez. A bias correction regression calibration approach in generalized linear mixed measurement error models. Communications in Statistics: Theory and Methods, 28(1):217–232, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wilcox:1999:THA

[WM99] Rand R. Wilcox and Jan Muska. Tests of hypotheses about regression parameters when using a robust estimator. Communications in Statistics: Theory and Methods, 28(9):2201–2212, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wei:1992:OWD

[WN92] Feifei Wei and William I. Notz. D-optimal weighing designs and optimal chemical balance weighing designs for estimating the total weight. Communications in Statistics: Theory and Methods, 21(3):667–678, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Williamson:1997:BAT

[WN97] Patricia Pepple Williamson and Sharon E. Navard. A Bayesian approach for testing periodicity of dichotomous observations. Communications in Statistics: Theory and Methods, 26(2):449– 468, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wong:1992:COL

[Won92] Augustine C. M. Wong. Converting observed likelihood to levels of significance for transformation models. Communications in Statistics: Theory and Methods, 21(10):2809–2823, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wong:1993:ESL

[Won93] M. Y. Wong. Estimation of the slope in linear non-normal structural relationships. Communications in Statistics: Theory and Methods, 22(9):1–32, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wong:1995:SAI

[Won95] A. C. M. Wong. A simple and accurate inference procedure for the difference of failure intensities from two exponential distributions. Communications in Statistics: Theory and Methods, 24(11):2865–2879, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Woodall:1992:NMS

[Woo92] William H. Woodall. A note on maximum Z-scores for control charts for individuals. Communications in Statistics: Theory and Methods, 21(11):3211–3217, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Weldon:1991:NRD

[WP91] K. L. Weldon and D. Potvin. Nonparametric recovery of duration distributions from cross-sectional sample surveys. Communications in Statistics: Theory and Methods, 20(12):3943–3973, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Williams:1992:AEM

[WP92] Paige L. Williams and Christopher J. Portier. Analytic expressions for maximum likelihood estimators in a nonparametric model of tumor incidence and death. Communications in Statistics: Theory and Methods, 21(3):711–732, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wright:1990:PPS

[Wri90] Tommy Wright. Probability proportional to size (πps) sampling using ranks. Communications in Statistics: Theory and Methods, 19(1):347–362, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wright:1998:PDF

[Wri98] Peter A. Wright. The probability density function of process capability index c_{pmk}. Communications in Statistics: Theory and Methods, 27(7):1781–1789, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Weissfeld:1991:MDG

[WS91a] Lisa A. Weissfeld and Susan M. Sereika. A multicollinearity diagnostic for generalized linear models. Communications in Statistics: Theory and Methods, 20(4):1183–1198, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Welch:1991:SQI

[WS91b] William J. Welch and Jerome Sacks. A system for quality improvement via computer experiments. Communications in Statistics: Theory and Methods, 20(2):477–495, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Weissfeld:1994:RAP

[WS94] L. A. Weissfeld and H. Schneider. Residual analysis for parametric models fit to censored data. Communications in Statistics: Theory and Methods, 23(8):2283–2297, 1994. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Walter:1999:CNN

[WS99] Gilbert G. Walter and Xiaoping Shen. Continuous non-negative wavelets and their use in density estimation. Communications in Statistics: Theory and Methods, 28(1):1–17, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Weissfeld:1991:CIC

[WSM91] Lisa A. Weissfeld, R. I. St. Laurent, and L. H. Moulton. Confidence intervals for the comparison of two Poisson distributions. Communications in Statistics: Theory and Methods, 20(10):307, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wu:1996:ANM

[Wu96] Dane W. Wu. Asymptotic normality of the multiscale wavelet density estimator. Communications in Statistics: Theory and Methods, 25(9):1957–1970, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wun:1991:RAA

[Wun91] Lap-Ming Wun. Regression analysis of autocorrelated Poissondistribution data. Communications in Statistics: Theory and Methods, 20(10):3083–3091, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wahba:1990:WOR

[WW90] Grace Wahba and Yonghua Wang. When is the optimal regularization parameter insensitive to the choice of the loss function? Communications in Statistics: Theory and Methods, 19(5):1685– 1700, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Watelet:1991:NEN

[WW91a] L. Watelet and B. B. Winter. Nonparametric estimation of nonincreasing densities and use of data from renewal processes. Communications in Statistics: Theory and Methods, 20(7):2073– 2094, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wilson:1991:ATC

[WW91b] Jeffrey R. Wilson and William D. Warde. An adjusted test for cross-classified cluster sample data. Communications in Statistics: Theory and Methods, 20(9):3029–3050, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wu:1999:LST

[WX99] Yanhong Wu and Yongxia Xu. Local sequential testing procedures in a normal mixture model. Communications in Statistics:

Theory and Methods, 28(8):1777–1792, 1999. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Wu:1994:OSH

[WZ94] Xishi Wu and Bojuan Zhao. Optimal sampling of hierarchical screening with inspection errors. Communications in Statistics: Theory and Methods, 23(3):803–814, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Xiaolong:1997:ETP

[XB97] Luo Xiaolong and James M. Boyett. Estimations of a threshold parameter in Cox regression. Communications in Statistics: Theory and Methods, 26(10):2329–2346, 1997. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Xue:1992:ONB

[XD92] Dixi Xue and James A. Deddens. Overdispersed negative binomial regression models. Communications in Statistics: Theory and Methods, 21(8):2215–2226, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Xiang:1995:SLL

[Xia95] Xiaojing Xiang. A strong law of large number for L-statistics in the non-i.d. case. Communications in Statistics: Theory and Methods, 24(7):1813–1819, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Xu:1990:RWA

[Xu90] Darning Xu. On random walks on affine group. Communications in Statistics: Theory and Methods, 19(8):2925–2942, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yatawara:1991:KFP

[YAM91] Nihal Yatawara, Bovas Abraham, and John F. MacGregor. A Kalman filter in the presence of outliers. Communications in Statistics: Theory and Methods, 20(5–6):1803–1820, 1991. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yamato:1997:GDT

[Yam97a] Hajime Yamato. A generalization of the Donnelly–Tavaré– Griffiths formula. Communications in Statistics: Theory and Methods, 26(8):2009–2019, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yamato:1997:DTG

[Yam97b] Hajime Yamato. On the Donnelly–Tavaré–Griffiths formula associated with the coalescent. Communications in Statistics: Theory and Methods, 26(3):589–599, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yang:1990:BPG

[Yan90a] Hu Yang. A brief proof on the generalized variance bound of the relative efficiency in statistics. Communications in Statistics: Theory and Methods, 19(12):4587–4590, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yang:1990:CBG

[Yan90b] Ming-Chung Yang. Compromise between generalized Bayes and Bayes estimators of Poisson means under entropy loss. Communications in Statistics: Theory and Methods, 19(3):935–951, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yang:1996:EMP

[Yan96a] Hu Yang. Efficiency matrix and the partial ordering of estimate. Communications in Statistics: Theory and Methods, 25(2):457– 468, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yang:1996:SAR

[Yan96b] Zhenlin Yang. Some asymptotic results on Box–Cox transformation methodology. Communications in Statistics: Theory and Methods, 25(2):403–414, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yang:1998:AMD

[Yan98] Ying Yang. Asymptotics of minimum distance estimator in linear regression models under strong mixing. Communications in Statistics: Theory and Methods, 27(7):1619–1637, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Young:1993:EEG

[YB93] David H. Young and Saad T. Bakir. Estimation efficiency of grouped and censored observation schemes for a generalised regression model. Communications in Statistics: Theory and Methods, 22(2):403–423, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yu:1994:CID

[YB94a] Qing-Ling Yu and Richard K. Burdick. Confidence intervals on the difference of failure intensities from two exponential distributions. Communications in Statistics: Theory and Methods, 23 (3):735–745, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yuan:1994:TLT

[YB94b] Ke-Hai Yuan and P. M. Bentler. Test of linear trend in eigenvalues of k covariance matrices with applications in common principal components analysis. Communications in Statistics: Theory and Methods, 23(11):3141–3156, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yu:1995:CIV

[YB95] Qing-Ling Yu and Richard K. Burdick. Confidence intervals on variance components in regression models with balanced (q −1)- fold nested error structure. Communications in Statistics: Theory and Methods, 24(5):1151–1167, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yi:1999:SDW

[YC99a] Seongbaek Yi and Byoung Seon Choi. Solving 2-D Wiener filtering problem using the block Cholesky decomposition. Communications in Statistics: Theory and Methods, 28(5):1059–1072, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yue:1999:CNS

[YC99b] Jack C. Yue and Murray K. Clayton. Comparing the number of species in two populations: group sampling case. Communications in Statistics: Theory and Methods, 28(2):467–477, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yeh:1992:NIG

[Yeh92] Lam Yeh. Nonparametric inference for geometric processes. Communications in Statistics: Theory and Methods, 21(7):2083– 2105, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yafune:1996:KLI

[YIK96] Akifumi Yafune, Makio Ishiguro, and Genshiro Kitagawa. Kullback–Leibler information approach to the optimum measurement point for Bayesian estimation. Communications in Statistics: Theory and Methods, 25(3):519–536, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yip:1991:CIM

[Yip91] Paul Yip. Conditional inference on a mixture model for the analysis of count data. Communications in Statistics: Theory and Methods, 20(7):2045–2057, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yip:1996:EPC

[Yip96] Paul S. F. Yip. Effect of plant-capture in a capture-recapture experiment. Communications in Statistics: Theory and Methods, 25(9):2025–2038, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yip:1998:UMM

[Yip98] Paul S. F. Yip. The use of a mixture model for estimating vaccine efficacy. Communications in Statistics: Theory and Methods, 27(12):3113–3119, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yuan:1991:EMB

[YJF91] F. Yuan, M. Jimbo, and R. Hara Fuji. An extension method for balanced arrays. Communications in Statistics: Theory and Methods, 20(3):1073–1085, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yum:1990:DLT

[YK90] Bong-Jin Yum and Sun-Ho Kim. Development of life-test sampling plans for exponential distributions based on accelerated life testing. Communications in Statistics: Theory and Methods, 19(7):2735–2743, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yuan:1997:BES

[YK97] Lin Yuan and H. K. Kesavan. Bayesian estimation of Shannon entropy. Communications in Statistics: Theory and Methods, 26 (1):139–148, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yu:1991:TSC

[YL91] P. L. H. Yu and K. Lam. Tightness of some confidence and predictive intervals related to selection. Communications in Statistics: Theory and Methods, 20(4):1401–1408, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yip:1992:CNP

[YL92] Paul Yip and K. F. Lam. A class of non-parametric tests for the equality of failure rates in a competing risks model. Communications in Statistics: Theory and Methods, 21(9):2541–2556, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yip:1993:MNT

[YL93] Paul Yip and K. F. Lam. A multivariate nonparametric test for the equality of failure rates in a competing risks model. Communications in Statistics: Theory and Methods, 22(11):3199–3222, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yoshida:1998:RPC

[YLB98] Olga S. Yoshida, José G. Leite, and Heleno Bolfarine. The roles of prior catchability assumptions and sample features on Bayes estimates of the number of classes in a population. Communications in Statistics: Theory and Methods, 27(8):1895–1904, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yadohisa:1999:VFR

[YN99] Hiroshi Yadohisa and Naoto Niki. Vector field representation of asymmetric proximity data. Communications in Statistics: Theory and Methods, 28(1):35–48, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yokoyama:1996:EGC

[Yok96] Takahisa Yokoyama. Extended growth curve models with random-effects covariance structures. Communications in Statistics: Theory and Methods, 25(3):571–584, 1996. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yong:1995:BKT

[Yon95] Zhou Yong. Bahadur–Kiefer theorems for kernel smooth product-limit quantile estimator. Communications in Statistics: Theory and Methods, 24(11):2815–2828, 1995. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Young:1992:IGO

[You92] K. D. S. Young. Influence of groups of observations on Bayes factors. Communications in Statistics: Theory and Methods, 21(5):1405–1426, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Young:1993:BML

[You93] David H. Young. Bias of maximum likelihood estimators of the response standard deviation in regression. Communications in Statistics: Theory and Methods, 22(7):1879–1891, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yu:1993:ABI

[YP93] Qiqing Yu and Eswar Phadia. Admissibility of the best invariant estimator of a distribution function under the Kolmogorov– Smirnov loss. Communications in Statistics: Theory and Methods, 22(8):2103–2124, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yuh:1994:REU

[YR94] Lianng Yuh and Stephen J. Ruberg. On robust estimation using individual ratios in bioavailability trials. Communications in Statistics: Theory and Methods, 23(2):545–554, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Young:1992:CID

[YSM92] Dean M. Young, John W. Seaman, Jr., and Laurie M. Meaux. A characterization of the independence-distribution-preserving covariance structure for the multivariate maximum squared-radii statistic. Communications in Statistics: Theory and Methods, 21(6):1605–1613, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yuan-Tsung:1995:TSJ

[YT95] Chang Yuan-Tsung. Two-stage James–Stein estimators of the mean based on prior knowledge. Communications in Statistics: Theory and Methods, 24(9):2211–2227, 1995. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yu:1991:NEM

[Yu91] Kai F. Yu. A note on the estimation of the mixing parameter in a mixture of two distributions. Communications in Statistics: Theory and Methods, 20(2):595–609, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yu:1997:UCR

[Yu97] Jun Yu. Uniform convergence rates for a nearest neighbor density estimator under dependence assumptions. Communications in Statistics: Theory and Methods, 26(3):601–616, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yu:1998:ESM

[Yu98] Kai Fun Yu. On estimating subpopulation means. Communications in Statistics: Theory and Methods, 27(1):235–273, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yuan:1992:CSB

[Yua92] Fuzhi Yuan. Characterization of singular balanced fractional s^{m} factorial designs derivable from balanced arrays with maximum number of constraints. Communications in Statistics: Theory and Methods, 21(7):1891–1907, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yue:1999:GTS

[Yue99] Jack C. Yue. Generalized two-stage bandit problem. Communications in Statistics: Theory and Methods, 28(9):2261–2276, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yang:1995:BPR

[YV95] Jim J. Yang and John W. Van Ness. Breakdown points for redescending M-estimates of location. Communications in Statistics: Theory and Methods, 24(7):1769–1787, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yu:1998:CSC

[YW98] Qiqing Yu and George Y. C. Wong. Consistency of self-consistent estimators of a discrete distribution function with bivariate right-censored data. Communications in Statistics: Theory and Methods, 27(6):1547–1556, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yanagimoto:1993:CSE

[YY93] Takemi Yanagimoto and Eiji Yamamoto. A criterion of sensitivity of an estimating function. Communications in Statistics: Theory and Methods, 22(2):451–460, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Yamamoto:1995:MLR

[YY95] Eiji Yamamoto and Takemi Yanagimoto. A modified likelihood ratio test for the mean direction in the von Mises distribution. Communications in Statistics: Theory and Methods, 24(10):2659–2678, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zacks:1991:IBT

[Zac91] S. Zacks. Inference based on Taguchi’s saturated designs. Communications in Statistics: Theory and Methods, 20(2):497–510, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zahn:1990:CCS

[Zah90] Douglas A. Zahn. Current challenges in statistics: large-lecture courses. Communications in Statistics: Theory and Methods, 19(11):4383–4418, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zhu:1998:MOO

[ZAW98] Wei Zhu, Hongshik Ahn, and Weng Kee Wong. Multipleobjective optimal designs for the logit model. Communications in Statistics: Theory and Methods, 27(6):1581–1592, 1998.

CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zhang:1993:THA

[ZB93] Ji Zhang and Dennis D. Boos. Testing hypotheses about covariance matrices using bootstrap methods. Communications in Statistics: Theory and Methods, 22(3):723–739, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zhang:1997:GCM

[ZB97] Jie Zhang and Dennis D. Boos. Generalized Cochran–Mantel– Haenszel test statistics for correlated categorical data. Communications in Statistics: Theory and Methods, 26(8):1813–1837, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zografos:1990:DSS

[ZFP90] K. Zografos, K. Ferentinos, and T. Papaioannou. Divergence statistics: sampling properties and multinomial goodness of fit and divergence tests. Communications in Statistics: Theory and Methods, 19(5):1785–1802, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zhang:1990:GDP

[Zha90] Nien Fan Zhang. Group delay, partial group delay and index-lag relationship for multidimensional processes. Communications in Statistics: Theory and Methods, 19(9):3137–3145, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zhang:1992:RAR

[Zha92] Zhiyi Zhang. The robustness of ANOVA with respect to interactions in some orthogonal block designs. Communications in Statistics: Theory and Methods, 21(1):233–240, 1992. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zhang:1993:CRM

[Zha93a] Ping Zhang. On the convergence rate of model selection criteria. Communications in Statistics: Theory and Methods, 22(10): 2765–2775, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zhao:1993:CPP

[Zha93b] Ming Zhao. Change-point problems in software and hardware reliability. Communications in Statistics: Theory and Methods, 22 (3):757–768, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zhang:1994:SAR

[Zha94a] Biao Zhang. Some asymptotic results for the residual maximum likelihood estimation of the parameters of a spatial process. Communications in Statistics: Theory and Methods, 23(6): 1617–1649, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zhang:1994:EPE

[Zha94b] Ping Zhang. On the estimation of prediction errors in logistic regression models. Communications in Statistics: Theory and Methods, 23(10):2881–2894, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zhang:1995:DCP

[Zha95a] Heping Zhang. Detecting change points and monitoring biomedical data. Communications in Statistics: Theory and Methods, 24(5):1307–1324, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zhang:1995:ARL

[Zha95b] Jian Zhang. Asymptotics of the run lengths of two control charts. Communications in Statistics: Theory and Methods, 24(9):2375– 2403, 1995. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zhang:1996:SBR

[Zha96] Jian Zhang. On the sample breakdown robustness of some nonparametric tests. Communications in Statistics: Theory and Methods, 25(6):1185–1198, 1996. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zhang:1998:NKD

[Zha98] Biao Zhang Biao Zhang. A note on kernel density estimation with auxiliary information. Communications in Statistics: Theory and Methods, 27(1):1–11, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zheng:1999:SHF

[Zhe99] Qi Zheng. Solution to the hazard function of some two-stage carcinogenesis models when normal cell growth is piecewise linear. Communications in Statistics: Theory and Methods, 28(8): 1921–1929, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zontek:1990:LAE

[ZK90] Stefan Zontek and Witold Klonecki. Limiting admissible estimators for variance components. Communications in Statistics: Theory and Methods, 19(7):2485–2507, 1990. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zheng:1994:SCE

[ZK94] Ming Zheng and John P. Klein. A self-consistent estimator of marginal survival functions based on dependent competing risk data and an assumed copula. Communications in Statistics: Theory and Methods, 23(8):2299–2311, 1994. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zhu:1991:CES

[ZL91] Li-Xing Zhu and Guo-Ying Li. A consistent estimator for the supremum of the projection index based on sliced inverse regression. Communications in Statistics: Theory and Methods, 20(4): 1211–1223, 1991. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zaman:1990:RSS

[ZM90] Arif Zaman and George Marsaglia. Random selection of subsets with specified element probabilities. Communications in Statistics: Theory and Methods, 19(11):4419–4434, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zheng:1993:ADM

[ZM93a] Qi Zheng and James H. Matis. Approximating discrete multivariate distributions from known moments. Communications in Statistics: Theory and Methods, 22(12):3553–3567, 1993. CO-DEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zheng:1993:SAP

[ZM93b] Qi Zheng and James H. Matis. Some applications, properties and conjectures for higher order cumulants of a Markovian steppingstone model. Communications in Statistics: Theory and Methods, 22(12):3305–3319, 1993. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zografos:1998:MDB

[Zog98] K. Zografos. On a measure of dependence based on Fisher’s information matrix. Communications in Statistics: Theory and Methods, 27(7):1715–1728, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zoubeidi:1994:OAS

[Zou94] Toufik Zoubeidi. Optimal allocations in sequential tests involving two populations with covariates. Communications in Statistics: Theory and Methods, 23(4):1215–1225, 1994. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zayed:1997:NDR

[ZQ97] Hany Zayed and Dana Quade. A note on the distribution of the resistance of rank tests. Communications in Statistics: Theory and Methods, 26(8):1867–1875, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zhou:1998:ERC

[ZR98] Xiao-Hua Zhou and C. A. Rodenberg. Estimating an ROC curve in the presence of non-ignorable verification bias. Communications in Statistics: Theory and Methods, 27(3):635–657, 1998. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zoroa:1990:CBC

[ZRM90] P. Zoroa, J. M. Ruiz, and J. Marín. A characterization based on conditional expectations. Communications in Statistics: Theory and Methods, 19(8):3127–3135, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zhang:1997:BEE

[ZS97] Shuanglin Zhang and Qiuying Sha. On the best equivariant estimator of covariance matrix of a multivariate normal population. Communications in Statistics: Theory and Methods, 26(8):

2021–2034, 1997. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zhao:1998:SMC

[ZS98] Hongyu Zhao and Terence P. Speed. Stochastic modeling of the crossover process during meiosis. Communications in Statistics: Theory and Methods, 27(6):1557–1580, 1998. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zhang:1990:IEP

[ZSW90] N. F. Zhang, G. A. Stenback, and D. M. Wardrop. Interval estimation of process capability index c_{pk}. Communications in Statistics: Theory and Methods, 19(12):4455–4470, 1990. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zhang:1999:NCR

[ZW99] Shunpu Zhang and Laisheng Wei. A note on the convergence rates for empirical Bayes estimators of parameters in multipleparameter exponential families. Communications in Statistics: Theory and Methods, 28(6):1273–1291, 1999. CODEN CST-MDC. ISSN 0361-0926 (print), 1532-415X (electronic).

Zhang:1999:ANE

[ZZ99] Shuanglin Zhang and Zhongguo Zheng. On the asymptotic normality for L_{2}-error of wavelet density estimator with application. Communications in Statistics: Theory and Methods, 28(5): 1093–1104, 1999. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).